A Bibliography of Papers in Lecture Notes in Computer Science (1997), Part 1 of 2

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 February 2014 Version 1.26

Title word cross-reference

((2^{n})^{m}) [843]. + [197, 2156]. 1 [205]. 12 [2283]. 2 [1255, 56, 2077, 291, 289, 1207, 1833, 1774, 2245, 290]. 3 [762, 104, 326, 295, 56, 681, 1212, 758, 929, 291, 754, 67, 289, 1742, 757, 305, 1770, 286, 742, 310, 751, 2259, 2295, 2278, 2244, 2289, 1745, 2283, 2286, 2263, 1766, 290]. 4 [56, 283]. 5 [2258]. 6 [2258]. << [2156]. β [959, 870]. χ^{2} [1510]. D [799, 1978, 1992, 2219]. E [812]. K [250, 1968, 1819, 143, 380, 248, 1702, 1341, 1706, 846, 848, 1844]. k = 3, 4, 5 [248]. l [848]. F_{2^{n}} [844]. μ [1130]. n [2072, 384, 143, 1702, 2217, 2171]. o [856]. O(n^{2}) [56]. O(n log_{2}(n)) [2083]. O(n log n)

[1577]. π [2078, 1672]. q [1283]. → [2245]. wL + wR [1284]. wL, wR [1284]. y^{2}x^{3} + bx^{2} mod n [1703]. Z [87]. Zm [204].

- [1819, 56]. -adic [1255, 2171]. -ary [1283]. -Axis [2258]. -bases [799]. -branching [2230]. -Calculus [2078, 1672, 1130]. -Categorical [2077]. -Categories [2072]. -Cells [959]. -Center [250]. -colorable [56]. -cube [143]. -D [2286]. -Dimensional [205, 1833, 1745, 1978]. -minimal [856]. -partitioning [143]. -Paths [1341]. -programs [1844]. -reductions [870]. -set [1968]. -Subiteration [2283]. -term [380]. -th [2217]. -trees [1706]. -unification [812]. -vertex [248]. -widths [384]. -wise [846].

1 [505, 1837, 1580]. 1-Mismatch [1580].

1-Way [1837]. 1024 [665]. 1st [2297].

2 [64, 1582]. 2-Opt [1582]. 2-Visibility [64]. 2.5D [299]. 2dmin [1284]. 2nd [2297].

3 [158]. 3-Trees [158]. 3.0 [1757, 1789].

5 [2093].

6R [1146].

81/2 [2116].

95 [1128, 1188, 1196, 1186, 2386, 1204, 1197].

a-D [2288]. A0* [595]. A5 [835]. Abdominal [413, 297]. Abduction [2342]. Abductive [520]. ABEL [1097]. Ability [697]. Absolute [2256]. Absorption [767]. Abstract [9, 783, 73, 1294, 2114, 1339, 45, 2215, 2327, 274, 403, 1241, 818]. Abstraction [1503, 170, 1966, 2349, 2119]. Abstractions [1879, 1230, 656, 1627, 1628]. AC [801]. ACACIA [2035]. Academic [435]. Acceleration [1735]. Acceptance [1832, 927]. Accepted [1727]. Accepting [2198, 544]. Access [1041, 1363, 1684, 1593, 502, 1302, 1646, 1659, 301, 1949, 1964]. Accessible [1639]. accessing [1982]. Accounting [918, 1359]. Accumulators [851]. Accuracy [1671, 323, 1791]. Accurate [971, 1805]. acetabular [321]. Achieve [2156]. Achieving [242]. ACLP [1476]. Acoustic [739, 354]. Acoustic-labial [354]. Acoustics [1002]. Acquiring [412]. Acquisition [1350, 609, 2375, 414, 1365, 2029, 2019]. across [1209]. Acting [103]. Action [782, 894, 93, 1932, 1674, 1908, 2229]. Action-Based [782]. Actions [520, 1954, 1942]. Activated [734]. Activation [698]. activations [1212]. Active [1563, 277, 755, 1535, 2324, 276, 2264, 2166, 1032, 9, 10, 306]. Activities [1047, 652, 1963]. Activity [959, 1808]. ACTL [1253]. Actor [1921]. Actors [100]. ACTS [1051, 1047, 1053]. Acute [413]. Acyclic [1686, 1822, 1446, 58]. Ada [1187, 1128, 1195, 1201, 1189, 1203, 1188, 1196, 1202, 1186, 1191, 2386, 1204, 1192, 1197, 1194]. Ada-Based [1191]. Ada-specific [1189]. Adaptable [928]. Adaptation [906, 703, 1641, 1495]. Adapted [410]. Adaptive [1385, 465, 2007, 825, 2034, 696, 2316, 591, 928, 588, 2268, 1658, 714, 2008, 1566, 712, 2297, 1976, 1221]. Adding [1075, 18, 929]. Addition [653, 2251]. Additional [864, 388]. additive [1985]. Additives [2068]. adequate [399]. adic [1255, 2171]. adjacent [319]. Adjusted [719]. Admissibility [1487]. Admissible [866, 2058]. Adoption [892, 1641]. Adopts [990]. Advanced [572, 1617, 1555, 1648, 1516]. Advances [348, 285, 1797]. Adversaries [2165, 2166]. Advice [2226]. Advocating [1966]. AFAN [983]. Affine [1711, 2263]. Affinity [742, 284]. affinity-based [284]. after [326, 322]. AG [500]. Against [2165, 825, 1704, 2166, 824]. Agencies [243]. Agency [2035]. Agent [97, 567, 90, 894, 244, 92, 242, 85, 1638, 575, 897, 565, 164, 570, 389, 2033, 573, 896, 568, 89, 895, 576, 2037, 162, 84, 578, 2030, 2042, 2040, 2036, 2035, 903, 902, 350, 620, 2031, 236, 1621, 2038, 571, 233, 900, 94]. Agent-Based [389, 233]. Agents [99, 238, 101, 95, 428, 235, 893, 1065, 580, 569, 586, 392, 161, 96, 2034, 93, 590, 84, 898, 566, 393, 1555, 1649, 2032, 160, 891, 564, 86, 1530, 1438, 241, 1949]. Aggregates [2087]. Aglets [1649]. Agreement [833, 1552]. AI [239, 1813, 26]. AI-Systems [26]. aid [299]. Aided [1189]. Air [1491, 112, 688, 1630]. Airborne [2315]. Aircraft [733]. Airspace [457]. Airways [763]. aitch [611, 1813]. Ajtai [2152]. AL220 [692]. Alarm [1136]. Alarms [1012]. Algebra [1311, 110, 206, 220, 859, 121, 1325, 197, 788, 1560, 1163]. Algebra-Based [1311]. Algebraic [483, 1372, 2365, 1273, 2311, 2104, 867, 499, 462, 1277, 1814, 1832, 157, 2215, 1966, 2359, 1085, 1286, 1078, 383]. Algebras [2307, 1380]. Algorithm [2303, 258, 2083, 1417, 753, 109, 979, 1687, 629, 1757, 2235, 1291, 1272, 570, 2027, 986, 256, 1540, 1517, 645, 1590, 65, 68, 411, 1498, 1733, 1510, 2098, 636, 2020, 1232, 1725, 1515, 2011, 1514, 2010, 2160, 2283, 2008, 1652, 1608, 2170, 1580, 1765, 2285, 719, 2023, 1423, 1723, 1389, 640, 1146, 1000, 1721, 137, 1972, 402, 1978, 2182]. Algorithmic [2186, 262, 189, 863, 2318, 2220, 193]. Algorithmics [2073]. Algorithms [1771, 720, 690, 1780, 721, 1227, 1695, 211, 673, 1269, 475, 2164, 1399, 246, 723, 1329, 1607, 643, 2047, 1312, 2309, 114, 1711, 2307, 970, 2007, 736, 689, 1447, 207, 1891, 593, 684, 1231, 121, 248, 152, 259, 525, 1705, 61, 63, 935, 184, 711, 344, 699, 1327, 2169, 460, 1307, 1830, 1589, 1499, 1709, 1704, 1775, 1783, 1576, 2219, 1651, 2313, 843, 1531, 1588, 1875, 1715, 2017, 2232, 2003, 1876, 1767, 732, 2023, 1734, 135, 463, 381, 1982, 11, 1991, 1997, 815, 812]. Aliasing [1198]. Aligned [932]. Aligning [1449, 1711]. Alignment [1452, 1803, 1805, 1448]. All-Optical [1321]. All-or-Nothing [1528]. Alleged [835, 834]. Allegories [2073]. Allocation [650, 150, 2091, 1760, 1232, 1860, 21, 1995]. Allometry [960]. Almost [188, 846]. Alogtime [2047]. Alphabet [1724]. ALT [1777, 1790]. ALT-models [1777]. Alternating [1727, 2180, 1776, 1455]. Alternative [1195, 461, 769, 973, 828]. Alternatives [909, 1118, 1221]. ALU [1803]. ALWAN [648]. always [1456]. Amalgamated [2036]. amidst [1277]. AMN [442]. Among [238]. Amp [1398]. Amplification [841, 2166]. Amplifications [43]. Amplitude [770]. AMUSE [1068]. Analog [730]. Analog-Digital [730].

Analogical [1506, 37, 1504, 1100]. Analogue [2054]. Analogy [1168]. Analyses [933, 1498, 1551]. Analysing [1137]. Analysis [994, 964, 76, 934, 1543, 1772, 1922, 1134, 1195, 1756, 2181, 216, 1757, 1324, 428, 677, 740, 1731, 1176, 1175, 2253, 1666, 2356, 337, 1044, 930, 1858, 2033, 232, 1540, 82, 606, 341, 659, 78, 676, 1853, 1302, 515, 1303, 2314, 1884, 951, 731, 1880, 1809, 263, 2254, 1801, 944, 1885, 1622, 2295, 1744, 1886, 1860, 1815, 1717, 2288, 1883, 1892, 1864, 1865, 1899, 1889, 1874, 1751, 1861, 2240, 1765, 1862, 1818, 786, 1435, 1353, 2293, 640, 463, 1863, 9, 1219, 395, 1246, 1076, 1122, 283, 328, 1267, 1224, 1997, 2003]. Analytic [1356]. analytical [1121, 322]. ANALYZE [288]. Analyzer [77]. Analyzing [709, 446, 986, 582, 2340]. Anatomical [741, 304]. Ancient [1588]. Aneurysms [761, 297]. Angiograms [731, 295, 291]. Angiography [2247, 769, 761, 291]. animal [320, 300]. Animated [99, 101]. Animating [1166]. Animation [100]. Animator [436]. anisotropic [1220, 1225]. Anisotropy [2295]. ANN [728]. ANN-based [728]. Anna [1193]. Anna/Ada [1193]. Annealing [596]. Annotated [1795, 1862]. Annotation [1795, 1049]. Announcement [952]. Annuli [1604]. Annulus [1603]. Anonymous [1822, 2173, 826]. Answer [87]. Answering [1493, 860, 1471]. Antenatal [972]. Antennas [631]. Anthropological [244]. Anticipation [573]. Anticlique [145]. Any [838]. Anytime [1109, 1110]. aortic [297]. APEmille [675]. Aperture [2274]. Appear [1017]. Appfications [1596]. Applicability [614]. Application [1177, 1010, 1879, 702, 1615, 1007, 2332, 985, 930, 416, 999, 649, 705, 1145, 163, 452, 648, 1759, 2092, 1196, 2248, 1390, 1743, 47, 1652, 1602, 1661, 1803, 279, 2023, 2071, 2041, 658, 733, 1076, 312]. Application-Level [649, 1196].

Applications [1060, 167, 584, 661, 677, 655, 685, 2188, 632, 1813, 501, 68, 1320, 197, 2057, 1067, 1202, 687, 1589, 1658, 176, 1319, 1915, 2259, 846, 1024, 1647, 2236, 1605, 133, 1045, 1722, 1139, 1393, 318, 384, 381, 1977, 1941, 2001]. Applicative [489]. Applied [420, 511, 1585, 2291, 678, 700, 967, 415, 1956]. Applying [736, 1364, 1480, 459]. Approach [1920, 683, 1366, 1141, 974, 2365, 1089, 1088, 1502, 345, 1030, 1090, 2321, 540, 37, 442, 1665, 947, 938, 189, 1317, 1182, 474, 1406, 1173, 602, 648, 1308, 1337, 1104, 1105, 1481, 2108, 2042, 757, 1760, 1392, 1109, 2129, 195, 1915, 2310, 1644, 2024, 1966, 1882, 2089, 1864, 1865, 1808, 1803, 2123, 488, 2266, 1566, 1485, 1734, 1642, 763, 1863, 1894, 1895, 1896, 1226, 1543, 8, 1264, 1222, 1465, 324, 1237, 2383, 290, 2004, 1455]. Approaches [509, 66, 2234, 304]. Approaching [243]. Approximate [745, 1584, 465, 933, 1984]. Approximating [268, 1955, 2002, 1973]. Approximation [113, 1688, 1447, 248, 1337, 1335, 1232, 1282, 1576, 1608, 1987]. Approximations [140, 107, 1671, 2345, 1629, 1931, 1994, 1998]. Approximators [715, 1221]. Arachne [173]. Arbitrary [2245, 1732]. arc [1972]. archimedean [2253]. Architectural [446, 486, 496, 1648, 2284, 1938]. Architecture [1391, 669, 1064, 1563, 1099, 1065, 985, 2012, 908, 432, 973, 389, 1785, 1027, 1637, 926, 2374, 1046, 917, 1024, 571, 1939, 1944, 1941, 179]. Architectures [842, 1768, 503, 698, 359, 1762, 236, 1777, 843, 1945]. Arcs [1558, 2233, 1077]. Area [1970, 57, 1601, 1652, 733]. Area-efficient [1970]. Argument [1096, 368]. Argument-Based [1096]. Argumentation [424, 1505]. Arguments [838]. Arithmetic [2172, 843, 2170, 1260, 798, 1081]. Arity [1381]. Arrangement [1632, 1575]. Array [947, 933, 544, 1328, 1745]. Arrays [943].

Art [450]. artery [299]. arthroplasty [325, 322]. arthroscopic [309]. Arthur [2192]. articular [286]. Artifacts [753]. Artificial [994, 1007, 553, 1016, 431, 992, 1013, 698, 968, 1397, 1011, 1015, 2024, 1883, 160]. ary [1283]. Ascription [244]. Ash [634]. ASICs [1237]. Asked [449]. Aspect [57, 1026]. Aspect-Oriented [1026]. Aspects [1350, 922, 1381, 2220, 1471]. Assembly [1450, 1597]. Assertion [495]. assertional [1261]. Assertions [272]. assessing [326]. Assessment [1086, 1416, 1881]. Asset [1065]. Assigning [1492]. Assignment [935]. Assistance [417]. Assistant [1153, 1855]. Assisted [991, 419, 312, 325, 324, 303, 300]. Associate [1311]. Association [1428, 2335]. Associations [1030]. Associative [2307, 451]. Assumption [1097]. Assumption-Based [1097]. Assumptions [254, 1938]. Asteroidal [1338]. Astrid [671]. Asymetrical [1001]. Asymmetric [203]. Asynchronous [1920, 901, 1822, 51, 81, 936, 2189, 1958, 1950, 1964]. ATLAS [172]. ATLAS-I [172]. Atlases [429]. ATM [1041, 1322, 1126, 920, 1067, 1121, 686, 687, 909, 172, 1735]. ATM-based [1041, 686, 687]. Atmospheric [633]. Atomic [629]. ATP [1145]. ATR’s [1397]. Attack [2159, 2180, 1512, 2162, 2160]. Attacks [819, 823, 2163]. Attention [95, 595]. Attribute [615]. Attributed [1836, 72]. Attributes [615]. Auctions [1704]. Audio [355, 1137]. Audio-visual [355]. Audio/Video [1137]. Augmented [298, 2259, 318, 311, 332]. Augmenting [1180]. Authenticated [833]. Authentication [1526, 351, 356, 353, 1670, 1525, 340, 2167, 338, 1550]. Author [853, 1040, 2183]. Authoring [1050]. authorisation [1543]. Authority [827]. autoepistemic [1458]. Autofocus [753]. Automata [222, 229, 1399, 1400, 216, 500, 204, 264, 2022, 777, 1618, 212, 776, 187, 474, 1773, 535, 1401, 1727, 1397, 1402, 198, 228, 1326, 2212, 183, 217, 545, 1833, 2220, 1840, 2371, 1625, 1846, 1847, 1750, 1629, 1753, 779, 1457, 1076, 1247, 776]. Automata-Theoretic [474]. Automated [1903, 650, 1143, 221, 413, 1480, 2040, 742, 1673, 2259, 741, 1139, 301]. Automatic [2303, 889, 753, 345, 342, 536, 572, 983, 117, 363, 358, 1193, 2131, 344, 765, 286, 1242, 2139, 49, 338, 308, 328, 1267]. Automatically [480, 940, 660, 802]. Automating [497, 1866, 437]. Automaton [205, 2218]. Autonomous [101, 894, 95, 1638, 84, 1634, 898, 1387, 393, 2318, 2238, 1389, 241]. Autonomous-Objects-Based [1634]. Available [1556]. Average [1815, 1842, 2003]. Average-Case [1815]. Avionics [1617]. Avoid [1148]. Avoidance [171]. Avoided [1198]. Avoiding [664]. Aware [583, 891]. Awareness [2137]. Axiom [2070]. Axiomatic [1376, 1915]. axiomatisation [808]. axiomatization [398]. Axiomatizations [1325]. Axis [2258, 1999]. axis-parallel [1999].

B [1227, 1645, 442, 299, 912]. B-ISDN [912]. B-spline [299]. B-tree [1645]. Back [1536]. Background [1094, 2140, 412]. Backgrounds [744]. Backpropagation [979, 716]. backward [1453]. Bacterial [2019]. Bad [2168, 1252]. Badly [2267, 2257]. bag [22]. bags [1074]. Balance [959, 1609, 1993]. Balanced [265, 1593, 261, 1652]. Balancing [1579, 1233, 1231, 645, 176]. Ballast [1906]. Balloon [1018]. Banded [106]. BANG [1899, 1137]. BANG-Clustering [1899]. Banishing [1451]. Base [242, 861, 696, 1497, 1104, 1782, 318]. Based [887, 1662, 618, 721, 641, 722, 1177, 1096, 1501, 2106, 993, 1366, 1097, 782, 679, 838, 1089, 1502, 1311, 1789, 95, 1136, 1503, 522, 91, 2096, 2172, 434, 1422, 1420, 1478, 1450, 2022, 1409, 432, 1703, 2308, 610, 981, 975, 363, 430, 389, 839, 2276, 1493, 168, 492, 1048, 718, 696, 940, 507, 497, 1507, 589, 992, 755, 1508, 1425, 514, 535, 2189, 1479, 74, 1480, 1634, 1009, 1365, 601, 2154, 1759, 1397, 1866, 100, 1481, 1557, 1572, 240, 1742, 231, 1382, 1390, 1174, 2035, 1482, 419, 1191, 996, 1487, 1505, 2254, 944, 2026, 2138, 2179, 1621]. Based [2318, 2272, 2238, 2028, 1644, 2289, 2162, 1524, 1651, 1784, 1855, 2009, 2158, 2063, 1720, 2229, 2319, 1678, 1864, 1865, 1899, 1660, 2267, 279, 488, 88, 1635, 1484, 875, 1485, 767, 2023, 233, 1509, 1389, 1931, 463, 1194, 2163, 421, 1550, 326, 1551, 298, 1255, 1238, 1125, 312, 296, 317, 1119, 1540, 328, 383, 907, 284, 336, 1963, 815, 338, 293, 12, 1547, 1041, 1756, 292, 728, 686, 687, 1149, 1646, 136]. Bases [1370, 1104, 866, 2036, 700, 2059, 1889, 799, 817]. Basic [988, 167, 1502, 110, 1318, 559, 588, 1015, 820, 1551, 2382]. Basis [2073, 691, 2356]. Batch [702]. Batches [617]. Bayes [964]. Bayesian [1213, 744, 351, 770, 576, 2278, 2280, 1901, 2266, 2237]. BDDs [668, 774, 1252]. Be [85, 539, 87, 1198, 1029, 2353]. bear [1998]. Beat [992, 1697]. before [326]. begin [2123]. behaved [791]. Behavior [735, 100, 1015]. Behavior-Based [100]. Behaviors [99, 1352]. Behaviour [1183, 494]. Behaviour-Refinement [494]. Behavioural [1852]. Belief [1099, 89, 895, 1093, 1104, 1110]. Belief-Function [89]. Beliefs [90, 1107, 1458]. Bends [1599, 1601]. benefits [1939]. BeoLink [1137]. Berlinda [1965]. Best [1513, 1227, 1059]. Best-Effort [1059]. Best-First [1227]. beta [408]. Bets [42]. Better [211, 208]. Between [955, 959, 989, 857, 963, 1005, 754, 2345, 1220, 1457, 2203, 2055, 1433, 1642]. Beyond [618, 1039, 2269]. bi [400]. bi-domains [400]. Bias [1490, 1451]. Bibliography [448]. Bicategory [1675]. Biconjugate [1228]. Bidding [1704]. Bidirectional [938, 2254]. Bifurcation [131]. Big [1343, 1856]. Big-Oh [1856]. Billing [918]. Binarization [2312]. Binary [2199, 144, 1541, 1444, 1827, 477, 2282, 1106, 1915, 2217, 193, 1931, 2248]. Binding [273, 828]. Bindings [1187]. Bins [1608]. Bio [1000]. Bio-inspired [1000]. Bioinformatics [1794, 1806]. Biological [1795, 990, 427, 997, 1001, 1389, 1808]. Biologically [975]. Biology [452, 1394, 1810]. Biomedicine [1806]. Biometric [339, 352]. Bipartite [52, 1336, 1740, 257, 1319, 1575]. Biplanar [1575]. Bisection [208]. Bisimulation [481, 1317, 1323, 191, 1243]. Bisimulations [1157]. bit [1240]. bit-vectors [1240]. Bits [837]. Bitwise [1783]. Bivariate [132, 1429]. Black [261]. blackboards [1949]. Blind [2155]. Block [1523, 1512, 1524, 1514, 1547]. blocks [1123]. Blood [993, 750]. Boards [1610]. bodies [319]. Boehm [1933]. Boiler [1010]. Boltzmann [644]. Bolu [714]. BONAPARTE [1067]. bone [282]. BONITA [1964]. Boolean [265, 1255, 668, 840, 1151, 848]. Borel [1297]. Bottleneck [1495]. Bottom [522, 1150]. Bottom-Up [522, 1150]. Bound [1313, 208, 1723, 382]. Boundary [113, 255, 754, 133]. Bounded [140, 473, 1074, 1342, 1291, 186, 1299, 2165, 1824, 1975, 539, 1375, 1829, 1307, 2059, 267, 1971, 1976]. Bounded-Variable [1375]. Bounding [2052, 1412, 1126]. Bounds [977, 1295, 1607, 1740, 65, 1830, 125, 558, 2191, 2192, 2194, 1849, 836, 2071, 1548]. boxes [373]. Brain [31, 954, 758, 431, 755, 1397, 429, 756, 741, 759, 1210, 314, 315, 313, 1397]. Brainstorming [1643]. Brainware [1391]. Branch [1723]. Branch-and-Bound [1723]. Branches [874]. Branching [1300, 2214, 1844, 379, 2230]. Branchwidth [1329]. Breadth [182]. Breadth-First [182]. Breaking [2033]. Breakout [1610]. Breakpoints [1710]. Brief [1092, 1281, 532]. Broadband [920]. Broadcast [1052]. Broadcasting [1824, 192, 150, 1976]. bronchoscopy [312]. Browser [167, 27, 1165]. Browsing [1425]. Bruijn [208]. BSP [1312, 657, 636]. Bucket [830]. Build [1928]. Building [166, 1502, 1872, 902, 1763, 1123]. Bulk [110]. Bumpless [1612]. Bunch [2290]. bundles [1207]. Business [1185, 1172]. Butterflies [1736]. Butting [2378]. Button [1382]. BVP [124]. Bytes [1527].

C [1187, 1789, 504, 280, 670, 217, 1404]. C/ E [217]. C98 [671]. Cache [943, 930, 664, 663, 1784, 1650]. Cache-Coherence [664]. Cache-Coherent [943]. Caches [1235]. CAD [2004]. Calculating [1541]. Calculation [108, 2311, 131]. Calculator [1166]. Calculi [1911, 1304, 1932, 1380, 157, 1908, 406, 408]. Calculus [2367, 1315, 2078, 2202, 2104, 489, 1172, 2189, 1560, 1672, 1158, 2191, 2087, 1917, 2344, 2360, 2050, 226, 1071, 1266, 397, 401, 1080, 404, 1130, 1959, 870]. Calibrated [2257]. Calibration [2322, 2258, 2259, 2256]. Call [1140, 1303]. Call-by-value [1303]. CAM [1397]. CAM-Brain [1397]. Cambridge [1062]. Camera [2322, 2258, 2321, 2259, 2257]. Cameras [2263]. Can [519, 365, 224, 1198, 1029, 1692, 2353, 94]. Candidate [2352]. CANM [105, 108]. Canonicity [1348]. Canonization [2047]. Capabilities [1191]. Capability [28]. Capacity [111, 557, 961]. Capture [1825]. Carbohydrates [1806]. Carcinogenesis [2351]. Cardiac [740, 751]. cardinality [1994]. cardiopulmonary [310]. Care [916, 411, 1862]. Carlo [593, 121]. Carriers [2298]. Carroll [1302]. Cartographic [1577]. Case [618, 1489, 1501, 1298, 1350, 1490, 882, 2164, 1502, 600, 1491, 1757, 1503, 434, 1422, 1478, 1493, 1497, 1507, 1179, 1508, 1479, 1480, 418, 1759, 1481, 917, 437, 1482, 419, 1872, 1505, 1815, 2221, 2079, 1488, 1484, 1509, 900, 1944, 1946, 1945, 2003, 1179, 1180]. Case-Base [1497]. Case-Based [618, 1501, 1502, 1503, 434, 1422, 1478, 1493, 1507, 1508, 1479, 1480, 1481, 1482, 419, 1505, 1484, 1509]. Case-Study [882]. Cases [1500, 1492, 235, 1494, 1495, 1484]. Cash [41, 40]. Catalogues [1562]. Categorical [2077, 2079, 399]. Categories [2072, 2080, 1923, 1874, 406]. Categorization [32]. Category [2197]. Catenation [1835]. Causal [897, 420]. Causal-Functional [420]. Cause [1351]. caused [314]. Cautious [2329]. CAVE [360]. CBR [1491, 24, 1483, 1487]. CCD [2315]. CD [547]. Cell [1735]. Cells [959, 428]. Cellular [1771, 204, 264, 2022, 2276, 205, 1773, 1397, 1770, 2218, 2220, 1777, 1840, 1790, 1747, 1750, 1753, 1457, 2383, 2182]. Cellular-Neural [1747]. Center [250, 2236]. Center-Surround [2236]. Central [107]. Centre [333]. Centred [1173]. Centroid [1602]. Cerebral [754, 750, 761, 1212]. Certain [549, 1278, 554, 1743]. Certainty [994]. Certificates [266]. Certification [39, 1248]. Certified [1928]. CFD [510]. CG [1368, 1356]. CGLex [1366]. CGLS [640]. Chain [593, 1275]. Chaining [700, 1453]. Challenge [1405]. Challenges [2373, 976, 1656, 1386]. Chameleon [1519]. Change [874, 883, 1101, 1107, 875, 2266, 1114, 387]. Change-Based [875]. Change-Oriented [883]. Changes [955, 739, 697, 754]. Changing [1483]. Channel [1052, 490, 1324]. Channels [839, 1256]. Chaos [264]. Chaotic [1288, 969, 2021]. Char [1139]. Char-Set [1139]. Character [2308, 103]. Characterisation [2376, 431].

Characterisations [1852]. Characteristic [1513, 2023, 1218]. Characterization [551, 141, 481, 252, 547, 1322, 2215, 1724, 2204, 2306, 1125, 1085]. Characterizations [1333]. Characterize [496]. Characterizing [756, 1212]. Charging [918]. Charts [534, 12]. Cheating [1704]. checker [1265]. Checkers [845, 1545]. Checking [527, 1316, 475, 822, 1315, 1301, 1586, 538, 1129, 75, 476, 1938, 1918, 191, 1931, 1249, 1268, 1266, 1076, 1260, 1257, 1251, 1254, 1261]. Chemical [109, 262]. Cheminformatics [1794]. Chervonenkis [372]. Children [413]. Chip [172]. Choice [989, 534, 635, 769]. Chordal [207, 1728, 1973]. Chromatic [1335, 249]. Chromosome [1802]. chrysanthemum [1216]. Church [804]. Cinema [102]. Cipher [1519, 1533, 1523, 835, 1524, 1547]. Ciphering [1517]. Ciphers [1512, 1522]. Circal [220]. Circle [131, 1715]. Circuit [2054, 1610]. Circuits [2195, 1692, 379]. Circular [59, 1444]. Citizens [1806]. CL [394]. CLAM [178]. Class [776, 1813, 966, 1277, 231, 1619, 1622, 1623, 1636, 1806, 1257, 1548]. Classes [210, 252, 1825, 154, 1682, 1286]. Classical [1922, 407]. Classification [618, 1854, 977, 1857, 971, 991, 194, 1160, 612, 990, 978, 992, 621, 1866, 32, 1900, 2347, 135, 1216, 1213]. Classifier [421]. Classifiers [1885]. Clausal [2334]. Clause [1142, 2340]. Clause-Diffusion [1142]. Clauses [616]. ClearCase [882]. Climbing [458]. Clinical [410, 418, 411, 320, 305]. Clique [145]. Clocks [80, 1078]. Cloning [2034]. close [2000]. closed [804]. Closest [2285]. closet [1936]. Closing [155]. Closure [1448, 148, 1254, 793]. Clover [1534]. CLP [368]. Cluster [177, 647, 168, 1855]. Cluster-Based [1855]. Clustered [60, 61]. Clustering [1432, 2336, 704, 1898, 2219, 2296, 1882, 2236, 1899, 1803, 1887]. Clusters [175, 654, 1431, 710, 1784].

Clusters-Based [1784]. CM [2093]. CM-5 [2093]. CMIP [921]. CMIP/SNMP [921]. CNAPS [982]. CNNs [981]. Co [1769, 2029, 2018, 1964, 2214, 1848]. Co-evolution [2029]. Co-evolutionary [2018]. co-NP [2214, 1848]. co-ordination [1964]. Co-processor [1769]. Coalgebraic [1317, 494]. Coarse [1312, 651]. Coarse-Grain [651]. Coarsening [940]. Cocycles [1280]. Code [562, 510, 1201, 938, 582, 506, 935, 588, 2131, 563, 671, 630, 1029, 1550, 396, 1154]. Code-Data [1201]. Codes [1269, 1271, 1291, 1274, 1275, 1278, 1279, 1280, 1283, 1284, 2179, 1817, 1285, 1286]. codesign [1237]. Coding [1449, 705, 998, 2275, 2019, 308]. Coevolving [2012]. Coexistence [655]. Cognitive [974]. Coherence [1089, 664, 2270]. Coherence-Based [1089]. Coherent [943]. Coincide [1560, 6]. Coinductive [398, 1157, 494]. coli [1798]. Collaborating [564]. Collaboration [2035]. Collaborative [679, 1046, 574]. Collage [1826]. Collecting [1990, 301]. Collection [710, 1862]. Collections [1425, 2087, 1877]. Collective [96]. Collision [851, 2178, 831]. Collision-Free [851, 831]. Collision-Resistant [2178]. Collisions [1025]. Collusion [203]. Collusion-Secure [203]. Colonies [550]. colonoscopy [331]. Color [2246, 1336, 1478, 2316, 2236, 2016, 598, 308]. Color-Segmentation [2236]. colorable [56]. Colored [1158]. Coloring [1319, 1996]. Coloured [1138, 1820, 1137, 533, 1135, 532]. Colouring [1320]. Combination [1141, 1088, 617, 815, 812]. Combinations [1785, 123, 805]. Combinatorial [1771, 1696, 820]. Combinators [1167, 813]. Combined [1504, 1863]. Combining [1141, 1257, 277, 526, 615, 1356, 1930, 2082, 1112, 2274, 1259]. Commands [1624]. Commerce [39, 927]. Commitment [391].

Communicating [783, 552].

Communication [1313, 2185, 1236, 943, 567, 783, 780, 642, 586, 577, 168, 896, 170, 93, 656, 645, 187, 578, 1308, 662, 178, 2190, 1943, 944, 942, 1950, 1948, 1990]. Communication-Minimal [943, 942]. Communication-passing [1943]. Communications [1768, 660, 1722]. Commutation [472]. Commutative [1834, 799]. Commute [2086]. Commuting [1331]. CoMoMAS [2030]. Comp.specification.z [449]. Compact [1446, 1720, 330]. Compactly [132, 207]. Comparative [994, 288, 1666, 1668, 1005, 1710, 1683]. Comparing [454, 603, 1924, 217, 2061, 1629, 2251, 875]. Comparison [988, 487, 177, 417, 2047, 317, 44, 758, 2043, 595]. COMPASS [495]. Compatible [2136]. COMPEL [1801]. Compensated [2247]. Compensation [1557]. Compensation-Based [1557]. Competition [963, 459]. Competitive [913, 2029, 1997]. Compilation [949, 2116, 272, 2129, 948, 493, 1081]. Compiler [946, 246, 273, 935, 271, 1966]. Compilers [941, 659, 1748]. Compiling [947]. Complementarity [232]. Complete [144, 1858, 1322, 787, 1302, 1809, 1356, 1724, 2219, 1848, 808, 4, 1375]. Completeness [251, 1718, 2064]. Completion [1291, 2277, 793, 792]. Complex [956, 735, 1792, 860, 968, 623, 126, 130, 1476]. Complexity [1313, 2185, 2052, 1443, 265, 2199, 2054, 210, 552, 186, 206, 252, 1825, 262, 860, 450, 862, 187, 1813, 1827, 1698, 2145, 2326, 1130, 155, 869, 560, 2190, 558, 2213, 201, 1834, 2234, 1594, 2196, 1842, 2194, 1817, 1850, 387, 2083, 1083]. Compliant [917]. Component [910, 748, 1637, 2211, 528, 515, 2240, 1635, 1938, 1959]. Component-Based [1635]. Component-Oriented [1637]. Components [725, 750, 592]. Componentwise [125]. Comportment [2110]. Composable [2361, 1921]. Composed [1777]. Composing [1469]. Composite [1693]. Composition [1509]. Compositional [1103, 1239]. Compositions [370]. Comprehensible [495]. Comprehension [865]. Compressed [1443, 1055]. Compressible [1778]. Compressing [1339]. Compression [627, 356, 588, 2268, 372]. Computability [1701, 1297]. Computable [1699, 194, 1623, 1816, 2232, 1700]. Computation [1448, 551, 762, 988, 522, 2366, 579, 939, 1741, 523, 452, 2124, 1314, 1870, 1011, 1303, 2190, 1403, 1640, 1778, 1749, 1750, 2360, 1871, 14]. Computational [1295, 552, 206, 1570, 2012, 506, 1585, 701, 254, 558, 1381, 1622, 1914, 2234, 2192, 2061, 1435]. Computations [1739, 2355, 2200, 635, 118, 1140, 187, 2212, 1743, 1790, 1692, 2121, 1753, 1791, 405]. Computed [763]. Computer [171, 675, 1795, 467, 529, 2005, 533, 1189, 2056, 325, 982, 764, 419, 1648, 1659, 311, 312, 324, 321, 2387, 300]. Computer-Annotated [1795]. Computer-Assisted [419, 312, 300]. Computer-Generated [2056]. Computer-Generation [1795]. computer-vision-enabled [311]. Computerised [1178]. Computers [1756, 110, 633, 515, 625, 1229, 1754, 639]. Computing [166, 177, 1342, 1599, 600, 572, 627, 1464, 1657, 1228, 857, 500, 1578, 1740, 180, 1792, 2242, 168, 1604, 192, 502, 1462, 501, 514, 703, 1698, 1828, 2145, 178, 1277, 693, 686, 687, 1782, 2244, 1656, 1812, 581, 2277, 1723, 179, 165, 1961, 817, 2003]. Concentrated [111]. Concept [953, 710, 2345, 425, 1353, 330, 910]. Concepticons [2374]. Conception [165]. Concepts [1414, 167, 567, 1645, 613, 278, 2382]. Conceptual [706, 1366, 1354, 1352, 1357, 1346, 1347, 1363, 1364, 1182, 1349, 1365, 1361, 1360, 1353, 1369]. Concerning [1830]. Conciliation [351]. Concordia [564]. Concrete [1487]. Concurrency [782, 789, 1373, 862, 1829, 16]. Concurrent [1666, 77, 1540, 645, 899, 74, 2105, 1401, 2108, 1962, 98, 779, 1267, 1082, 1083, 1537]. Condensed [1154]. Condition [1275, 2339, 804]. Conditional [2354, 1098, 2291, 541, 195, 794, 2363, 471, 818]. Conditional-Logical [195]. Conditionals [1111]. Conditioning [576]. Conditions [614, 2142, 133]. Conference [46, 1552]. Conference-Key [46]. Configuration [882, 878, 873, 880, 917, 231, 962, 879, 1959]. Configuration-Based [231]. Configurations [1324, 1732]. Conflict [1684, 1593, 2210]. Conflict-Free [1684, 1593, 2210]. Conflicts [904, 2037]. Confluence [561, 1922]. Confocal [2248]. Conformational [1792]. Conforming [129]. Congestion [1059, 457]. congruence [793]. Conic [2256]. Conjecture [1691, 1332, 806]. conjunctions [3]. Conjunctive [2, 364, 4]. Connected [840, 248, 1837, 592, 141]. Connection [1141, 178, 1841, 1747]. Connection-less [178]. Connectionist [973, 976, 1864, 1865, 1894, 1895]. Connections [2055, 1220]. Connective [1927]. Connectivity [211, 919, 920, 1991]. Connectors [486]. Conquer [2109, 1043]. Consecutive [1709]. Consensus [1451, 1047]. conservative [1467]. Conservativeness [1383]. considerations [311]. Consistency [925, 2046, 1201, 1416, 20]. Consistent [604]. Constant [964, 1295, 1824, 2191]. Constant-Depth [2191]. Constants [1383]. Constrained [1578, 1604, 987, 69, 1319, 2265, 2009, 459, 2068]. Constraint [1288, 1408, 1019, 219, 2115, 1409, 523, 458, 1129, 697, 948, 2317, 2327, 1951, 1257, 1082, 1453]. Constraint-Based [1409]. Constraints [2370, 713, 894, 484, 2356, 1568, 2334, 1115, 362, 479, 1807, 1151, 231, 2138, 518, 397, 1454, 1257, 4]. Constructing [1689, 1152, 613, 1343, 1606, 876, 23]. Construction [690, 202, 1714, 707, 426, 1359, 2202, 1446, 1758, 1322, 1188, 270, 1241, 615, 1403, 1777, 1582, 1983]. Constructions [553, 1672, 1286, 404]. Constructive [1329, 1667, 1868]. constructor [805]. constructor-sharing [805]. Constructs [1199, 1172, 1956]. Consultation [410]. Consumer [1064, 1062]. Contacts [1839]. Containment [1245, 1834, 2, 3]. Content [2134, 1051, 2376, 1061]. Contest [62]. Context [472, 552, 553, 148, 1826, 541, 561, 2362, 1719, 1488, 803]. Context-Free [552, 148, 541, 561]. Context-Freeness [1826]. Context-Sensitive [2362, 803]. Contextual [2140, 1835, 2221, 2142]. Continuation [104, 969, 1923]. Continuity [2084]. Continuous [746, 785, 1627, 1628, 1629, 1701, 283]. Continuously [232]. Contour [2275]. Contours [752, 591]. Contraction [150]. Contractions [269]. Contracts [781]. contradiction [1466]. Contrast [1702, 764, 2267, 295]. Contrast-Optimal [1702]. Control [1060, 722, 993, 702, 246, 101, 1059, 1491, 1615, 692, 386, 180, 736, 1617, 1540, 718, 119, 726, 544, 223, 1419, 727, 734, 961, 276, 218, 1326, 513, 231, 445, 1646, 1620, 1621, 1735, 1024, 1625, 1626, 1963, 1764, 1627, 1659, 1630, 2019, 712, 1944, 357, 1946, 16, 1945]. Control-based [1963]. Control-Driven [1764]. Control-Flow [1540]. Controllable [441]. Controlled [542]. Controller [984, 714]. Controllers [983, 1619, 230, 719]. Controlling [1018, 543]. Convection [683, 124, 764]. Convergence [111, 115, 908, 127, 133, 2023]. Conversational [1489]. Conversion [935, 131, 813]. Convex [643, 1689, 682, 123, 2251, 384, 296, 676]. Convexity [1586]. Conveyer [1608]. Convolution [2311, 1775]. Cooling [722].

Cooperating [549, 548]. Cooperation [91, 242, 880, 237]. Cooperative [721, 1060, 92, 913, 1493, 2036, 236, 1640, 179, 1963]. Coordinate [2322]. Coordinate-Free [2322]. Coordinating [1538, 1949, 1954, 1942]. Coordination [97, 575, 586, 389, 896, 684, 2189, 1961, 1764, 1032, 1935, 1950, 1960, 1952, 1943, 1962, 1936, 1956, 1940, 1941, 1963, 1965]. Copyright [1055, 1554]. Copyrighted [1632]. CoQ [1666, 1683]. CORALI [1346]. CORBA [908, 921, 237]. Core [661, 2145, 2372]. Cores [1574]. Corner [2261]. Coronary [417]. Corpora [360]. Corporate [1360]. Corporations [1641]. Correct [2208, 1248]. Correcting [1817]. Correction [753, 314]. Correctness [867, 494]. CORRELATE [902]. Correlation [847, 2180, 2272]. Correlation-Immune [847]. Correspondence [19, 607, 765, 2274]. Correspondences [2287]. Cortex [960, 759]. Cortical [995, 999]. Cost [831, 1048, 1433, 68, 1335, 249, 1652, 1875, 1876, 1721, 2156, 307]. Cost-Delay [1652]. Cost-Effectiveness [1433]. Could [87]. Count [1322]. Counter [1168, 1323]. Counter-examples [1168]. counters [15]. Counting [546, 147, 2222, 1520]. Coupled [109, 511]. Coupling [959, 1496]. Course [1135]. Coverability [1834]. covering [1969, 384]. Coverings [1838]. COVERS [1757, 1789]. Crank [111]. Crash [1044]. CRAY [2095, 508, 671]. Creating [164, 1369]. Creation [1051]. Creativity [1506]. Criteria [1516, 1826, 1832, 848, 1998]. Criterion [867, 2264, 812]. Critic [1157]. critical [1248]. Critique [1058]. Cropped [2015]. CROQUE [1560]. Cross [930, 247, 195, 795, 2133]. Cross-Entropy [195]. Cross-Loop [930]. Cross-sections [795]. Crossing [475]. Crossings [1575]. Crossover [454, 462]. cryosection [294].

Cryptanalysis [1521, 835, 1510, 1511, 2158, 1520, 2182]. Cryptographic [2164, 822, 1546, 839, 848, 1530, 1071, 1244]. Cryptography [2188, 845, 2149, 1702, 45, 824]. Cryptologic [846]. Cryptosystem [2159, 2161, 44, 2152, 821, 2158]. Cryptosystems [2181, 2153, 2169, 844, 47, 2171, 2163]. CSG [1577]. CSP [1670]. CT [288, 327, 294, 430, 292, 2014, 290]. CT-based [292]. CT-retinography [327]. cube [143]. Cubic [268]. Cues [1002]. Cultural [2007]. Culture [1642]. Cumulative [939]. curly [212, 183]. Current [976]. curvature [1216]. curvatures [1218]. Curve [2169, 45]. Curves [107, 764, 38, 844, 2170]. Customer [916]. Cut [1371, 2056, 865, 1380]. Cut-Elimination [1371]. Cut-Free [1380]. Cycle [1794, 2015, 1252, 1991]. Cyclic [1911, 1275, 1283, 2281, 1286, 406]. Cyclically [1286]. Cylinders [2249].

D [295, 1212, 291, 754, 289, 290, 762, 104, 326, 681, 758, 283, 929, 632, 67, 289, 1742, 757, 305, 1770, 286, 742, 1207, 310, 751, 1774, 2259, 2295, 2278, 2244, 2289, 2245, 2283, 2286, 2288, 2263, 1766, 290]. D/ [289, 290]. D0L [2216]. D0L-Languages [2216]. DAC [2195]. Dag [147]. DAGs [156, 274]. DAL [767]. Dangers [1200]. Data [1391, 2246, 1018, 1755, 1427, 1086, 229, 931, 167, 27, 943, 1879, 433, 1408, 234, 1857, 246, 1090, 1905, 1903, 2332, 740, 672, 1441, 224, 1201, 1450, 2307, 180, 2381, 1858, 986, 416, 1891, 1593, 1233, 504, 861, 1294, 1230, 1596, 1426, 1406, 663, 1567, 359, 1859, 1440, 1881, 1853, 1742, 2212, 1884, 951, 1655, 2074, 176, 1880, 630, 1915, 2278, 1867, 1886, 1860, 2059, 1856, 1888, 1647, 2315, 1883, 1892, 1864, 1865, 1899, 2044, 1874, 1919, 1861, 617, 1862, 274, 761, 1902, 932, 1863, 1411, 21, 19, 1, 396, 18, 291, 1979, 5, 794]. Data-Dependent [2212]. Data-Flow [224]. Data-Independent [2212]. Data-Intensive [1647]. Data-Parallel [504, 176]. Data-Structures [229]. Database [925, 355, 1019, 1563, 1559, 1556, 2332, 1781, 1056, 1406, 1564, 750, 1644, 1648, 1796, 2044, 2352, 1439, 1475]. Databases [1795, 1087, 1408, 1409, 860, 2325, 1565, 1646, 1801, 1901, 1890, 1569, 1894, 1895, 20, 1471]. Dataflow [1751, 932]. Datalog [2195, 6, 1454]. datapath [1242]. Datasets [1802]. Dataspace [1764]. Datatypes [2085, 2086]. Dates [939]. Datum [1893]. DAVIC [908]. DCT [2307]. Deadlines [225, 1592]. Deadlock [171, 1920, 1558, 200]. Deadlock-Free [200]. Deadlock-Freedom [1920]. Deakin [1365]. Dealing [1093, 1900]. Debugging [156, 1941, 1763, 1786]. Decentralized [562]. Decibel [1398]. Decidability [1616, 561, 2064]. Decidable [1140, 1323, 2216, 1245, 795]. Deciding [2083, 2226]. Decimation [287]. Decision [1141, 1429, 1147, 2199, 724, 1928, 416, 2336, 1507, 35, 2341, 418, 1155, 422, 1872, 2214, 1673, 2280, 2196, 193, 1931, 139, 1240, 374]. Decision-Support [418]. Declarative [1490, 2128, 66, 2108, 1465]. Declustering [1721]. Decoding [1269]. Decomposable [1327]. Decomposition [1854, 971, 1272, 510, 1974, 613, 511, 676, 1759, 1174, 129, 1775, 2224, 2305, 932, 412, 1121]. Decompositions [1626]. Decryption [2152]. Deduction [701, 1480, 364]. Deductive [772, 189, 861, 1566, 1569, 1474, 1471, 1475]. Deductive-Algorithmic [189]. Default [1089, 371, 1102, 1108, 1113, 1114]. Defaults [1349, 1459]. Defect [2262, 621]. Defects [2255, 1753]. deferred [16]. Definability [1823, 1341, 158, 2060]. Defined [951, 844]. Definition [283, 1172, 2144]. Definitions [2348]. Deformable [606, 591, 743, 604, 742, 759, 287, 284].

Deformation [748, 2252]. Deformations [738]. degeneracy [254]. Degree [140, 1824, 1601, 1971, 1975]. Delaunay [2001, 1690]. Delay [1291, 992, 1652]. DeleGate [1661]. Delegation [904]. Deletion [1337]. Delivery [1055, 1653]. delta [2367]. Demand [905]. Demonstration [1041, 840, 1063, 1626]. Dempster [1099]. Dempster-Shafer [1099]. Dendritic [2248]. Denesting [1971]. Deniable [2151]. Denotational [2068]. Density [1571, 764, 1896]. Dependability [1117]. Dependable [649]. Dependence [246, 1107, 942]. Dependency [1087, 2229]. Dependency-Based [2229]. Dependent [1350, 472, 1657, 220, 1378, 2212, 347, 404]. Deployment [1650, 165]. Depth [2324, 2191, 1600, 379]. Depth-First [1600]. deques [1545]. Derandomization [1298]. Deregulating [913]. Derivation [546, 1680, 1766]. Derivations [157]. Derivative [1295, 2260]. Derived [995]. Deriving [1783]. Described [2130, 1833]. Describing [1905]. Description [735, 409, 554, 1487, 1875, 279, 1876, 1944, 1947]. Descriptions [244]. Descriptive [193]. Desgin [1626]. Design [624, 1551, 97, 311, 780, 1506, 441, 842, 1007, 442, 983, 533, 1044, 506, 685, 1637, 1039, 734, 429, 1866, 1481, 218, 155, 24, 714, 2031, 848, 1515, 1937, 1745, 1717, 1038, 1751, 165, 1393, 1543, 1249, 382, 910]. Design/CPN [533]. Designing [673, 888, 1531]. Designs [901, 1255]. Desired [1180]. Destroying [2298]. Detachment [1154]. Detectable [828]. Detecting [2321, 1905, 2037, 156]. Detection [2302, 744, 534, 709, 1136, 2262, 770, 1558, 2291, 737, 2270, 32, 2101, 1410, 336, 2296, 2317, 2261, 2266, 763, 1000, 2297, 1253, 1466, 1252, 1218, 286]. Detector [1009]. determinants [2003]. Determination [2286, 2352]. Determining [936]. Determinism [187]. Deterministic [258, 596, 597, 2355]. Develop [1799].

Developing [445, 1191, 915]. Development [887, 1177, 355, 1189, 632, 328, 697, 1479, 2030, 271, 2031, 1868, 498, 1566, 331]. Developments [1795, 693, 1069, 814]. device [314]. Diabetic [1879]. Diagnoses [1370]. Diagnosis [417, 1176, 73, 420, 151, 419, 1886, 1906, 415, 298, 1473]. Diagnostics [1576]. diagram [1245]. Diagrams [1587, 1354, 2199, 189, 1904, 193, 1931]. Diagraphs [182]. Dialog [887, 1169]. Dialog-Based [887]. Dialogue [889, 888, 884, 92, 885, 1169, 886, 1855]. Diameters [1306]. Diamond [156]. DIANE [1049]. dichotomous [1607]. Difference [473, 111, 127, 133, 1212]. Differences [610]. Different [1041, 1756, 1091, 509, 1115, 615, 1348, 2000]. Differential [1513, 2181, 2321, 1616, 113, 1273, 767]. Differentiation [117]. Difficulties [991]. Diffie [45]. Diffusion [1142, 116, 764, 1213, 1210, 1220, 1221, 1225, 1205, 1214]. Diffusive [1231]. Digit [1008]. Digital [1052, 2247, 1181, 2157, 429, 976, 730, 731, 2155, 2284, 49, 2156, 2306, 1237]. Digraphs [152]. Dilation [2209]. Dilemmas [393]. dimension [373, 372, 383]. Dimensional [1443, 116, 205, 63, 2248, 1833, 2218, 1836, 1745, 21, 318, 306, 288, 299, 1978, 1983, 1224, 1988, 807]. Dimensionality [2235]. Dimensions [2219, 1992]. Diophantine [1987, 811]. Direct [1780, 2307, 1446, 125, 1774]. Directed [1446, 1374, 58, 1320, 47, 493, 1972, 792]. Direction [1776]. Directional [503, 2260]. Directions [693]. disagreement [377]. Disambiguation [2134]. disbeliefs [1458]. Discipline [881]. Disciplines [2097]. Discontinuous [993]. Discovered [1434]. Discovering [434, 1422, 1859, 11, 1887]. Discovery [1168, 2325, 1419, 1421, 1794, 1888, 1890, 1893, 2352, 1423, 241]. Discrepancy [1570]. Discrete [735, 509, 232, 118, 1326, 623, 1904, 2252, 2162, 2265, 1628, 836, 2163, 2306, 2004]. Discrete-Log [2163]. Discrete-Ordinates [509]. Discrete-Time [1326, 2265]. Discretization [2331, 511]. Discriminant [341]. Discussion [833]. Disease [417, 739, 422, 763]. Disease-Induced [739]. diseases [327]. Disjoint [1606, 1972]. Disjunctive [523, 872, 1114, 1472, 1454, 1471]. disk [1119]. disk-based [1119]. Disks [1597, 2306]. DisLoP [1472]. Disorder [450]. Disparity [2273]. Dispatch [1028]. Dispatching [2105]. Display [1380]. Disprove [1826]. Dissimilarity [1877]. Distance [1587, 1269, 30, 1973, 2180, 1797, 2345, 1864, 1865, 2350, 267]. Distance-Based [1864, 1865]. Distances [857, 2203]. Distinction [1093]. Distinctions [2379]. Distinguishes [251]. Distributed [925, 641, 239, 943, 426, 175, 1757, 661, 442, 1822, 945, 1135, 878, 725, 81, 684, 895, 645, 589, 1027, 899, 1637, 1643, 784, 178, 1202, 1565, 628, 1742, 686, 687, 1633, 1760, 879, 156, 849, 1762, 625, 2385, 1835, 1746, 1651, 1636, 1749, 1045, 548, 581, 1635, 2102, 779, 1229, 1754, 639, 556, 2039, 1125, 1960, 1990, 1940, 1941, 12]. Distributed-Memory [943, 661]. Distribution [931, 46, 1053, 1151, 2267, 1554]. Distributions [745, 1133]. Divergence [534]. Divide [985, 2109, 1043]. Division [345, 1099]. DNA [1449, 1799, 1812, 2019]. DNA-Regions [1799]. DNF [380]. Do [101, 278, 276, 2086, 393, 2123]. DOA [1566]. Document [1425, 710, 1361, 281]. Documentation [499, 1037, 1038]. Documents [2141, 2137, 2138, 2140, 2144, 2142, 2133]. Dodgson [1302]. DOF [2258]. DOLMEN [906]. Domain [1350, 855, 345, 919, 918, 2376, 1916, 1359, 2115, 510, 511, 414, 1376, 676, 926, 1759, 1421, 129, 2089, 1889, 915].

Domain-Decomposition [1759]. Domain-Free [855]. Domains [485, 856, 857, 48, 416, 1339, 422, 1487, 400]. Dominance [460]. Dominant [1396]. Dominating [140]. Domination [1708]. Done [874]. DOSFOP [499]. down [1150, 1149]. Downward [199]. Draft [1516]. Drainage [604]. DRASTIC [1027]. Drawing [55, 56, 58, 62, 61, 63, 68, 1601]. Drawings [1599, 57, 149, 257, 64, 1970]. Drift [253, 376]. Driven [1364, 1868, 1764, 741]. DrScheme [2126]. Drug [1794]. DSS [2164]. DTS [985]. Dual [1337, 1278]. Duality [259]. dually [1973]. Due [769]. Dungeon [2128]. Duration [226]. durations [1995]. durative [1942]. During [2330, 326, 319]. Dwork [2152]. Dyck [1842]. Dynamic [587, 1306, 76, 464, 1134, 390, 1571, 536, 181, 975, 921, 1618, 458, 1596, 697, 514, 1979, 182, 588, 1307, 1742, 422, 951, 1025, 1654, 1995, 1763, 1818, 1070, 1991]. Dynamical [546, 656, 1401, 1622, 2220, 1627, 133]. Dynamically [1981, 1889]. Dynamics [1305, 678, 2003, 1938].

E-Unification [1296, 2069]. Ear [1002]. Early [945, 1000]. Ease [1495]. Easy [266]. ECG [992, 1878]. echography [299]. Economic [720, 653, 2022]. Eden [1537]. Edge [2211, 737, 32, 1319, 71, 1713, 1000, 159, 2297]. Edge-Rankings [159]. Edges [147, 32]. Edit [2180]. Editor [2127]. Education [1066]. EEG [709]. Effect [76, 1351, 951]. Effective [659, 813]. effectively [1256]. Effectiveness [1433, 2300, 2084]. Effects [319, 764, 2120, 1914]. Efficiency [465, 1433, 1888]. Efficient [1526, 1418, 762, 505, 202, 1236, 1253, 1342, 1270, 1120, 2175, 1928, 46, 1312, 850, 2174, 242, 1407, 650, 43, 827, 839, 168, 506, 1334, 170, 685, 152, 2173, 1426, 1590, 1828, 1866, 1327, 2169, 154, 1133, 1610, 4, 1328, 1708, 2111, 2261, 191, 1229, 1754, 1156, 1970, 1266, 1240, 1119, 910, 1242, 1263]. Efficiently [774]. Effort [1059, 1479]. Ego [2257]. Ego-Motion [2257]. Egomotion [2259]. Eigensolver [507]. Eigenspace [2234]. Eigenvalue [130, 133]. Eigenvalues [1738]. Eigenvectors [1738]. Elaboration [1481]. ELAN [815]. Elastic [739, 726, 2290, 285]. Elastically [2247]. Election [1757, 827]. Elections [1302]. Electrical [959]. Electronic [41, 1007, 39, 927]. Electronic-Cash [41]. Element [107, 1994]. Elementary [264, 1314, 1282, 385, 1102]. Elements [132]. ElGamal [828]. Eliminating [940, 2152]. Elimination [1371, 614, 865, 1144]. Elliptic [2169, 38, 45, 844, 2170]. Elliptic-Curve [45]. embedded [1554, 1248]. Embedding [144, 1670, 247]. Embeddings [52]. Emergent [2005, 392]. Emerging [1641, 1631]. EMG [420]. Emotion [95]. Emotion-Based [95]. EMPA [1311]. Empirical [619, 741, 382]. enabled [311]. Enabling [1553]. Enantioseparations [2332]. Encapsulation [1030, 1054]. Encoded [359, 1553]. Encoding [337, 1022, 2223]. Encodings [1924]. Encryption [1516, 2150, 2151, 1333, 1510, 1515, 1529, 1514, 1528, 1531, 2182, 1530, 1527, 2156, 820]. Encryption/MAC [1527]. Encyclopedic [1485]. End [1536, 997, 890, 927]. End-to-End [890]. Ended [2115]. Endgame [1890]. Endoscopy [762, 302, 304]. Energy [603, 600, 607, 672, 2310, 2300, 592]. enforced [1467]. Engagements [2033]. Engineering [1178, 1231, 1145, 2006, 698, 1635, 233]. Enhance [1501, 609]. Enhanced [453]. Enhancement [2267, 1221]. ENT [332]. Entailment [2353]. Enterprise [236]. Entire [1802]. Entrenchment [1105]. Entropic [2011]. Entropy [832, 2308, 2334, 195]. Enumerating [1732]. Enumeration [1284]. Enumerations [1816]. Enumerative [1290]. Environment [641, 1050, 889, 882, 679, 627, 677, 1657, 913, 878, 685, 89, 895, 1162, 535, 622, 2126, 1643, 2030, 1572, 1483, 1636, 1763, 1659, 498, 423, 311, 1116, 310]. Environmental [955, 1880]. Environments [464, 1571, 642, 568, 1046, 1923]. EOL [549]. Epidemics [434, 1422]. Epidemiology [1880]. Episode [1442]. EPNet [2021]. EpoDB [1796]. epsilon [212, 183]. epsilon-Free [183]. epsilon-Transitions [212]. Equality [989, 398]. Equals [1341, 158]. Equation [514]. Equational [469, 2358, 794]. Equations [111, 113, 1273, 511, 1309, 699, 1292, 133, 811]. Equipped [1832]. Equivalence [210, 1131, 1323, 1834, 1846, 2348, 1286, 1991]. Equivariant [2260]. Ergodic [204]. ergonomic [328]. Erlang [2129]. Error [2311, 151, 125, 1709, 1817]. Error-Correcting [1817]. Error-Free [2311]. Error-Tolerant [1709]. Errors [1449, 979, 884, 1340, 2152, 135]. Erythropoiesis [1796]. Escape [461]. Escrow [48, 1333, 40, 828, 1551]. essays [2387]. Esterel [524]. Estet [671]. Estet-Astrid [671]. Estimated [605]. Estimates [745, 127]. Estimating [663, 1479, 1897, 987, 1748]. Estimation [505, 2235, 2287, 593, 1897, 2273, 313, 2323, 2259, 2257, 1896, 288]. Estimation-Propagation [2257]. Estimations [764]. Estimators [2271]. ET-LOTOS [790]. Eta [404, 1379, 870]. Eta-Expansions [1379, 404, 870]. EURESCOM [912]. Europe [1047]. European [916]. Evaluate [623]. evaluating [1124, 292]. Evaluation [1526, 1516, 889, 1385, 1091, 2096, 769, 622, 1433, 890, 982, 1560, 729, 1279, 1110, 321, 2264, 2196, 2117, 1945, 2102, 1488, 2119, 1642, 658, 320, 1122, 1117]. Evaluations [896, 2111]. Evasion [1572]. Event [232, 1172, 623, 1904, 2210, 2344, 49]. Event-Stream [49]. Evidence [1696, 1112]. Evolution [721, 1398, 713, 723, 575, 1811, 271, 1395, 1396, 1025, 2383, 459, 456, 2029]. Evolutionary [1696, 600, 465, 466, 2022, 2027, 2007, 1811, 458, 452, 711, 2006, 703, 1698, 1870, 2020, 451, 1390, 269, 1388, 2009, 2014, 2313, 2008, 1749, 2016, 719, 1871, 463, 2018]. Evolvable [1392, 1388, 1386, 1393]. Evolving [1027, 1167, 2013, 1389]. Evolware [1387]. Exact [933, 1302, 1978]. Exactly [1293, 1980]. Examining [1499]. Example [724, 1047, 422]. Examples [707, 1168]. Exception [1020]. Exceptions [1349, 2135]. Exchange [918, 680]. Exclusion [344]. Executable [272]. Executing [1663, 1103]. Execution [426, 958, 936, 1633, 585, 786, 12]. Exemplar [643, 682, 676]. Existence [2339, 266]. Existential [2232]. Existentially [2370]. Existing [1830]. Exogenous [1901]. Expansion [2377, 2112, 2171]. Expansions [1379, 404, 870]. Expectation [602]. Expectation-Maximisation [602]. Experience [945, 883, 1390, 445, 303]. Experiences [670, 329, 436]. Experiment [672, 164, 895, 576, 1361, 320]. Experimental [1091, 2096, 58, 303, 305, 322]. Experimentation [315]. Experiments [946, 1231, 1382, 360]. Expert [351, 414, 413, 1855, 2039]. experts [378]. Explaining [2046]. Explanation [1496]. Explanations [410, 424]. Explicit [2367, 1486, 492, 1101, 2358, 1909]. Exploiting [562, 609, 651, 1873]. exploits [1251]. Exploration [1234, 1258]. Exploratory [1874]. Exploring [465, 1600]. Exponential [858, 2191, 2194]. Exponentials [1282]. Exponentiation [1787, 1276]. Exposing [1936]. Expressed [1032]. Expressibility [1292]. Expression [1796]. Expressions [183, 1310]. Expressive [15, 1344, 1314]. Expressiveness [2195]. Extended [783, 1023, 226]. Extendible [2032]. Extending [789, 1909, 1378, 1927]. Extensible [369]. Extension [1376, 1006, 1285]. Extensional [1304]. Extensionality [865]. Extent [1573]. Exterior [741]. External [1445, 1002, 3]. Extract [2283, 1439]. Extracting [27, 1367, 1894, 1895]. Extraction [2246, 2375, 333, 234, 1615, 2308, 973, 2374, 2373, 2379, 1224, 336, 2024, 2380, 1625, 2372, 1226, 1242]. Extractor [1367]. Extrema [464]. Eye [333, 327].

F [1925]. F0omega [1379]. Face [357, 353, 341, 2291, 358, 335, 344, 336, 2290, 2293]. Faces [2293]. Facial [2309, 334, 343]. FACILE [82]. Facilitating [1176]. Facility [565, 1636]. Factor [947]. Factor-Join [947]. Factorisation [2075]. Factorization [1270]. Factorizations [2158]. Factors [1355]. Facts [554, 385]. Factual [1107]. Fail [480, 851]. Fail-Stop [851]. Failure [2159, 1496, 1958]. Failures [1486, 38]. Fair [1433, 784, 1084]. fairness [1251]. Families [549, 2215]. Family [549, 1522, 2170, 1791, 459, 1982, 814]. Fanin [1692]. Farmer [649]. Farmer-Worker [649]. Farms [173]. FAs [1404]. Fast [1526, 1385, 980, 1371, 1532, 1099, 114, 1711, 2287, 1533, 825, 1593, 1517, 1221, 2270, 2179, 2296, 843, 1720, 1531, 2171, 456, 1690, 1805, 381, 312, 1775, 2305]. Fast-Division [1099]. Fat [1596]. Fault [2181, 1136, 1202, 2042, 1989, 250]. Fault-Tolerant [1202, 1989]. Faults [822, 1824, 1828, 2211]. faulty [1976]. FDT [1845]. FEAL [1513]. feasability [300]. Feature [618, 608, 1014, 346, 612, 1870, 2298, 1034, 2319, 1871, 1208, 1209, 1224]. Feature-Based [2319]. Feature-Oriented [1034]. Features [1199, 334, 928, 1002, 336]. Fed [702]. Fed-Batch [702]. Federated [919]. FEDGE [32]. Feed [963]. Feed-Forward [963]. feedback [309]. Feedbacked [691]. FEM [682]. Fermat [2311]. Fermentation [702]. Fetal [972]. Few [707, 1606, 1601]. Field [108, 594, 593]. Fields [1270, 2172, 847, 2311, 2253, 969, 982, 756, 1006, 843, 2236, 1276, 2277]. FIFO [1324]. FIFO-Channel [1324]. Figure [600, 1218]. Figure-Ground [600]. File [1755, 664]. Files [1427]. Filled [2246, 769]. Filled-In [2246]. Filter [1518, 1412, 2279, 2261]. Filter-Refine [1412]. Filtering [466, 999, 163, 240, 1210, 1205]. Filters [691, 334, 427, 2294, 2304, 1221, 1223, 1214]. Find [1157, 184, 698, 154, 1856]. Finding [1574, 115, 1424, 248, 1431, 2213, 844, 1808, 759, 381]. Finds [1395]. Fine [945, 1762]. Fine-Grain [945]. Fingerprinting [203, 826]. Fingerprints [340]. Finitary [1916]. Finite [1739, 107, 132, 1293, 1270, 1400, 2115, 859, 119, 187, 541, 1897, 1727, 1402, 866, 2212, 1186, 183, 1933, 127, 1833, 2223, 2196, 1846, 1847, 2044, 2231, 814, 1276, 1285, 1994]. Finite-Horizon [2196]. Finite-State [1400, 1933]. Finitely [1845, 795, 796]. FIR [691]. FIRE [1404]. Firmware [1524]. Firmware-Oriented [1524]. First [1227, 1147, 219, 1669, 895, 182, 1600, 1023, 2064, 2347, 2133, 2350, 2352, 1310, 871, 303, 809, 810]. First-Order [1147, 1669, 2347, 2352, 1310, 810]. Fish [276]. Fitness [1882]. Fits [1642]. fixation [321]. Fixed [1447, 696, 1827, 961, 1745, 1722, 1240]. Fixed-Size [1745]. fixed-sized [1240]. Fixed-Topology [1447]. Fixing [1986]. Fixpoint [1375, 1455]. Flat [996]. Flattened [2249]. Flexible [562, 1476, 531]. FLIP [25]. FLORID [1477]. Florist [1187]. Flow [109, 111, 224, 1540, 644, 78, 223, 68, 750, 1735, 2265, 2113, 2317, 1778, 2015, 493, 2023, 12]. Flow-Directed [493]. Fluid [644, 678]. fluoroscopies [290]. fluoroscopy [326]. fly [1251]. Focus [595]. Focusing [1093, 72]. Fold [1805]. Folding [2354, 1804]. Folding/unfolding [2354]. Following [1209]. Forall [953]. Force [629]. Forced [769]. Forced-Choice [769]. Forecasting [720, 1400, 434, 1422, 1012]. Foreign [680]. Forests [1722]. Forgery [852]. Forging [682]. fork [1127]. Form [1152, 2038]. Formal [439, 1663, 90, 529, 1666, 1135, 497, 1331, 1947, 2379, 435, 1237, 2130, 2043, 1744, 1812, 1353, 779, 1543, 1264, 1171, 12, 1259]. Formalisation [91, 1672]. Formalised [1199]. Formalising [74]. Formalism [470, 1618, 89, 1349, 778, 2138]. Formalisms [450]. Formalization [1031, 1675]. formation [307]. Formatting [2143]. Forming [1874]. Forms [2246, 1309, 2338, 2227, 1851, 872]. Formulae [1577, 872, 385]. Formulas [869, 2224, 1253]. Formulation [1669, 574]. Fortran [638, 658]. FORUM [449]. Forward [963, 1453, 2346]. forwarding [1956]. Found [1870, 1871, 2353]. foundation [12]. Foundations [2149, 2384]. Founded [522, 523, 6]. Four [1370]. Four-Valued [1370]. Fourier [1775, 2292]. Fractal [627, 1750]. Fractal/Wavelet [627]. fraction [1976]. fractures [321]. Fragment [862]. Frame [1449, 317]. frame-based [317]. Frame-shift [1449]. Framework [2106, 483, 1800, 1175, 416, 784, 2341, 1564, 1164, 1028, 2035, 1110, 1744, 1927, 2319, 1219, 294, 1468]. Frameworks [36, 1037]. Franklin [86]. Fraud [828]. Fraud-Detectable [828]. Free [2322, 851, 855, 2053, 831, 1338, 552, 2311, 553, 148, 1684, 1593, 863, 541, 200, 1380, 477, 183, 561, 2210, 2217, 1735, 1310, 812, 800]. Freedom [1920]. Freeness [1826]. Freezable [274]. Frege [2191]. Frequency [345, 2273, 2217]. Frequently [449]. Fresh [1034]. Front [997]. Front-End [997]. Frontier [1436]. Full [1315, 506, 865, 1672]. Fully [984, 2236, 1243, 1991]. Function [838, 668, 2172, 2007, 89, 1001, 1882, 2062, 412, 1221, 812, 1547]. Functional [2355, 483, 1538, 749, 1536, 1534, 1358, 2122, 420, 1541, 2357, 2124, 1870, 1539, 2111, 2121, 2360, 2123, 1871, 79, 396, 1212, 315, 1193, 289]. Functionality [653, 1180]. Functionally [2125]. Functions [202, 1385, 1587, 132, 265, 1099, 37, 2177, 847, 613, 192, 698, 1279, 951, 38, 848, 2110, 1694, 1887, 1701, 1700, 1217, 375, 296, 1548]. Fundamental [132, 1093]. Furnace [1010, 634, 138, 137]. Fusing [353]. Fusion [439, 1086, 1090, 1092, 352, 987, 2237, 327]. Future [467, 468, 693, 976, 910]. Fuzzy [28, 706, 722, 542, 167, 27, 735, 433, 707, 702, 36, 30, 25, 692, 29, 1357, 983, 2027, 2308, 31, 984, 725, 708, 718, 458, 1507, 701, 697, 727, 34, 734, 716, 699, 704, 693, 757, 26, 32, 1884, 24, 700, 731, 742, 714, 1885, 1886, 2028, 2010, 1882, 1887, 719, 1004, 759, 1509, 23, 2019, 165, 421, 712, 733]. Fuzzy-Morphology [1004]. Fuzzy-Neural [693].

GA [1866, 2028, 2019]. Gabor [2279, 2300, 2274]. Gabor- [2274]. Gaining [466]. Gait [339]. Galois [847, 843]. Game [2184, 1303, 2029]. Games [407, 1620, 2192, 1625]. Gang [2100, 2099, 2102]. Gap [1711, 362, 155]. Gap-Order [362]. garbage [405, 1990]. Gardens [271]. Gas [632]. Gate [1392, 1388]. Gate-Level [1392, 1388]. Gated [740]. Gateway [1661]. Gaussian [745, 746, 461, 756]. Gaze [2323]. Gbit [1525]. Gbit/second [1525]. GCHQ [829]. Gene [1798, 1796]. General [508, 1645, 517, 1407, 150, 1028, 2317, 2089, 1661, 1127, 1469, 1978, 1468]. Generalisation [1416, 1157, 1875, 1876]. Generalization [1359, 2357, 1421]. Generalized [258, 472, 2249, 850, 2311, 1823, 232, 953, 474, 1133, 561, 1706, 336, 1842, 1803, 1471, 159, 1470]. Generated [549, 2056, 660]. Generating [1686, 2330, 465, 83, 1785, 201, 751, 1974]. Generation [2370, 1428, 1795, 2164, 2175, 2206, 538, 2131, 886, 437, 275, 2218, 1630, 425]. Generative [280, 557, 1948]. Generator [1177, 2180, 834]. Generator-Based [1177]. Generators [1518]. Generic [856, 2048, 913, 1093, 965]. Genericity [1035]. Genes [1802]. Genetic [720, 690, 2303, 721, 673, 1398, 643, 2309, 455, 2027, 736, 453, 689, 698, 699, 460, 1733, 2026, 1576, 2011, 2010, 1882, 2013, 2015, 2017, 1887, 732, 2023]. Genetic-Based [2023]. Genetic-Entropic [2011]. Genome [1811, 1809]. Genomic [1802]. Genomics [1710]. Gentle [1135]. Gentzen [2046]. Geo [1414]. Geo-Opera [1414]. Geodesic [30]. Geographic [687, 2044]. Geometric [1407, 1591, 462, 2264, 1581, 2245, 1849, 1605, 2240, 1765, 1979, 1211, 811]. Geometrical [297]. Geometry [1570, 2122, 1585, 254, 1797, 1330, 2044, 1146]. GeoStore [1413]. GeoToolKit [1413]. GF} [843]. Giant [2211]. Gibbs [2267]. GIOTTO3D [67]. Given [2302, 1322, 1628]. Glacial [934]. Glass [138, 137]. Glaucoma [421]. GLCM [294]. Global [938, 689, 924, 754, 920, 2101, 2084, 1066]. Glucose [993]. Goal [390, 428, 792]. Goal-directed [792]. Goal-Oriented [428]. Goals [1496]. Goedel [2063]. Goes [1145]. Going [1039]. Gonzalo [960]. Good [2168, 1021, 580, 844]. GOODS [1017]. Gossiping [192]. Government [2022]. GP [280]. gpo [479]. Gradient [1228, 2317]. Graduality [1100]. Graesser [86]. Grain [945, 2282, 651, 1762]. Grained [1312]. Gram [1788]. Grammar [551, 549, 546, 552, 553, 547, 557, 554, 2136]. Grammar-Compatible [2136]. Grammars [859, 1811, 544, 2130, 1835, 2221, 2227, 1946].

Grammatical [540, 1811]. Grand [1405]. Graph [55, 1312, 1357, 2077, 859, 1364, 59, 62, 63, 602, 68, 2341, 601, 154, 69, 2205, 476, 157, 2363, 2228, 1738, 597, 2290, 72, 1970, 1977, 1946]. Graph-Drawing [62]. Graphical [1789, 1162, 1094, 157]. Graphlet [70]. Graphs [268, 140, 1366, 1305, 141, 1336, 1352, 54, 144, 1338, 145, 56, 1346, 1347, 146, 1740, 1363, 1446, 207, 58, 153, 60, 61, 1349, 257, 64, 1365, 1361, 1360, 2211, 154, 1307, 1421, 1729, 1884, 1600, 1708, 2113, 1601, 530, 1353, 1369, 267, 1345, 1972, 1973, 1981, 1241, 1985, 792]. Gravitational [631]. gray [1218]. gray-level [1218]. Greedy [1448, 1875, 1876]. Green [1221]. Grey [2302]. Grey-Valued [2302]. Grid [1755, 1899, 1369]. Grid-Based [1899]. Grids [247]. Grin [1536]. Gröbner [817]. Ground [600, 523, 2319, 1462, 801, 809]. Group [850, 2174, 1274, 1775, 2158]. Grouping [2296]. Groups [2174, 863]. Groupware [1641, 1642]. Growing [990]. growth [282]. Guide [1363]. Guided [1399, 31, 748, 1488, 330, 319, 305, 323]. Guidelines [1546, 410, 411]. gynecologic [307].

Hadamard [2305, 2305]. Hamiltonian [1598]. Handling [1657, 913, 1133, 2042, 2120, 1020, 1208]. haptic [309]. Hard [1592, 2145, 2225, 1697, 1691, 2225]. harder [1456]. Hardness [268, 1298]. Hardness-Randomness [1298]. Hardware [1385, 524, 441, 175, 985, 970, 981, 1016, 2294, 1396, 1770, 1392, 976, 730, 1673, 1394, 1388, 2043, 1386, 1393, 1544]. Harmonic [349]. Harmonisation [912]. Harmonisation/Integration [912]. Harmonised [916]. Hartmanis [1691]. Hash [2177, 1547]. Hashing [2178, 831, 830, 2179, 2240]. Haskell [2105, 1537]. Having [1808]. HCT [132].

HDD [1242]. Head [430, 2212, 407]. Heads [1846, 2378]. Health [1422]. Healthcare [411]. Heart [417, 972, 422]. Heat [109, 111, 138, 137]. Heat-Mass [109]. Heat-Transfer [138, 137]. heavy [1984]. heavy-traffic [1984]. Heilbronn [1687]. Hellman [45, 2158]. Help [1750]. Helping [2127]. HEMATOOL [414]. Herbrand [2061, 2345]. Heredity [1396]. Hermite [132, 2304]. Heterogeneous [1854, 234, 173, 568, 1802, 2098, 2035, 243, 237, 19]. Heuristic [1838, 1252]. Heuristics [1771, 1582]. Hexagonal [1716]. Hidden [65]. Hide [2177]. Hierarchical [1030, 1432, 61, 67, 1022, 2261, 1488, 1219]. Hierarchies [390, 1503, 1381]. Hierarchy [252, 1287, 478, 199, 1831, 185, 1836, 1844, 2230]. High [683, 177, 475, 635, 627, 672, 117, 506, 631, 502, 877, 501, 621, 2324, 359, 1440, 686, 687, 623, 243, 1601, 1791, 786, 1194, 373, 638, 658]. High-Accuracy [1791]. High-Level [475, 786]. High-Order [117]. High-Performance [506]. Higher [505, 1663, 481, 1540, 1304, 1669, 1932, 1909, 1164, 2358, 1908, 2113, 1680, 1681, 2359, 2360, 1682, 403]. Higher-Order [505, 481, 1540, 1669, 1932, 1909, 1164, 2358, 1908, 2113, 1680, 2359, 2360, 1682, 403]. Highly [234, 1443, 1536, 1556, 1799, 1788, 1548]. Highway [714, 1621]. Hilbert [865]. Hill [458]. Hill-Climbing [458]. Hints [998]. hip [322]. HIPERCLASS [621]. HIPERCOMBATS [624]. histogram [296]. histogram-based [296]. Histopathology [419]. Historical [2092]. Hitting [202]. HOL [1161, 1671, 1674, 1681, 1683]. holding [328]. Homogeneous [2249]. Homomorphic [2227]. Homotopy [1277]. Honest [2147, 2148]. Hop [1322]. Hopfield [343, 451, 967]. HORB [1633]. Horizon [119, 2196]. Horizontal [1079, 1590].

Hormonal [955]. Horn [2348, 862, 1150]. Host [1658]. hosts [1956]. Hough [333]. HP [1804, 643, 682, 676]. HP-Convex [676]. HP/ [643, 682]. HPF [945, 1759]. Human [1664, 355, 884, 1352, 1178, 1504, 1048, 427, 341, 1802, 886, 620, 2323, 1963]. Human-Agent [620]. Human-Machine [884, 886, 2323]. Human-Style [1664]. Humanities [1362]. HW [1237]. HW/SW [1237]. Hybrid [222, 1612, 854, 216, 225, 219, 1614, 1613, 1616, 627, 1665, 2027, 432, 1617, 189, 1618, 1161, 223, 33, 972, 218, 1326, 26, 231, 1619, 1760, 230, 1622, 1623, 1644, 1625, 1626, 1847, 1628, 1629, 1630, 227, 763, 1072, 1246, 1265]. Hybrids [1802]. Hydrodynamic [1743]. Hydrodynamics [124]. Hydroelectric [1012]. Hyper [2200]. Hyperbolic [133]. Hypercube [673]. Hypercubes [1684]. Hyperedge [148, 2206]. Hypergraphs [142, 139]. Hypermedia [928]. Hyperset [2059]. Hypertext [1583, 444, 1180, 1037]. Hypertextual [2139]. Hypotheses [2353, 377]. Hypothesis [2339, 386]. HYTECH [1265].

I/O [2099]. IBL [1498]. IBM [1268]. ICE [1515]. Iconic [1384]. Idea [580, 819, 1170]. Ideal [2172, 1272]. ideals [799]. Identification [355, 342, 691, 1903, 1707, 725, 752, 352, 357, 427, 349, 765, 1002, 350, 2028, 2241, 2167, 717, 138, 387]. identifier [1554]. Identifying [1725]. Identity [340]. IEEE [45]. If [935, 87, 700]. If-Conversion [935]. IF-THEN [700]. Ignorance [1112]. II [980, 1116]. IIR [466]. ILF [1143, 1144]. ILF-SETHEO [1144]. iliosacral [323]. Illuminated [2267]. ILP [2331, 2332, 2351]. Image [1554, 594, 749, 426, 1054, 627, 296, 2309, 2308, 430, 748, 705, 743, 1211, 601, 605, 757, 2248, 2268, 1283, 128, 2280, 2014, 2313, 2315, 2284, 2016, 592, 1000, 652, 330, 282, 1221, 319, 305, 1218, 323, 286]. Image-Guided [748, 330, 305]. Imagery [2266]. Images [762, 753, 629, 30, 2249, 740, 2247, 343, 752, 755, 341, 704, 2282, 1833, 2298, 2310, 2295, 2227, 2267, 767, 741, 763, 1210, 298, 316, 294, 1206, 1212, 1207, 290]. Imaging [739, 766, 306]. IMGS [136]. IMGS-based [136]. Immune [847]. Immunization [825]. Impersonation [825]. Implementable [776]. Implementation [508, 980, 1532, 1537, 1007, 643, 512, 1407, 1044, 981, 1016, 685, 507, 589, 676, 1193, 1401, 2294, 477, 1164, 1015, 1776, 730, 1180, 2117, 1681, 79, 1229, 1754, 1543, 1617, 285, 1187]. Implementations [970, 1131, 1624, 875]. Implemented [2125]. Implementing [92, 77, 1402, 1196, 917, 2040, 657, 1174, 2097, 2041, 1965]. Implication [2114]. Implications [2099]. Implicit [1080, 1776, 637, 630]. Implicitly [949]. Importance [822]. IMPROV [100]. Improve [1491, 465]. Improved [140, 258, 1559, 681, 248, 427, 755, 443, 2100, 1131, 1994, 2170, 1229, 1754, 639, 1246]. Improvement [2247, 1897, 316]. Improving [970, 662, 1410, 2279, 1520, 2023, 1455]. IN/ B [911, 905, 1042]. IN/B-ISDN [911, 905, 1042]. Inaccessible [1565]. inaccuracy [314]. Including [1193]. inclusion [808]. Inclusions [1616]. Incomparability [549]. Incomparable [1300]. Incomplete [484, 1901, 1788]. Incompressibility [1815]. Incorporate [1150]. Incorporating [13]. Increases [769]. Increasing [1888]. Incremental [1448, 970, 845, 1970]. Incrementality [831]. Indecision [989]. Independence [188]. Independent [140, 2321, 1556, 1338, 2212, 1728, 846, 1595, 571, 1967, 347]. Index [1418, 853, 1040, 2183, 541]. Indexed [1923]. Indexing [1645, 681, 1181, 1508, 1029, 1485]. Indexings [1843]. Indices [1492]. INDIE [608]. Indirect [1897]. Indistinguishability [708]. Individual [391, 2291]. Induce [955]. Induced [739, 1198]. Inducing [2201]. Induction [1501, 2329, 608, 2053, 1891, 2337, 2340, 1868, 2350, 1680]. Inductive [2329, 577, 2338, 2326, 2343, 2327, 2349, 794, 2384]. Industrial [882, 624, 621, 623, 1238]. Industry [1441, 621, 1463]. inequalities [383]. Inference [25, 1363, 273, 1150, 2010, 2025, 2237, 402]. Inferences [577, 2193, 518]. Inferring [368, 1696]. Infinitary [1916, 2046]. Infinite [1315, 1668, 119, 558, 1260, 1081, 388]. Infinity [2256, 134]. Influence [2223, 1747, 386]. Informal [2049, 1259]. Information [1313, 2375, 1088, 918, 2168, 235, 2177, 1175, 1092, 1057, 31, 963, 1494, 161, 1364, 2034, 2141, 353, 1181, 209, 2324, 928, 864, 1643, 2374, 2373, 2379, 240, 687, 1107, 2264, 236, 1555, 1653, 833, 1639, 2380, 2044, 2378, 237, 2372, 396, 1542, 1463, 388, 1963]. Information-Theoretic [209]. Information-Theoretically [833]. Informations [234]. Infrastructure [564]. Infusion [993]. Inheritance [1200, 278, 2105, 1028]. inhibitor [1077]. Initial [934, 134, 307]. Initiated [1653]. Injuries [415]. Ink [849]. Inline [2112]. innermost [802]. Innocuous [805]. Input [959, 572, 2130, 2010]. Input/Output [572]. Insertions [560]. Insight [466]. Inspection [1016, 621, 1482, 1873]. Inspiration [923]. Inspired [439, 975, 999, 1000]. Instances [201, 2225]. Instantaneous [2273]. Instruction [937, 1403, 652]. instruments [308]. Insulin [993]. Insulin-Pump [993]. Insurance [724]. Integer [181, 1752, 1974]. integral [798]. Integrated [1800, 672, 610, 911, 905, 622, 920, 1042, 915, 321, 1463]. Integrating [1096, 1501, 937, 754, 1859, 1802, 2205, 1171, 1569]. Integration [433, 1147, 1014, 1065, 1199, 1568, 650, 1143, 922, 66, 2379, 1644, 281]. Integrator [957]. Integrity [1019, 1568, 1115, 1194]. Intel [1531]. Intelligence [1435]. Intelligent [166, 27, 417, 1638, 736, 1858, 2034, 538, 703, 972, 1853, 909, 1880, 236, 1555, 1889, 1862, 165]. Intensional [244]. Intensive [572, 1647, 2193, 1862]. Intentions [91]. Inter [918, 926, 2038, 1659]. Inter-Agent [2038]. Inter-Domain [918, 926]. Inter-Organizational [1659]. Interaction [1052, 974, 92, 590, 568, 620, 2080]. Interactions [955, 1781, 310]. Interactive [721, 1050, 102, 1432, 1044, 1143, 1061, 1334, 1056, 1161, 332, 100, 1053]. Interactivity [1062]. Intercellular [959]. Interconnection [247]. Interest [240]. Interface [111, 642, 913, 1504, 654, 514]. Interfaces [887, 1366, 1187, 2323, 1566]. Interfacing [26]. Intermediate [613, 2377, 493, 395]. intermodality [293]. Internal [2221, 3]. International [1650]. Internet [1405, 923, 908, 1122, 1567, 1058, 2041]. Interoperability [919, 1409, 243]. Interoperable [916, 1053, 950]. Interpolation [132, 769, 1512, 127]. Interpretation [426, 1432, 755, 2114, 2063, 9]. Interpretations [1926, 2345]. Interpreting [1879]. Interreflections [767]. Intersect [1410]. intersection [2214, 1848]. Intersections [1693]. Interval [1125, 775, 1740, 200, 2209, 1297]. Intervals [612]. interventional [332]. Interventions [748]. Intracerebral [768]. Intractability [1597, 812]. intranasal [305]. intraoperative [314, 313]. Intravascular [606]. Intrinsic [2235, 205]. Introducing [1534, 2128, 1396]. Introduction [166, 2072, 1135, 2056, 1440, 790, 532]. Introspective [1491]. Intuitionistic [867, 2114, 2058, 2065, 1383]. Invariance [1699]. Invariant [2201, 996, 2281, 2060, 2245, 191, 1004, 816]. Invariants [2242, 2244, 2243]. invasive [302, 303, 305]. Invention [2343]. Inverse [1445, 1774, 2050, 2353, 1146]. Inverted [1615]. Invertible [1207, 2220]. Investigation [105, 466]. Involving [1781]. IP [2041]. IPC [174]. Irregular [661, 638, 504, 1759, 662, 626]. Irrigation [2015]. Isabelle [1678, 1679, 1662, 437, 1681]. Isabelle-Based [1662]. Isabelle/HOL [1678]. Ischemic [422]. ISDN [911, 905, 1042, 912]. ISIS [1063]. Isn’t [50]. iso [287]. iso-surfaces [287]. Isochronous [1655]. Isoefficiency [640]. Isomorphism [2083, 2047, 267]. Isosurfaces [759]. Isothetic [256]. Issue [1058]. Issues [925, 928, 1193, 1853, 1180, 1020, 1648, 233, 10]. Itemsets [1420]. Iterated [2085, 560]. Iteration [1288, 667, 22]. Iterative [106, 115, 121, 1452, 588, 615, 2285, 2266].

J [1332]. Jacket [722]. Jackknifing [750]. Jada [586]. Jam [876]. JANUS [890]. Jape [1166]. JAVA [908, 586, 590, 1033, 1633, 879, 1649, 1937, 1035, 1965, 1631]. Job [2090, 2098]. Join [947, 1917, 1721, 1411, 1127, 1080]. Join-Calculus [1917, 1080]. Joining [1695]. Joins [1410, 1561]. Joint [91]. Josefsson [131]. Just [1736, 84, 275, 1095]. Just-in-Time [275]. Justo [960].

K0*BMDs [525]. Kannel [1957]. kappa [2130, 2230]. KDD [1433, 1430, 1438]. Keeping [2147, 2148]. keg [1552]. Kernel [92, 646, 1655]. Kernels [506, 2260, 1654, 658]. Key [2159, 2181, 2161, 46, 48, 44, 2153, 830, 2162, 833, 821, 843, 828, 1551, 910]. Key-Escrow [828]. Keyed [1529]. Keys [2175]. Keystream [834]. Kind [1519]. Kinds [1115]. Kinematic [289]. Kinematics [427, 1146, 326]. Kleptographic [2163]. Kleptography [824]. Klimov [1984]. KMOV [2161]. knapsack [1978, 1987]. knee [326, 325, 309, 329]. Knowledge [1500, 1350, 1370, 27, 609, 838, 1434, 1359, 432, 31, 1367, 1334, 752, 414, 1093, 755, 1355, 2325, 1419, 2146, 1365, 1360, 1495, 2040, 2036, 1106, 962, 1421, 1108, 1362, 236, 2140, 1868, 2024, 1901, 2193, 1890, 1889, 98, 741, 2019, 1894, 421, 241, 1895, 1439, 412, 1465]. Knowledge-Based [432, 755, 421]. Knowledge-Bases [1370]. Knowledge-Driven [1868]. Knowledge-Intensive [2193]. Knowledge-Theoretic [98]. known [1995]. KOGGE [1179]. Kohn [1615]. Kohn-Nerode [1615]. Kohonen [980, 986]. Kolmogorov [186, 2194]. KYDONIA [1057].

L [1793, 548]. L-Systems [1793]. Label [71]. Labeled [156]. Labeling [430, 1602, 2266]. Labelled [1346, 1340, 1131]. Labelling [601]. Labels [260]. labial [354]. Labyrinth [1395]. Ladder [1521]. Ladder-DES [1521]. Lag [134, 323]. Lagrangian [2009]. Lagrangian-Based [2009]. Lambda [1911, 2119, 397, 1304, 406, 1158, 2368, 870, 408, 394]. lambda-Calculi [1304, 408]. lambda-Calculus [1158, 870]. lambda-sigma [2368]. lambda-to-CL [394]. lambda0s0y0mPA [1922]. lambdacap [2367]. Lanczos [508, 1754]. Land [688]. Lander [1368]. Landmark [765]. Landscape [466]. Language [583, 949, 97, 1534, 1031, 572, 774, 555, 619, 1618, 499, 589, 409, 75, 74, 886, 554, 1676, 1842, 1848, 2378, 2360, 425, 493, 2372, 1566, 1144, 387, 1128, 396, 1123, 271, 1962, 815, 2386, 1959]. language-oriented [1123]. Languages [2184, 549, 1293, 475, 1019, 1536, 529, 1021, 194, 224, 947, 553, 2203, 1189, 582, 953, 1826, 541, 2131, 1727, 270, 1028, 561, 1292, 1725, 2226, 1812, 2204, 79, 556, 395, 22, 7, 1245, 1943, 1261, 1965, 2216]. LANL [2093]. laparoscopic [308]. laparoscopy [307]. Large [1772, 112, 2174, 1046, 513, 1539, 1746, 1692, 1484, 876, 1118, 1944, 1119, 1979]. Large-Scale [1772, 1539]. Largest [1575].

Late [273]. Lately [874]. later [1250]. Lateral [963]. Latin [1550]. Lattice [1804, 823, 644, 2153, 1278, 2051]. Lattice-Boltzmann [644]. Lattices [2188, 800]. Law [1615]. Laws [1107]. layer [1957]. Layered [715, 963, 247, 1109, 1581]. Layout [59, 1322, 69, 72]. Lazy [2364, 1536, 1539, 2110, 2121, 2360]. LBA [540]. Lead [1808]. Learn [224, 1901, 2127, 94]. learnability [372]. Learned [1385, 1187, 657]. Learners [1850]. Learning [28, 618, 1384, 2328, 901, 385, 373, 433, 1491, 1091, 1492, 375, 380, 1014, 575, 2012, 2027, 432, 610, 975, 1496, 619, 573, 958, 696, 576, 1497, 892, 1013, 727, 864, 376, 578, 703, 2341, 1498, 374, 240, 1495, 378, 2314, 615, 2020, 1392, 1390, 620, 2342, 1867, 2344, 2380, 1873, 2348, 2193, 2015, 279, 415, 2018, 412, 379, 381, 386, 377, 388]. Least [128, 640, 136]. Leather [2262]. Leaves [1290, 1216]. Lee [1274]. Leech [1278]. Left [2368, 2371, 804]. Left-Linear [2368, 804]. Left-to-Right [2371]. Legacy [1175]. Legal [1351, 710, 1619, 1632, 1484]. Lemma [1149, 1673]. Length [1574, 2307, 1127]. less [178, 1988]. Lessons [1187, 657]. Letter [1724, 2217]. Level [988, 475, 937, 370, 432, 650, 649, 877, 1215, 601, 1196, 1392, 1388, 2221, 1784, 1786, 1661, 245, 786, 652, 101, 1218, 1938]. Levels [1563, 1438]. Lewis [1302]. LexBFS [153]. LexBFS-Orderings [153]. Lexical [2375, 1108]. Lexicographical [1309]. Lexicons [2374, 425]. Libraries [1489, 1128, 2386]. Library [1562, 25, 643, 173, 670, 1202]. Lie [2022, 1282]. Life [164, 2024]. Lifetime [939]. Liftable [1160]. Lifting [2088]. Light [767, 2081, 1164]. Lighted [1382]. Lightweight [642, 178]. Like [1771, 2301, 38, 2059, 917]. Likelihood [745, 593, 336]. Limitations [666, 2192]. Limited [1574, 978, 1735, 548]. Limits [674, 893]. Linda [492, 1947]. Linda’s [1935]. Line [1321, 1579, 57, 2287, 696, 61, 1592, 1557, 1577, 2296, 1595, 1849, 1969, 1220, 219, 1983, 1997]. Linear [1780, 106, 2083, 202, 1417, 1772, 1269, 713, 110, 1329, 840, 1616, 206, 1731, 204, 970, 2055, 1274, 573, 506, 232, 511, 120, 121, 862, 1005, 1869, 197, 834, 383, 1130, 513, 125, 1281, 2250, 1838, 2220, 2368, 2139, 2226, 2064, 1520, 1575, 2001, 1580, 1752, 2277, 1285, 1951, 1075, 1246, 1972, 1257, 1212, 1773, 1070, 2268, 804, 2300, 811, 810]. Linear-Time [1130, 1838, 1580, 2277, 2001]. Linearly [1824]. Lines [2242, 2244, 2283, 2263, 760, 1209]. Linguistic [2377]. Linguistics [1435]. Link [444, 1277]. Linkage [2219]. Linked [951, 1801]. Linking [587, 1795]. Links [1606, 1870, 1871]. Lip [337]. Lipreading [2292]. LISP [252]. list [1982]. Lists [274, 374]. Lite [1404]. Literals [2330]. Liveness [1665]. Living [428]. LJ [1148]. LK [1148, 2068]. LK-to-LJ [1148]. LLL [2160]. Llull [771]. Load [1579, 1593, 1233, 1231, 645, 1322, 176, 2023]. Loadable [1654]. Local [1010, 80, 7, 1636, 1918, 2299, 1968, 534, 4, 1993]. Localising [334]. Localities [492]. Locality [681, 492, 1843, 1958]. Localization [34, 1002, 2288, 932, 598, 317, 284]. Locally [1499, 1816]. Locating [2291]. Location [1556, 1657, 2272]. Loci [743, 1211]. Lock [940, 357]. Lock-control [357]. Locking [1657]. LOD [1646]. LOD-based [1646]. Log [2321, 2162, 2163]. Log-Based [2162]. Log-Polar [2321]. Logarithmic [551]. Logarithmic-Space [551]. Logarithms [836]. Logic [491, 1663, 2329, 2355, 483, 1385, 167, 1098, 517, 25, 1916, 469, 2107, 370, 2048, 775, 2027, 1092, 2055, 2356, 984, 73, 1473, 860, 718, 523, 516, 1669, 862, 701, 1169, 2125, 1813, 2338, 367, 2340, 865, 369, 1378, 866, 2326, 2124, 1314, 757, 361, 2205, 869, 1106, 26, 2120, 1108, 531, 2058, 2138, 1620, 2343, 2064, 2065, 2327, 2349, 2350, 1680, 1719, 2360, 1682, 1113, 1310, 165, 712, 733, 1472, 1075, 1469, 1470, 1465, 1070, 1254, 2387, 2384]. Logical [1384, 92, 400, 1823, 2336, 1115, 2341, 195, 1915, 1927, 2082, 2089, 2081, 1383, 871]. Logicality [471]. Logically [1111]. Logics [2052, 517, 2055, 1825, 190, 1155, 196, 1164, 1487, 2063, 1852, 2382]. Logs [874]. LOKI91 [1520]. LOLA [1475]. Longitudinal [1879]. Look [1034, 1793]. Lookahead [2331]. Looking [2054]. Loop [931, 930, 1325, 952, 666, 932]. LooPo [952]. Loops [943, 1293, 1758, 2109, 942]. Loose [1496]. Loss [1735, 1126]. Lossless [2268]. Lossy [490]. LOTOS [789, 496, 790]. Low [1571, 1578, 343, 1048, 307]. low-cost [307]. Low-Power [1578]. Lower [1607, 1740, 558, 2191, 1849, 836, 1701, 382, 1548]. LP [520, 1666, 349]. LP-residual [349]. LPs [1998]. LR [2130]. LTL [1129]. Lu [507]. Lucas [38]. Lukasiewicz [2055]. lumbar [289]. LV [283]. Lyapunov [214].

M [740, 893, 2271]. M-A [893]. M-Estimators [2271]. M-Mode [740]. M2L [1263]. Mach [1654]. Machine [674, 884, 2258, 1492, 1504, 906, 886, 1392, 445, 620, 2323, 1867, 2380, 1163, 415, 412, 1073, 1969, 818]. Machine-Supported [1163]. Machines [673, 1579, 947, 1162, 1832, 2192, 2223, 2232, 2383]. Macrame [665]. Macro [932]. Macro-Dataflow [932]. Macromutation [454]. Magic [1150, 849]. Magnetic [108, 431, 757, 315]. Magnetohydrodynamics [630]. Magnitude [339]. Maintaining [1573, 1562, 1613, 255, 1889]. Maintenance [733, 2385]. Major [1641]. Making [2178, 2198, 2033, 1047, 422, 900]. Mallat [2274]. Mallat-Wavelets [2274]. Mammograms [760]. Man [825]. Man-in-the-Middle [825]. Management [994, 882, 1051, 562, 679, 1019, 919, 923, 672, 878, 873, 916, 1558, 921, 929, 184, 880, 926, 920, 1365, 917, 1028, 879, 881, 1651, 1841, 1630, 1045, 915, 1650, 1979]. Managing [1855]. Maneuvers [221]. MANIFOLD [1749]. manifolds [756]. Manipulating [1408]. Manipulation [1051, 668, 525]. Manipulators [514, 1146]. manual [2386]. Manufacturing [1581]. Many [1859, 251, 1685, 2225]. Many-One [251]. Many-to-One [1685]. Map [695, 1206]. MAPLE [120, 500]. Mapping [1593, 1802, 1709, 315]. Maps [1014, 758, 1008, 34, 694]. Margins [939]. Mark [1016]. Marker [1358]. Markers [1802, 751]. Market [680, 1043, 891]. Market-Aware [891]. Marketplace [164]. Marketplaces [1405]. Markov [594, 1119, 593, 2196]. Markov’s [2058]. Mars [1368]. MAS [816, 2041]. Mass [109, 631, 726]. Massively [572, 1228, 633, 1229, 1754, 639]. Match [1021]. Matched [334]. Matching [1583, 603, 1584, 1443, 397, 1478, 1442, 611, 1168, 291, 602, 1445, 2379, 1149, 1833, 2310, 2271, 2371, 597, 2290, 285, 324]. Matchings [1685]. Matchmaking [162]. Material [751, 1632]. materialize [5]. materialized [16]. Mathematical [1153, 1885, 2049]. mathematics [1069]. Matrices [1772, 1707, 118, 123, 126, 2305]. Matrices-Nonsingularity [123]. Matrix [638, 1712, 1734, 375]. Matroidal [1337]. Matter [903]. max [197, 185]. Maxima [255]. Maximal [1271, 381]. Maximisation [602]. Maximum [745, 2334, 593]. mcd [1142]. McEliece [2159]. MDL [964]. Me [874]. Means [1100, 644, 1003, 998, 1887, 872, 2068]. Measure [1299, 1882, 2345, 1877]. Measurement [36, 1124, 754, 767]. Measurement-Theoretic [36]. Measurements [993, 288]. Measures [1091, 210, 1420, 1297, 296, 1212]. Measuring [198]. mechanical [310]. Mechanism [1490, 1733, 1735, 1868].

Mechanisms [168, 958, 278, 544, 585, 22]. mechanization [1244]. Media [1952, 1774, 1644, 2132, 2133]. Media-Based [1644]. Medial [2283]. Median [1710]. Medical [762, 426, 416, 424, 409, 743, 422, 425, 306]. Meets [239]. MEGA [632]. MEGA-D [632]. Melting [634, 138, 137]. Membership [469, 1842, 2348, 1887]. Memory [1234, 931, 943, 2096, 661, 2165, 29, 945, 655, 343, 1684, 1593, 654, 2093, 184, 845, 1828, 628, 451, 1390, 1762, 2013, 2231, 1229, 1754, 639, 1545, 1990]. Memory-Based [1390]. Memory-Bounded [2165]. Menger [1972]. Menu [1366]. Menu-Based [1366]. Merges [874]. Mergesort [263]. Merging [1563, 1327]. Merkle [2158]. Merlin [2192]. Mersenne [2311]. Mesh [1819, 681, 24, 1843]. Mesh-Indexings [1843]. Meshes [116, 1742, 1988, 1994]. Message [1526, 534, 2159, 655, 685, 1525, 2182, 1550]. Message-Passing [655, 685]. Message-Resend [2159]. Messages [566]. MESSENGERS [1634]. Meta [1563, 370, 432, 1182, 1860, 1438, 1179]. Meta-CASE [1179]. Meta-Level [370, 432]. Meta-Levels [1563]. Meta-Modelling [1182]. Meta-Reasoning [1860]. Metabolic [1800]. Metacomputing [641, 674, 1647]. Metadatabase [239]. METAFrame [530]. Metaphor [575]. Metaspaces [1656]. METATEM [98, 1962]. Method [171, 439, 104, 508, 1177, 1428, 1298, 214, 470, 603, 1141, 2052, 1134, 1311, 441, 495, 1228, 509, 507, 2316, 1799, 2105, 670, 1155, 765, 1028, 126, 419, 731, 127, 130, 1111, 2295, 1763, 1520, 2299, 2312, 2025, 2050, 1139, 1569, 779, 1931, 1229, 639, 2019, 1554, 1246, 1238, 292]. Methodologies [2039]. Methodology [1502, 1170, 429, 2030, 915]. Methods [1772, 709, 461, 113, 465, 1199, 1135, 116, 1858, 1496, 122, 1897, 197, 435, 1776, 128, 2043, 134, 1215]. Meticulous [263]. Metric [1271, 2201, 2084]. Metrical [423]. Metrics [616]. Micro [1655]. Micro-kernel [1655]. Microcircuits [954]. Microelectrode [995]. Microprocessor [692, 1249]. microrobot [331]. Microscope [629]. Microscopy [2248]. Middle [825]. Migrant [173]. Migration [922, 914, 1762]. Migratory [584]. Mill [1858]. MIMD [1766]. Min [1998, 185]. Min-Sum [1998]. Mind [903, 387]. Minder [96]. Minimal [943, 1203, 2217, 942, 1114, 856, 286, 1229, 639]. Minimally [1619, 302, 303]. Minimax [119, 382]. Minimisation [2310]. Minimization [603, 600, 607, 592]. Minimizing [1306, 1721]. Minimum [1516, 211, 1336, 1269, 396, 1599, 1604, 939, 68, 1277, 195, 1992, 1838, 1723, 377, 1994]. Minimum-Link [1277]. Minimum-Width [1604]. Mining [1436, 1087, 1437, 2332, 1441, 1891, 2335, 1406, 1440, 1564, 1888, 1874, 1435]. Minkowski [2251]. MIPS [1752]. Mismatch [1580]. Missing [1900]. Mistrusting [48]. MISTY [1514]. Mixed [722, 730]. Mixin [2106]. Mixin-Based [2106]. Mixture [746, 2291]. Mixtures [968]. ML [1667, 1080, 1909, 2127]. MLP [971, 998]. MLPs [982]. MMH [1525]. Mobile [587, 583, 567, 1657, 579, 565, 570, 580, 569, 582, 80, 568, 535, 588, 34, 563, 1658, 566, 1015, 1390, 585, 1649, 2238, 2032, 1656, 1956, 564, 2019, 1530, 1949]. Mobility [562, 924, 1658, 907, 1958]. Modal [481, 520, 2052, 351, 1315, 1813, 1155, 869, 1106, 1620, 1927, 227]. Mode [740, 51, 736, 2318]. Model [745, 2370, 1041, 956, 722, 1177, 527, 1227, 949, 1316, 1249, 475, 1268, 749, 707, 1804, 522, 91, 1344, 1315, 1065, 680, 739, 873, 859, 646, 420, 444, 343, 1811, 911, 910, 1048, 633, 755, 748, 538, 1129, 965, 1397, 960, 2124, 476, 1254, 2279, 1006, 2010, 1037, 1855, 2029, 2317, 1918, 1786, 1678, 1660, 2267, 1752, 2360, 617, 1144, 1931, 1958, 8, 1266, 1076, 399, 282, 1960, 1260, 1257, 304, 1985, 1251, 1265, 1261, 2004]. Model-Based [1678]. Model-Checking [475, 1931, 1076]. Modeled [222]. Modeling [706, 1234, 1097, 746, 1502, 1789, 1352, 1031, 1800, 1950, 29, 220, 455, 2381, 2334, 1351, 656, 1507, 877, 223, 688, 1946, 2094, 1793, 1810, 1765, 1509, 712, 1116, 1124, 1121, 297, 310, 1959]. Modelization [1010]. Modelling [1185, 1563, 242, 1172, 93, 1182, 928, 604, 1881, 1904, 903, 1774, 2032, 2079, 767, 1902, 760, 1542]. Models [762, 2134, 594, 433, 112, 1373, 2078, 1798, 995, 1376, 1799, 604, 1932, 1094, 1378, 1866, 2314, 1872, 742, 550, 2210, 2061, 1883, 1921, 1630, 1907, 759, 23, 747, 1119, 287, 1462, 406, 284, 1777]. Modern [2149, 1058]. Modification [2025]. Modified [979]. MODp [188]. MODp-tests [188]. Modular [487, 1304, 1470, 698, 2146, 2076, 1524, 1966, 1787, 498]. Modularity [1072, 1718]. Modulated [1896]. Module [489, 1541, 1782, 401]. Modules [1593, 1746]. Modulo [1758, 1693]. Modulus [2084]. Molecular [1342, 1792, 452, 1797]. Moments [505, 2306, 1127]. Monadic [1296, 531]. Monads [2076]. Monitor [780, 421]. Monitoring [1879]. Monocular [2289]. monoidal [406]. Monoids [1845, 795, 796]. Monotone [872, 380, 374]. Monotonic [516, 518, 1456, 1474, 1477, 1468, 1471]. Monte [593, 121]. MORAL [2238]. Morphisms [559, 1827, 1724]. Morphological [2294, 2282, 2285]. Morphology [1004, 312]. morphology-based [312]. Mosaicing [2249]. MOSEL [531, 1263]. MOSIX [647]. Most [2339, 11]. Motion [753, 1571, 2321, 607, 2247, 601, 2319, 2257, 2274, 283, 322]. Motion-Based [601]. Motivational [96]. Motor [957, 958, 718, 1013]. Mouse [2289]. Mouth [336]. Movement [2321, 180, 958, 313]. Movements [337].

Movie [163]. Moving [1573]. MPEG [1553]. MPEG-encoded [1553]. MPI [670]. MPPs [2094]. MR [753, 295, 316, 294, 769, 752, 755, 291, 754, 761, 741]. MR/CT [316]. MRA [768]. MRF [2276]. MRI [749, 751]. Mu [1315, 1266]. Mu-Calculus [1315, 1266]. mucke [1266]. Müller [1284]. Multi [1686, 1213, 1819, 97, 90, 351, 101, 244, 1665, 242, 1175, 575, 897, 985, 827, 573, 896, 1785, 89, 895, 2037, 162, 578, 2030, 271, 2042, 2040, 2036, 2212, 1218, 2035, 902, 2031, 915, 245, 2237, 571, 1423, 900, 1226]. Multi-Agent [97, 90, 244, 242, 575, 897, 573, 896, 89, 895, 2037, 162, 578, 2030, 2042, 2040, 2036, 2035, 902, 2031, 571, 900]. Multi-Architecture [1785]. Multi-Authority [827]. Multi-domain [915]. Multi-Head [2212]. Multi-language [271]. Multi-Level [245, 101]. Multi-Mesh [1819]. Multi-neural [985]. Multi-perspective [1175]. Multi-Processor [1665]. Multi-relational [1423]. Multi-scale [1218, 1226]. Multi-sensor [2237]. Multi-spectral [1213]. Multi-target [271]. Multi-tree [1686]. Multiagent [243, 245, 891, 2385]. Multicast [2173, 1046, 1652]. multiclass [1984]. Multicomputers [1312, 513]. Multicriteria [729]. multidatabases [17]. Multidimensional [1418, 1645, 986, 999, 1880]. Multigrid [2276]. Multilevel [2012, 60]. Multilingual [2374]. Multimedia [1041, 1060, 1755, 1050, 1064, 1051, 1049, 1059, 1065, 1057, 1044, 410, 1047, 1056, 1068, 924, 1637, 1067, 1053, 2137, 907, 1042, 1058, 2138, 1653, 1644, 1045, 1066, 13]. Multimodal [2133]. Multiobjective [2303]. Multiparty [907]. Multiple [1584, 2335, 1452, 2320, 1704, 1505, 2286, 2025, 617, 1078]. Multiple-Data-Batches [617]. Multiple-Object [1704]. Multiplication [1607, 1734]. Multiplications [1524]. Multiplicative [2160]. Multiprocessor [1782, 2097, 2017, 1967]. Multiprocessors [931, 943, 2096, 1784]. Multiprogramming [2102]. Multiresolution [1688]. Multiscale [1054, 1212]. Multisensor [1086]. Multiset [75]. Multispectral [2266]. Multisplitting [1426]. multistage [1126]. Multitape [1727]. Multiterminal [1716]. Multithreaded [949, 1769]. Multithreading [1761]. Multivariate [1270]. Multiway [1714, 178]. Munich [1068]. Muscle [427]. Must [539]. Mutation [461]. Mutual [1737]. Mutually [48]. Myocardial [752].

Naccache [44]. Naccache-Stern [44]. NaDSyl [2057]. Name [1025, 1636]. Narrowing [2364, 2354, 2103, 2360]. Natural [553, 619, 2239, 886, 554, 2378, 425, 1566, 1144, 1206, 798]. Navier [511, 637]. Navigation [1388, 2319, 1861, 306]. NCCA [179]. Near [1022, 1595]. Nearest [1499]. Nearly [1714, 380]. Necessary [959]. Necklaces [156]. Need [1140, 276]. Needed [2362]. needs [1463]. negation [4]. Negotiate [94]. Negotiation [2038]. Neighbor [1695, 1499]. Neighbor-Joining [1695]. Neighborhood [463]. Neighbourhoods [1181]. Nelson [1141]. Nelson-Oppen [1141]. Neonatal [994]. Neonates [991]. Nerode [1615]. Nested [1347, 1558, 2109, 17]. Net [980, 1136, 2366, 1839, 1746, 1632, 1085]. NetCard [41]. Nets [1538, 1132, 1821, 1134, 1138, 1820, 1137, 533, 1135, 867, 223, 1131, 367, 1130, 1133, 532, 1839, 1930, 1716, 1226, 1077, 1123, 405]. Network [956, 993, 901, 562, 923, 653, 1364, 921, 929, 170, 173, 990, 906, 920, 1009, 965, 688, 961, 1046, 987, 1881, 665, 962, 2101, 615, 1042, 1012, 1390, 730, 967, 176, 2278, 2281, 2024, 1651, 1784, 1747, 1652, 1883, 2231, 1659, 1722, 338, 1004, 1439, 179]. Network-Based [993]. Network-wide [179]. Networked [643, 677]. Networks [171, 1041, 1384, 994, 1385, 177, 974, 1321, 977, 1010, 1059, 691, 1091, 1136, 1007, 1822, 512, 1768, 2309, 985, 51, 1824, 715, 1792, 963, 555, 2276, 1016, 431, 1618, 192, 458, 978, 992, 1322, 538, 247, 697, 208, 2291, 1003, 698, 968, 920, 716, 1870, 1055, 737, 1011, 909, 2020, 1015, 1001, 1110, 996, 1761, 1735, 1648, 1901, 1750, 1730, 717, 2275, 1871, 1066, 2018, 1127, 384, 1984, 1121]. Neumann [973]. Neural [1771, 956, 994, 988, 957, 993, 974, 977, 1010, 1007, 512, 2309, 983, 2301, 715, 963, 2276, 1016, 958, 431, 990, 978, 992, 726, 2291, 1003, 698, 968, 1009, 965, 716, 961, 987, 693, 1870, 737, 1011, 962, 615, 2020, 1015, 1012, 1001, 1390, 730, 967, 996, 2281, 2024, 1747, 1883, 1750, 717, 2275, 1871, 1004, 2018, 1439, 384, 985]. Neural-Like [1771, 2301]. Neuro [982, 1886, 421]. Neuro-Computer [982]. Neuro-Fuzzy [1886]. Neuro-Fuzzy-Classifier [421]. Neurodynamics [966]. Neuroendocrine [955]. Neurofuzzy [1881]. Neurons [1384, 2275]. neurosurgery [302]. neurosurgical [304, 303]. Neutrality [1807]. Newton [130]. Newtonian [104]. Next [1630]. NIC [964]. Nice [257]. Nicolson [111]. NL [1691]. NL-hard [1691]. NLP [1435]. Node [1417, 142, 1337, 665, 1713, 1723]. Node-Deletion [1337]. Nodes [65, 1976]. Noetherian [799, 810]. Noise [745, 614, 2317, 1692, 377]. Noisy [839, 2298, 1902]. Nominal [2337]. Non [2355, 1772, 1922, 534, 1503, 282, 512, 2201, 2253, 553, 83, 1160, 1334, 523, 516, 1102, 1773, 1193, 254, 1150, 305, 998, 765, 2268, 637, 1477, 2300, 2240, 2068, 1257, 1456, 1462, 1474, 1245, 1468, 1471]. Non-archimedean [2253]. Non-confluence [1922]. Non-constrained [2068]. Non-Context-Free [553]. Non-degeneracy [254]. Non-deterministic [2355].

Non-elementary [1102]. Non-functional [1193]. Non-Ground [523, 1462]. Non-Horn [1150]. Non-Interactive [1334]. Non-invasive [305]. Non-Liftable [1160]. Non-linear [1773, 2268, 2300, 1257]. Non-local [534]. Non-Monotonic [516, 1477, 1456, 1474, 1468, 1471]. Non-overlapping [637]. Non-recurrent [512]. Non-refinable [1503]. non-regular [1245]. Non-rigid [282, 765, 2240]. Non-Standard [83]. Non-Symmetric [1772]. Non-trivial [2201]. noncommutative [817]. Nondeterminism [1300, 1342, 198]. Nondeterministic [2199, 1401, 183, 1844]. None [2187]. Noninvasive [766]. Nonlinear [713, 1616, 1518, 726, 1005, 1851, 1221, 1205, 1214, 1548]. nonlinearity [1548]. nonlocal [1221]. Nonmonotonic [1103, 2382, 518]. Nonmonotonicity [1105]. Nonperiodic [1286]. Nonregular [1324]. Nonsingularity [123]. Nonsymmetric [511]. Norm [2356]. Normal [1152, 1309, 2338, 750, 1851, 872]. Normalisation [347, 335, 1908, 802]. normalisations [408]. Normalization [1669, 2369, 394]. Normalized [2253]. Norms [894]. Notarization [49]. Note [549, 461, 1824, 260, 581]. Notes [135]. Nothing [1528]. Nourishment [1733]. Novel [570]. Novices [2127]. NOWs [172]. NP [251, 1302, 2214, 2225, 1848]. NP-hard [2225]. NSM [175]. NTRU [823]. NUMA [2096, 654]. Number [2164, 1599, 2172, 1447, 65, 1693]. Numbers [708, 1294, 384, 1255]. Numerical [104, 105, 2330, 109, 110, 113, 122, 1426, 124, 2008, 134, 135, 137]. Nuprl [1161, 1164]. Nuprl-Light [1164].

OBDD [2223]. OBDD-Representations [2223]. OBDDs [1694]. Oberon [277, 276, 281]. Object [491, 439, 2303, 603, 1019, 1534, 1054, 1800, 1924, 1021, 643, 1199, 2287, 2104, 940, 1036, 588, 74, 2131, 1957, 2137, 1028, 599, 1704, 2252, 2278, 2238, 1037, 1025, 1020, 1654, 2241, 1676, 1636, 2286, 2288, 1660, 2240, 1635, 1566, 574, 598, 1960, 309, 20, 1965, 1646]. Object- [1037]. Object-Based [940, 74]. Object-Oriented [491, 1019, 1534, 1800, 1021, 643, 1199, 1036, 2131, 1028, 1020, 1676, 1636, 1566, 1957, 20, 1965]. Object-Stacking [1660]. Object-Z [439]. Objective [2098]. Objective-Oriented [2098]. Objects [1418, 2302, 782, 735, 1030, 484, 650, 504, 277, 1596, 743, 1634, 1039, 276, 585, 1023, 1034, 1032, 1877, 397, 1544, 1224]. Oblique [1869]. Oblivious [841, 1988]. Observable [422, 2196]. Observables [1318]. Observation [536]. Obstacle [1571]. Obstacles [1277]. Obtained [750]. Obtaining [2324]. Occam [614]. Occluded [1004]. Occurrence [1839]. ocular [327]. Oculo [957]. Oculo-Motor [957]. Odd [1278]. ODE [117]. ODP [781]. Off [1841]. Offline [2352]. offs [1298]. Offset [1603, 1587]. Offset-Polygon [1603]. Oh [101, 1856]. Oil [632]. Olan [1959]. Olfactory [1007]. Olufsen [1137]. omega [1287]. omega-rational [1287]. omegaMEGA [1153]. OMG [565]. Omnidirectional [279]. On-Line [1321, 1579, 696, 1592, 1557, 1595, 1849, 1969, 219, 1983, 1997]. On-the-fly [1251]. Once [1300, 2214]. Oncology [418]. One [838, 2172, 2307, 116, 256, 187, 2340, 2377, 251, 266, 1323, 130, 2218, 1685, 809, 810]. One-Counter [1323]. One-Dimensional [2218]. One-Separation [256]. One-Step [2307, 809]. One-Way [838, 266, 187]. One-Word [2377]. Ones [1709, 1988]. Online [1387, 281]. Onset [991]. onto [1608]. Ontological [2379]. Ontologies [244]. OO [439]. Op [1398]. Opal [499, 817]. Open [1064, 1293, 2115, 1065, 1937, 1940]. Open-Ended [2115]. Opera [1414]. Operating [170, 1544]. operation [1948]. Operational [1926, 2033, 1340, 688, 2132, 2079, 1369]. Operations [482, 1358, 820, 794]. Operator [122, 790]. Operators [2106, 840, 1407, 2282, 185, 2026, 2297]. ophthalmic [311]. Oppen [1141]. Opportunities [1853]. Opt [1582]. Optical [1321, 2265, 2317]. Optimal [21, 1910, 2355, 1714, 838, 345, 635, 1688, 146, 256, 119, 1976, 645, 477, 1985, 1589, 1702, 1704, 1022, 2219, 1595, 2371, 1843, 1848, 1531, 1715]. optimality [1984]. Optimally [827, 2260]. Optimisation [2313, 1875, 1876, 1561]. Optimising [1536]. Optimization [1771, 524, 1756, 713, 572, 2007, 457, 684, 633, 892, 588, 2273, 599, 967, 1581, 2009, 2008, 2016, 719, 732, 459, 595, 1117]. Optimizations [937, 930, 935]. Optimized [1559]. Optimizing [723, 57, 346, 1533, 694, 1029]. Optimum [248, 1335, 249]. Oracles [2177]. orbital [288]. Order [2364, 505, 1663, 481, 1147, 913, 117, 450, 1540, 1304, 1669, 1355, 526, 362, 1932, 1909, 865, 1378, 1327, 1831, 1164, 1279, 243, 1718, 531, 2358, 848, 2162, 1023, 1908, 2113, 2064, 2347, 2350, 1680, 1681, 2359, 2360, 1852, 2233, 2352, 1682, 1310, 134, 870, 871, 2297, 1258, 1951, 403, 809, 818, 810]. Order-Sorted [1718, 818]. Ordered [1300, 1739, 53, 1159, 1628, 193]. Ordering [479, 1694, 801]. Orderings [1096, 480, 153]. Orders [1160, 2228, 808, 1075]. Ordinal [387]. Ordinary [113, 1273]. Ordinates [509]. ordination [1964]. Organic [1640]. Organisations [1176]. Organization [957, 975, 1355, 2379, 1782]. Organizational [29, 1659, 1642]. Organized [1438, 241]. Organizing [1014, 758, 1008, 695, 694]. Orientation [1822, 770, 1207]. Orientations [1686]. Oriented [491, 1019, 428, 1534, 2262, 1800, 1021, 643, 1199, 883, 1036, 1637, 2131, 2098, 1028, 1026, 1524, 1037, 1020, 1676, 1636, 1034, 1566, 1430, 1123, 1957, 20, 1965]. Orthogonal [1599, 1138, 57, 63, 1412, 1330, 1601, 1978].

Orthomodular [2197]. orthopedic [330]. Osaka [279]. OSAM [910]. Oscillations [955]. Oscillatory [959]. Other [154, 848, 1963]. Out-of-Core [661]. Outliers [1898, 1902]. Outlook [711]. Output [1144, 1948]. Overcome [674, 2274]. Overcoming [666]. Overhead [940, 1748, 1990]. Overlapped [660]. Overlapping [1626, 637]. Overlaps [1851]. overload [1120]. Overloading [1682]. Overriding [2106]. Overshootings [727]. Overview [1092, 2325, 1814].

P [2214, 134]. P-stable [134]. P1363 [45]. P4 [141]. P4-connected [141]. P506 [912]. PA [1930]. PA-Processes [1930]. PAC [1498]. Package [1528, 816]. Packages [500, 875]. Packet [1059, 1685, 1956]. packing [1968]. Pagenumber [53]. pages [1990]. Pain [413]. Pair [768]. Pairwise [1606]. Pallets [1608]. Palliating [2377]. Pan [916]. Pan-European [916]. Panacea [435]. Pancreatic [959]. Panel [1546]. Paper [1166]. Parabolic [116, 133]. Paradigm [831, 985, 1362, 1783]. Paradigms [988, 2337, 563]. Parallel [1739, 675, 2303, 1780, 508, 1234, 669, 53, 106, 949, 542, 1236, 177, 624, 647, 943, 783, 780, 1756, 668, 110, 842, 572, 552, 642, 643, 677, 1228, 1312, 512, 1768, 947, 680, 146, 180, 682, 655, 650, 664, 555, 207, 1891, 1684, 1593, 168, 649, 1785, 638, 953, 504, 644, 1741, 1781, 150, 507, 645, 633, 632, 941, 1773, 663, 2090, 514, 622, 1235, 676, 678, 1828, 670, 662, 2092, 626, 628, 2098, 1742, 1589, 1302, 2109, 1421, 2120, 1776, 637, 634, 948, 176, 950, 630, 1743, 1775, 1783, 1744, 1717, 1594, 1763, 1778, 2283]. Parallel [1748, 1786, 1749, 1750, 1779, 1766, 1791, 2102, 1767, 79, 1734, 1788, 640, 1411, 652, 808, 1961, 1984, 804, 1999, 2383]. Parallel/Distributed [664]. Parallelisation [660]. Parallelism [937, 651, 1762]. Parallelization [931, 629, 572, 661, 510, 509, 656, 671, 666].

Parallelized [940]. Parallelizer [952]. Parallelizing [941, 659, 2357, 1748]. Parameter [2303, 2053, 966, 1198, 2313, 1210]. Parameter-Induced [1198]. Parameterised [2359]. Parameters [593, 2016]. Parametric [2364, 216, 299, 939, 998, 1594]. Parametrically [591]. Paramodulation [2045]. Parasitic [961]. Parity [193]. Parsimony [1697]. Part [2333]. Part-of-Speech [2333]. Partial [1258, 527, 1951, 2177, 1168, 526, 364, 1341, 158, 1706, 2272, 2204, 1852, 2233, 1112, 808, 1075, 6, 794]. Partial-order [1258]. Partializing [485]. Partially [1739, 1331, 422, 2196]. Partially-Observable [2196]. Particle [670, 683]. Partition [601, 1335, 249]. Partitioned [2199, 1561]. Partitioning [943, 1559, 1181, 1511, 1328, 2014, 1738, 1715, 1411, 143]. Partitions [145, 592]. Partly [1565]. Parzen [1896]. Parzen-Windows [1896]. Passing [655, 685, 1918, 1943]. Passive [2323]. Past [467]. Path [1591, 152, 1322, 1277, 1602, 801]. Paths [1740, 1363, 1590, 182, 1320, 1606, 1314, 1341]. Pathway [411]. Patient [762, 2247, 305]. Patient-Specific [762]. Patients [1879]. Pattern [1583, 1443, 1792, 973, 1425, 1445, 1833, 2371, 2299, 2025]. Pattern-Matching [1833, 2371]. Patterns [1905, 1535, 1036, 1039, 496, 996, 2031, 1037, 2142, 1038, 2025, 1961]. paxaconsistent [1466]. PC [553]. PCCG [1369]. PCF [399, 407]. PC’s [647]. Pebbles [1846]. Pedagogic [2126]. Pedestrian [2239]. pedicle [319]. Peers [1142]. Peers-mcd [1142]. PEFPT [667]. Peirce [1345]. Penalties [1711]. [289, 290, 643, 682]. 2 [2116]. Ada [1193]. B-ISDN [911, 905, 1042]. CPN [533]. CT [316]. Distributed [664]. E [217]. HOL [1678].

Integration [912]. MAC [1527]. O [2099]. Output [572]. PC [848]. Resume [790]. second [1525]. SNMP [921]. SW [1237]. unfolding [2354]. Video [1137]. Wavelet [627]. Pendulum [1615]. PENELOPE [1117]. Pentium [1531]. PEP [1262, 528]. Perceptual [1008]. percutaneous [301, 321]. Perform [1692, 2193]. performability [1116, 1117]. Performance [683, 1695, 1562, 647, 1756, 1178, 635, 627, 677, 970, 453, 506, 190, 1558, 631, 638, 502, 633, 663, 659, 501, 621, 1634, 359, 1440, 1881, 686, 687, 623, 2279, 1765, 2102, 658, 1122, 1124, 1121]. Performance-Enhanced [453]. PERicardial [300]. Permutable [1286]. Permutation [1740, 2242, 1708, 2243]. Permutations [266]. Perpetual [1325]. Perpetuality [2369]. Persistent [1027, 376]. Person [351, 356, 352, 353, 338]. Personal [1643]. Personalities [99]. Personalized [928]. Perspective [1051, 789, 1419, 886, 1348, 1935, 1175]. Perspectives [1490, 77, 1094]. Persuasion [2038]. PET [1212, 747]. Petri [1538, 1132, 1821, 1134, 1138, 1820, 1136, 1137, 533, 1135, 367, 1123, 1130, 1133, 532, 1839, 1746, 1930, 1085]. pH [949]. Phantom [2293, 317]. Phase [109, 339, 1133, 134, 226, 295]. Phase-Lag [134]. Phase-Type [1133]. Phase-weighted [339]. Phenomena [2005, 1773]. phone [1122]. Phoneme [1003]. Phonetic [2328]. Photos [344]. Phrasal [1436]. Phylogenetic [1447, 990, 1712]. Phylogeny [1695]. Physical [1709]. Physics [672]. Pi [2189]. Pi-Calculus [2189]. Picture [2276, 1310]. Pictures [2302]. Piecewise [1295, 970]. piercing [1999]. PIERS [622]. Pilot [1063]. Pipeline [1758, 936]. Pipelining [1752]. Pitfalls [1148]. PiZA [436]. Placement [1603, 71, 944, 323]. Plan [426, 424]. Planar [53, 152, 257, 64, 154, 1972, 1981]. Planarity [1611, 1586]. Plane [52, 1597, 1277, 2256, 2319]. Planetary [2318]. Planning [99, 1571, 610, 1504, 2286, 423, 302, 327, 325]. Plans [1638]. Plant [734, 1012, 717]. Plants [1010, 1793]. Plastic [955]. Platform [1637, 1965]. Platoon [221]. Plausibility [1173]. Plausibility-Centred [1173]. Play [2150]. PLC [776]. PLC-Automata [776]. PLTL [872]. Plug [2150]. Point [1573, 1604, 271, 2243, 2285]. Pointer [951]. Pointer-Linked [951]. Pointers [174]. Points [964, 1689, 2242, 1827, 961, 1589, 2244, 1732]. Polar [2321, 283]. Polarized [2068]. Polaron [104]. Policies [1172]. Policy [1546, 2196, 1543]. Political [2095]. Pollution [112]. Polygon [1603, 1587, 1577]. Polygon-Offset [1587]. Polygonal [1688, 1572]. Polygons [256, 1412, 1595, 1598]. Polymorphic [1910, 76, 483, 870]. Polymorphism [78, 1909, 1912, 530]. Polynomial [1526, 1271, 252, 115, 259, 187, 2146, 1671, 199, 1831, 185, 1832, 2216, 2191, 2232, 799]. Polynomial-time [187]. Polynomials [1301, 1282, 1549]. Polyploidy [460]. Polytopes [1586]. Polyvariance [78]. pomset [1085]. Pool [2096]. Pool-Based [2096]. Poppycock [435]. Population [1897]. PORT [1362]. Portable [888, 173]. Portfolio [635]. Pose [2287, 2286]. Posedness [2250]. Posets [2197, 1605]. Position [1657, 692, 770, 1595, 806]. Position-Independent [1595]. Positive [1347, 2188]. Possibilistic [1088, 1091, 35, 1094]. Possibilities [978]. Possibly [1668, 87]. Posteriori [745]. Postmodernism [1095]. Potential [914, 651]. Power [674, 1010, 1295, 1344, 1578, 1331, 501, 728, 1012, 2215, 2217, 530, 193, 1459, 15]. Power-Free [2217]. Powerful [986, 1154, 1988]. Powers [153]. Practical [41, 2178, 1147, 441, 970, 1545, 40, 1104, 1029, 88, 1146]. Practice [1179, 2090, 364, 1276, 88, 1980]. Pragmatic [2129, 1237]. Pre [1008, 1055]. Pre-compressed [1055]. Pre-processing [1008]. precise [301]. Precision [978]. Preconditioner [1788]. Preconditioners [511, 637, 136]. Precorrected [747]. Predecessor [181]. Predefined [2298]. Predicate [1019, 935, 1378, 2343]. Predicated [935]. Predicates [1407, 705, 2145]. Predicting [727, 968]. Prediction [117, 1508, 1005, 2021]. Predictions [2091, 2351]. Predictive [1003]. Predictor [2268]. Preemption [2101]. Preference [1096]. preferences [13]. Prefix [869]. preliminary [330]. preorders [1239]. Preprocessing [709, 972]. Preprocessor [678]. Presburger [1260, 1081]. Presence [1449, 1914, 1692]. Present [467]. Presentation [2077, 1679]. Presented [1845, 795, 796]. Preserve [365]. Preserving [1183, 797]. Presheaf [1373, 2078]. Pressing [1382]. Prevalence [2163]. Priced [2041]. Primal [1337]. Primal-Dual [1337]. Prime [2162]. Primes [2218]. Primitive [403, 174, 1023]. primitives [1964]. Principal [515, 750, 2240, 1218]. Principle [1433, 2070]. Principles [887, 975, 2085]. Printed [1610]. Prior [752, 962]. priori [316]. Priorities [546]. Prioritised [1104]. Prioritized [23]. Priority [1134, 518, 1984]. Privacy [841, 928, 2166]. Private [1313]. Proactive [2176]. Probabilistic [1227, 1316, 1091, 1317, 1508, 213, 787, 788, 1082, 490, 2347, 1520, 1213, 1243]. Probabilistic-Timed [788]. Probabilities [461, 1998]. Probability [188, 1297, 1896]. Probably [1033]. Probe [1016]. Problem [104, 721, 994, 1501, 1141, 475, 1687, 540, 1615, 635, 664, 1447, 119, 1322, 899, 729, 156, 561, 126, 1335, 1232, 249, 71, 130, 1622, 1576, 2250, 1842, 1710, 1588, 1575, 1580, 2123, 2274, 1383, 1972, 1215, 1987, 791, 800].

Problem-Solving [721, 899]. Problems [1771, 268, 105, 829, 1854, 211, 1603, 977, 246, 206, 114, 465, 1591, 116, 77, 232, 638, 121, 259, 978, 863, 122, 452, 1154, 1337, 678, 1198, 2153, 2326, 1327, 1296, 1832, 599, 976, 967, 1833, 1774, 1743, 250, 1834, 1581, 2222, 1845, 1713, 836, 2017, 133, 2143, 134, 640, 136, 733, 1978, 1980, 20, 1476, 1991, 795, 2000, 2002]. Procedural [948, 1782, 1471]. Procedure [520, 1615, 1357, 1785, 1356, 1240]. Procedures [1141, 1147, 1928, 371, 409, 1155, 1673, 1919, 741, 303]. Process [534, 1502, 175, 220, 170, 877, 1325, 727, 788, 694, 1314, 1024, 1430, 1754, 1958, 1964]. Process-Oriented [1430]. Processes [1414, 1086, 1538, 1132, 1316, 481, 974, 783, 780, 1318, 1800, 1315, 682, 777, 80, 535, 223, 1123, 787, 411, 1323, 2196, 1918, 1818, 72, 1243, 1930]. Processing [1391, 669, 723, 2168, 1559, 2301, 1792, 963, 619, 995, 1412, 1013, 997, 1365, 359, 1557, 1565, 1742, 1770, 1362, 1783, 2313, 2315, 1892, 1786, 2378, 1767, 1144, 932, 1411, 652, 1220, 1008]. Processor [505, 25, 2096, 1665, 2091, 1531, 1752, 1769]. Processors [1533, 555, 506, 1837, 1745]. Product [144, 247, 697, 881, 876, 1070]. Production [1016, 886]. Products [1732]. Profile [342, 1218]. Profiler [2092]. Profiles [240]. Profiling [1539]. PROforma [409]. Progification [1676]. Prognosis [994]. Progol [2332, 2333, 2339, 2346]. Program [937, 84, 369, 270, 1202, 445, 488, 498, 652]. Programmable [1952]. Programming [2184, 1663, 2329, 483, 1922, 1398, 1019, 2107, 642, 1731, 465, 466, 455, 2007, 453, 1667, 1189, 655, 2122, 646, 492, 1618, 499, 523, 516, 1230, 654, 280, 1129, 2126, 2338, 270, 2326, 2108, 2124, 451, 1281, 950, 1026, 1649, 2111, 2343, 2009, 1640, 1764, 1034, 2327, 2013, 2349, 2016, 1752, 2123, 732, 1472, 1473, 1470, 1465, 1082, 1453, 2384, 815, 2383]. Programming-with-Proofs [1922].

Programs [491, 1300, 587, 583, 519, 1195, 524, 521, 517, 536, 370, 366, 857, 1357, 1256, 1666, 2356, 73, 363, 77, 649, 81, 1540, 82, 1781, 940, 663, 1193, 864, 367, 2340, 784, 2357, 361, 1633, 651, 2120, 1539, 948, 2214, 263, 1744, 2344, 1786, 2230, 2067, 1751, 1766, 1192, 379, 1951, 8, 1457, 1469, 1974, 1844]. Progress [225, 1796, 1247]. ProHos [505]. ProHos-1 [505]. Project [1063, 1346, 1190, 1068, 1397, 360, 1249, 1464, 912, 1536, 271]. Projection [1371, 768, 999, 2020]. Projections [1356, 316]. Projective [2242, 2244, 2243, 1730]. Projectivity [996]. Prolegomena [516]. Prolog [366, 363]. Promises [1656, 1386]. Proof [1166, 469, 2046, 2366, 1357, 867, 1669, 371, 1162, 1308, 213, 787, 405, 1814, 2085, 1356, 2191, 2066, 1848, 1679]. Proof-Nets [867]. Proof-on-Paper [1166]. Proofs [1922, 1168, 1157, 2056, 837, 479, 2213, 494, 2194, 2067, 1163, 1144, 383]. Propagation [1288, 1136, 626, 1774, 2257]. Proper [1740]. Properties [2361, 543, 148, 392, 1301, 73, 1304, 701, 1337, 2110, 820, 7, 1081]. Property [1372, 1152, 1709, 797]. Proposal [911, 276]. Proposals [828]. Propositional [1098, 367, 865, 2065, 1848, 1156]. Prospecting [632]. Prospectives [1542]. PROT [1795]. Protection [1055, 1554]. Protective [728]. ProtectOS [1544]. Protein [1805]. Proteins [1804]. Protocol [829, 220, 539, 38, 1244, 1957, 1239]. Protocols [1934, 203, 822, 839, 1670, 409, 2146, 1058, 1071, 1551]. proton [327]. Prototype [907]. Prototyping [74, 815]. Provable [43]. provably [1248]. Prove [2146]. Prover [1662, 1142, 1154, 1149, 1156]. Provers [1668]. Provide [1065]. Providers [913]. Providing [502, 2377]. Proving [1664, 480, 802, 1613, 73, 1143, 1161, 899, 1081, 1150, 1746, 1464, 801, 794]. Provision [928, 1042]. Provisional [409]. Provisioning [1068]. Proximate [1589].

Przymusinski [1460]. PScheD [2095]. Pseudo [2164, 630, 2019]. Pseudo-Bacterial [2019]. Pseudo-Random [2164]. Pseudo-spectral [630]. PSPACE [1375]. PSPACE-complete [1375]. Psychological [160]. PTTP [1149]. PTTP-based [1149]. Public [2159, 2161, 44, 2153, 833, 821, 843]. Public-Key [2159, 44, 2153, 821, 843]. Publishing [2133]. Pulse [981, 2275]. Pulse-Coding [2275]. Pump [993]. punctures [300]. Pure [855, 366, 401, 396]. Purpose [1391, 1661]. Pursuit [1572]. Pursuit-Evasion [1572]. Pushdown [198, 1076]. PVM [510]. PVS [1664, 1670, 1241]. Pyramid [2277].

QCD [105]. QRS [1009]. Qu [2158]. Qu-Vanstone [2158]. quadrangulations [1994]. Quadratic [738]. Qualitative [1500, 609, 433, 946, 986, 72, 1907]. Quality [215]. Quantal [958]. Quantification [1098, 2065]. Quantificational [1719]. Quantified [2370, 1111, 1848]. Quantifier [1113]. Quantifiers [2107, 1470]. Quantiles [1897]. Quantitative [986, 474, 1264]. Quantity [215]. quantizer [382]. Quantum [2168]. Quasi [683, 856, 2248, 1991, 1229, 639, 1256]. Quasi- [856]. Quasi-binary [2248]. Quasi-fully [1991]. Quasi-minimal [1229, 639]. Quasi-particle [683]. quasi-stable [1256]. Quasiperiodicity [1289]. Queries [14, 856, 181, 1657, 152, 1375, 2377, 2348, 2232, 13, 3, 20]. Query [1559, 1591, 1493, 860, 1412, 2377, 1557, 1560, 1565, 1831, 1892, 2, 7, 4, 1471]. Querying [1, 1427, 2381]. Question [87, 1485]. Question-Based [1485]. Questions [449, 595]. Queue [2091, 1127]. queueing [1984]. Queues [1011, 1074, 1545]. queuing [1127]. Quick [685, 2010]. Quotients [1681].

R [368, 1137, 1417]. R-trees [1417]. R0S0N1 [251]. R0S0N1-tt [251]. R8000 [1752]. Rabin [1703, 837]. Rabin-Type [1703]. Race [1188]. Radial [691]. Radiation [631, 1802, 445]. radicals [1971]. radiographies [289]. Radon [2254]. Rail [1482]. Rail-Inspection [1482]. Railroad [475]. railway [1946]. RALL [1163]. Random [258, 594, 744, 2164, 2177, 593, 209, 2211, 756, 2026, 844, 2267, 1907, 286]. Random-Self-Reducibility [209]. Randomization [1300]. Randomized [377, 1588, 1982, 11, 1998]. Randomness [1298, 1334]. Randomness-Efficient [1334]. Randoms [747]. Randoms-Precorrected [747]. Range [322, 2278, 767, 1690]. Rankings [1706, 159]. Rapid [840]. Rat [431]. Rate [972, 127]. Rated [1886]. Rates [1525, 1126]. Ratio [57, 336]. Ratio-based [336]. Rational [1290, 92, 1823, 117, 1331, 1113, 2041, 1287, 798]. rationale [1128]. Rationality [893]. Ray [764, 291, 290]. Razor [614]. RBF [338]. RBFNN [969]. RC4 [834]. RDBMS [1562]. Re [1632, 547, 556]. Re-use [1632]. Reach [1623]. Reachability [1076, 1324, 476, 1622, 1246]. Reaction [109, 116]. Reaction-Diffusion [116]. Reactive [99, 238, 536, 96, 537, 443, 1015, 550, 1674, 1073, 1947, 1267, 1069]. Reactor [722]. Read [1300, 2214, 2233]. Read-Once [1300, 2214]. Real [1185, 772, 894, 164, 646, 304, 708, 613, 2239, 1294, 776, 778, 418, 100, 785, 451, 2323, 2218, 2289, 1654, 279, 1701, 1931, 1700, 1120, 1264, 309, 1989]. Real-Life [164]. Real-Scale [1185]. Real-Time [772, 894, 2239, 776, 778, 100, 785, 2323, 2218, 2289, 1654, 1931, 1120, 1264, 309, 1989]. Real-Valued [451]. Realisation [909]. Reality [2259, 1810, 318, 311, 298, 332]. Realization [2356, 1643]. Realizing [2177, 2346]. Really [276]. Reals [1825, 2222]. Realtime [1203, 1979]. Rearrangeable [1729]. reason [2385].

Reasoner [1484]. Reasoning [1500, 1071, 1096, 1501, 520, 1089, 1502, 1147, 37, 529, 1903, 1503, 434, 2048, 897, 1478, 1486, 1504, 1493, 1100, 516, 1507, 371, 1101, 1102, 474, 1103, 1479, 1480, 33, 2040, 2036, 1106, 1482, 419, 1505, 1111, 1860, 2346, 518, 1139, 1902, 1509, 1114, 1456, 1474, 1458, 1477]. rebalancing [1993]. Reception [2173]. Receptive [982, 2236]. Recessive [1396]. Recipe [443]. Recognition [2303, 2328, 345, 2309, 1792, 343, 861, 1008, 341, 705, 358, 997, 347, 335, 1003, 348, 344, 359, 599, 996, 2279, 1006, 1801, 2310, 2278, 2238, 1595, 2286, 2299, 2240, 2290, 1004, 1805]. Recognizability [1341, 158]. Recognizable [1331, 1256]. Recognize [1293]. Recognizing [1434]. Recombination [1812]. Recommendation [163]. Recomputation [2120]. Reconfigurable [571]. Reconstructed [440]. Reconstructing [2004]. Reconstruction [1695, 1217, 768, 2320, 2014, 2263, 327, 316, 400, 1985, 2001]. Record [1925]. Recordings [995]. Recovery [1556, 2162, 2319]. Rectangle [54, 2296]. rectangles [1999]. Rectangular [1326, 1623, 1838]. Rectification [2300]. rectilinear [1980]. Recurrent [482, 2275, 512]. Recursion [1342, 858, 406, 2085, 403]. Recursive [1926, 681, 2340, 2116, 961, 868, 1744, 1214, 2297, 398]. Recursive-Parallel [1744]. Red [261]. Red-Black [261]. Redex [2121]. Reduce [718, 697]. Reduced [831, 819, 83, 359, 556, 1990]. Reducibility [209]. Reduction [1372, 2153, 704, 951, 1330, 2234, 1258, 407, 813]. Reductions [1340, 1374, 526, 367, 2362, 870]. Redundancy [852, 2160]. Reed [1284]. Reed-Müller [1284]. Reference [528, 2386]. refinable [1503]. Refine [1412]. Refinement [2330, 783, 780, 1820, 773, 447, 784, 785, 494, 1915, 1966, 2229, 1919, 227, 1957].

Refinement-Algebraic [1966]. Refining [1489, 775, 446, 1339, 361, 1674]. Reflection [2053, 1928, 1643, 2070]. Reflections [2118]. Reflective [1563, 1024, 2232]. Reflexive [219]. Reflexivity [2203]. Region [1054, 933, 601]. Region-Level [601]. Regional [754, 750]. Regions [1415, 1799]. Register [939, 150, 1232]. Registered [754]. Registers [547]. Registration [605, 2310, 316, 312, 282, 294, 296, 292, 324, 289, 305, 290, 293]. Regular [1293, 777, 2211, 183, 1933, 176, 1245]. Regularities [2352]. regularity [797]. Regularization [964, 2269]. Regulated [1940]. Regulation [1798, 2207]. Regulatory [1799, 1801]. Rehearsal [762]. Reinforcement [1014, 1497]. Reinventing [1527]. Related [1579, 2159, 246, 1198, 1296, 836, 1991]. Related-Message [2159]. Relating [225, 2210, 384]. Relation [1358, 1380, 1163]. Relational [1087, 1926, 699, 1864, 1865, 2352, 597, 1423]. Relations [2197, 472, 840, 2203, 1823, 2335, 1056, 2146, 1105, 477, 1292, 1915]. Relationships [368, 2033, 1582]. Relative [478, 1377]. relaxations [1986]. Relaxed [261, 1993, 1609]. relevance [812]. Reliability [1434, 190, 576, 622, 987, 1886, 733]. Reliable [1216, 472, 2287, 1617, 1692]. Reluctance [718]. Remote [276, 174, 1919]. Remotely [1529, 2266]. Removing [212, 254]. renal [301]. Rendering [628, 309]. Rényi [832]. Reorganization [898]. Repair [1496]. Repeated [1275]. Repeated-Root [1275]. Repeats [1803]. Repetitive [1724]. Repetitiveness [2216]. Replacement [365, 148, 2206]. Replication [1556]. Report [62, 418, 1464]. Repositories [1360]. Represent [1354]. Representation [1755, 1136, 2307, 689, 1486, 1115, 1349, 601, 24, 2279, 2140, 2139, 1712, 1488, 299].

Representations [1910, 603, 1054, 1867, 2061, 1462, 309, 2223]. Represented [207]. Representing [774, 986, 2082, 1694, 1465]. Republication [1632]. Requirement [959, 778]. Requirements [1060, 1516, 1385, 1173, 1481, 881, 1032, 1171]. REs [1404]. Resemblance [37]. Resend [2159]. Residential [51]. Residual [1229, 639, 349]. Resilient [847]. Resistant [2178]. Resistence [959]. Resolution [1461, 2107, 1691, 1357, 343, 1160, 2324, 1159, 2213, 2044, 2194, 2050, 2277]. Resonance [431, 757]. Resonant [631]. Resonant-Mass [631]. Resource [583, 1060, 186, 1299, 906, 917, 1151, 944, 1045]. Resource-Aware [583]. Resource-Based [944]. Resource-Bounded [186, 1299]. Resource-Distribution [1151]. Respect [2258, 2230, 135]. Respiratory [991]. Response [86]. Responses [971]. Responsiveness [2100]. Restart [2062]. Restarting [545]. Restoration [2248, 128]. Restricted [1711, 324, 1693, 2064, 779]. Restrictive [1619]. Result [238, 384]. Resultant [1273]. Results [934, 2186, 2199, 1091, 1299, 895, 1813, 462, 196, 1335, 1539, 967, 2318, 1917, 1841, 330, 305, 796, 2382]. Retino [999]. Retino-cortical [999]. retinography [327]. Retransmission [539]. Retrieval [1313, 1490, 1491, 1494, 2141, 1361, 2280, 1720, 2378, 1488]. Retrospective [711, 293]. Return [1052]. reusable [1549]. Reuse [1177, 930, 1145, 648]. Reversible [205]. review [1205]. REVISE [1473]. revising [1468]. Revision [521, 424, 89, 895, 1093, 413, 1104]. Revisited [186, 740, 1611, 1840, 2158, 2]. Rewards [1311]. Rewrite [798, 2187, 2208, 478, 471, 801]. Rewriting [2361, 542, 142, 543, 2104, 2077, 1140, 75, 1377, 545, 1673, 1718, 2362, 2076, 2075, 2363, 2228, 1851, 807, 809, 818, 797, 804, 810, 803].

Rewritings [2117]. RICAD [1493]. Rice [2054]. rich [1261]. Ridge [1395]. ridges [1211]. Right [101, 941, 2371, 809]. rights [1949]. Rigid [294, 1296, 2069, 282, 765, 2240]. Rigidity [2265]. Rigidity-Constrained [2265]. Rigor [2049]. Rigorously [441]. Rings [847, 1728, 1285, 799, 375, 1999]. Ripple [718]. RISC [1756, 506, 652]. RISC-based [1756]. Risk [151]. Risks [1422]. RL [547]. RNA [1807]. Road [2316]. Robbins [791]. ROBDDs [527]. RoboCup [279]. Robodoc [320]. Robot [514, 34, 303, 1390, 1388, 279, 1389, 330]. Robot-assisted [303]. robotic [329]. Robotics [1014, 1015]. Robots [703, 2019]. Robust [333, 1220, 228, 251, 1658, 388, 2272, 1677, 2266]. Robustly [1741]. Robustness [1898, 1223]. Role [1555]. Roles [165]. ROM [1527]. Root [115, 1275, 1851]. Root-Finding [115]. Rorschach [86]. Rosser [804]. Rotation [2260]. Rotation-Equivariant [2260]. Rotations [2281, 1609, 2245]. Rough [1437]. Round [838]. Round-Optimal [838]. Roundoff [135]. Route [1830]. Routers [2041]. Routing [1686, 1321, 146, 200, 1320, 2209, 155, 1610, 1319, 1685, 2231, 1716, 1988]. Rover [2318]. Rovers [236]. RPC [174]. RSA [2175, 44, 837, 2176, 2154, 852, 38, 45, 2160, 2171]. RSA-Based [2154]. RSA-Like [38]. RSA-Type [2171]. Rule [1501, 521, 1731, 2308, 430, 696, 700, 2058, 1888, 10, 8, 1984, 815]. Rule-Based [2308, 430, 815]. RuleBase [1268]. Rules [2328, 238, 971, 2335, 547, 413, 478, 866, 700, 2079, 2019, 9]. Run [1050, 1741, 1027, 2131]. Run-Time [1050, 1741, 1027, 2131]. Runtime [572, 1340].

S [1418]. S-tree [1418]. S0Z [2201]. S52C [866]. SAC [848]. SAC/PC [848]. Safe [2354, 858, 1203, 1033, 585, 1624]. Safer [1953]. Safety [221, 231, 1248, 1081].

Salesman [465]. Sample [846, 372]. Sampling [258, 2011, 286]. SAR [605]. Satisfaction [219]. satisfiability [2002]. satisfiable [2002]. SATO [1156]. Saturations [2340]. Saw [1494]. Scalability [1734]. Scalable [625, 1485]. Scalar [713, 1207]. Scale [1772, 1185, 1206, 770, 1046, 1223, 1216, 1226, 1217, 1212, 1209, 1222, 1218, 1224, 1539]. Scale-space [1206, 1223, 1217, 1212, 1222, 1224]. Scales [2245, 1208]. Scaling [191]. Scaling-Invariant [191]. Scan [503]. Scan-Directional [503]. Scanner [2315]. Scans [754, 747]. Scenario [911, 905, 1626, 617]. Scenarios [1173]. Scene [2308]. Scenes [2239, 1979]. Schedufing [1592]. Schedule [1741]. Scheduled [1758, 2099]. Schedulers [2092, 1196]. Schedules [253]. Scheduling [1967, 937, 2096, 938, 1781, 2100, 2090, 1984, 729, 2098, 2095, 2097, 1588, 2017, 2102, 1767, 652, 1989, 1998]. Scheduling-LPs [1998]. Schema [1568, 975]. Schema-Based [975]. Schemas [2142]. Schemata [690, 369, 2349]. Schematizations [482]. Scheme [1686, 1059, 111, 1703, 858, 689, 827, 39, 1758, 1233, 644, 2126, 1763, 1720, 2261, 1554]. Schemes [851, 2174, 200, 1333, 1702, 129, 2162, 1680, 133, 871, 1214]. Schmidt [1788]. Schur [123]. Science [467, 2387]. Scientific [180]. SCM [883, 877, 881, 875]. ScmEngine [878]. Scope [589]. Score [347]. Scott [857, 806]. screw [319]. screws [323]. Scripting [1191]. SDH [920]. Sea [688]. SEAL [1510]. Seamless [281]. Search [1513, 1227, 2366, 570, 149, 1541, 182, 2026, 1600, 1644, 1737, 597, 574, 1968, 1993, 791]. Searching [1236, 29, 1181, 451, 1595, 1713, 1849, 1690]. Second [1413, 865, 1378, 531, 1784, 1806, 134, 870]. Second-Class [1806]. Second-Level [1784]. Second-Order [531, 870]. Secrecy [1934].

Secret [2181, 1549, 1702, 833]. Secret-Key [833]. sections [795]. Secure [203, 46, 43, 825, 827, 585, 2179, 833, 2166]. Securing [51]. Security [1934, 1939, 2161, 2165, 928, 837, 926, 2155, 1529, 488, 245, 927, 1542]. Seed [1494]. Segmentation [594, 2247, 31, 999, 2276, 758, 755, 2316, 743, 344, 605, 757, 596, 2314, 731, 742, 284, 2304, 2236, 2288, 2284, 2016, 1878, 592, 2275, 761, 2274, 741, 759, 1000, 1219]. Segments [1590]. Seismic [1774]. Select [448]. Selected [1707, 452]. Selection [2303, 461, 1193, 5, 1870, 2010, 2062, 1602, 1871, 1210, 2387]. Selective [2105, 852, 2120]. Selector [1029]. Self [957, 1014, 465, 2007, 758, 1008, 259, 209, 695, 1278, 276, 694, 2264, 2256, 1747, 2070, 1217]. Self-Active [276]. Self-Adaptive [465, 2007]. Self-Calibration [2256]. Self-Connection [1747]. Self-Dual [1278]. Self-Duality [259]. Self-Evaluation [2264]. Self-Organization [957]. Self-Organizing [1014, 758, 1008, 695]. Self-Organizing-Process [694]. self-similar [1217]. Self-Verifying [2070]. Semantic [483, 2356, 2379, 1174, 1921]. Semantic-Based [1174]. Semantics [2184, 76, 854, 475, 517, 1344, 2366, 1358, 83, 81, 523, 486, 2116, 3, 788, 1382, 1644, 1839, 2063, 1923, 2068, 2233, 98, 226, 1194, 1951, 1460, 1461, 1948, 1466, 1085]. Semantics-Based [1194]. Semi [2381, 863, 182, 1909, 1277, 630, 1701, 1171, 1]. Semi-Algebraic [1277]. Semi-continuous [1701]. Semi-Dynamic [182]. Semi-explicit [1909]. Semi-formal [1171]. Semi-groups [863]. Semi-implicit [630]. Semi-structured [2381, 1]. semidefinite [1986]. Semigroups [500, 1834]. Sensed [2266]. Sensing [1846]. Sensitive [2362, 1875, 1876, 806, 803]. Sensitivity [1731]. Sensor [1903, 352, 987, 1109, 2237]. Sensorial [960]. Sensors [987]. sentences [11]. Sentential [2227]. Sentinel [2042]. separable [1214]. Separate [272]. Separating [1458, 2074]. Separation [600, 970, 256]. separators [1977]. Sequence [534, 1450, 1496, 990, 1452, 1013, 1805]. Sequence-Structure [1805]. Sequences [1290, 1451, 749, 1711, 1668, 1343, 606, 1904, 38, 1801]. Sequencing [1597]. Sequent [2050]. Sequential [1315, 207, 663, 2116, 651, 2190]. Serial [1756, 754]. Serializability [17]. Series [53, 1427, 1437, 1400, 1424, 1331, 2215, 2021, 808]. Series-Parallel [53, 808]. Seriously [438]. Server [1057, 1196, 1653, 1124]. servers [1984]. Service [55, 1065, 913, 916, 1068, 906, 920, 907, 915, 1127]. Services [1041, 1052, 1050, 1637, 917, 1053, 909, 1042, 281, 1463]. SESAM [352]. SESP [360]. Session [1546]. Set [1573, 140, 482, 36, 500, 346, 256, 1188, 1693, 2058, 1623, 1717, 1856, 1628, 2067, 1907, 1139, 1968, 1215]. Set-Term [1717]. SETHEO [1145, 1144]. Sets [1739, 53, 202, 1437, 1324, 181, 1338, 1691, 1287, 416, 1604, 1150, 266, 2225, 1750, 875, 1297, 1929, 1974, 1987, 1999]. Setting [1926]. Seven [2000]. Several [1563]. SFP [485]. SGML [2134, 2135]. SHA [842]. Shading [2250]. Shafer [1099]. Shallit [1332]. Shape [2302, 755, 335, 743, 2074, 2279, 2250, 598, 395, 1222]. Shape-from-Shading [2250]. Shapes [1726, 2251]. Share [1420]. Shared [1234, 931, 2175, 2096, 945, 655, 1230, 654, 628, 1784, 1764, 1990]. Shared-Memory [1234, 931, 2096, 655, 654]. Sharing [1018, 1305, 238, 1702, 1639, 1632, 1549, 805, 406]. Shedding [2081]. Shell [1660]. Shift [95, 2201, 1004, 1449, 314, 1618]. Shift-Invariant [2201, 1004]. Shifts [754, 2281]. Short [2307, 1343, 253, 2377, 2213]. Short-Length [2307]. Shortest [1740, 152, 1590, 182, 1830]. Shostak [793]. Should [2022]. Shoulder [427]. SHRIMP [169]. Shuffle [1835]. Side [2120, 951]. Side-Effects [2120]. sigma [2368]. Sign [2157]. Signal [723, 359]. Signals [744, 972]. Signature [1427, 851, 2174, 49, 2156]. Signatures [850, 2154, 852, 849, 2155, 47, 2160, 2241]. Signcryption [2156]. signs [2003]. Silicon [1395]. SIMD [1533]. Similar [1424, 1217]. Similarity [696, 1498, 1219, 296, 1222]. Simple [955, 946, 1804, 1785, 170, 240, 1872, 1598, 1632, 1146, 326, 381, 399, 1600]. Simplices [1732]. Simplification [480, 2045, 1577, 2228]. Simplify [985]. Simulate [629, 514]. simulated [304]. Simulating [1352, 309, 1790, 870]. Simulation [683, 109, 1789, 1614, 680, 682, 631, 895, 576, 1773, 2208, 965, 670, 982, 626, 764, 1884, 634, 967, 2318, 1779, 302, 1122, 1084]. Simulations [1457, 2022, 644, 623]. Simulator [1191, 307, 322]. Simultaneous [1452, 1296, 2069, 1987]. Sine [118]. Single [674, 478, 2124, 172]. Single-Chip [172]. Singularities [2262]. Situation [1086]. Size [977, 368, 696, 830, 1745, 1784, 1694]. sized [1240]. Sizes [463]. skeleton [1936]. skeptical [1461]. Sketch [2066]. Sketch-as-Proof [2066]. Skill [609]. Skin [766]. skull [318]. Sleep [709]. Sliding [736, 1530, 1239]. Sliding-Mode [736]. Slope [216]. Slope-Parametric [216]. Slovene [2337]. Slowed [1225]. Sludge [734]. Small [188, 864, 183, 830, 2192, 1625, 1601, 763, 379, 1544]. SmallEiffel [273]. Smallest [151]. Smalltalk [1029]. Smart [1203, 566]. Smooth [832]. Smoothing [457, 2298]. SNARKS [1165]. SNMP [922]. Sobolev [122]. Social [974, 391, 93, 163, 393]. Societies [898, 160]. Society [428, 1438, 241]. Sockets [169]. Soft [166, 739, 444, 1159, 2294, 165]. Soft-Link [444]. Software [1177, 1064, 1502, 524, 1532, 467, 985, 878, 873, 389, 1758, 1944, 1479, 1145, 1235, 648, 626, 1525, 1555, 1531, 1752, 233, 1194, 1939, 1248, 1267, 1945]. Soils [2015]. Solution [1780, 638, 124, 130, 1743, 2318, 134, 1509, 2039, 811]. Solutions [465, 916, 1476]. Solvable [1845]. Solve [393]. Solver [2115, 637]. Solvers [106, 117, 506]. Solving [1771, 721, 1501, 1772, 1309, 120, 458, 899, 1980, 699, 125, 126, 2123, 811, 2017, 733, 1464, 1257, 1119]. Some [107, 1091, 1295, 1247, 1607, 206, 553, 259, 1148, 462, 2326, 1250, 967, 1917, 1581, 384, 2057, 1394]. Sort [1188]. Sorted [2364, 1718, 818]. Sorting [1819, 662]. SOS [1951]. Sound [787, 1002, 1356, 1111, 1914, 1263]. SOUR [792]. Source [234]. Space [551, 1234, 2103, 749, 1359, 1274, 1233, 966, 2192, 1636, 2025, 667, 1216, 1258, 1217, 1206, 1212, 1222, 1223, 1224, 1964, 811]. Space-Time [749]. Spaces [706, 485, 188, 1271, 122, 869, 1330, 846, 2084, 386, 1118, 1211, 1953]. Spanning [248, 1728, 1722, 1975, 1992]. Sparse [1780, 946, 1691, 506, 638, 154]. Sparsed [1772]. SPARTH [1791]. Spatial [1405, 1414, 1755, 1408, 740, 1406, 1431, 601, 33, 1410, 1646, 2044, 2266, 1721, 1411, 1208]. Spatial-Temporal [740]. Spatiotemporal [1409]. Speaker [345, 346, 2291, 347, 348, 359, 349, 360, 354, 350]. Special [1391, 116, 134]. Special-Purpose [1391]. Specialising [370]. Specialization [1914]. Specializing [1912]. Specific [762, 1799, 2339, 1189, 11]. Specification [90, 2376, 469, 363, 444, 190, 1364, 497, 785, 2205, 1677, 237, 779]. Specification-Based [363, 497]. Specifications [1663, 918, 536, 1613, 775, 778, 361, 437, 1814, 494, 1678, 2359, 794]. Specifying [2200, 443, 2080]. SPECT [1210, 316, 750]. Spectra [339]. Spectral [1002, 126, 129, 2266, 1213, 630]. Spectral-Spatial [2266]. Spectroscopy [431]. Spectrum [418, 349]. Speculation [1673]. Speculative [935]. Speculators [455]. Speech [2328, 2333, 353, 1005, 997, 890, 886, 998]. Speech-to-speech [890]. Speed [177, 1497, 1102, 2191, 1531]. Speed-Up [1497, 2191]. Speed-Ups [1102]. Speeds [2231]. Spert [980]. Spert-II [980]. Spi [1071]. SPIF [1190]. spine [289]. Spliced [1725]. Splicing [557, 556]. Spline [132, 299]. Split [1281, 1284]. Splitting [1417, 116, 1327]. SPNL [1123]. Spoken [889, 888, 884, 92, 1008]. Spoken-Digit [1008]. Sporadic [1196]. Sport [415]. Spring [1705]. SQUARE [1523]. Squares [1444, 128, 640, 136, 1550]. Stability [861, 123, 1743, 1747, 135]. Stabilizing [513]. Stable [1344, 2210, 1461, 1256, 1462, 134]. stack [1997]. stack-up [1997]. Stacking [1608, 1660]. stacks [1074, 1545]. Stage [1017]. Stalmarck [1238]. Stand [2198]. Standard [971, 83, 670, 45, 1609, 2127, 2269, 1128, 2386, 1516]. Standardisation [1768]. Standardizations [2071]. Standards [1768]. Star [146, 1595, 1310]. Star-Free [1310]. Start [1748]. State [1234, 1018, 1400, 911, 119, 1162, 2101, 1904, 1933, 2223, 2013, 1258, 1260, 1118, 1081, 1241, 12]. State-Space [1234]. statecharts [1079, 1250]. States [131, 2218]. Static [1019, 77, 921, 458, 1970, 1460]. Statistical [994, 964, 735, 749, 2146, 834, 599, 1855, 2288, 747, 760]. Statistics [690, 351, 434, 2022, 1220]. Steady [109]. Steel [621]. Steiner [1980, 1722, 1723]. Steklov [133]. Steklov-Type [133]. Step [1413, 780, 2307, 2180, 134, 463, 809, 810, 772]. Steps [219, 914, 2133]. Stepwise [1966]. Stereo [2273, 2271, 2315, 1215]. stereotactic [314]. stereotaxy [317]. Stereovision [2270]. Stern [44]. Stiff [117]. Still [2198, 2378]. stimulation [315]. Stimuli [1172]. Stochastic [1234, 984, 457, 119, 1133, 422, 2192, 1779, 1818, 2277, 1123]. Stokes [511, 637]. Stone [485]. Stop [851]. Storage [672, 1118, 1720, 1995]. Story [1170]. Str [1263]. Straight [2249, 57, 61]. Straight-Line [57, 61]. Strategic [893]. Strategically [569]. Strategies [713, 723, 1412, 597, 459, 456, 595, 813]. Strategy [1600, 1888, 2029, 2117]. Stratified [1370, 1821, 1503]. Stream [1519, 1533, 981, 169, 1517, 835, 49]. Streams [2157, 2116]. Strengthened [122]. Strengthening [424]. Striatal [956]. Strictly [1310]. Strictness [1913, 402]. String [1584, 797]. string-rewriting [797]. Strings [543]. Strong [1696, 1454, 1908, 394, 408]. Stronger [837]. Strongly [1689, 207, 1246]. Structural [988, 10, 2255, 1199, 611, 1131, 1807, 1782, 2254, 1810, 2079]. Structural-Procedural [1782]. Structure [559, 739, 450, 2294, 1859, 1733, 1742, 349, 1805, 18, 1224, 811]. Structured [603, 2141, 1104, 1814, 2281, 1678, 1, 1118, 2381]. Structures [229, 1354, 246, 1450, 1593, 1596, 1412, 622, 698, 67, 2377, 868, 1832, 951, 2387, 1808, 760, 386, 1979, 1218]. Structuring [2056, 1037, 1079]. STS [1802]. Stub [275]. Studies [1641, 2079]. Study [104, 1350, 882, 1562, 1502, 600, 417, 1491, 1757, 1668, 58, 1005, 1759, 917, 1872, 1683, 2018, 1460, 300, 1944, 289, 1946, 1945]. Studying [863, 137]. Stupid [1095]. Style [1664, 446, 1943]. Styles [496]. Stylesheets [2136]. Sub [756]. Sub-manifolds [756]. subarrays [381]. Subband [345]. Subclasses [1700]. Subdivisions [1586]. Subgraphs [248, 1575]. Subgroup [2162]. Subgroups [1423]. Subiteration [2283]. Subject [1372, 231]. sublinear [1996]. Submission [565]. Subregion [1720]. Subset [869, 1106, 1403, 286]. Subspace [2299]. Substitutions [2367, 2358]. Substract [731]. Substrate [949, 180]. Subsumption [1340, 611]. Subsystems [1030]. Subtraction [2247]. Subtree [454]. Subtyping [1910, 76, 484, 1021, 2202, 1925, 2051, 398].

Subword [1827]. success [1238]. Successes [1641]. Suffices [1298]. Sufficient [2339]. Suffix [260, 1983]. Sugiyama [65]. Suitable [1228]. Sum [1336, 2307, 1998]. Sumatra [583]. Superhull [1689]. Superimposed [705]. Superposition [2045]. Superscalar [1533]. Supervised [1210]. Supervisory [1624]. Supplement [1795]. Support [1500, 918, 175, 724, 529, 572, 646, 1493, 278, 659, 1036, 648, 418, 178, 1770, 1058, 1880, 236, 1654, 1862, 1430, 1194, 1466, 1544]. Supported [132, 1163]. Supporting [1177, 1504, 173, 920, 460, 1766]. Supports [1658]. Surface [2241, 759, 324, 313, 293]. surface-based [293]. Surfaces [1497, 756, 2252, 287, 286]. Surgery [762, 318, 326, 330, 332, 309, 319, 324, 329, 305]. surgical [292, 328, 321]. Surround [2236]. Surveillance [1422]. Survey [213, 1393]. Suspend [790]. Suspend/Resume [790]. Suspensions [624]. SWISS [1795, 1061]. SWISS-PROT [1795]. Switch [172, 1126]. Switchbox [155]. Switched [718]. Switching [1612, 1059, 911, 665, 1981]. Syllogism [700]. symbol [865]. Symbolic [1316, 971, 1324, 1260, 500, 2356, 659, 362, 1410, 476, 1261, 1873, 1931, 1146, 1258, 1247, 1257]. Symbols [543, 812]. Symmetric [1772, 1665, 507, 1320, 1776]. Symmetry [957, 989, 1274, 1705, 526, 1251]. Synaptic [992, 451]. Synchronised [1044]. Synchronization [782, 175, 1291, 656, 1837, 2204]. Synchronized [780, 1122]. Synchronous [110, 224, 1255]. Synergies [2012]. Syntactic [2365, 534, 1088, 252, 1199]. Syntactical [1090, 695]. Syntax [588, 403]. Synthesis [524, 536, 1077, 371, 231, 1619, 2039]. Synthesizing [230]. Synthetic [1376, 100, 2089]. System [2328, 41, 647, 355, 888, 1563, 1049, 668, 1903, 92, 1422, 692, 672, 242, 1063, 1137, 736, 1044, 664, 189, 410, 352, 151, 1056, 357, 632, 1169, 1103, 726, 501, 2324, 1013, 358, 928, 727, 40, 1634, 67, 890, 418, 2030, 100, 787, 972, 728, 686, 687, 2040, 2036, 1302, 1396, 70, 125, 24, 1482, 1782, 714, 1388, 2323, 2191, 2238, 1886, 1023, 1651, 1639, 1625, 2313, 1912, 1925, 1628, 1848, 1899, 279, 1862, 767, 556, 281, 712, 1472, 306, 301, 331, 317, 1474, 328, 315, 1947, 817, 1544, 1945, 1475]. Systematic [780, 1745, 2343, 1808, 1875, 1876, 1192, 1465]. Systems [28, 1920, 1780, 549, 106, 202, 2361, 889, 222, 1612, 854, 166, 1772, 546, 735, 97, 855, 1372, 2365, 567, 433, 90, 2197, 351, 772, 1916, 2200, 691, 244, 225, 1324, 1295, 219, 1614, 572, 661, 1007, 1616, 552, 442, 1175, 893, 1638, 575, 897, 553, 1135, 148, 1617, 401, 1684, 573, 537, 83, 1364, 232, 1317, 2187, 170, 511, 120, 645, 547, 2037, 877, 1027, 443, 2208, 676, 588, 578, 1401, 557, 699, 213, 785, 1557, 1693, 693, 2042, 218, 628, 2098, 687, 1814, 513, 231, 1619, 1760, 623, 636, 217, 490, 2035]. Systems [243, 554, 1043, 948, 902, 230, 1718, 550, 2031, 1555, 174, 1621, 1622, 1623, 2191, 2043, 1674, 2238, 2028, 1644, 1746, 1020, 2082, 1648, 1636, 1763, 1627, 1793, 1847, 1906, 2044, 1779, 1861, 1629, 1766, 1791, 1851, 237, 915, 548, 425, 245, 1635, 1767, 227, 2070, 571, 2081, 226, 1509, 471, 1389, 779, 1931, 900, 165, 1393, 2039, 733, 1072, 10, 1120, 1246, 1125, 1960, 1260, 1264, 1257, 798, 1944, 910, 1081, 1069, 1083, 329, 1954, 1265, 1946, 1237, 801, 2385, 1937, 1940, 818, 797, 804, 1963, 810, 26]. Systolic [1787]. SZK [2147, 2148]. SZK-Verifier [2148, 2147].

T [1203]. T-Smart [1203]. T3D [508]. T3E [2095]. Tableau [1165]. Tableaux [1356, 2062]. Tabular [2143]. Tactics [1162]. Tagged [752, 751]. Tagging [2333]. Tags [752]. Taking [438]. Tally [1727]. Tangibility [2070]. TANGRAM [1116]. TANGRAM-II [1116]. Tank [1906].

TAPSOFT [468]. Target [941, 2345, 1004, 317, 271]. Task [1350, 1203, 1760, 651, 1762, 1786, 1197]. Task-Dependent [1350]. Task-Safe [1203]. Tasks [2253, 619, 1394, 1967]. Taxonomy [885, 84, 1393]. Teaching [1135, 1189, 2124, 2123]. Team [1638, 279, 574]. Teams [901, 392]. Teamwork [1638]. Technical [1350, 2376]. Technique [2308, 649, 26, 127, 1743, 1602, 2023, 326, 321]. Techniques [994, 487, 1756, 110, 2186, 1559, 510, 981, 1493, 676, 1433, 1866, 2377, 1859, 2373, 1565, 2205, 1867, 1900, 1204, 72, 1255, 1124, 1079, 2382, 293]. Technologies [502, 588, 563, 1779]. Technology [1751, 1631, 1635, 927, 2372, 1553]. Telecommunication [1136, 910]. Telecommunications [923, 1441, 913, 625]. Telelearning [1066]. Telemedicine [1067]. telepresence [332]. Telescript [590]. Template [670]. Templates [1684, 1593, 606]. Temporal [1879, 740, 775, 610, 190, 1056, 474, 2205, 2138, 2064, 1561, 1075, 1208, 1070]. Teraflop [675]. Term [482, 2361, 2077, 1140, 1377, 1718, 1717, 2076, 2117, 2363, 2228, 1851, 380, 374, 818, 804]. Terminals [1064]. Terminating [2227]. Termination [480, 2365, 365, 2356, 479, 478, 1197, 803, 9, 801]. Terminology [2376]. Terms [608, 1311, 616]. Terrain [604, 2320]. Terrestrial [1052]. Test [538, 437, 1987, 1709, 201, 86]. Testbed [1362]. Testing [262, 81, 497, 1155, 788, 972, 2300, 1204, 1192]. Tests [1044, 1120, 4, 188]. Text [1367, 424, 1351, 347, 2377, 240, 1889, 1850]. Text-dependent [347]. Text-independent [347]. Text-Learners [1850]. Texts [1443, 2139, 1485]. Textual [234, 1492]. Texture [2253, 999, 596, 2304, 2312, 598]. Textures [2255, 2254]. th [2217]. The-Based [1450]. Their [475, 1558, 1172, 1596, 846, 1582, 135, 1823, 1977, 1223, 1832, 1681]. Them [122, 1148]. Theorem [1662, 1664, 1142, 2054, 1143, 1668, 1161, 899, 1150, 1149, 2075, 794]. Theorema [1464]. Theorems [1285]. Theoretic [1920, 36, 474, 209, 1303, 98, 8]. Theoretical [467, 2199, 416, 1308, 2341, 967, 2067, 595]. Theoretically [833]. Theories [1103, 1921, 871, 1468]. Theory [1506, 36, 2054, 92, 2046, 1273, 1570, 1707, 1667, 859, 1152, 1364, 35, 1162, 2090, 1376, 1419, 1378, 868, 1672, 2058, 1620, 1023, 2059, 2075, 1675, 2132, 2089, 1678, 1276, 88, 1907, 2269, 381, 1078, 1240, 1980, 285, 404, 16, 809, 810]. Theory-Based [88]. Therapies [409]. Therapy [445]. Thermal [683, 1010, 114]. Thermocapillary [124]. theta [611]. theta-subsumption [611]. theta0p2 [1813]. Thin [1226]. Thing [101]. Thinning [2301, 2283]. thoughts [1250]. Threads [173, 950, 1935]. Three [318, 177, 1948, 1798, 63, 748, 2248, 1988, 848, 1626, 1719, 306, 288, 1224]. Three-Dimensional [2248, 1988, 1224]. Three-Valued [1719]. Threshold [262, 2211, 47, 1694]. Ticketing [916]. Tight [977, 1740]. Tile [2189]. Tile-Based [2189]. Tileworld [392]. Tiling [942]. Tilings [1289]. Time [1427, 1050, 1132, 749, 2198, 772, 1437, 1329, 894, 1400, 225, 789, 1905, 773, 774, 220, 116, 858, 1044, 646, 539, 1424, 2239, 776, 1741, 259, 2091, 1101, 1027, 778, 2131, 359, 100, 785, 1130, 1830, 2040, 1106, 1326, 1577, 1832, 275, 764, 2323, 2216, 2218, 1838, 2289, 2265, 1654, 1580, 2232, 1878, 786, 2277, 1931, 2021, 2292, 1120, 1972, 1264, 187, 309, 1070, 1989, 1996, 2001]. Time-Dependent [220]. Time-Series [1427, 2021]. Time-Synchronised [1044]. Time-Varying [1905]. Time-Warping [1878]. Timed [222, 229, 475, 777, 212, 788, 228, 217, 1629, 191, 779, 1072, 1247]. Times [936, 1127]. timing [1245]. TINA [917, 914, 907]. Tissue [739, 431, 748]. TM [288]. TMN [918, 916, 922, 926, 917]. TMN-like [917]. TOAS [1435]. Toeplitz [118]. Tolerant [1202, 1709, 250, 1989]. Tolerate [979]. Tomographic [740, 763]. Tomography [114, 332]. Tool [166, 624, 679, 2258, 1789, 529, 366, 1264, 983, 533, 986, 389, 1367, 499, 414, 660, 1036, 880, 1860, 1766, 1430, 1530, 1116, 1119, 1267, 1262, 1121, 1263]. Toolbox [1857, 1613]. Toolkit [59, 965]. Tools [1051, 102, 650, 883, 758, 409, 1365, 218, 1180, 1801, 1194, 1238, 1124]. Toolset [1195, 1203, 1365, 531, 571]. Top [1537, 2006, 1150, 1149]. Top-down [1150, 1149]. Top-Evolutionary [2006]. topographic [1206]. Topological [854, 264]. Topologies [1625]. Topology [2201, 1447, 990, 694, 1722, 2004]. TopVision [1061]. Torque [718]. Torus [1730]. Total [1344, 6, 128, 325, 322, 801]. Totality [1913]. Totally [1628]. Tour [1707, 1281]. Toxicity [968]. Trace [1309]. traced [406]. Tracing [2121]. Tracked [1369]. Trackies [279]. Tracking [464, 2239, 606, 2291, 601, 318, 1208, 283, 319, 308]. Tractable [22]. Trade [1298, 1841]. Trade-Off [1841]. Trade-offs [1298]. trading [1553]. Traditional [758, 666]. Traffic [1059, 1491, 1011, 1630, 1984]. Trails [2121]. Training [736, 716, 962, 307]. TRAM [818]. Trans [1607]. Trans-dichotomous [1607]. Transaction [1185, 1557, 1767]. Transactions [1558, 1174, 42, 17]. Transcopyright [1632]. transcranial [315]. Transcription [1801]. Transducer [955]. Transducers [473, 1933, 1836]. TRANSFAC [1801]. Transfer [109, 51, 1655, 138, 137]. Transfers [841]. Transform [333, 118, 2254, 2304, 1528, 2292, 1775, 2305]. Transform-Based [2254]. Transformation [780, 522, 366, 2208, 1433, 2205, 2109, 2142, 2245, 283, 1455]. Transformation-Based [522]. Transformations [937, 1138, 2111, 652]. Transformers [1019]. Transforming [2110]. Transforms [2311]. transient [1120]. Transition [2197, 1916, 2200, 83, 1317, 1904, 2210, 226, 1083]. Transitions [212]. Transitive [1448, 50, 1254]. transitive-closure [1254]. Translating [183, 2067]. Translation [1663, 55, 890, 199, 2130, 19, 394]. Translations [1148, 2130, 2245]. Transmission [1824, 1053]. Transnational [1042]. Transparency [1184]. Transparent [174]. Transputer [646]. Transputers [514]. transverse [321]. Trapdoor [1522]. Trapezoid [1708]. Traveling [465]. Travois [1527]. Treating [413]. Treatment [112, 209, 174, 717, 327]. Tree [1399, 2047, 57, 150, 990, 260, 1869, 2341, 1733, 269, 1836, 1803, 1418, 1686, 1645, 1980]. treeNets [1110]. Trees [52, 1306, 473, 851, 1290, 1714, 1696, 1429, 144, 416, 149, 1684, 2336, 1541, 2206, 1343, 588, 1320, 1698, 261, 1106, 1726, 2248, 1933, 1728, 1872, 2214, 269, 1713, 1712, 1652, 1602, 1722, 1723, 139, 159, 1417, 1973, 1975, 1983, 1706, 1992, 1993, 158]. Treewidth [2186, 1307]. TREMBL [1795]. Trends [468, 1066]. Triadic [1354]. Trial [1061, 1068]. Trials [1063]. Triangular [132]. Triangulations [1598, 1732, 1690, 2001]. Tridiagonal [636]. Triple [1338]. Triple-Free [1338]. trivial [2201]. Trouble [916]. TrPTL [2224]. TRRD [1801]. True [789, 806]. Truncating [610]. Trust [50, 378, 1542]. TSP [1578, 1582]. TTP [1551]. TTP-based [1551]. Tumorous [431]. Tumors [31]. Tuning [1784]. Tunnel [714]. Tuple [1742, 1964, 1953]. Tuple-Based [1742]. tuple-space [1964]. Turing [251, 2192]. Tutorial [214, 1440, 1790]. Twenty [595]. Twin [242]. Twin-Base [242]. Two [988, 53, 624, 109, 1443, 461, 819, 769, 2055, 509, 2187, 547, 63, 1275, 726, 588, 196, 558, 636, 1724, 1836, 807, 2221, 2063, 1732, 2069, 134, 21, 1551, 299, 1122, 1983, 1988, 1991]. two-connectivity [1991]. Two-Dimensional [1443, 1836, 21, 299, 1983, 1988]. Two-Letter [1724]. Two-Level [2221]. Two-Mass [726]. Two-Phase [109]. Two-Step [134]. Two-Variable [196]. Two-Way [1275, 636, 1122]. Two-Wheeler [624]. TWOPRIME [1517]. Type [1920, 132, 855, 1372, 2365, 1703, 273, 1667, 401, 1540, 1294, 990, 1033, 589, 75, 1378, 1133, 20, 1912, 2067, 2171, 488, 133, 274, 1682, 712, 398, 404]. Type-Based [589, 488, 1540]. Type-consistency [20]. Type-Theoretic [1920]. Typed [395, 2141, 493, 408]. Typelab [498]. Types [1910, 1018, 2364, 2083, 483, 783, 1926, 1932, 615, 1022, 1032, 2044, 1035, 1929, 396, 794]. Typing [1934, 1159, 1080].

Ugly [2168]. Ultrafilter [2048]. Ultrasonic [1482]. Ultrasound [606, 318, 298, 314]. UMTS [925]. unary [15]. Uncertain [1088, 416, 1111, 1863]. Uncertainties [120]. Uncertainty [770, 1741, 2257, 1907]. Unconditional [2165]. Unconditionally [2147, 2148]. Undecidability [478, 1377, 196, 809, 2069, 1383, 796]. Undecidable [1845, 810]. Undeniable [2154]. Underlying [958]. Understanding [2049]. Unfolding [367]. Unification [1357, 1296, 2358, 1717, 2069, 815, 812]. Unified [1031, 1758]. Uniform [873, 1115, 1827, 1105, 2369, 942]. Uniformly [548]. Unifying [530]. Union [184, 154]. Union-Find [184, 154]. Unions [1186]. Unique [1344, 947, 1851]. Uniqueness [2263]. Unit [1597, 728, 1862, 1192, 1297]. Unit-Testing [1192]. UNITY [1955, 1956]. Universal [1645, 2107, 1358, 715, 859, 1233, 1850, 1073]. Universality [205, 1840]. University [279]. Unknown [1600, 2028]. Unpredicted [1101]. Unsteady [1778]. unstructured [18]. Unsupervised [594, 990, 596, 1012, 2304, 2236, 2266]. Unsymmetric [1754]. untangling [1935]. Unwrapping [295]. UOWHFs [2178]. Update [519, 521, 1114, 8]. Update-Programs [519]. Updates [255, 1467]. Updating [260, 1925, 1889]. Upper [1313, 1607, 208, 65, 1701, 2071]. Ups [1102]. Upward [149]. Usage [2093, 1242]. Use [762, 735, 506, 118, 2092, 962, 24, 1901, 1680, 9, 1551, 1632]. Used [629, 299]. User [1060, 884, 650, 410, 1364, 928, 951, 1037, 1855, 927, 13]. User-Adapted [410]. User-Defined [951]. User-Driven [1364]. User-Level [650]. Using [1041, 1526, 720, 2303, 683, 485, 994, 1427, 993, 1686, 1097, 1664, 1538, 27, 594, 333, 1398, 1502, 238, 1857, 668, 713, 772, 1437, 2200, 1491, 691, 1757, 1928, 1905, 1492, 740, 29, 2309, 510, 220, 1065, 465, 2311, 908, 1137, 2007, 2308, 1363, 31, 2356, 1496, 339, 2333, 343, 1494, 2336, 638, 1016, 1157, 593, 1008, 352, 1309, 1231, 1741, 357, 2091, 458, 1497, 414, 547, 755, 1670, 748, 1355, 538, 1129, 1161, 514, 1173, 698, 743, 604, 1897, 34, 1361, 479, 852, 605, 460, 240, 2273, 349, 496, 1107, 1421, 765, 340, 2314, 2101]. Using [750, 1410, 1884, 1108, 451, 243, 625, 742, 38, 1239, 1505, 751, 1388, 174, 2140, 1674, 2162, 2024, 2010, 2271, 1037, 2160, 2304, 2313, 1038, 1626, 2317, 2110, 2236, 2288, 1793, 2284, 1964, 2349, 49, 2016, 2121, 2351, 2240, 2171, 2025, 2251, 717, 1878, 1488, 2352, 338, 42, 2277, 759, 932, 2019, 824, 2292, 598, 1439, 412, 1226, 1213, 1219, 282, 294, 314, 1260, 317, 921, 309, 289, 1252, 1946, 1218, 284, 792, 1941, 811]. Utilization [108, 2100]. utilizing [288].

vacuity [1253]. Vague [1415]. Valences [2207]. Validate [1173]. Validating [412]. Validation [1172, 1109]. Validity [1927, 1882]. Value [113, 2113, 1918, 134, 1303]. Value-Passing [1918]. Valued [2302, 1370, 451, 1719, 1700]. Values [860, 1900, 1877]. Vanstone [2158]. Vapnik [372]. Variable [934, 232, 1375, 196, 1694]. Variables [2370, 1331, 2010, 2069, 1986]. Variant [1228, 2368]. Variation [738]. Varying [1905, 1499]. Vasculature [768]. VC [383]. VC-dimension [383]. VDM [1662]. VDM-SL [1662]. Vector [505, 1769, 515]. Vectors [618, 460, 1240]. Vegas [187]. Vehicle [1044, 1626, 2319]. vein [299]. Velocity [764, 1621]. Ventilation [991, 714]. Verbose [2226]. Verification [487, 229, 524, 772, 346, 1666, 363, 189, 537, 221, 535, 362, 784, 359, 360, 354, 1673, 2043, 1746, 1966, 1677, 226, 779, 872, 1244, 1238, 1247, 1264, 1237, 322, 1239, 1259]. Verified [120]. Verifier [2147, 1248, 2148]. Verify [1670]. Verifying [222, 1665, 1671, 2070, 1255, 1083]. VeriSoft [1267]. Versatile [1428]. Version [873, 1158, 1070]. Versus [521, 553, 187, 892, 2280, 875, 1206, 933, 192, 1452, 2212, 2214, 293]. vertebra [290]. vertebral [319]. Vertex [211, 1706, 248]. Vertex-Connectivity [211]. Vertex-Rankings [1706]. Vertical [1590, 1079]. vertices [1981]. Verus [774, 1264]. very [1119]. Vessel [2247, 769]. Vessels [761]. Via [1812, 1019, 600, 1950, 2332, 857, 1638, 1346, 117, 339, 1949, 247, 1131, 1402, 1151, 879, 2358, 1685, 1815, 1639, 592]. Viability [1761]. Video [669, 723, 356, 905, 1055, 1655, 338, 1553, 332]. Video-based [338]. Video-Data [1655]. View [964, 2189, 711, 1935]. Viewing [966]. Viewpoints [1505]. Views [768, 2320, 1646, 2286, 5, 16]. VINES [1651]. Virtual [762, 302, 102, 1374, 1233, 1046, 2144, 1810, 1035, 299, 310]. Visibility [54, 769, 1572, 64]. Visibility-Based [1572]. Vision [2324, 2264, 2318, 2238, 2289, 279, 1000, 311].

Vision-Based [2318, 2238, 2289, 279]. VISLAN [317]. Visual [537, 1016, 1048, 995, 778, 1702, 1886, 2144, 2167, 1763, 1873, 1793, 1766, 355]. Visualization [466, 60, 1809, 72]. Visualizing [67, 667]. Vitro [431]. VLIW [1533]. VLSI [1016, 155]. voice [1122]. Volume [56, 754, 628, 309, 293]. volume-based [293]. volumetric [309]. Voronoi [1587]. Voting [612, 1302, 2280]. VQ [347]. VRML [929, 1646]. VRML-object [1646]. vs [921, 895, 1093, 2374, 1784, 1594].

W [440]. W.I.S.E. [1857]. WAAC [2315]. Wadge [1287]. Wadge-Wegner [1287]. Wait [477]. Wait-Free [477]. Walk [2267]. Want [1494]. Warehouse [1567, 5]. Warping [1878]. Was [771]. Waste [717]. Watchman [1830]. Water [2015, 717]. Watermarking [1048, 1055]. Watersheds [2284]. Wave [626, 1774]. Wavelength [1319]. Wavelet [2288, 1720]. Wavelet-Based [1720]. Wavelets [1905, 2274]. Way [838, 2172, 1275, 1395, 266, 636, 1122, 187, 1837]. WDM [1320]. Weak [1243, 481, 447, 1444, 1382, 227, 408, 1454, 407]. weak-head [407]. Weaker [1310]. Weakness [834]. weaving [1935]. WEB [1124, 14, 641, 679, 1935, 161, 1181, 879, 2059, 1651, 2144, 1639]. Web-Based [641, 679]. Web-like [2059]. WebRC [880]. Wegner [1287]. Weight [211, 979, 1284]. Weighted [1977, 1723, 1968, 339]. weighting [296]. Weights [978, 1499, 451, 1747, 806]. Well [722, 522, 523, 278, 2250, 791, 6]. Well-Founded [522, 523, 6]. Well-Posedness [2250]. Wheeler [624]. Where [480, 2123]. Which [2353, 1825, 378]. Wide [161, 879, 2144, 1652, 1660, 179]. Width [1604, 208, 1925, 267]. Width-Subtyping [1925]. widths [384]. WIND [1567]. window [1239]. Windows [1896]. wise [846]. WISH [1639]. within [2164, 2046]. Without [851, 752, 2298, 2288, 2198, 1607, 1025]. Word [473, 1141, 1446, 710, 2377, 1832, 1292, 1845, 795, 800]. Word-Concept [710]. Word-into-Trees [473]. Words [2198, 1444, 558, 2217]. Work [1053, 12]. Worker [649]. Workflow [1414, 1185, 1184]. Workflows [1178]. Workload [2093, 2094, 1125]. Workloads [2099]. Workstation [175, 176]. Workstations [643, 677, 1768, 682, 941, 1761]. World [418, 1660, 891, 2081, 161, 879, 2144]. World-Wide-Web [2144]. Worlds [2005]. Worldwide [1636]. Worst [1298]. Worst-Case [1298]. Would [2022]. Wrist [2258]. WSIS [1402]. WWW [55, 908, 1639, 1650].

x [2003, 291, 1802, 764, 290]. X-Ray

[764, 291, 290]. X.500 [878]. XMX [1524].

YACS [1802]. Yau [507]. years [1250].

Yeast [702, 1809]. YOHO [360].

Z [439, 449, 448, 444, 447, 443, 438, 437, 436, 445]. Z0nq [1274]. Z4 [1278]. Zeno [2198]. Zero [838, 769, 1334, 2146]. Zero-filled [769]. Zero-Knowledge [838]. Zeroes [708]. Zm [2220]. Zqn [1274]. Zqn-Linear [1274].

References

Abiteboul:1997:QSS

[1] S. Abiteboul. Querying semistructured data. Lecture Notes in Computer Science, 1186:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chekuri:1997:CQC

[2] C. Chekuri and A. Rajaraman. Conjunctive query containment revisited. Lecture Notes in Computer Science, 1186:56–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grahne:1997:SCQ

[3] G. Grahne, N. Spyratos, and D. Sramate. Semantics and containment of queries with internal and external conjunctions. Lecture Notes in Computer Science, 1186:71–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huyn:1997:ECL

[4] N. Huyn. Efficient complete local tests for conjunctive query constraints with negation. Lecture Notes in Computer Science, 1186:83–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1997:SVM

[5] H. Gupta. Selection of views to materialize in a data warehouse. Lecture Notes in Computer Science, 1186:98– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flum:1997:TPW

[6] J. Flum, M. Kubierschky, and B. Ludaescher. Total and partial wellfounded datalog coincide. Lecture Notes in Computer Science, 1186:113– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dong:1997:LPQ

[7] G. Dong, L. Libkin, and L. Wong. Local properties of query languages. Lecture Notes in Computer Science, 1186: 140–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1997:MTA

[8] N. Bidoit and S. Maabout. A model theoretic approach to update rule programs. Lecture Notes in Computer Science, 1186:173–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bailey:1997:AIA

[9] J. Bailey, L. Crnogorac, K. Ramamohanarao, and H. Soendergaard. Abstract interpretation of active rules and its use in termination analysis. Lecture Notes in Computer Science, 1186: 188–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bailey:1997:SIA

[10] J. Bailey, G. Dong, and K. Ramamohanarao. Structural issues in active rule systems. Lecture Notes in Computer Science, 1186:203–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunopulos:1997:DAM

[11] D. Gunopulos, H. Mannila, and S. Saluja. Discovering all most specific sentences by randomized algorithms. Lecture Notes in Computer Science, 1186:215–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wodtke:1997:FFD

[12] D. Wodtke and G. Weikum. A formal foundation for distributed work flow execution based on state charts. Lecture Notes in Computer Science, 1186: 230–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fagin:1997:IUP

[13] R. Fagin and E. L. Wimmers. Incorporating user preferences in multimedia queries. Lecture Notes in Computer Science, 1186:247–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abiteboul:1997:QCW

[14] S. Abiteboul and V. Vianu. Queries and computation on the Web. Lecture Notes in Computer Science, 1186:262– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benedikt:1997:EPU

[15] M. Benedikt and H. J. Keisler. Expressive power of unary counters. Lecture Notes in Computer Science, 1186: 291–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawaguchi:1997:CCT

[16] A. Kawaguchi, D. Lieuwen, I. Mumick, and D. Quass. Concurrency control theory for deferred materialized views. Lecture Notes in Computer Science, 1186:306–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halici:1997:SNT

[17] U. Halici, B. Arpinar, and A. Dogac. Serializability of nested transactions in multidatabases. Lecture Notes in Computer Science, 1186:321–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buneman:1997:ASU

[18] P. Buneman, S. Davidson, M. Fernandez, and D. Suciu. Adding structure to unstructured data. Lecture Notes in Computer Science, 1186:336–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abiteboul:1997:CTH

[19] S. Abiteboul, S. Cluet, and T. Milo. Correspondence and translation for heterogeneous data. Lecture Notes in Computer Science, 1186:351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishihara:1997:TCP

[20] Y. Ishihara, H. Seki, and M. Ito. Type-consistency problems for queries in object-oriented databases. Lecture Notes in Computer Science, 1186:364– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdel-Ghaffar:1997:OAT

[21] K. Abdel-Ghaffar and A. El Abbadi. Optimal allocation of two-dimensional data. Lecture Notes in Computer Science, 1186:409–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colby:1997:TIM

[22] L. Colby and L. Libkin. Tractable iteration mechanisms for bag languages. Lecture Notes in Computer Science, 1186:461–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yager:1997:CPF

[23] R. R. Yager. Constructing prioritized fuzzy models. Lecture Notes in Computer Science, 1188:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hurley:1997:UFR

[24] N. Hurley. The use of fuzzy representation in a CBR system for mesh design. Lecture Notes in Computer Science, 1188:29–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonner:1997:FFL

[25] M. Bonner, S. Mayer, A. Raggl, and W. Slany. FLIP++: A fuzzy logic inference processor library. Lecture Notes in Computer Science, 1188:44–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herrmann:1997:FLI

[26] C. S. Herrmann. Fuzzy logic as interfacing technique in hybrid AI-Systems. Lecture Notes in Computer Science, 1188:69–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldwin:1997:EKD

[27] J. F. Baldwin and T. P. Martin. Extracting knowledge from data using an intelligent fuzzy data browser. Lecture Notes in Computer Science, 1188:81– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abe:1997:FSL

[28] S. Abe. Fuzzy systems with learning capability. Lecture Notes in Computer Science, 1188:101–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cannavacciuolo:1997:SOM

[29] A. Cannavacciuolo, G. Capaldo, A. Ventre, and A. Volpe. Searching for the organizational memory using fuzzy modeling. Lecture Notes in Computer Science, 1188:144–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloch:1997:FGD

[30] I. Bloch. Fuzzy geodesic distance in images. Lecture Notes in Computer Science, 1188:153–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clark:1997:UFI

[31] M. C. Clark, L. O. Hall, D. B. Goldgof, and M. S. Silbiger. Using fuzzy information in knowledge guided segmentation of brain tumors. Lecture Notes in Computer Science, 1188:167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ho:1997:FFE

[32] K. H. L. Ho and N. Ohnishi. FEDGE — fuzzy edge detection by fuzzy categorization and classification of edges. Lecture Notes in Computer Science, 1188: 182–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guesgen:1997:THS

[33] H. W. Guesgen. Towards hybrid spatial reasoning. Lecture Notes in Computer Science, 1188:197–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasos:1997:MRL

[34] J. Gasos and A. Martin. Mobile robot localization using fuzzy maps. Lecture Notes in Computer Science, 1188:207– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1997:TPD

[35] D. Dubois and H. Prade. Towards possibilistic decision theory. Lecture Notes in Computer Science, 1188:240– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bilgic:1997:MTF

[36] T. Bilgic and I. B. Tuerksen. Measurement-theoretic frameworks for fuzzy set theory. Lecture Notes in Computer Science, 1188:252–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouchon-Meunier:1997:RAA

[37] B. Bouchon-Meunier and L. Valverde. A resemblance approach to analogical reasoning functions. Lecture Notes in Computer Science, 1188:266–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joye:1997:PFR

[38] Marc Joye and Jean-Jacques Quisquater. Protocol failures for RSA-like functions using Lucas sequences and elliptic curves. Lecture Notes in Computer Science, 1189:93–100, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crispo:1997:CSE

[39] B. Crispo and M. Lomas. A certification scheme for electronic commerce. Lecture Notes in Computer Science, 1189:19–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujisaki:1997:PEC

[40] E. Fujisaki and T. Okamoto. Practical escrow cash system. Lecture Notes in Computer Science, 1189:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1997:NPE

[41] R. Anderson, C. Manifavas, and C. Sutherland. NetCard — A practical electronic-cash system. Lecture Notes in Computer Science, 1189:49– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wheeler:1997:TUB

[42] D. Wheeler. Transactions using bets. Lecture Notes in Computer Science, 1189:89–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cramer:1997:EPS

[43] R. Cramer, I. Damgaard, and T. P. Pedersen. Efficient and provable secure amplifications. Lecture Notes in Computer Science, 1189:101–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cusick:1997:CRN

[44] T. W. Cusick. A comparison of RSA and the Naccache-Stern publickey cryptosystem. Lecture Notes in Computer Science, 1189:111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaliski:1997:IPS

[45] B. S. Kaliski. IEEE P1363: A standard for RSA, Diffie–Hellman and ellipticcurve cryptography [abstract]. Lecture Notes in Computer Science, 1189:117– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burmester:1997:ESC

[46] M. Burmester and Y. G. Desmedt. Efficient and secure conference-key distribution. Lecture Notes in Computer Science, 1189:119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lim:1997:DSA

[47] Chae Hoon Lim and Pil Joong Lee. Directed signatures and application to threshold cryptosystems. Lecture Notes in Computer Science, 1189:131–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:KEM

[48] L. Chen, D. Gollmann, and C. J. Mitchell. Key escrow in mutually mistrusting domains. Lecture Notes in Computer Science, 1189:139–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneier:1997:AES

[49] B. Schneier and J. Kelsey. Automatic event-stream notarization using digital signature. Lecture Notes in Computer Science, 1189:155–169, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christianson:1997:WIT

[50] B. Christianson and W. S. Harbison. Why isn’t trust transitive? Lecture Notes in Computer Science, 1189:171– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chuang:1997:SRA

[51] S.-C. Chuang and D. Greaves. Securing residential asynchronous transfer mode networks. Lecture Notes in Computer Science, 1189:177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abellanas:1997:BET

[52] M. Abellanas, J. Garcia, G. Hernandez, and M. Noy. Bipartite embeddings of trees in the plane. Lecture Notes in Computer Science, 1190:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alzohairi:1997:SPP

[53] M. Alzohairi and I. Rival. Seriesparallel planar ordered sets have pagenumber two. Lecture Notes in Computer Science, 1190:11–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bose:1997:RVG

[54] P. Bose, A. Dean, J. Hutchinson, and T. Shermer. On rectangle visibility graphs. Lecture Notes in Computer Science, 1190:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bridgeman:1997:GDT

[55] S. Bridgeman, A. Garg, and R. Tamassia. A graph drawing and translation service on the WWW. Lecture Notes in Computer Science, 1190:45–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calamoneri:1997:DCG

[56] T. Calamoneri and A. Sterbini. Drawing 2-, 3- and 4-colorable graphs in O(n^{2}) volume. Lecture Notes in Computer Science, 1190:53–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1997:OAA

[57] T. Chan, M. T. Goodrich, S. R. Kosaraju, and R. Tamassia. Optimizing area and aspect ratio in straightline orthogonal tree drawings. Lecture Notes in Computer Science, 1190:63– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiBattista:1997:DDA

[58] G. Di Battista, A. Garg, G. Liotta, and A. Parise. Drawing directed acyclic graphs: An experimental study. Lecture Notes in Computer Science, 1190:76–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dogrusoez:1997:CLG

[59] U. Dogrusoez, B. Madden, and P. Madden. Circular layout in the graph layout toolkit. Lecture Notes in Computer Science, 1190:92–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:MVC

[60] P. Eades and Q. W. Feng. Multilevel visualization of clustered graphs. Lecture Notes in Computer Science, 1190: 101–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:SLD

[61] P. Eades, Q. W. Feng, and X. Lin. Straight-line drawing algorithms for hierarchical graphs and clustered graphs. Lecture Notes in Computer Science, 1190:113–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:GDC

[62] P. Eades, J. Marks, and S. North. Graph-drawing contest report. Lecture Notes in Computer Science, 1190:129– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:TAT

[63] P. Eades, A. Symvonis, and S. Whitesides. Two algorithms for three dimensional orthogonal graph drawing. Lecture Notes in Computer Science, 1190: 139–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1997:VDP

[64] U. Foessmeier, G. Kant, and M. Kaufmann. 2-Visibility drawings of planar graphs. Lecture Notes in Computer Science, 1190:155–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frick:1997:UBN

[65] A. Frick. Upper bounds on the number of hidden nodes in Sugiyama’s algorithm. Lecture Notes in Computer Science, 1190:169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frick:1997:IDA

[66] A. Frick, C. Keskin, and V. Vogelmann. Integration of declarative approaches. Lecture Notes in Computer Science, 1190:184–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garg:1997:GSV

[67] A. Garg and R. Tamassia. GIOTTO3D: A system for visualizing hierarchical structures in 3D. Lecture Notes in Computer Science, 1190:193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garg:1997:NMC

[68] A. Garg and R. Tamassia. A new minimum cost flow algorithm with applications to graph drawing. Lecture Notes in Computer Science, 1190:201– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1997:CGL

[69] W. He and K. Marriott. Constrained graph layout. Lecture Notes in Computer Science, 1190:217–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Himsolt:1997:GS

[70] M. Himsolt. The graphlet system. Lecture Notes in Computer Science, 1190: 233–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kakoulis:1997:ELP

[71] K. G. Kakoulis and I. G. Tollis. On the edge label placement problem. Lecture Notes in Computer Science, 1190: 241–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wittenburg:1997:QVP

[72] K. Wittenburg and L. Weitzman. Qualitative visualization of processes: Attributed graph layout and focusing techniques. Lecture Notes in Computer Science, 1190:401–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comini:1997:PPL

[73] M. Comini, G. Levi, M. C. Meo, and G. Vitiello. Proving properties of logic programs by abstract diagnosis. Lecture Notes in Computer Science, 1192:22–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fredlund:1997:FPC

[74] L.-A. Fredlund, J. Koistinen, and F. Orava. Formalising and prototyping a concurrent object-based language. Lecture Notes in Computer Science, 1192:116–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fradet:1997:TCM

[75] P. Fradet and D. Le Metayer. Type checking for a multiset rewriting language. Lecture Notes in Computer Science, 1192:126–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amtoft:1997:PSE

[76] T. Amtoft, F. Nielson, H. R. Nielson, and J. Ammann. Polymorphic subtyping for effect analysis: The dynamic semantics. Lecture Notes in Computer Science, 1192:172–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cridlig:1997:ISA

[77] R. Cridlig. Implementing a static analyzer of concurrent programs: Problems and perspectives. Lecture Notes in Computer Science, 1192:244–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faxen:1997:PPF

[78] K.-F Faxen. Polyvariance, polymorphism, and flow analysis. Lecture Notes in Computer Science, 1192:260– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilhelm:1997:PIF

[79] R. Wilhelm, M. Alt, F. Martin, and M. Raber. Parallel implementation of functional languages. Lecture Notes in Computer Science, 1192:279–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degano:1997:MPL

[80] P. Degano, J.-V. Loddo, and C. Priami. Mobile processes with local clocks. Lecture Notes in Computer Science, 1192:296–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNicola:1997:TSA

[81] R. De Nicola and R. Pugliese. Testing semantics of asynchronous distributed programs. Lecture Notes in Computer Science, 1192:320–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degano:1997:AFP

[82] P. Degano, C. Priami, L. Leth, and B. Thomsen. Analysis of FACILE programs. Lecture Notes in Computer Science, 1192:345–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFrancesco:1997:NSS

[83] N. De Francesco, A. Santone, and G. Vaglini. A non-standard semantics for generating reduced transition systems. Lecture Notes in Computer Science, 1192:370–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franklin:1997:IAJ

[84] S. Franklin and A. Graesser. Is it an agent, or just a program?: A taxonomy for autonomous agents. Lecture Notes in Computer Science, 1193:21– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castelfranchi:1997:A

[85] C. Castelfranchi. To be or not to be an agent. Lecture Notes in Computer Science, 1193:37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wooldridge:1997:ART

[86] M. Wooldridge. Agents as a Rorschach test: A response to Franklin and Graesser. Lecture Notes in Computer Science, 1193:47–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1997:IAW

[87] M. Fisher. If Z is the answer, what could the question possibly be? Lecture Notes in Computer Science, 1193: 65–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1997:PTT

[88] G. Wagner. Practical theory and theory-based practice. Lecture Notes in Computer Science, 1193:67–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dragoni:1997:BRT

[89] A. F. Dragoni and P. Giorgini. Belief revision through the belief-function formalism in a multi-agent environment. Lecture Notes in Computer Science, 1193:103–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benerecetti:1997:FSB

[90] M. Benerecetti, A. Cimatti, E. Giunchiglia, and F. Giunchiglia. Formal specification of beliefs in multi-agent systems. Lecture Notes in Computer Science, 1193:117–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brazier:1997:FCM

[91] F. Brazier, C. Jonker, and J. Treur. Formalisation of a cooperation model based on joint intentions. Lecture Notes in Computer Science, 1193:141– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bretier:1997:RAK

[92] P. Bretier and D. Sadek. A rational agent as the kernel of a cooperative spoken dialogue system: Implementing a logical theory of interaction. Lecture Notes in Computer Science, 1193:189–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dignum:1997:MSA

[93] F. Dignum and B. Van Linder. Modelling social agents: Communication as action. Lecture Notes in Computer Science, 1193:205–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeng:1997:HCA

[94] D. Zeng and K. Sycara. How can an agent learn to negotiate? Lecture Notes in Computer Science, 1193:233– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botelho:1997:EBA

[95] L. M. Botelho and H. Coelho. Emotionbased attention shift in autonomous agents. Lecture Notes in Computer Science, 1193:277–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davis:1997:RMA

[96] D. Davis. Reactive and motivational agents: Towards a collective minder. Lecture Notes in Computer Science, 1193:309–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbuceanu:1997:DCL

[97] M. Barbuceanu and M. Fox. The design of a coordination language for multi-agent systems. Lecture Notes in Computer Science, 1193:341–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wooldridge:1997:KTS

[98] M. Wooldridge. A knowledge-theoretic semantics for concurrent METATEM. Lecture Notes in Computer Science, 1193:357–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badler:1997:TPA

[99] N. I. Badler, B. D. Reich, and B. L. Webber. Towards personalities for animated agents with reactive and planning behaviors. Lecture Notes in Computer Science, 1195:43–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldberg:1997:ISR

[100] A. Goldberg. IMPROV: A system for real-time animation of behavior-based interactive synthetic actors. Lecture Notes in Computer Science, 1195:58– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blumberg:1997:MLC

[101] B. Blumberg and T. Galyean. Multilevel control for animated autonomous agents: Do the right thing… oh, not that. Lecture Notes in Computer Science, 1195:74–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blair:1997:TIV

[102] D. Blair and T. Meyer. Tools for an interactive virtual cinema. Lecture Notes in Computer Science, 1195:83– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayes-Roth:1997:AC

[103] B. Hayes-Roth, R. Van Gent, and D. Huber. Acting in character. Lecture Notes in Computer Science, 1195: 92–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akishin:1997:NCM

[104] P. G. Akishin, I. V. Puzynin, and Y. S. Smirnov. The Newtonian continuation method for numerical study of 3D polaron problem. Lecture Notes in Computer Science, 1196:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amirkhanov:1997:CNI

[105] I. V. Amirkhanov, I. V. Puzynin, T. P. Puzynina, and T. A. Strizh. CANM in numerical investigation of QCD problems. Lecture Notes in Computer Science, 1196:9–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amodio:1997:PIS

[106] P. Amodio and F. Mazzia. Parallel iterative solvers for banded linear systems. Lecture Notes in Computer Science, 1196:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreev:1997:FEA

[107] A. B. Andreev and T. D. Todorov. Finite element approximations of some central curves. Lecture Notes in Computer Science, 1196:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayrian:1997:UCM

[108] E. A. Ayrian and A. V. Fedorov. Utilization of CANM for magnetic field calculation. Lecture Notes in Computer Science, 1196:34–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bazhlekov:1997:NAS

[109] I. B. Bazhlekov, M. G. Koleva, and D. P. Vasileva. Numerical algorithm for simulation of coupled heat-mass transfer and chemical reaction in twophase steady flow. Lecture Notes in Computer Science, 1196:38–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bisseling:1997:BTN

[110] R. H. Bisseling. Basic techniques for numerical linear algebra on bulk synchronous parallel computers. Lecture Notes in Computer Science, 1196:46– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braianov:1997:CCN

[111] I. Braianov. Convergence of a Crank– Nicolson difference scheme for heat equations with interface in the heat flow and concentrated heat capacity. Lecture Notes in Computer Science, 1196:58–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandt:1997:TLA

[112] J. Brandt, J. Christensen, I. Dimov, and K. Georgiev. Treatment of large air pollution models. Lecture Notes in Computer Science, 1196:66–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brugnano:1997:BVM

[113] L. Brugnano. Boundary value methods for the numerical approximation of ordinary differential equations. Lecture Notes in Computer Science, 1196:78– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1997:FAP

[114] R. H. Chan, C. Cheung, and H. Sun. Fast algorithms for problems on thermal tomography. Lecture Notes in Computer Science, 1196:90–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:CIP

[115] T. C. Chen. Convergence in iterative polynomial root-finding. Lecture Notes in Computer Science, 1196:98– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clavero:1997:STM

[116] C. Clavero, J. C. Jorge, F. Lisbona, and G. I. Shishkin. Splitting time methods and one dimensional special meshes for reaction-diffusion parabolic problems. Lecture Notes in Computer Science, 1196:106–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corliss:1997:HOS

[117] G. F. Corliss, A. Griewank, P. Henneberger, and G. Kirlinger. High-order stiff ODE solvers via automatic differentiation and rational prediction. Lecture Notes in Computer Science, 1196: 114–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiBenedetto:1997:UDS

[118] F. Di Benedetto. The use of discrete sine transform in computations with Toeplitz matrices. Lecture Notes in Computer Science, 1196:126–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMarco:1997:FSS

[119] S. C. Di Marco and R. L. V. Gonzalez. A finite state stochastic minimax optimal control problem with infinite horizon. Lecture Notes in Computer Science, 1196:134–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimitrova:1997:VSL

[120] N. S. Dimitrova and C. P. Ullrich. Verified solving of linear systems with uncertainties in MAPLE. Lecture Notes in Computer Science, 1196:142–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimov:1997:IMC

[121] I. T. Dimov and A. N. Karaivanova. Iterative Monte Carlo algorithms for linear algebra problems. Lecture Notes in Computer Science, 1196:150–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dyakonov:1997:OPS

[122] E. G. D’yakonov. Operator problems in strengthened Sobolev spaces and numerical methods for them. Lecture Notes in Computer Science, 1196:161– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elsner:1997:CCM

[123] L. Elsner and T. Szulc. Convex combinations of matrices-nonsingularity and Schur stability. Lecture Notes in Computer Science, 1196:170–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filippov:1997:NSB

[124] S. S. Filippov, Y. V. Sanochkin, A. Y. Sogomonyan, and A. V. Tygliyan. Numerical solution of a BVP in hydrodynamics (thermocapillary convection). Lecture Notes in Computer Science, 1196:176–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoermann:1997:CEB

[125] B. Hoermann and C. P. Ullrich. Componentwise error bounds and direct linear system solving. Lecture Notes in Computer Science, 1196:184–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ivanov:1997:MSS

[126] I. G. Ivanov. A method for solving the spectral problem for complex matrices. Lecture Notes in Computer Science, 1196:192–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jovanovic:1997:ITC

[127] B. S. Jovanovic. Interpolation technique and convergence rate estimates for finite difference method. Lecture Notes in Computer Science, 1196:200– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamm:1997:LST

[128] J. Kamm and J. G. Nagy. Least squares and total least squares methods in image restoration. Lecture Notes in Computer Science, 1196:212–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karageorghis:1997:CSD

[129] A. Karageorghis. Conforming spectral domain decomposition schemes. Lecture Notes in Computer Science, 1196: 220–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaschiev:1997:NMS

[130] M. Kaschiev and D. Koulova-Nenova. Newton’s method for solution of one complex eigenvalue problem. Lecture Notes in Computer Science, 1196:227– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaschieva:1997:CBS

[131] V. A. Kaschieva. Calculation of bifurcation states in the circle Josefsson’s conversion. Lecture Notes in Computer Science, 1196:236–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BarreraRosillo:1997:CSF

[132] D. Barrera Rosillo, M. J. Ibathez Perez, and A. Lopez Carmona. Compactly supported fundamental functions for bivariate Hermite spline interpolation and triangular finite elements of HCT type. Lecture Notes in Computer Science, 1196:414–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vulkov:1997:AST

[133] L. G. Vulkov. Applications of Steklovtype eigenvalue problems to convergence of difference schemes for parabolic and hyperbolic equations with dynamical boundary conditions. Lecture Notes in Computer Science, 1196:557–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1997:TSP

[134] P. S. Williams and T. E. Simos. Twostep P-stable methods with phase-lag of order infinity for the numerical solution of special second order initial value problems. Lecture Notes in Computer Science, 1196:565–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yalamov:1997:NCN

[135] P. Y. Yalamov. Notes on the classification of numerical algorithms with respect to their stability to roundoff errors. Lecture Notes in Computer Science, 1196:573–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:NIB

[136] T. Yang. New IMGS-based preconditioners for least squares problems. Lecture Notes in Computer Science, 1196: 581–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheleva:1997:NAS

[137] I. Zheleva, V. Kambourova, P. Dimitrov, and R. Rusev. Numerical algorithm for studying heat-transfer in a glass melting furnace. Lecture Notes in Computer Science, 1196:589–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheleva:1997:IHT

[138] I. Zheleva, V. Kambourova, P. Dimitrov, and R. Rusev. Identification of the heat-transfer in a glass melting furnace. Lecture Notes in Computer Science, 1196:596–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1997:HDT

[139] A. C. Yao. Hypergraphs and decision trees. Lecture Notes in Computer Science, 1197:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alimonti:1997:IAI

[140] P. Alimonti and T. Calamoneri. Improved approximations of independent dominating set in bounded degree graphs. Lecture Notes in Computer Science, 1197:2–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babel:1997:NCP

[141] L. Babel and S. Olariu. A new characterization of P4-connected graphs. Lecture Notes in Computer Science, 1197: 17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauderon:1997:NRH

[142] M. Bauderon and H. Jacquet. Node rewriting in hypergraphs. Lecture Notes in Computer Science, 1197:31– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bezrukov:1997:PC

[143] S. L. Bezrukov. On k-partitioning the n-cube. Lecture Notes in Computer Science, 1197:44–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broadwater:1997:ECB

[144] A. Broadwater, K. Efe, and A. Fernandez. Embedding complete binary trees in product graphs. Lecture Notes in Computer Science, 1197:56–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brys:1997:CAP

[145] K. Brys and Z. Lonc. Clique and anticlique partitions of graphs. Lecture Notes in Computer Science, 1197:67– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:OPR

[146] C. C. Chen and J. Chen. Optimal parallel routing in star graphs. Lecture Notes in Computer Science, 1197:73– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cicerone:1997:CED

[147] S. Cicerone, D. Frigioni, U. Nanni, and F. Pugliese. Counting edges in a dag. Lecture Notes in Computer Science, 1197:85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciotti:1997:CPC

[148] O. Ciotti and F. Parisi-Presicce. Closure properties of context-free hyperedge replacement systems. Lecture Notes in Computer Science, 1197:101– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crescenzi:1997:UDS

[149] P. Crescenzi and P. Penna. Upward drawings of search trees. Lecture Notes in Computer Science, 1197:114–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diks:1997:MGP

[150] K. Diks and T. Hagerup. More general parallel tree contraction: Register allocation and broadcasting in a tree. Lecture Notes in Computer Science, 1197:126–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diks:1997:SDS

[151] K. Diks and A. Pelc. System diagnosis with smallest risk of error. Lecture Notes in Computer Science, 1197:141– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djidjev:1997:EAS

[152] H. N. Djidjev. Efficient algorithms for shortest path queries in planar digraphs. Lecture Notes in Computer Science, 1197:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dragan:1997:LOP

[153] F. F. Dragan, F. Nicolai, and A. Brandstaedt. LexBFS-Orderings and powers of graphs. Lecture Notes in Computer Science, 1197:166–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gustedt:1997:EUF

[154] J. Gustedt. Efficient union-find for planar graphs and other sparse graph classes. Lecture Notes in Computer Science, 1197:181–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartmann:1997:SRV

[155] S. Hartmann, M. W. Schaeffter, and A. S. Schulz. Switchbox routing in VLSI design: Closing the complexity gap. Lecture Notes in Computer Science, 1197:196–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hurfin:1997:DDN

[156] M. Hurfin and M. Raynal. Detecting diamond necklaces in labeled DAGs (A problem from distributed debugging). Lecture Notes in Computer Science, 1197:211–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahl:1997:AGD

[157] W. Kahl. Algebraic graph derivations for graphical calculi. Lecture Notes in Computer Science, 1197:224–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaller:1997:DER

[158] D. Kaller. Definability equals recognizability of partial 3-Trees. Lecture Notes in Computer Science, 1197:239– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1997:GER

[159] X. Zhou, M. A. Kashem, and T. Nishizeki. Generalized edgerankings of trees. Lecture Notes in Computer Science, 1197:390–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watt:1997:ASP

[160] S. N. K. Watt. Artificial societies and psychological agents. Lecture Notes in Computer Science, 1198:27–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davies:1997:IAW

[161] N. J. Davies, R. Weeks, and M. C. Revett. Information agents for the World Wide Web. Lecture Notes in Computer Science, 1198:81–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foner:1997:MAM

[162] L. Foner and I. B. Crabtree. Multiagent matchmaking. Lecture Notes in Computer Science, 1198:100–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisk:1997:ASF

[163] D. Fisk. An application of social filtering to movie recommendation. Lecture Notes in Computer Science, 1198:

116–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chavez:1997:RLE

[164] A. Chavez, D. Dreilinger, R. Guttman, and P. Maes. A real-life experiment in creating an agent marketplace. Lecture Notes in Computer Science, 1198:160–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zadeh:1997:RFL

[165] L. A. Zadeh. The roles of fuzzy logic and soft computing in the conception, design and deployment of intelligent systems. Lecture Notes in Computer Science, 1198:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Azvine:1997:ISC

[166] B. Azvine, N. Azarmi, and K. C. Tsui. An introduction to soft computing — A tool for building intelligent systems. Lecture Notes in Computer Science, 1198:191–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldwin:1997:BCF

[167] J. F. Baldwin and T. P. Martin. Basic concepts of a fuzzy logic data browser with applications. Lecture Notes in Computer Science, 1198:211–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davis:1997:ECM

[168] A. Davis, M. Swanson, and M. Parker. Efficient communication mechanisms for cluster based parallel computing.

Lecture Notes in Computer Science, 1199:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damianakis:1997:SSS

[169] S. N. Damianakis, C. Dubnicki, and E. W. Felten. Stream sockets on SHRIMP. Lecture Notes in Computer Science, 1199:16–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiNucci:1997:SEP

[170] D. C. DiNucci. A simple and efficient process and communication abstraction for network operating systems. Lecture Notes in Computer Science, 1199:31–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abali:1997:DAM

[171] B. Abali. A deadlock avoidance method for computer networks. Lecture Notes in Computer Science, 1199: 61–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katevenis:1997:ASC

[172] M. G. H. Katevenis, P. Vatsolaki, D. Serpanos, and E. Markatos. ATLAS-I: A single-chip ATM switch for NOWs. Lecture Notes in Computer Science, 1199:88–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimitrov:1997:APT

[173] B. Dimitrov and V. Rego. Arachne: A portable threads library supporting migrant threads on heterogeneous network farms. Lecture Notes in Computer Science, 1199:102–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1997:TTR

[174] S. Kim, M. Hyun, and S. Lee. Transparent treatment of remote pointers using IPC primitive in RPC systems. Lecture Notes in Computer Science, 1199: 115–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonney:1997:DHS

[175] J. Bonney, R. Ramanujan, A. Ahamad, and S. Takkella. Distributed hardware support for process synchronization in NSM workstation clusters. Lecture Notes in Computer Science, 1199: 144–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaddoura:1997:LBR

[176] M. Kaddoura. Load balancing for regular data-parallel applications on workstation network. Lecture Notes in Computer Science, 1199:173–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bal:1997:CTH

[177] H. Bal, R. Hofman, and K. Verstoep. A comparison of three high speed networks for parallel cluster computing. Lecture Notes in Computer Science, 1199:184–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomez:1997:CCL

[178] J. C. Gomez, V. Rego, and V. S. Sunderam. CLAM: Connection-less, lightweight, and multiway communication support for distributed computing. Lecture Notes in Computer Science, 1199:227–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamashita:1997:NWC

[179] H. Yamashita, T. Suguri, S. Kinoshita, and Y. Okada. Network-wide Cooperative Computing Architecture (NCCA). Lecture Notes in Computer Science, 1199:241–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chrisochoides:1997:DMC

[180] N. Chrisochoides, I. Kodukula, and K. Pingali. Data movement and control substrate for parallel scientific computing. Lecture Notes in Computer Science, 1199:256–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodal:1997:PQD

[181] G. S. Brodal. Predecessor queries in dynamic integer sets. Lecture Notes in Computer Science, 1200:21–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franciosa:1997:SDS

[182] P. G. Franciosa, D. Frigioni, and R. Giaccio. Semi-dynamic shortest paths and breadth-first search in diagraphs. Lecture Notes in Computer Science, 1200:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1997:TRE

[183] J. Hromkovic, S. Seibert, and T. Wilke. Translating regular expressions into small curly epsilon-free nondeterministic finite automata. Lecture Notes in Computer Science, 1200:55–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiorio:1997:MMU

[184] C. Fiorio and J. Gustedt. Memory management for union-find algorithms. Lecture Notes in Computer Science, 1200:67–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hempel:1997:OMM

[185] H. Hempel and G. Wechsung. The operators min and max on the polynomial hierarchy. Lecture Notes in Computer Science, 1200:93–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buhrman:1997:RBK

[186] H. Buhrman and L. Fortnow. Resourcebounded Kolmogorov complexity revisited. Lecture Notes in Computer Science, 1200:105–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duris:1997:VVD

[187] P. Duris, J. Hromkovic, J. D. P. Rolim, and W. Schnitger. Las Vegas versus determinism for one-way communication complexity, finite automata, and polynomial-time computations. Lecture Notes in Computer Science, 1200:117–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertram-Kretzberg:1997:MTA

[188] C. Bertram-Kretzberg and H. Lefmann. MODp-tests, almost independence and small probability spaces. Lecture Notes in Computer Science, 1200:141–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeAlfaro:1997:HDD

[189] L. De Alfaro, A. Kapur, and Z. Manna. Hybrid diagrams: A deductivealgorithmic approach to hybrid system verification. Lecture Notes in Computer Science, 1200:153–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeAlfaro:1997:TLS

[190] L. De Alfaro. Temporal logics for the specification of performance and reliability. Lecture Notes in Computer Science, 1200:165–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weise:1997:ESI

[191] C. Weise and D. Lenzkes. Efficient scaling-invariant checking of timed bisimulation. Lecture Notes in Computer Science, 1200:177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietzfelbinger:1997:GBV

[192] M. Dietzfelbinger. Gossiping and broadcasting versus computing functions in networks. Lecture Notes in Computer Science, 1200:189–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waack:1997:DAP

[193] S. Waack. On the descriptive and algorithmic power of parity ordered binary decision diagrams. Lecture Notes in Computer Science, 1200:201–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Case:1997:CCL

[194] J. Case, E. Kinber, A. Sharma, and F. Stephan. On the classification of computable languages. Lecture Notes in Computer Science, 1200:225–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kern-Isberner:1997:CLA

[195] G. Kern-Isberner. A conditionallogical approach to minimum crossentropy. Lecture Notes in Computer Science, 1200:237–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graedel:1997:URT

[196] E. Graedel, M. Otto, and E. Rosen. Undecidability results on two-variable logics. Lecture Notes in Computer Science, 1200:249–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaubert:1997:MAM

[197] S. Gaubert and M. Plus. Methods and applications of (max,+ ) linear algebra. Lecture Notes in Computer Science, 1200:261–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldstine:1997:MNP

[198] J. Goldstine, H. Leung, and D. Wotschke. Measuring nondeterminism in pushdown automata. Lecture Notes in Computer Science, 1200:295–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemaspaandra:1997:DTP

[199] E. Hemaspaandra, L. A. Hemaspaandra, and H. Hempel. A downward translation in the polynomial hierarchy. Lecture Notes in Computer Science, 1200:319–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flaremini:1997:DFI

[200] M. Flaremini. Deadlock-free interval routing schemes. Lecture Notes in Computer Science, 1200:351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karg:1997:CGT

[201] C. Karg, J. Koebler, and R. Schuler. The complexity of generating test instances. Lecture Notes in Computer Science, 1200:375–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrew:1997:ECH

[202] A. E. Andrew, A. E. F. Clementi, and J. D. P. Rolim. Efficient construction of hitting sets for systems of linear functions. Lecture Notes in Computer Science, 1200:387–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biehl:1997:PCS

[203] I. Biehl and B. Meyer. Protocols for collusion-secure asymmetric fingerprinting. Lecture Notes in Computer Science, 1200:399–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattaneo:1997:ELC

[204] G. Cattaneo, E. Formenti, G. Manzini, and L. Margara. On ergodic linear cellular automata over Zm. Lecture Notes in Computer Science, 1200:427– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durand-Lose:1997:IUD

[205] J. O. Durand-Lose. Intrinsic universality of a 1-dimensional reversible cellular automaton. Lecture Notes in Computer Science, 1200:439–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buss:1997:CCS

[206] J. F. Buss, G. S. Frandsen, and J. O. Shallit. The computational complexity of some problems of linear algebra. Lecture Notes in Computer Science, 1200:451–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahlhaus:1997:SPA

[207] E. Dahlhaus. Sequential and parallel algorithms on compactly represented chordal and strongly chordal graphs. Lecture Notes in Computer Science, 1200:487–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldmann:1997:BUB

[208] R. Feldmann, B. Monien, P. Mysliwietz, and S. Tschoeke. A better upper bound on the bisection width of de Bruijn networks. Lecture Notes in Computer Science, 1200:511–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feigenbaum:1997:ITT

[209] J. Feigenbaum and M. Strauss. An information-theoretic treatment of random-self-reducibility. Lecture Notes in Computer Science, 1200:523–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breutzmann:1997:EMC

[210] J. M. Breutzmann and J. H. Lutz. Equivalence of measures of complexity classes. Lecture Notes in Computer Science, 1200:535–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auletta:1997:BAM

[211] V. Auletta and M. Parente. Better algorithms for minimum weight vertexconnectivity problems. Lecture Notes in Computer Science, 1200:547–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diekert:1997:RCE

[212] V. Diekert, P. Gastin, and A. Petit. Removing curly epsilon-transitions in timed automata. Lecture Notes in Computer Science, 1200:583–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldreich:1997:PPS

[213] O. Goldreich. Probabilistic proof systems — A survey. Lecture Notes in Computer Science, 1200:595–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artstein:1997:TLM

[214] Z. Artstein. Tutorial: The Lyapunov method. Lecture Notes in Computer Science, 1201:2–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1997:QQ

[215] T. A. Henzinger and O. Kupferman. From quantity to quality. Lecture Notes in Computer Science, 1201:48– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boniol:1997:ASP

[216] F. Boniol, A. Burgueño, O. Roux, and V. Rusu. Analysis of slope-parametric hybrid automata. Lecture Notes in Computer Science, 1201:75–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huuck:1997:CTC

[217] R. Huuck, Y. Lakhnech, L. Urbina, and S. Engell. Comparing timed C/E systems with timed automata. Lecture Notes in Computer Science, 1201:81– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hajji:1997:DTH

[218] M. S. Hajji, J. M. Bass, A. R. Browne, and P. J. Fleming. Design tools for hybrid control systems. Lecture Notes in Computer Science, 1201:87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branicky:1997:LRC

[219] M. S. Branicky. On-line, reflexive constraint satisfaction for hybrid systems: First steps. Lecture Notes in Computer Science, 1201:93–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerone:1997:MTD

[220] A. Cerone, A. J. Cowie, G. J. Milne, and P. A. Moseley. Modeling a timedependent protocol using the Circal process algebra. Lecture Notes in Computer Science, 1201:124–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dolginova:1997:SVA

[221] E. Dolginova and N. Lynch. Safety verification for automated platoon maneuvers. Lecture Notes in Computer Science, 1201:154–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Archer:1997:VHS

[222] M. Archer and C. Heitmeyer. Verifying hybrid systems modeled as timed automata. Lecture Notes in Computer Science, 1201:171–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flaus:1997:HFN

[223] J.-M. Flaus and G. Ollagnon. Hybrid flow nets for hybrid processes modeling and control. Lecture Notes in Computer Science, 1201:213–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caspi:1997:WCW

[224] P. Caspi. What can we learn from synchronous data-flow languages? Lecture Notes in Computer Science, 1201: 255–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bornot:1997:RTP

[225] S. Bornot and J. Sifakis. Relating time progress and deadlines in hybrid systems. Lecture Notes in Computer Science, 1201:286–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1997:SVE

[226] Q. Xu. Semantics and verification of extended phase transition systems in duration calculus. Lecture Notes in Computer Science, 1201:301–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weise:1997:WRM

[227] C. Weise and D. Lenzkes. Weak refinement for modal hybrid systems. Lecture Notes in Computer Science, 1201: 316–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1997:RTA

[228] V. Gupta, T. A. Henzinger, and R. Jagadeesan. Robust timed automata. Lecture Notes in Computer Science, 1201:331–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asarin:1997:DSV

[229] E. Asarin, M. Bozga, A. Kerbrat, and O. Maler. Data-structures for the verification of timed automata. Lecture Notes in Computer Science, 1201:346– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1997:SCH

[230] D. Kapur and R. K. Shyamasundar. Synthesizing controllers for hybrid systems. Lecture Notes in Computer Science, 1201:361–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heymann:1997:CSC

[231] M. Heymann, F. Lin, and G. Meyer. Control synthesis for a class of hybrid systems subject to configuration-based safety constraints. Lecture Notes in Computer Science, 1201:376–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSchutter:1997:GLC

[232] B. De Schutter and B. De Moor. Generalized linear complementarity problems and the analysis of continuously variable systems and discrete event systems. Lecture Notes in Computer Science, 1201:409–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wooldridge:1997:IAB

[233] M. Wooldridge. Issues in agent-based software engineering. Lecture Notes in Computer Science, 1202:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergamaschi:1997:EIH

[234] S. Bergamaschi. Extraction of informations from highly heterogeneous source of textual data. Lecture Notes in Computer Science, 1202:42–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkhard:1997:CIA

[235] H.-D. Burkhard. Cases, information, and agents. Lecture Notes in Computer Science, 1202:64–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerschberg:1997:KRC

[236] L. Kerschberg. Knowledge rovers: Cooperative intelligent agent support for enterprise information architectures. Lecture Notes in Computer Science, 1202:79–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vossen:1997:CSC

[237] G. Vossen. The CORBA specification for cooperation in heterogeneous information systems. Lecture Notes in Computer Science, 1202:101–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berndtsson:1997:RSA

[238] M. Berndtsson, S. Chakravarthy, and B. Lings. Result sharing among agents using reactive rules. Lecture Notes in Computer Science, 1202:126–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babin:1997:MMD

[239] G. Babin, Z. Maamar, and B. Chaib-Draa. Metadatabase meets distributed AI. Lecture Notes in Computer Science, 1202:138–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haneke:1997:LBF

[240] E. Haneke. ‘learning’ based filtering of text information using simple interest profiles. Lecture Notes in Computer Science, 1202:171–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhong:1997:OSA

[241] N. Zhong, Y. Kakemoto, and S. Ohsuga. An organized society of autonomous knowledge discovery agents. Lecture Notes in Computer Science, 1202:183– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cao:1997:AEC

[242] W. Cao, C.-G. Bian, and G. Hartvigsen. Achieving efficient cooperation in a multi-agent system: The twin-base modelling. Lecture Notes in Computer Science, 1202:210–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jamison:1997:AIH

[243] W. C. Jamison. Approaching interoperability for heterogeneous multiagent systems using high order agencies. Lecture Notes in Computer Science, 1202:222–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bordini:1997:AIO

[244] R. H. Bordini, J. A. Campbell, and R. Vieira. Ascription of intensional ontologies in anthropological descriptions of multi-agent systems. Lecture Notes in Computer Science, 1202:235– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1997:MLS

[245] G. Wagner. Multi-level security in multiagent systems. Lecture Notes in Computer Science, 1202:272–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bilardi:1997:ADS

[246] G. Bilardi. Algorithms and data structures for control dependence and related compiler problems. Lecture Notes in Computer Science, 1203:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Even:1997:EIN

[247] G. Even and S. Even. Embedding interconnection networks in grids via the layered cross product. Lecture Notes in Computer Science, 1203:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dinitz:1997:FOV

[248] Y. Dinitz and Z. Nutov. Finding optimum k-vertex connected spanning subgraphs: Improved approximation algorithms for k = 3, 4, 5. Lecture Notes in Computer Science, 1203:13–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1997:OCC

[249] K. Jansen. The optimum cost chromatic partition problem. Lecture Notes in Computer Science, 1203:25– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khuller:1997:FTC

[250] S. Khuller, R. Pless, and Y. J. Sussmann. Fault tolerant K-center problems. Lecture Notes in Computer Science, 1203:37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemaspaandra:1997:RTN

[251] E. Hemaspaandra, L. A. Hemaspaandra, and H. Hempel. R0S0N1-tt(NP) distinguishes robust many-one and Turing completeness. Lecture Notes in Computer Science, 1203:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caporaso:1997:SCL

[252] S. Caporaso, M. Zito, N. Galesi, and E. Covino. Syntactic characterization in LISP of the polynomial complexity classes and hierarchy. Lecture Notes in Computer Science, 1203:61–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feige:1997:DSS

[253] U. Feige and G. Rayzman. On the drift of short schedules. Lecture Notes in Computer Science, 1203:74–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomez:1997:RND

[254] F. Gomez, S. Ramaswami, and G. Toussaint. On removing non-degeneracy assumptions in computational geometry. Lecture Notes in Computer Science, 1203:86–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAmore:1997:MMU

[255] F. D’Amore, P. G. Franciosa, R. Giaccio, and M. Talamo. Maintaining maxima under boundary updates. Lecture Notes in Computer Science, 1203:100– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Datta:1997:OAO

[256] A. Datta, K. Krithivasan, and T. Ottmann. An optimal algorithm for one-separation of a set of isothetic polygons. Lecture Notes in Computer Science, 1203:110–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1997:NDP

[257] U. Foessmeier and M. Kaufmann. Nice drawings for planar bipartite graphs. Lecture Notes in Computer Science, 1203:122–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amir:1997:IDA

[258] A. Amir and E. Dar. An improved deterministic algorithm for generalized random sampling. Lecture Notes in Computer Science, 1203:159–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domingo:1997:PTA

[259] C. Domingo. Polynomial time algorithms for some self-duality problems. Lecture Notes in Computer Science, 1203:171–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferragina:1997:NUS

[260] P. Ferragina, R. Grossi, and M. Montangero. A note on updating suffix tree labels. Lecture Notes in Computer Science, 1203:181–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanke:1997:RBR

[261] S. Hanke, T. Ottmann, and E. Soisalon-Soininen. Relaxed balanced red-black trees. Lecture Notes in Computer Science, 1203:193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damaschke:1997:ACC

[262] P. Damaschke. The algorithmic complexity of chemical threshold testing. Lecture Notes in Computer Science, 1203:205–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katajainen:1997:MAM

[263] J. Katajainen and J. L. Traeff. A meticulous analysis of mergesort programs. Lecture Notes in Computer Science, 1203:217–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattaneo:1997:TCE

[264] G. Cattaneo, M. Finelli, and L. Margara. Topological chaos for elementary cellular automata. Lecture Notes in Computer Science, 1203:241–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernasconi:1997:CBB

[265] A. Bernasconi. On the complexity of balanced Boolean functions. Lecture Notes in Computer Science, 1203:253– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemaspaandra:1997:SEC

[266] L. A. Hemaspaandra, J. Rothe, and G. Wechsung. On sets with easy certificates and the existence of one-way permutations. Lecture Notes in Computer Science, 1203:264–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamazaki:1997:IGB

[267] K. Yamazaki, H. L. Bodlaender, B. De Fluiter, and D. M. Thilikos. Isomorphism for graphs of bounded distance width. Lecture Notes in Computer Science, 1203:276–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alimonti:1997:HAP

[268] P. Alimonti and V. Kann. Hardness of approximating problems on cubic graphs. Lecture Notes in Computer Science, 1203:288–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kao:1997:TCE

[269] M.-Y Kao. Tree contractions and evolutionary trees. Lecture Notes in Computer Science, 1203:299–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goos:1997:PLP

[270] G. Goos. From programming languages to program construction. Lecture Notes in Computer Science, 1204:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gough:1997:MLM

[271] K. J. Gough. Multi-language, multitarget compiler development: Evolution of the Gardens Point Compiler Project. Lecture Notes in Computer Science, 1204:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gough:1997:EAS

[272] K. J. Gough and H. Klaeren. Executable assertions and separate compilation. Lecture Notes in Computer Science, 1204:41–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collin:1997:TIL

[273] S. Collin, D. Colnet, and O. Zendra. Type inference for late binding: The SmallEiffel compiler. Lecture Notes in Computer Science, 1204:67–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weck:1997:ADT

[274] W. Weck. An abstract data type for freezable lists and DAGs. Lecture Notes in Computer Science, 1204:112– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hof:1997:JTS

[275] M. Hof. Just-in-time stub generation. Lecture Notes in Computer Science, 1204:197–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gutknecht:1997:DFR

[276] J. Gutknecht. Do the fish really need remote control? A proposal for selfactive objects in Oberon. Lecture Notes in Computer Science, 1204:207– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Disteli:1997:COA

[277] A. R. Disteli and P. Reali. Combining Oberon with active objects. Lecture Notes in Computer Science, 1204: 221–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evered:1997:HWD

[278] M. Evered, J. L. Keedy, A. Schmolitzky, and G. Menger. How well do inheritance mechanisms support inheritance concepts? Lecture Notes in Computer Science, 1204:252–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suzuki:1997:AVB

[279] Sho’ji Suzuki, Tatsunori Kato, Hiroshi Ishizuka, Yasutake Takahashi, Eiji Uchibe, and Minoru Asada. An application of vision-based learning in RoboCup for a real robot with an omnidirectional vision system and the team description of Osaka University “Trackies”. Lecture Notes in Computer Science, 1204:316–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1604/16040316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1604/16040316.
pdf.

Eisenecker:1997:GPG

[280] U. W. Eisenecker. Generative programming (GP) with C++. Lecture Notes in Computer Science, 1204:351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeller:1997:SIO

[281] E. Zeller. Seamless integration of online services in the Oberon document system. Lecture Notes in Computer Science, 1204:366–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bro-Nielsen:1997:NRI

[282] M. Bro-Nielsen, C. Gramkow, and S. Kreiborg. Non-rigid image registration using bone growth model. Lecture Notes in Computer Science, 1205:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Declerck:1997:DCP

[283] J. Declerck, J. Feldmar, and N. Ayache. Definition of a 4D continuous polar transformation for the tracking and the analysis of LV motion. Lecture Notes in Computer Science, 1205:33–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1997:SUD

[284] T. N. Jones and D. N. Metaxas. Segmentation using deformable models with affinity-based localization. Lecture Notes in Computer Science, 1205: 53–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gee:1997:AEM

[285] J. C. Gee, D. R. Haynor, L. Le Briquer, and R. K. Bajcsy. Advances in elastic matching theory and its implementation. Lecture Notes in Computer Science, 1205:63–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacq:1997:ADA

[286] J. J. Jacq and C. Roux. Automatic detection of articular surfaces in 3D image through minimal subset random sampling. Lecture Notes in Computer Science, 1205:73–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delingette:1997:DIS

[287] H. Delingette. Decimation of isosurfaces with deformable models. Lecture Notes in Computer Science, 1205: 83–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bite:1997:CEO

[288] U. Bite, P. Yugueros, S. TerKonda, and D. D’Souza. Comparative estimation of orbital three dimensional CT measurements utilizing ANALYZE[TM]. Lecture Notes in Computer Science, 1205: 93–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamadeh:1997:KSL

[289] A. Hamadeh and P. Cinquin. Kinematic study of lumbar spine using functional radiographies and 3D/2D registration. Lecture Notes in Computer Science, 1205:109–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weese:1997:ARV

[290] J. Weese, T. M. Buzug, C. Lorenz, and C. Fassnacht. An approach to 2D/3D registration of a vertebra in 2D X-ray fluoroscopies with 3D CT images. Lecture Notes in Computer Science, 1205:119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldmar:1997:MMA

[291] Y. Feldmar, G. Malandain, N. Ayache, and S. Fernandez-Vidal. Matching 3D MR angiography data and 2D X-ray angiograms. Lecture Notes in Computer Science, 1205:129–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellis:1997:MEC

[292] R. E. Ellis, D. J. Fleet, J. T. Bryant, and J. Rudan. A method for evaluating CT-based surgical registration. Lecture Notes in Computer Science, 1205: 141–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

West:1997:RIR

[293] J. West, J. M. Fitzpatrick, M. Y. Wang, and B. M. Dawant. Retrospective intermodality registration techniques: surface-based versus volumebased. Lecture Notes in Computer Science, 1205:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bro-Nielsen:1997:RRC

[294] M. Bro-Nielsen. Rigid registration of CT, MR and cryosection images using a GLCM framework. Lecture Notes in Computer Science, 1205:171–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhalerao:1997:UPM

[295] A. Bhalerao, C.-F. Westin, and R. Kikinis. Unwrapping phase in 3D MR phase contrast angiograms. Lecture Notes in Computer Science, 1205:193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buzug:1997:IRC

[296] T. M. Buzug, J. Weese, C. Fassnacht, and C. Lorenz. Image registration: convex weighting functions for histogram-based similarity measures. Lecture Notes in Computer Science, 1205:203–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juhan:1997:GMA

[297] V. Juhan, B. Nazarian, K. Malkani, and R. Bulot. Geometrical modeling of abdominal aortic aneurysms. Lecture Notes in Computer Science, 1205: 243–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berlage:1997:ARD

[298] T. Berlage. Augmented reality for diagnosis based on ultrasound images. Lecture Notes in Computer Science, 1205: 253–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Champleboux:1997:PTD

[299] G. Champleboux and D. Henry. Parametric two-dimensional B-spline representation of vein and artery from 2.5D echography used to aid virtual echography. Lecture Notes in Computer Science, 1205:263–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chavanon:1997:CAP

[300] O. Chavanon, C. Barbe, J. Troccaz, and L. Carrat. Computer-ASsisted PERicardial punctures: animal feasability study. Lecture Notes in Computer Science, 1205:285–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bzostek:1997:ASP

[301] A. Bzostek, S. Schreiner, A. C. Barnes, and J. A. Cadeddu. An automated system for precise percutaneous access of the renal collecting system. Lecture Notes in Computer Science, 1205:299– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auer:1997:VEP

[302] L. M. Auer, D. Auer, and J. F. Knoplioch. Virtual endoscopy for planning and simulation of minimally invasive neurosurgery. Lecture Notes in Computer Science, 1205:315–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorodia:1997:RAM

[303] T. M. Gorodia, R. H. Taylor, and L. M. Auer. Robot-assisted minimally invasive neurosurgical procedures: first experimental experience. Lecture Notes in Computer Science, 1205:319–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darabi:1997:RSE

[304] K. Darabi, K. D. M. Resch, J. Weinert, and U. Jendrysiak. Real and simulated endoscopy of neurosurgical approaches in an anatomical model. Lecture Notes in Computer Science, 1205: 323–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hauser:1997:NIP

[305] R. Hauser, B. Westermann, and R. Probst. Non-invasive 3D patient registration for image-guided intranasal surgery — experimental and clinical results. Lecture Notes in Computer Science, 1205:327–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellemare:1997:TAT

[306] M. E. Bellemare, P. Haigron, and J. L. Coatrieux. Toward an active three dimensional navigation system in medical imaging. Lecture Notes in Computer Science, 1205:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jambon:1997:LCT

[307] A.-C. Jambon, P. Dubois, and S. Karpf. A low-cost training simulator for initial formation in gynecologic laparoscopy. Lecture Notes in Computer Science, 1205:347–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wei:1997:ATL

[308] G.-Q. Wei, K. Arbter, and G. Hirzinger. Automatic tracking of laparoscopic instruments by color coding. Lecture Notes in Computer Science, 1205:357– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gibson:1997:SAK

[309] S. Gibson, J. Samosky, A. Mor, and C. Fyock. Simulating arthroscopic knee surgery using volumetric object representations, real-time volume rendering and haptic feedback. Lecture Notes in Computer Science, 1205:369–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaye:1997:VEM

[310] J. Kaye, D. N. Metaxas, and F. P. Primiano. A 3D virtual environment for modeling mechanical cardiopulmonary interactions. Lecture Notes in Computer Science, 1205:389–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berger:1997:DCC

[311] J. W. Berger, M. E. Leventon, N. Hata, and W. Wells. Design considerations for a computer-vision-enabled ophthalmic augmented reality environment. Lecture Notes in Computer Science, 1205:399–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bricault:1997:FMB

[312] I. Bricault. A fast morphologybased registration — application to computer-assisted bronchoscopy. Lecture Notes in Computer Science, 1205: 417–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hill:1997:EIB

[313] D. L. G. Hill, C. R. Maurer, M. Y. Wang, and R. J. Maciunas. Estimation of intraoperative brain surface movement. Lecture Notes in Computer Science, 1205:449–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bucholz:1997:CSI

[314] R. D. Bucholz, D. D. Yeh, J. Trobaugh, and L. L. McDurmont. The correction of stereotactic inaccuracy caused by brain shift using an intraoperative ultrasound device. Lecture Notes in Computer Science, 1205:459–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ettinger:1997:ETM

[315] G. J. Ettinger, M. E. Leventon, W. E. L. Grimson, and R. Kikinis. Experimentation with a transcranial magnetic stimulation system for functional brain mapping. Lecture Notes in Computer Science, 1205:477–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouchard:1997:PRS

[316] V. Bouchard, P. Cinquin, L. Desbat, and A. François Joubert. A priori registration of SPECT projections and MR/CT images for SPECT reconstruction improvement. Lecture Notes in Computer Science, 1205:491–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colchester:1997:CPT

[317] A. C. F. Colchester, J. Zhao, N. Dorward, and G. Subsol. Comparison of phantom target localization by frame-based stereotaxy and using the VISLAN system. Lecture Notes in Computer Science, 1205:501–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alusi:1997:TDT

[318] G. Alusi, A. C. Tan, A. D. Linney, and K. Raoof. Three dimensional tracking with ultrasound for augmented reality applications in skull base surgery. Lecture Notes in Computer Science, 1205: 511–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glossop:1997:ETA

[319] N. Glossop and R. Hu. Effects of tracking adjacent vertebral bodies during image guided pedicle screw surgery. Lecture Notes in Computer Science, 1205:531–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauer:1997:RAE

[320] A. Bauer, M. Boerner, and A. Lahmer. Robodoc — animal experiment and clinical evaluation. Lecture Notes in Computer Science, 1205:561–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahler:1997:ECI

[321] D. M. Kahler and R. Zura. Evaluation of a computer integrated surgical technique for percutaneous fixation of transverse acetabular fractures. Lecture Notes in Computer Science, 1205:565–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaramaz:1997:RMA

[322] B. Jaramaz, C. Nikou, D. A. Simon, and A. M. DiGioia. Range of motion after total hip arthroplasty: experimental verification of the analytical simulator. Lecture Notes in Computer Science, 1205:573–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1997:AIG

[323] R. Hu, N. Glossop, D. Steven, and J. Randle. Accuracy of image guided placement of iliosacral lag screws. Lecture Notes in Computer Science, 1205: 593–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gong:1997:RSM

[324] J. Gong, R. Baechler, M. Sati, and L.- P. Nolte. Restricted surface matching: a new approach to registration in computer assisted surgery. Lecture Notes in Computer Science, 1205:597–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fadda:1997:CAP

[325] M. Fadda, D. Bertelli, S. Martelli, and M. Marcacci. Computer assisted planning for total knee arthroplasty. Lecture Notes in Computer Science, 1205:619–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banks:1997:SFB

[326] S. A. Banks, A. Z. Banks, T. V. S. Klos, and F. F. Cook. A simple fluoroscopy based technique for assessing 3D knee kinematics before, during, and after surgery. Lecture Notes in Computer Science, 1205:639–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bondiau:1997:ERC

[327] P.-Y. Bondiau and G. Malandain. Eye reconstruction and CT-retinography fusion for proton treatment planning of ocular diseases. Lecture Notes in Computer Science, 1205:705–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erbse:1997:DAS

[328] S. Erbse, K. Radermacher, M. Anton, and G. Rau. Development of an automatic surgical holding system based on ergonomic analysis. Lecture Notes in Computer Science, 1205:737–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harris:1997:ERS

[329] S. J. Harris, W. J. Lin, K. L. Fan, and R. D. Hibberd. Experiences with robotic systems for knee surgery. Lecture Notes in Computer Science, 1205:

757–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandt:1997:CRI

[330] G. Brandt, K. Radermacher, S. Lavallee, and H.-W. Staudte. A compact robot for image-guided orthopedic surgery: concept and preliminary results. Lecture Notes in Computer Science, 1205: 767–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carrozza:1997:DMS

[331] M. C. Carrozza, L. Lencioni, B. Magnani, and S. D’Attanasio. The development of a microrobot system for colonoscopy. Lecture Notes in Computer Science, 1205:779–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freysinger:1997:ITA

[332] W. Freysinger, M. J. Truppe, A. R. Gunkel, and W. L. Thumfart. Interactive telepresence and augmented reality in ENT surgery: interventional video tomography. Lecture Notes in Computer Science, 1205:817–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benn:1997:REC

[333] D. E. Benn, M. S. Nixon, and J. N. Carter. Robust eye centre extraction using the Hough transform. Lecture Notes in Computer Science, 1206:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1997:LFF

[334] K. N. Choi, A. D. J. Cross, and E. R. Hancock. Localising facial features with matched filters. Lecture Notes in Computer Science, 1206:11–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1997:SNF

[335] S. Fischer and B. Duc. Shape normalisation for face recognition. Lecture Notes in Computer Science, 1206: 21–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kervrann:1997:GLR

[336] C. Kervrann, F. Davoine, P. Perez, and H. Li. Generalized likelihood ratiobased face detection and extraction of mouth features. Lecture Notes in Computer Science, 1206:27–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coianiz:1997:AEL

[337] T. Coianiz and L. Torresani. Analysis and encoding of lip movements. Lecture Notes in Computer Science, 1206: 53–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wechsler:1997:AVB

[338] H. Wechsler, V. Kakkad, J. Huang, and S. Gutta. Automatic video-based person authentication using the RBF network. Lecture Notes in Computer Science, 1206:85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cunado:1997:UGB

[339] D. Cunado, M. S. Nixon, and J. N. Carter. Using gait as a biometric, via phase-weighted magnitude spectra. Lecture Notes in Computer Science, 1206:95–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hong:1997:IAU

[340] L. Hong, A. Jain, S. Pankanti, and R. Bolle. Identity authentication using fingerprints. Lecture Notes in Computer Science, 1206:103–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Etemad:1997:DAR

[341] K. Etemad and R. Chellappa. Discriminant analysis for recognition of human face images. Lecture Notes in Computer Science, 1206:127–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beumier:1997:API

[342] C. Beumier and M. Acheroy. Automatic profile identification. Lecture Notes in Computer Science, 1206:145– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dai:1997:RFI

[343] Y. Dai and Y. Nakano. Recognition of facial images with low resolution using a Hopfield memory model. Lecture Notes in Computer Science, 1206:153– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gebhart:1997:EPN

[344] A. Gebhart and T. Eppler. Exclusion of photos and new segmentation algorithms for the automatic face recognition. Lecture Notes in Computer Science, 1206:161–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besacier:1997:SAA

[345] L. Besacier and J.-F. Bonastre. Subband approach for automatic speaker recognition: Optimal division of the frequency domain. Lecture Notes in Computer Science, 1206:195–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charlet:1997:OFS

[346] D. Charlet and D. Jouvet. Optimizing feature set for speaker verification. Lecture Notes in Computer Science, 1206:203–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finan:1997:VSN

[347] R. A. Finan, A. T. Sapeluk, and R. I. Damper. VQ score normalisation for text-dependent and text-independent speaker recognition. Lecture Notes in Computer Science, 1206:211–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furui:1997:RAS

[348] S. Furui. Recent advances in speaker recognition. Lecture Notes in Computer Science, 1206:237–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayakawa:1997:SIU

[349] S. Hayakawa, K. Takeda, and F. Itakura. Speaker identification using harmonic structure of LP-residual spectrum. Lecture Notes in Computer Science, 1206:253–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Julia:1997:SIA

[350] L. E. Julia, L. P. Heck, and A. J. Cheyer. A speaker identification agent. Lecture Notes in Computer Science, 1206:261–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biguen:1997:ECM

[351] E. S. Biguen, J. Biguen, B. Duc, and S. Fischer. Expert conciliation for multi modal person authentication systems by Bayesian statistics. Lecture Notes in Computer Science, 1206:291–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dieckmann:1997:SBP

[352] U. Dieckmann, P. Plankensteiner, R. Schamburger, and B. Froeba. SESAM: A biometric person identification system using sensor fusion. Lecture Notes in Computer Science, 1206: 301–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duc:1997:PAF

[353] B. Duc, G. Maitre, S. Fischer, and J. Bigun. Person authentication by fusing face and speech information. Lecture Notes in Computer Science, 1206: 311–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jourlin:1997:ALS

[354] P. Jourlin, J. Luettin, D. Genoud, and H. Wassner. Acoustic-labial speaker verification. Lecture Notes in Computer Science, 1206:319–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bargale:1997:DAV

[355] C. B. Bargale, S. Chaudhuri, and P. Bhattacharyya. Development of an audio-visual database system for human identification. Lecture Notes in Computer Science, 1206:345–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davoine:1997:VCP

[356] F. Davoine, H. Li, and R. Forchheimer. Video compression and person authentication. Lecture Notes in Computer Science, 1206:353–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doi:1997:LCS

[357] M. Doi, Q. Chen, K. Sato, and K. Chihara. Lock-control system using face identification. Lecture Notes in Computer Science, 1206:361–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferralasco:1997:SAF

[358] D. Ferralasco, R. Ferrari, and M. Tistarelli. A system for automatic face recognition. Lecture Notes in Computer Science, 1206:369–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

George:1997:TES

[359] M. H. George and R. A. King. Time encoded signal processing and recognition for reduced data, high performance speaker verification architectures. Lecture Notes in Computer Science, 1206:377–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

James:1997:CSV

[360] D. James, H.-P. Hutter, and F. Bimbot. The CAVE speaker verification project — experiments on the YOHO and SESP corpora. Lecture Notes in Computer Science, 1206:385–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayes:1997:RSL

[361] I. J. Hayes, R. G. Nickson, and P. A. Strooper. Refining specifications to logic programs. Lecture Notes in Computer Science, 1207:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fribourg:1997:SVG

[362] L. Fribourg and J. Richardson. Symbolic verification with gap-order constraints. Lecture Notes in Computer Science, 1207:20–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cortesi:1997:SBA

[363] A. Cortesi, B. Le Charlier, and S. Rossi. Specification-based automatic verification of Prolog programs. Lecture Notes in Computer Science, 1207: 38–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joergensen:1997:CPD

[364] J. Joergensen, M. Leuschel, and B. Martens. Conjunctive partial deduction in practice. Lecture Notes in Computer Science, 1207:59–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bossi:1997:RCP

[365] A. Bossi and N. Cocco. Replacement can preserve termination. Lecture Notes in Computer Science, 1207: 104–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brunekreef:1997:TTP

[366] J. Brunekreef. A transformation tool for pure Prolog programs. Lecture Notes in Computer Science, 1207:130– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fribourg:1997:RPN

[367] L. Fribourg and H. Olsen. Reductions of Petri nets and unfolding of propositional logic programs. Lecture Notes in Computer Science, 1207:187–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benoy:1997:IAS

[368] F. Benoy and A. King. Inferring argument size relationships with CLP(R). Lecture Notes in Computer Science, 1207:204–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gegg-Harrison:1997:ELP

[369] T. S. Gegg-Harrison. Extensible logic program schemata. Lecture Notes in Computer Science, 1207:256–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1997:SML

[370] A. Brogi and S. Contiero. Specialising meta-level compositions of logic programs. Lecture Notes in Computer Science, 1207:275–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dung:1997:SPP

[371] P. M. Dung, R. A. Kowalski, and F. Toni. Synthesis of proof procedures for default reasoning. Lecture Notes in Computer Science, 1207:313–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Warmuth:1997:SCL

[372] M. Warmuth. Sample compression, learnability, and the Vapnik– Chervonenkis dimension. Lecture Notes in Computer Science, 1208:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beimel:1997:LBH

[373] A. Beimel and E. Kushilevitz. Learning boxes in high dimension. Lecture Notes in Computer Science, 1208:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guijarro:1997:LMT

[374] D. Guijarro, V. Lavin, and V. Raghavan. Learning monotone term decision lists. Lecture Notes in Computer Science, 1208:16–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bshouty:1997:LMF

[375] N. H. Bshouty, C. Tamon, and D. K. Wilson. Learning matrix functions over rings. Lecture Notes in Computer Science, 1208:27–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freund:1997:LUP

[376] Y. Freund and Y. Mansour. Learning under persistent drift. Lecture Notes in Computer Science, 1208:109–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cesa-Bianchi:1997:RHM

[377] N. Cesa-Bianchi, P. Fischer, E. Shamir, and H. U. Simon. Randomized hypotheses and minimum disagreement hypotheses for learning with noise. Lecture Notes in Computer Science, 1208: 119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helmbold:1997:LWT

[378] D. Helmbold, S. Kwek, and L. Pitt. Learning when to trust which experts. Lecture Notes in Computer Science, 1208:134–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergadano:1997:LBP

[379] F. Bergadano, N. H. Bshouty, C. Tamon, and S. Varricchio. On learning branching programs and small depth circuits. Lecture Notes in Computer Science, 1208:150–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castro:1997:LNM

[380] J. Castro, D. Guijarro, and V. Lavin. Learning nearly monotone k-term DNF. Lecture Notes in Computer Science, 1208:162–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birkendorf:1997:FSA

[381] A. Birkendorf. On fast and simple algorithms for finding maximal subarrays and applications in learning theory. Lecture Notes in Computer Science, 1208:198–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bartlett:1997:MLB

[382] P. Bartlett, T. Linder, and G. Lugosi. A minimax lower bound for empirical quantizer design. Lecture Notes in Computer Science, 1208:210–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gurvits:1997:LAP

[383] L. Gurvits. Linear algebraic proofs of VC-dimension based inequalities. Lecture Notes in Computer Science, 1208: 238–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baxter:1997:RRC

[384] J. Baxter and P. Bartlett. A result relating convex n-widths to covering numbers with some applications to neural networks. Lecture Notes in Computer Science, 1208:251–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barzdins:1997:LFE

[385] J. Barzdins, R. Freivalds, and C. H. Smith. Learning formulae from elementary facts. Lecture Notes in Computer Science, 1208:272–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Case:1997:CSH

[386] J. Case, S. Jain, and M. Suraj. Control structures in hypothesis spaces: the influence on learning. Lecture Notes in Computer Science, 1208:286–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambainis:1997:OMC

[387] A. Ambainis, S. Jain, and A. Sharma. Ordinal mind change complexity of language identification. Lecture Notes in Computer Science, 1208:301–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufmann:1997:RLI

[388] S. Kaufmann and F. Stephan. Robust learning with infinite additional information. Lecture Notes in Computer Science, 1208:316–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cranefield:1997:ABA

[389] S. Cranefield and M. Purvis. An agentbased architecture for software tool coordination. Lecture Notes in Computer Science, 1209:44–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bell:1997:DGH

[390] J. Bell and Z. Huang. Dynamic goal hierarchies. Lecture Notes in Computer Science, 1209:88–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cavedon:1997:SIC

[391] L. Cavedon, A. Rao, and G. Tidhar. Social and individual commitment. Lecture Notes in Computer Science, 1209:152–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clark:1997:EPT

[392] M. Clark, K. Irwig, and W. Wobcke. Emergent properties of teams of agents in the tileworld. Lecture Notes in Computer Science, 1209:164–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1997:HDA

[393] A. Ito. How do autonomous agents solve social dilemmas? Lecture Notes in Computer Science, 1209:177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akama:1997:LCT

[394] Y. Akama. A lambda-to-CL translation for strong normalization. Lecture Notes in Computer Science, 1210:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belle:1997:TIL

[395] G. Belle and E. Moggi. Typed intermediate languages for shape analysis. Lecture Notes in Computer Science, 1210:11–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berardi:1997:MIC

[396] S. Berardi and L. Boerio. Minimum information code in a pure functional language with data types. Lecture Notes in Computer Science, 1210:30–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bono:1997:MCL

[397] V. Bono and M. Bugliesi. Matching constraints for the lambda calculus of objects. Lecture Notes in Computer Science, 1210:46–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandt:1997:CAR

[398] M. Brandt and F. Henglein. Coinductive axiomatization of recursive type equality and subtyping. Lecture Notes in Computer Science, 1210:63– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brauener:1997:SAC

[399] T. Brauener. A simple adequate categorical model for PCF. Lecture Notes in Computer Science, 1210:82– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bucciarelli:1997:LRB

[400] A. Bucciarelli. Logical reconstruction of bi-domains. Lecture Notes in Computer Science, 1210:99–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courant:1997:MCP

[401] J. Courant. A module calculus for pure type systems. Lecture Notes in Computer Science, 1210:112–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damiani:1997:IAS

[402] F. Damiani and P. Giannini. An inference algorithm for strictness. Lecture Notes in Computer Science, 1210: 129–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Despeyroux:1997:PRH

[403] J. Despeyroux, F. Pfenning, and C. Schuermann. Primitive recursion for higher-order abstract syntax. Lecture Notes in Computer Science, 1210:147– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghani:1997:EED

[404] N. Ghani. Eta-expansions in dependent type theory — the calculus of constructions. Lecture Notes in Computer Science, 1210:164–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerrini:1997:PNG

[405] S. Guerrini, S. Martini, and A. Masini. Proof nets, garbage, and computations. Lecture Notes in Computer Science, 1210:181–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasegawa:1997:RCS

[406] M. Hasegawa. Recursion from cyclic sharing: traced monoidal categories and models of cyclic lambda calculi. Lecture Notes in Computer Science, 1210:196–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herbelin:1997:GWH

[407] H. Herbelin. Games and weak-head reduction for classical PCF. Lecture Notes in Computer Science, 1210:214– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xi:1997:WSB

[408] H. Xi. Weak and strong beta normalisations in typed lambda-calculi. Lecture Notes in Computer Science, 1210: 390–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1997:PMP

[409] J. Fox, N. Johns, and A. Rahmanzadeh. Protocols for medical procedures and therapies: A provisional description of the PROforma language and tools. Lecture Notes in Computer Science, 1211:21–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCarolis:1997:UAM

[410] B. De Carolis, G. Rumolo, and V. Cavallo. User-adapted multimedia explanations in a clinical guidelines consultation system. Lecture Notes in Computer Science, 1211:62–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gordon:1997:ACP

[411] C. Gordon, P. Johnson, C. Waite, and M. Veloso. Algorithm and care pathway: Clinical guidelines and healthcare processes. Lecture Notes in Computer Science, 1211:66–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zupan:1997:AVB

[412] B. Zupan and S. Dzeroski. Acquiring and validating background knowledge for machine learning using function decomposition. Lecture Notes in Computer Science, 1211:86–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzeroski:1997:ARE

[413] S. Dzeroski, G. Potamias, V. Moustakis, and G. Charissis. Automated revision of expert rules for treating acute abdominal pain in children. Lecture Notes in Computer Science, 1211:98– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DuPlessis:1997:KAD

[414] J. P. Du Plessis and C. J. Tolmie. Knowledge acquisition by the domain expert using the tool HEMATOOL. Lecture Notes in Computer Science, 1211:122–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zelic:1997:MLA

[415] I. Zelic, I. Kononenko, N. Lavrac, and V. Vuga. Machine learning applied to diagnosis of sport injuries. Lecture Notes in Computer Science, 1211:138– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cremilleux:1997:TFD

[416] B. Cremilleux and C. Robert. A theoretical framework for decision trees in uncertain domains: Application to medical data sets. Lecture Notes in Computer Science, 1211:145–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bologna:1997:IAC

[417] G. Bologna, A. Rida, and C. Pellegrini. Intelligent assistance for coronary heart disease diagnosis: A comparison study. Lecture Notes in Computer Science, 1211:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geissbuehler:1997:CSD

[418] A. Geissbuehler, R. A. Miller, and W. W. Stead. The clinical spectrum of decision-support in oncology with a case report of a real world system. Lecture Notes in Computer Science, 1211:235–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaulent:1997:CBR

[419] M.-C. Jaulent, C. Le Bozec, E. Zapletal, and P. Degoulet. A case-based reasoning method for computer-assisted diagnosis in histopathology. Lecture Notes in Computer Science, 1211:239– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cruz:1997:CFM

[420] J. Cruz and P. Barahona. A causalfunctional model applied to EMG diagnosis. Lecture Notes in Computer Science, 1211:249–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zahlmann:1997:NFC

[421] G. Zahlmann, M. Scherf, and A. Wegner. A neuro-fuzzy-classifier for a knowledge-based glaucoma monitor. Lecture Notes in Computer Science, 1211:273–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hauskrecht:1997:DDM

[422] M. Hauskrecht. Dynamic decision making in stochastic partially observable medical domains: Ischemic heart disease example. Lecture Notes in Computer Science, 1211:296–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ziebelin:1997:MPE

[423] D. Ziebelin and A. Vila. Metrical planning environment. Lecture Notes in Computer Science, 1211:343–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeRosis:1997:SAM

[424] F. De Rosis, F. Grasso, and D. C. Berry. Strengthening argumentation in medical explanations by text plan revision. Lecture Notes in Computer Science, 1211:349–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1997:MCS

[425] J. C. Wagner, C. Lovis, R. H. Baud, and J.-R. Scherrer. Medical concept systems, lexicons and natural language generation. Lecture Notes in Computer Science, 1211:398–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bianchi:1997:DPC

[426] N. Bianchi, P. Bottoni, C. Garbay, and P. Mussio. Distributed plan construction and execution for medical image interpretation. Lecture Notes in Computer Science, 1211:405–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Draye:1997:IIH

[427] J.-P. Draye, G. Cheron, D. Pavisic, and G. Libert. Improved identification of the human shoulder kinematics with muscle biological filters. Lecture Notes in Computer Science, 1211: 417–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boucher:1997:SGO

[428] A. Boucher, A. Doisy, X. Ronot, and C. Garbay. A society of goaloriented agents for the analysis of living cells. Lecture Notes in Computer Science, 1211:429–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gibaud:1997:MDD

[429] B. Gibaud, S. Garlatti, C. Barillot, and E. Faure. Methodology for the design of digital brain atlases. Lecture Notes in Computer Science, 1211:441–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cosic:1997:RBL

[430] D. Cosic and S. Loncaric. Rule-based labeling of CT head image. Lecture Notes in Computer Science, 1211:453– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derr:1997:CTT

[431] T. Derr, T. Els, M. Gyngell, and D. Leibfritz. Characterisation of tumorous tissue in rat brain by in vitro magnetic resonance spectroscopy and artificial neural networks. Lecture Notes in Computer Science, 1211:457– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christodoulou:1997:MLL

[432] E. Christodoulou and E. T. Keravnou. Meta-level learning in a hybrid knowledge-based architecture. Lecture Notes in Computer Science, 1211:467– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellazzi:1997:LDT

[433] R. Bellazzi, L. Ironi, R. Guglielmann, and M. Stefanelli. Learning from data through the integration of qualitative models and fuzzy systems. Lecture Notes in Computer Science, 1211:501– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bull:1997:CBR

[434] M. Bull, G. Kundt, and L. Gierl. Case-based reasoning and statistics for discovering and forecasting of epidemics. Lecture Notes in Computer Science, 1211:513–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heitmeyer:1997:FMP

[435] C. L. Heitmeyer. Formal methods: A panacea or academic poppycock? Lecture Notes in Computer Science, 1212: 3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hewitt:1997:EPA

[436] M. A. Hewitt, C. M. O’Halloran, and C. T. Sennett. Experiences with PiZA, an animator for Z. Lecture Notes in Computer Science, 1212:37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helke:1997:ATC

[437] S. Helke, T. Neustupny, and T. Santen. Automating test case generation from Z specifications with Isabelle. Lecture Notes in Computer Science, 1212:52–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1997:TZS

[438] J. A. Hall. Taking Z seriously. Lecture Notes in Computer Science, 1212: 89–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Achatz:1997:FOM

[439] K. Achatz and W. Schulte. A formal OO method inspired by fusion and object-Z. Lecture Notes in Computer Science, 1212:92–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1997:WR

[440] J. Hall and A. Martin. W reconstructed. Lecture Notes in Computer Science, 1212:115–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boerger:1997:PMR

[441] E. Boerger and S. Mazzanti. A practical method for rigorously controllable hardware design. Lecture Notes in Computer Science, 1212:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butler:1997:ADD

[442] M. Butler. An approach to the design of distributed systems with B AMN. Lecture Notes in Computer Science, 1212:223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evans:1997:IRS

[443] A. S. Evans. An improved recipe for specifying reactive systems in Z. Lecture Notes in Computer Science, 1212: 275–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DInverno:1997:ZSS

[444] M. D’Inverno and M. Hu. A Z specification of the soft-link hypertext model. Lecture Notes in Computer Science, 1212:297–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacky:1997:EZD

[445] J. Jacky, J. Unger, M. Patrick, and D. Reid. Experience with Z developing a control program for a radiation therapy machine. Lecture Notes in Computer Science, 1212:317–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciancarini:1997:ARA

[446] P. Ciancarini and C. Mascolo. Analyzing and refining an architectural style. Lecture Notes in Computer Science, 1212:349–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derrick:1997:WRZ

[447] J. Derrick, E. Boiten, H. Bowman, and M. Steen. Weak refinement in Z. Lecture Notes in Computer Science, 1212:369–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bowen:1997:SZB

[448] J. P. Bowen. Select Z bibliography. Lecture Notes in Computer Science, 1212:391–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bowen:1997:CSZ

[449] J. P. Bowen. Comp.specification.z and Z FORUM frequently asked questions. Lecture Notes in Computer Science, 1212:425–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davis:1997:CFO

[450] M. W. Davis. Complexity formalisms, order and disorder in the structure of art. Lecture Notes in Computer Science, 1213:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imada:1997:SRV

[451] A. Imada and K. Araki. Searching real-valued synaptic weights of Hopfield’s associative memory using evolutionary programming. Lecture Notes in Computer Science, 1213:13–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fogel:1997:AEC

[452] G. B. Fogel. The application of evolutionary computation to selected problems in molecular biology. Lecture Notes in Computer Science, 1213:23– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clack:1997:PEG

[453] C. Clack and T. Yu. Performanceenhanced genetic programming. Lecture Notes in Computer Science, 1213: 87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Angeline:1997:CSC

[454] P. J. Angeline. Comparing subtree crossover with macromutation. Lecture Notes in Computer Science, 1213:101– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:MSG

[455] S.-H. Chen and C.-H Yeh. Modeling speculators with genetic programming. Lecture Notes in Computer Science, 1213:137–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1997:FES

[456] X. Yao and Y. Liu. Fast evolution strategies. Lecture Notes in Computer Science, 1213:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delahaye:1997:ACS

[457] D. Delahaye and A. R. Odoni. Airspace congestion smoothing by stochastic optimization. Lecture Notes in Computer Science, 1213:163–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dozier:1997:SSD

[458] G. Dozier, J. Bowen, A. Homaifar, and A. Esterline. Solving static and dynamic fuzzy constraint networks using evolutionary hill-climbing. Lecture Notes in Computer Science, 1213:189– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:AFC

[459] J.-M. Yang, Y.-P. Chen, J.-T. Horng, and C.-Y. Kao. Applying family competition to evolution strategies for constrained optimization. Lecture Notes in Computer Science, 1213:201–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hadad:1997:SPG

[460] B. S. Hadad and C. F. Eick. Supporting polyploidy in genetic algorithms using dominance vectors. Lecture Notes in Computer Science, 1213:223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beyer:1997:NEP

[461] H.-G. Beyer and D. B. Fogel. A note on the escape probabilities for two alternative methods of selection under Gaussian mutation. Lecture Notes in Computer Science, 1213:265–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

English:1997:SGA

[462] T. M. English. Some geometric and algebraic results on crossover. Lecture Notes in Computer Science, 1213:285– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1997:AEA

[463] X. Yao, G. Lin, and Y. Liu. An analysis of evolutionary algorithms based on neighborhood and step sizes. Lecture Notes in Computer Science, 1213: 297–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Angeline:1997:TED

[464] P. J. Angeline. Tracking extrema in dynamic environments. Lecture Notes in Computer Science, 1213:335–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chellapilla:1997:ESA

[465] K. Chellapilla and D. B. Fogel. Exploring self-adaptive methods to improve the efficiency of generating approximate solutions to traveling salesman problems using evolutionary programming. Lecture Notes in Computer Science, 1213:361–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chellapilla:1997:GIE

[466] K. Chellapilla, D. B. Fogel, and S. S. Rao. Gaining insight into evolutionary programming through landscape visualization: An investigation into IIR filtering. Lecture Notes in Computer Science, 1213:407–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boehm:1997:TCS

[467] C. Boehm. Theoretical computer science and software science: The past, the present and the future. Lecture Notes in Computer Science, 1214:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1997:FTT

[468] H. Ehrig and B. Mahr. Future trends of TAPSOFT. Lecture Notes in Computer Science, 1214:6–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouhoula:1997:SPM

[469] A. Bouhoula, J.-P. Jouannaud, and J. Meseguer. Specification and proof in membership equational logic. Lecture Notes in Computer Science, 1214: 67–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Astesiano:1997:FM

[470] E. Astesiano and G. Reggio. Formalism and method. Lecture Notes in Computer Science, 1214:93–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamada:1997:LCR

[471] T. Yamada, J. Avenhaus, C. Loria-Saenz, and A. Middeldorp. Logicality of conditional rewrite systems. Lecture Notes in Computer Science, 1214:141– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biermann:1997:RGC

[472] I. Biermann and B. Rozoy. Reliable generalized and context dependent commutation relations. Lecture Notes in Computer Science, 1214:165– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andre:1997:WTT

[473] Y. Andre and F. Bossut. Word-into-trees transducers with bounded difference. Lecture Notes in Computer Science, 1214:177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emerson:1997:GQT

[474] E. A. Emerson and R. J. Trefler. Generalized quantitative temporal reasoning: An automata-theoretic approach. Lecture Notes in Computer Science, 1214:189–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beauquier:1997:RCP

[475] D. Beauquier and A. Slissenko. The railroad crossing problem: Towards semantics of timed algorithms and their model-checking in high-level languages. Lecture Notes in Computer Science, 1214:201–212, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ilie:1997:MCT

[476] J. M. Ilie and K. Ajami. Model checking through symbolic reachability graph. Lecture Notes in Computer Science, 1214:213–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goubault:1997:OIW

[477] E. Goubault. Optimal implementation of wait-free binary relations. Lecture Notes in Computer Science, 1214:225– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geser:1997:RUTa

[478] A. Geser, A. Middeldorp, E. Ohlebush, and H. Zantema. Relative undecidability in the termination hierarchy of single rewrite rules. Lecture Notes in Computer Science, 1214:237–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Genet:1997:TPU

[479] T. Genet and I. Gnaedig. Termination proofs using gpo ordering constraints. Lecture Notes in Computer Science, 1214:249–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arts:1997:APT

[480] T. Arts and J. Giesl. Automatically proving termination where simplification orderings fail. Lecture Notes in Computer Science, 1214:261–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldamus:1997:MCW

[481] M. Baldamus and J. Dingel. Modal characterization of weak bisimulation for higher-order processes. Lecture Notes in Computer Science, 1214:285– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amaniss:1997:SOR

[482] A. Amaniss, M. Hermann, and D. Lugiez. Set operations for recurrent term schematizations. Lecture Notes in Computer Science, 1214:333–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arenas-Sanchez:1997:SFF

[483] P. Arenas-Sanchez and M. Rodriguez-Artalejo. A semantic framework for functional logic programming with algebraic polymorphic types. Lecture Notes in Computer Science, 1214:453– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bono:1997:SCI

[484] V. Bono, M. Bugliesi, M. Dezani-Ciancaglini, and L. Liquori. Subtyping constraints for incomplete objects. Lecture Notes in Computer Science, 1214:465–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alessi:1997:PSS

[485] F. Alessi, P. Baldan, and F. Honsell. Partializing Stone spaces using SFP domains. Lecture Notes in Computer Science, 1214:478–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiadeiro:1997:SAC

[486] J. L. Fiadeiro and A. Lopes. Semantics of architectural connectors. Lecture Notes in Computer Science, 1214: 505–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1997:CMV

[487] H. R. Andersen, J. Staunstrup, and N. Maretti. A comparison of modular verification techniques. Lecture Notes in Computer Science, 1214:550– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Volpano:1997:TBA

[488] D. Volpano and G. Smith. A typebased approach to program security. Lecture Notes in Computer Science, 1214:607–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courant:1997:AMC

[489] J. Courant. An applicative module calculus. Lecture Notes in Computer Science, 1214:622–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iyer:1997:PLC

[490] P. Iyer and M. Narasimha. Probabilistic lossy channel systems. Lecture Notes in Computer Science, 1214:667– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:1997:LOO

[491] M. Abadi and K. R. M. Leino. A logic of object-oriented programs. Lecture Notes in Computer Science, 1214:682– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNicola:1997:LBL

[492] R. De Nicola, G. Ferrari, and R. Pugliese. Locality based Linda: Programming with explicit localities. Lecture Notes in Computer Science, 1214: 712–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wells:1997:TIL

[493] J. B. Wells, A. Dimock, R. Muller, and F. Turbak. A typed intermediate language for flow-directed compilation. Lecture Notes in Computer Science, 1214:757–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:1997:BRC

[494] B. Jacobs. Behaviour-refinement of coalgebraic specifications with coinductive correctness proofs. Lecture Notes in Computer Science, 1214:787–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonnier:1997:CCA

[495] S. Bonnier and T. Heyer. COMPASS: A comprehensible assertion method. Lecture Notes in Computer Science, 1214:803–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heisel:1997:ULP

[496] M. Heisel and N. Levy. Using LO-TOS patterns to characterize architectural styles. Lecture Notes in Computer Science, 1214:818–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donat:1997:AFS

[497] M. R. Donat. Automating formal specification-based testing. Lecture Notes in Computer Science, 1214:833– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonHenke:1997:TEM

[498] F. W. Von Henke, M. Luther, and M. Strecker. Typelab: An environment for modular program development. Lecture Notes in Computer Science, 1214:851–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Didrich:1997:DDT

[499] K. Didrich and T. Klein. DOS-FOP — A documentation tool for the algebraic programming language Opal. Lecture Notes in Computer Science, 1214:875–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caron:1997:ASM

[500] P. Caron. AG: A set of Maple packages for symbolic computing of automata and semigroups. Lecture Notes in Computer Science, 1214:879–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Falcao:1997:HPC

[501] D. M. Falcao. High performance computing in power system applications. Lecture Notes in Computer Science, 1215:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1997:PAH

[502] J. Dongarra, S. Browne, and H. Casanova. Providing access to high performance computing technologies. Lecture Notes in Computer Science, 1215:24–34, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coltuc:1997:SDA

[503] D. Coltuc and J.-M. Becker. Scandirectional architectures. Lecture Notes in Computer Science, 1215:35– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dekeyser:1997:IDP

[504] J.-L. Dekeyser, B. Kokoszko, J.-L. Levaire, and P. Marquet. Irregular dataparallel objects in C++. Lecture Notes in Computer Science, 1215:65– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alves:1997:PVP

[505] J. C. Alves, A. Puga, L. Corte-Real, and J. S. Matos. ProHos-1: A vector processor for the efficient estimation of higher-order moments. Lecture Notes in Computer Science, 1215:96– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dayde:1997:UCK

[506] M. J. Dayde and I. S. Duff. The use of computational kernels in full and sparse linear solvers, efficient code design on high-performance RISC processors. Lecture Notes in Computer Science, 1215:108–139, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domas:1997:PIS

[507] S. Domas and F. Tisseur. Parallel implementation of a symmetric eigensolver based on the Yau and Lu method. Lecture Notes in Computer Science, 1215:140–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aliaga:1997:PIG

[508] J. I. Aliaga, V. Hernandez, and J. L. Perez. A parallel implementation of the general Lanczos method on the Cray T3D. Lecture Notes in Computer Science, 1215:168–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coelho:1997:PDO

[509] P. J. Coelho, J. Goncalves, and P. Novo. Parallelization of the discreteordinates method: Two different approaches. Lecture Notes in Computer Science, 1215:222–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carvalho:1997:PCC

[510] L. M. R. Carvalho and J. M. L. M. Palma. Parallelization of a CFD code using PVM and domain decomposition techniques. Lecture Notes in Computer Science, 1215:247–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiasDAlmeida:1997:PNL

[511] F. Dias D’Almeida and P. B. Vasconcelos. Preconditioners for nonsymmetric linear systems in domain decomposition applied to a coupled discretization of Navier–Stokes equations. Lecture Notes in Computer Science, 1215:295–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calonge:1997:PIN

[512] T. Calonge, L. Alonso, R. Ralha, and A. L. Sanchez. Parallel implementation of non-recurrent neural networks. Lecture Notes in Computer Science, 1215:313–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hernandez:1997:SLC

[513] V. Hernandez and E. S. Quintana-Orti. Stabilizing large control linear systems on multicomputers. Lecture Notes in Computer Science, 1215:338–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1997:IBT

[514] J. C. Fernandez, L. Peñalver, M. Sobejano, and J. Martinez. An interface based on transputers to simulate the dynamic equation of robot manipulators using parallel computing. Lecture Notes in Computer Science, 1215:365–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heras:1997:PCA

[515] D. B. Heras, J. C. Cabaleiro, V. Blanco, and P. Costas. Principal component analysis on vector computers. Lecture Notes in Computer Science, 1215:416– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dix:1997:PLP

[516] J. Dix, L. M. Peteira, and T. Przymusinski. Prolegomena to logic programming for non-monotonic reasoning. Lecture Notes in Computer Science, 1216:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bochman:1997:LSG

[517] A. Bochman. On logics and semantics for general logic programs. Lecture Notes in Computer Science, 1216: 37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:NRM

[518] X. Wang, J.-H. You, and L.-Y. Yuan. Nonmonotonic reasoning by monotonic inferences + priority constraints. Lecture Notes in Computer Science, 1216:

91–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alferes:1997:UPC

[519] J. J. Alferes and L. M. Pereira. Updateprograms can update programs. Lecture Notes in Computer Science, 1216: 110–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldoni:1997:APR

[520] M. Baldoni, L. Giordano, A. Martelli, and V. Patti. An abductive procedure for reasoning about actions in modal LP. Lecture Notes in Computer Science, 1216:132–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1997:URP

[521] N. Bidoit and S. Maabout. Update rule programs versus revision programs. Lecture Notes in Computer Science, 1216:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brass:1997:TBB

[522] S. Brass, S. Zukowski, and H. Freitag. Transformation-based bottomup computation of the well-founded model. Lecture Notes in Computer Science, 1216:171–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dix:1997:CNG

[523] J. Dix and F. Stolzenburg. Computation of non-ground disjunctive well-founded semantics with constraint logic programming. Lecture Notes in Computer Science, 1216:202–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berry:1997:HSS

[524] G. Berry. Hardware and software synthesis, optimization, and verification from Esterel programs. Lecture Notes in Computer Science, 1217:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drechsler:1997:MAK

[525] R. Drechsler, B. Becker, and S. Ruppertz. Manipulation algorithms for K0*BMDs. Lecture Notes in Computer Science, 1217:4–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emerson:1997:CPO

[526] E. A. Emerson, S. Jha, and D. Peled. Combining partial order and symmetry reductions. Lecture Notes in Computer Science, 1217:19–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1997:PMC

[527] H. R. Andersen, J. Staunstrup, and N. Maretti. Partial model checking with ROBDDs. Lecture Notes in Computer Science, 1217:35–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grahlmann:1997:RCP

[528] B. Grahlmann. The reference component of PEP. Lecture Notes in Computer Science, 1217:65–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulton:1997:TSF

[529] R. J. Boulton. A tool to support formal reasoning about computer languages. Lecture Notes in Computer Science, 1217:81–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonderBeeck:1997:GMU

[530] M. Von der Beeck, V. Braun, A. Classen, and A. Dannecker. Graphs in METAFrame: The unifying power of polymorphism. Lecture Notes in Computer Science, 1217:112–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelb:1997:MFT

[531] P. Kelb, T. Margaria, M. Mendler, and C. Gsottberger. MOSEL: A flexible toolset for monadic second-order logic. Lecture Notes in Computer Science, 1217:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jensen:1997:BIC

[532] K. Jensen. A brief introduction to coloured Petri nets. Lecture Notes in Computer Science, 1217:203–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:1997:DCC

[533] S. Christensen, J. B. Joergensen, and L. M. Kristensen. Design/CPN — A computer tool for coloured Petri nets. Lecture Notes in Computer Science, 1217:209–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-Abdallah:1997:SDP

[534] H. Ben-Abdallah and S. Leue. Syntactic detection of process divergence and non-local choice in message sequence charts. Lecture Notes in Computer Science, 1217:259–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1997:ABV

[535] G. Ferrari, G. Ferro, S. Gnesi, and U. Montanari. An automata based verification environment for mobile processes. Lecture Notes in Computer Science, 1217:275–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boigelot:1997:ASS

[536] B. Boigelot and P. Godefroid. Automatic synthesis of specifications from the dynamic observation of reactive programs. Lecture Notes in Computer Science, 1217:321–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeAlfaro:1997:VVR

[537] L. De Alfaro, Z. Manna, H. B. Sipma, and T. E. Uribe. Visual verification of reactive systems. Lecture Notes in Computer Science, 1217:334–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engels:1997:TGI

[538] A. Engels, L. Feijs, and S. Mauw. Test generation for intelligent networks using model checking. Lecture Notes in Computer Science, 1217:384–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DArgenio:1997:BRP

[539] P. R. D’Argenio, J.-P. Katoen, T. C. Ruys, and J. Tretmans. The bounded retransmission protocol must be on time! Lecture Notes in Computer Science, 1217:416–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bordihn:1997:GAL

[540] H. Bordihn. A grammatical approach to the LBA problem. Lecture Notes in Computer Science, 1218:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernau:1997:CCF

[541] H. Fernau and M. Holzer. Conditional context-free languages of finite index. Lecture Notes in Computer Science, 1218:10–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asveld:1997:CFP

[542] P. R. J. Asveld. Controlled fuzzy parallel rewriting. Lecture Notes in Computer Science, 1218:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bottoni:1997:CRP

[543] P. Bottoni, G. Mauri, and P. Mussio. On controlling rewriting by properties of strings and symbols. Lecture Notes in Computer Science, 1218:71– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernau:1997:AAG

[544] H. Fernau and R. Freund. Accepting array grammars with control mechanisms. Lecture Notes in Computer Science, 1218:95–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jancar:1997:RAR

[545] P. Jancar, F. Mraz, M. Platek, and J. Vogel. On restarting automata with rewriting. Lecture Notes in Computer Science, 1218:119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balanescu:1997:GSC

[546] T. Balanescu, H. Georgescu, and M. Gheorghe. Grammar systems with counting derivation and dynamical priorities. Lecture Notes in Computer Science, 1218:150–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dumitrescu:1997:CRU

[547] S. Dumitrescu. Characterization of RE using CD grammar systems with two registers and RL rules. Lecture Notes in Computer Science, 1218:167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waetjen:1997:CDU

[548] O. Waetjen. On cooperating distributed uniformly limited L systems. Lecture Notes in Computer Science, 1218:178–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amin:1997:NIE

[549] M. Amin. A note on the incomparability of the EOL family with certain families of languages generated by cooperating grammar systems. Lecture Notes in Computer Science, 1218:217–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelemen:1997:CMR

[550] J. Kelemen. Colonies as models of reactive systems. Lecture Notes in Computer Science, 1218:220–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abrahamson:1997:GCL

[551] K. Abrahamson, L. Cai, and S. Gordon. A grammar characterization of logarithmic-space computation. Lecture Notes in Computer Science, 1218: 247–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruda:1997:CCC

[552] S. Bruda. On the computational complexity of context-free parallel communicating grammar systems. Lecture Notes in Computer Science, 1218:256– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chitu:1997:PGS

[553] A. Chitu. PC grammar systems versus some non-context-free constructions from natural and artificial languages. Lecture Notes in Computer Science, 1218:278–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jimenez-Lopez:1997:GSD

[554] M. D. Jimenez-Lopez and C. Martin-Vide. Grammar systems for the description of certain natural language facts. Lecture Notes in Computer Science, 1218:288–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Csuhaj-Varju:1997:NPL

[555] E. Csuhaj-Varju and A. Salomaa. Networks of parallel language processors. Lecture Notes in Computer Science, 1218:299–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zandron:1997:RDS

[556] C. Zandron, C. Ferretti, and G. Mauri. A reduced distributed splicing system for RE languages. Lecture Notes in Computer Science, 1218:319–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgescu:1997:GCS

[557] G. Georgescu. On the generative capacity of splicing grammar systems. Lecture Notes in Computer Science, 1218: 330–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1997:TLB

[558] J. Hromkovic and J. Karhumaeki. Two lower bounds on computational complexity of infinite words. Lecture Notes in Computer Science, 1218:366–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cazanescu:1997:SBM

[559] V. E. Cazanescu. The structure of the basic morphisms. Lecture Notes in Computer Science, 1218:422–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzer:1997:CII

[560] M. Holzer and K.-J Lange. On the complexity of iterated insertions. Lecture Notes in Computer Science, 1218: 440–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ilie:1997:DGC

[561] L. Ilie. The decidability of the generalized confluence problem for contextfree languages. Lecture Notes in Computer Science, 1218:454–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldi:1997:ECM

[562] M. Baldi, S. Gai, and G. Picco. Exploiting code mobility in decentralized and flexible network management. Lecture Notes in Computer Science, 1219: 13–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghezzi:1997:MCP

[563] C. Ghezzi and G. Vigna. Mobile code paradigms and technologies. Lecture Notes in Computer Science, 1219:39– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1997:CIC

[564] D. Wong, N. Paciorek, T. Walsh, and J. DiCelie. Concordia: An infrastructure for collaborating mobile agents. Lecture Notes in Computer Science, 1219:86–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1997:OMA

[565] D. Chang and S. Covaci. The OMG mobile agent facility: A submission. Lecture Notes in Computer Science, 1219:98–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hurst:1997:MAS

[566] L. Hurst, P. Cunningham, and F. Somers. Mobile agents — smart messages. Lecture Notes in Computer Science, 1219:111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumann:1997:CCM

[567] J. Baumann, F. Hohl, N. Radouniklis, and K. Rothermel. Communication concepts for mobile agent systems. Lecture Notes in Computer Science, 1219: 123–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doemel:1997:MAI

[568] P. Doemel, A. Lingnau, and O. Drobnik. Mobile agent interaction in heterogeneous environments. Lecture Notes in Computer Science, 1219:136–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chia:1997:SMA

[569] T. Chia and S. Kannapan. Strategically mobile agents. Lecture Notes in Computer Science, 1219:149–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:NMA

[570] W. E. Chen and C. R. Leng. A novel mobile agent search algorithm. Lecture Notes in Computer Science, 1219:162– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1997:TAI

[571] M. Williams and A. Bendiab. A toolset for architecture independent, reconfigurable, multi-agent systems. Lecture Notes in Computer Science, 1219:210– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brezany:1997:IOI

[572] Peter Brezany. Input/output intensive massively parallel computing: Language support, automatic parallelization, advanced optimization, and runtime systems. Lecture Notes in Computer Science, 1220:vi + 288, 1997. CODEN LNCSD9. ISBN 3-540-62840-1. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davidsson:1997:LLA

[573] P. Davidsson. Learning by linear anticipation in multi-agent systems. Lecture Notes in Computer Science, 1221: 62–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ye:1997:COS

[574] Y. Ye and J. K. Tsotsos. On the collaborative object search team: A formulation. Lecture Notes in Computer Science, 1221:94–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

CetertickBazzan:1997:ECM

[575] A. L. Cetertick Bazzan. Evolution of coordination as a metaphor for learning in multi-agent systems. Lecture Notes in Computer Science, 1221:117– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dragoni:1997:LAR

[576] A. F. Dragoni and P. Giorgini. Learning agents’ reliability through Bayesian conditioning: A simulation experiment. Lecture Notes in Computer Science, 1221:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davies:1997:CII

[577] W. Davies and P. Edwards. The communication of inductive inferences. Lecture Notes in Computer Science, 1221: 223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Friedrich:1997:LCM

[578] H. Friedrich, M. Kaiser, O. Rogalla, and R. Dillmann. Learning and communication in multi-agent systems. Lecture Notes in Computer Science, 1221:259–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardelli:1997:MC

[579] L. Cardelli. Mobile computation. Lecture Notes in Computer Science, 1222: 3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chess:1997:MAT

[580] D. Chess, C. Harrison, and A. Kershenbaum. Mobile agents: Are they a good idea? Lecture Notes in Computer Science, 1222:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waldo:1997:NDC

[581] J. Waldo, G. Wyant, A. Wollrath, and S. Kendall. A note on distributed computing. Lecture Notes in Computer Science, 1222:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cugola:1997:AMC

[582] G. Cugola, C. Ghezzi, G. P. Picco, and G. Vigna. Analyzing mobile code languages. Lecture Notes in Computer Science, 1222:93–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Acharya:1997:SLR

[583] A. Acharya, M. Ranganathan, and J. Saltz. Sumatra: A language for resource-aware mobile programs. Lecture Notes in Computer Science, 1222: 111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bharat:1997:MA

[584] K. Bharat and L. Cardelli. Migratory applications. Lecture Notes in Computer Science, 1222:131–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kato:1997:SSE

[585] K. Kato. Safe and secure execution mechanisms for mobile objects. Lecture Notes in Computer Science, 1222: 201–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciancarini:1997:JCC

[586] P. Ciancarini and D. Rossi. Jada — coordination and communication for Java agents. Lecture Notes in Computer Science, 1222:213–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Acharya:1997:DLM

[587] A. Acharya and J. Saltz. Dynamic linking for mobile programs. Lecture Notes in Computer Science, 1222:245– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franz:1997:ACS

[588] M. Franz. Adaptive compression of syntax trees and iterative dynamic code optimization: Two basic technologies for mobile object systems. Lecture Notes in Computer Science, 1222:263– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duggan:1997:TBI

[589] D. Duggan. A type-based implementation of a language with distributed scope. Lecture Notes in Computer Science, 1222:277–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doemel:1997:IJT

[590] P. Doemel. Interaction of Java and Telescript agents. Lecture Notes in Computer Science, 1222:295–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Figueiredo:1997:APD

[591] M. A. T. Figueiredo, J. M. N. Leitao, and A. K. Jain. Adaptive parametrically deformable contours. Lecture Notes in Computer Science, 1223:35– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:ISE

[592] J.-P. Wang. Image segmentation via energy minimization on partitions with connected components. Lecture Notes in Computer Science, 1223:101–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Descombes:1997:MLE

[593] X. Descombes, R. Morris, J. Zerubia, and M. Berthod. Maximum likelihood estimation of Markov random field parameters using Markov chain Monte Carlo algorithms. Lecture Notes in Computer Science, 1223:133–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barker:1997:UIS

[594] S. A. Barker and P. J. W. Rayner. Unsupervised image segmentation using Markov random field models. Lecture Notes in Computer Science, 1223: 165–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuille:1997:TQF

[595] A. L. Yuille and J. Coughlan. Twenty questions, focus of attention, and A0*: A theoretical comparison of optimization strategies. Lecture Notes in Computer Science, 1223:197–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1997:DAU

[596] T. Hofmann, J. Puzicha, and J. M. Buhmann. Deterministic annealing for unsupervised texture segmentation. Lecture Notes in Computer Science, 1223:213–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1997:DSS

[597] M. L. Williams, R. C. Wilson, and E. R. Hancock. Deterministic search strategies for relational graph matching. Lecture Notes in Computer Science, 1223:261–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhong:1997:OLU

[598] Y. Zhong and A. K. Jain. Object localization using color, texture and shape. Lecture Notes in Computer Science, 1223:279–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hornegger:1997:OPS

[599] J. Hornegger and H. Niemann. Optimization problems in statistical object recognition. Lecture Notes in Computer Science, 1223:311–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhandarkar:1997:FGS

[600] S. M. Bhandarkar and X. Zeng. Figureground separation: A case study in energy minimization via evolutionary computing. Lecture Notes in Computer Science, 1223:375–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gelgon:1997:RLM

[601] M. Gelgon and P. Bouthemy. A regionlevel motion-based graph representation and labelling for tracking a spatial image partition. Lecture Notes in Computer Science, 1223:409–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finch:1997:EMA

[602] A. M. Finch, R. C. Wilson, and E. R. Hancock. An expectationmaximisation approach to graph matching. Lecture Notes in Computer Science, 1223:425–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Azencott:1997:EMM

[603] R. Azencott and L. Younes. An energy minimization method for matching and comparing structured object representations. Lecture Notes in Computer Science, 1223:441–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fua:1997:CMT

[604] P. Fua. Consistent modelling of terrain and drainage using deformable models. Lecture Notes in Computer Science, 1223:459–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glasbey:1997:SIR

[605] C. A. Glasbey. SAR image registration and segmentation using an estimated DEM. Lecture Notes in Computer Science, 1223:507–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Escolano:1997:DTT

[606] F. Escolano, M. Cazorla, D. Gallardo, and R. Rizo. Deformable templates for tracking and analysis of intravascular ultrasound sequences. Lecture Notes in Computer Science, 1223:521–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branca:1997:MCT

[607] A. Branca, E. Stella, and A. Distante. Motion correspondence through energy minimization. Lecture Notes in Computer Science, 1223:535–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armengol:1997:IFT

[608] E. Armengol and E. Plaza. Induction of feature terms with INDIE. Lecture Notes in Computer Science, 1224:33– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baroglio:1997:EQK

[609] C. Baroglio. Exploiting qualitative knowledge to enhance skill acquisition. Lecture Notes in Computer Science, 1224:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cichosz:1997:ILP

[610] P. Cichosz. Integrated learning and planning based on truncating temporal differences. Lecture Notes in Computer Science, 1224:57–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeRaedt:1997:ATS

[611] L. De Raedt, P. Idestam-Almquist, and G. Sablon. aitch theta-subsumption for structural matching. Lecture Notes in Computer Science, 1224:73–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demiroez:1997:CVF

[612] G. Demiroez and H. Altay Guevenir. Classification by voting feature intervals. Lecture Notes in Computer Science, 1224:85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demsar:1997:CIC

[613] J. Demsar, B. Zupan, M. Bohanec, and I. Bratko. Constructing intermediate concepts by decomposition of real functions. Lecture Notes in Computer Science, 1224:93–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gamberger:1997:COR

[614] D. Gamberger and N. Lavrac. Conditions for Occam’s razor applicability and noise elimination. Lecture Notes in Computer Science, 1224:108–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1997:LDT

[615] Y.-J Hu. Learning different types of new attributes by combining the neural network and iterative attribute construction. Lecture Notes in Computer Science, 1224:124–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hutchinson:1997:MTC

[616] A. Hutchinson. Metrics on terms and clauses. Lecture Notes in Computer Science, 1224:138–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ting:1997:MCM

[617] Kai Ming Ting and Boon Toh Low. Model combination in the multipledata-batches scenario. Lecture Notes in Computer Science, 1224:250–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aha:1997:CBL

[618] D. W. Aha and D. Wettschereck. Casebased learning: Beyond classification of feature vectors. Lecture Notes in Computer Science, 1224:329–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daelemans:1997:ELN

[619] W. Daelemans, A. Van den Bosch, and T. Weijters. Empirical learning of natural language processing tasks. Lecture Notes in Computer Science, 1224:337– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaiser:1997:HAI

[620] M. Kaiser, V. Klingspor, and H. Friedrich. Human-agent interaction and machine learning. Lecture Notes in Computer Science, 1224:345–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Falessi:1997:HHP

[621] R. Falessi, M. P. Pensini, G. Rovigatti, and S. Taraglio. HIPERCLASS: High performance industrial inspection and defect classification in steel industry. Lecture Notes in Computer Science, 1225:21–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferri:1997:PPI

[622] C. Ferri, P. Contri, P. Salvaneschi, and G. Meloni. PIERS: A parallel integrated environment for reliability evaluation of structures. Lecture Notes in Computer Science, 1225:31–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoekstra:1997:HPD

[623] A. G. Hoekstra, L. Dorst, M. Bergman, and J. Lagerberg. High performance discrete event simulations to evaluate complex industrial systems. Lecture Notes in Computer Science, 1225:41– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldini:1997:HPI

[624] S. Baldini, L. Giraud, L. Hamel, and J. M. Jimenez. HIPERCOMBATS: A parallel industrial tool for two-wheeler suspensions design. Lecture Notes in Computer Science, 1225:51–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jezequel:1997:USD

[625] J.-M. Jezequel. Using scalable distributed computers in telecommunications. Lecture Notes in Computer Science, 1225:60–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guidec:1997:PIS

[626] F. Guidec, P. Calegari, and P. Kuonen. Parallel irregular software for wave propagation simulation. Lecture Notes in Computer Science, 1225:84– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruckmann:1997:HFW

[627] A. Bruckmann, J. Haemmerle, M. Reichl, and A. Uhl. Hybrid fractal/wavelet image compression in a high performance computing environment. Lecture Notes in Computer Science, 1225:117–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hancock:1997:DPV

[628] D. J. Hancock and R. J. Hubbold. Distributed parallel volume rendering on shared memory systems. Lecture Notes in Computer Science, 1225:157– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bigeard:1997:PAU

[629] R. Bigeard, M. Devel, and J. Julliand. Parallelization of an algorithm used to simulate atomic force microscope images. Lecture Notes in Computer Science, 1225:175–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keppens:1997:DPP

[630] R. Keppens, S. Poedts, P. M. Meijer, and J. P. Goedbloed. A data parallel pseudo-spectral semi-implicit magnetohydrodynamics code. Lecture Notes in Computer Science, 1225:190–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeRonde:1997:HPS

[631] J. F. De Ronde, G. D. Van Albada, and P. M. A. Sloot. High performance simulation for resonant-mass gravitational radiation antennas. Lecture Notes in Computer Science, 1225:200–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzwinel:1997:DPA

[632] W. Dzwinel, J. Dzwinel, and K. Dzwinel. Development of parallel applications for MEGA-D — system for oil and gas prospecting. Lecture Notes in Computer Science, 1225:223–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drummond:1997:POA

[633] L. A. Drummond, J. D. Farrara, C. R. Mechoso, and J. Z. Lou. Performance optimization of an atmospheric model in massively parallel computers. Lecture Notes in Computer Science, 1225:263–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaschke:1997:PSA

[634] L. Jaschke. Parallel simulation of an ash melting furnace. Lecture Notes in Computer Science, 1225:282–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breitler:1997:HPC

[635] M. Breitler, S. Hegi, J.-D. Reymond, and N. S. Tuchschmid. High performance computations for an optimal portfolio choice problem. Lecture Notes in Computer Science, 1225:293– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1997:TWB

[636] Y. Huang and W. F. McColl. A two-way BSP algorithm for tridiagonal systems. Lecture Notes in Computer Science, 1225:321–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Issman:1997:NOP

[637] E. Issman and G. Degrez. Nonoverlapping preconditioners for a parallel implicit Navier–Stokes solver. Lecture Notes in Computer Science, 1225:

351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSturler:1997:PSI

[638] E. De Sturler and D. Loher. Parallel solution of irregular, sparse matrix problems using High Performance Fortran. Lecture Notes in Computer Science, 1225:360–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:IQM

[639] T.-R. Yang and H.-X. Lin. The improved quasi-minimal residual method on massively distributed memory computers. Lecture Notes in Computer Science, 1225:389–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:IAC

[640] T.-R. Yang and H.-X. Lin. Isoefficiency analysis of CGLS algorithm for parallel least squares problems. Lecture Notes in Computer Science, 1225:452– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aloisio:1997:DWB

[641] G. Aloisio, M. Cafaro, R. Williams, and P. Messina. A distributed Web-based metacomputing environment. Lecture Notes in Computer Science, 1225:480– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brune:1997:LCI

[642] M. Brune, J. Gehring, and A. Reinefeld. A lightweight communication interface for parallel programming environments. Lecture Notes in Computer Science, 1225:503–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1997:OOL

[643] M. Bubak, W. Ciesla, and K. Sowa. Object-oriented library of parallel genetic algorithms and its implementation on networked workstations and HP/ convex exemplar. Lecture Notes in Computer Science, 1225:514–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derksen:1997:PFF

[644] J. J. Derksen, J. L. Kooman, and H. E. A. Van den Akker. Parallel fluid flow simulations by means of a lattice-Boltzmann scheme. Lecture Notes in Computer Science, 1225:524–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dralle:1997:DAO

[645] U. Dralle and A. Reinefeld. A distributed algorithm for optimal concurrent communication and load balancing in parallel systems. Lecture Notes in Computer Science, 1225:588–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cox:1997:RTK

[646] A. Cox and A. Hall. A real time kernel to support the Transputer programming model. Lecture Notes in Computer Science, 1225:615–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barak:1997:PMP

[647] A. Barak and O. La’adan. Performance of the MOSIX parallel system for a cluster of PC’s. Lecture Notes in Computer Science, 1225:624–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frank:1997:ASS

[648] R. Frank and H. Burkhart. Application support by software reuse: The ALWAN approach. Lecture Notes in Computer Science, 1225:636–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFlorio:1997:ALD

[649] V. De Florio, G. Decomnck, and R. Lauwereins. An application-level dependable technique for farmer-worker parallel programs. Lecture Notes in Computer Science, 1225:644–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradi:1997:IAU

[650] A. Corradi, L. Leonardi, and F. Zambonelli. Integration of automated and user-level tools toward efficient parallel objects allocation. Lecture Notes in Computer Science, 1225:654–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hordijk:1997:PEC

[651] J. Hordijk and H. Corporaal. The potential of exploiting coarse-grain task parallelism from sequential programs. Lecture Notes in Computer Science, 1225:664–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zingirian:1997:SIP

[652] N. Zingirian and M. Maresca. Scheduling image processing program activities on instruction level parallel RISC through program transformations. Lecture Notes in Computer Science, 1225: 674–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyd:1997:EAF

[653] S. Boyd and K. Marcus. The economic addition of functionality to a network. Lecture Notes in Computer Science, 1225:688–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dormanns:1997:PIN

[654] M. Dormanns, W. Sprangers, H. Ertl, and T. Bemmerl. A programming interface for NUMA shared-memory clusters. Lecture Notes in Computer Science, 1225:698–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordsen:1997:CSM

[655] J. Cordsen and W. Schroeder-Preikschat. On the coexistence of shared-memory and message-passing in the programming of parallel applications. Lecture Notes in Computer Science, 1225:718–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doroshenko:1997:MSC

[656] A. E. Doroshenko. Modeling synchronization and communication abstractions for dynamical parallelization. Lecture Notes in Computer Science, 1225:752–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hill:1997:LLI

[657] J. M. D. Hill and D. B. Skillicorn. Lessons learned from implementing BSP. Lecture Notes in Computer Science, 1225:762–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yau:1997:EHP

[658] H. W. Yau, G. C. Fox, and K. A. Hawick. Evaluation of High Performance Fortran through application kernels. Lecture Notes in Computer Science, 1225:772–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fahringer:1997:ESA

[659] T. Fahringer. Effective symbolic analysis to support parallelizing compilers and performance analysis. Lecture Notes in Computer Science, 1225:781– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evans:1997:OCA

[660] E. W. Evans, S. P. Johnson, P. F. Leggett, and M. Cross. Overlapped communications automatically generated in a parallelisation tool. Lecture Notes in Computer Science, 1225:801– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brezany:1997:PIC

[661] P. Brezany, A. Choudhary, and M. Dang. Parallelization of irregular out-of-core applications for distributedmemory systems. Lecture Notes in Computer Science, 1225:811–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Germain:1997:IIP

[662] C. Germain and D. Gautier de Lahaut. Improving irregular parallel communication through sorting. Lecture Notes in Computer Science, 1225:821– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fahringer:1997:ECP

[663] T. Fahringer. Estimating cache performance for sequential and data parallel programs. Lecture Notes in Computer Science, 1225:840–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cortes:1997:ACC

[664] T. Cortes, S. Girona, and J. Labarta. Avoiding the cache-coherence problem in a parallel/distributed file system. Lecture Notes in Computer Science, 1225:860–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haas:1997:MNS

[665] S. Haas, D. A. Thornley, M. Zhu, and R. W. Dobinson. The Macrame 1024 node switching network. Lecture Notes in Computer Science, 1225:870– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karkowski:1997:OLT

[666] I. Karkowski and H. Corporaal. Overcoming the limitations of the traditional loop parallelization. Lecture Notes in Computer Science, 1225:898– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:VIS

[667] Q. Wang, Y. Yijun, and E. D’Hollander. Visualizing the iteration space in PEFPT. Lecture Notes in Computer Science, 1225:908–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bianchi:1997:BFM

[668] F. Bianchi, E. Corno, M. Rebaudengo, and M. Sonza Reorda. Boolean function manipulation on a parallel system using BDDs. Lecture Notes in Computer Science, 1225:916–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altilar:1997:PAV

[669] D. T. Altilar, Y. Paker, and A. V. Sahiner. A parallel architecture for video processing. Lecture Notes in Computer Science, 1225:929–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerlach:1997:ECS

[670] J. Gerlach, M. Sato, and Y. Ishikawa. Experiences with the C++ standard template library and MPI for a parallel particle simulation method. Lecture Notes in Computer Science, 1225: 961–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gregoire:1997:PEA

[671] J. P. Gregoire, J. D. Mattei, and G. Simeoni. Parallelization of Estet-Astrid code on Cray C98. Lecture Notes in Computer Science, 1225:970– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calafiura:1997:ISD

[672] P. Calafiura, G. Wirrer, and B. Panzer-Steindel. An integrated storage and data management system for a high energy physics experiment. Lecture Notes in Computer Science, 1225:975– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baraglia:1997:DGA

[673] R. Baraglia, M. Bucci, D. Circelli, and R. Perego. On designing genetic algorithms for hypercube machines. Lecture Notes in Computer Science, 1225:980–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baraglia:1997:MOP

[674] R. Baraglia, R. Ferrini, D. Laforenza, and A. Lagana. Metacomputing to overcome the power limits of a single machine. Lecture Notes in Computer Science, 1225:982–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aglietti:1997:TPC

[675] F. Aglietti, A. Bartolini, C. Battista, and S. Cabasino. The teraflop parallel computer APEmille. Lecture Notes in Computer Science, 1225:991–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fornasari:1997:DDT

[676] N. Fornasari, G. Gazzaniga, S. Rovida, and G. Sacchi. Domain decomposition techniques: Analysis of a parallel implementation on HP-Convex exemplar systems. Lecture Notes in Computer Science, 1225:1000–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1997:PAE

[677] M. Bubak, W. Funika, and J. Moscinski. Performance analysis environment for parallel applications on networked workstations. Lecture Notes in Computer Science, 1225:1002–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galtier:1997:PPA

[678] J. Galtier and P. Klein. A parallel preprocessor applied to fluid dynamics problems. Lecture Notes in Computer Science, 1225:1006–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnes:1997:CME

[679] K. D. Barnes, N. Ivezic, and M. R. Leuze. Collaborative management environment: A Web-based management tool. Lecture Notes in Computer Science, 1225:1010–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chatagny:1997:PSF

[680] R. Chatagny and B. Chopard. Parallel simulation of a foreign exchange market model. Lecture Notes in Computer Science, 1225:1012–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chochia:1997:RMI

[681] G. Chochia and M. Cole. Recursive 3D mesh indexing with improved locality. Lecture Notes in Computer Science, 1225:1014–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chrobak:1997:PFS

[682] R. Chrobak, M. Bubak, J. Kitowski, and J. Moscinski. Parallel FEM simulation of forging processes on workstations and HP/ convex exemplar. Lecture Notes in Computer Science, 1225: 1018–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alda:1997:HPS

[683] W. Alda, W. Dzwinel, J. Kitowski, and J. Moscinski. High performance simulation of thermal convection using quasiparticle approach. Lecture Notes in Computer Science, 1225:1022–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMartino:1997:DCO

[684] B. Di Martino, A. Mazzeo, N. Mazzocca, and S. Russo. Distributed coordination in optimization algorithms. Lecture Notes in Computer Science, 1225:1024–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dillon:1997:EQD

[685] E. Dillon, C. Gamboa dos Santos, and J. Guyard. An environment for quick design and efficient implementation of message-passing applications. Lecture Notes in Computer Science, 1225:1026–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hawick:1997:ABD

[686] K. A. Hawick, H. A. James, K. J. Maciunas, and F. A. Vaughan. An ATMbased distributed high performance computing system. Lecture Notes in Computer Science, 1225:1032–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hawick:1997:GIS

[687] K. A. Hawick, H. A. James, K. J. Maciunas, and F. A. Vaughan. Geographic information systems applications on an ATM-based distributed high performance computing system. Lecture Notes in Computer Science, 1225:1035– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghiot:1997:SAL

[688] C. Ghiot, J.-M. Beckets, G. Carabin, and A. Dassargues. Sea air land modeling operational network. Lecture Notes in Computer Science, 1225:1040–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collins:1997:GRS

[689] J. J. Collins and M. Eaton. A global representation scheme for genetic algorithms. Lecture Notes in Computer Science, 1226:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agapie:1997:GAS

[690] A. Agapie and D. Caragea. Genetic algorithms, schemata construction and statistics. Lecture Notes in Computer Science, 1226:16–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boquete:1997:ISU

[691] L. Boquete, R. Barea, R. Garcia, and M. Mazo. Identification of systems using radial basis networks feedbacked with FIR filters. Lecture Notes in Computer Science, 1226:46–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butkiewicz:1997:PCS

[692] B. Butkiewicz. Position control system with fuzzy microprocessor AL220. Lecture Notes in Computer Science, 1226: 74–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1997:FNC

[693] M. M. Gupta. Fuzzy-neural computing systems: Recent developments and future directions. Lecture Notes in Computer Science, 1226:82–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haese:1997:OSO

[694] K. Haese. Optimizing the selforganizing-process of topology maps. Lecture Notes in Computer Science, 1226:92–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grigore:1997:SSO

[695] O. Grigore. Syntactical self-organizing map. Lecture Notes in Computer Science, 1226:101–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dias:1997:LLB

[696] J. M. Dias and A. D. Correia. Online learning based on adaptive similarity and fixed size rule base. Lecture Notes in Computer Science, 1226: 110–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eversheim:1997:AFC

[697] W. Eversheim, A. Roggatz, and R. E. Young. Ability of fuzzy constraint networks to reduce the dynamic of changes in product development. Lecture Notes in Computer Science, 1226: 122–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Friedrich:1997:UGE

[698] C. M. Friedrich and C. Moraga. Using genetic engineering to find modular structures and activation functions for architectures of artificial neural networks. Lecture Notes in Computer Science, 1226:150–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giuclea:1997:GAS

[699] M. Giuclea and A. Agapie. Genetic algorithms for solving systems of fuzzy relational equations. Lecture Notes in Computer Science, 1226:162–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Igel:1997:CSA

[700] C. Igel and K.-H. Temme. Chaining syllogism applied to fuzzy IF-THEN rules and rule bases. Lecture Notes in Computer Science, 1226:179–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dvorak:1997:CPF

[701] A. Dvorak. Computational properties of fuzzy logic deduction. Lecture Notes in Computer Science, 1226:189– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besli:1997:AFC

[702] N. Besli, E. Gul, and M. Tuerker. Application of fuzzy control to fed-batch yeast fermentation. Lecture Notes in Computer Science, 1226:207–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fukuda:1997:ALE

[703] T. Fukuda and K. Shimojima. Adaptation, learning and evolutionary computing for intelligent robots. Lecture Notes in Computer Science, 1226:217– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomez-Skarmeta:1997:FCI

[704] A. F. Gomez-Skarmeta, J. Piernas, and M. Delgado. Fuzzy clustering and images reduction. Lecture Notes in Computer Science, 1226:241–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Falkowski:1997:PSC

[705] B.-J. Falkowski. Predicates and superimposed coding: An application to image recognition. Lecture Notes in Computer Science, 1226:250–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adrian:1997:FMC

[706] S. Adrian. Fuzzy modeling of conceptual spaces. Lecture Notes in Computer Science, 1226:277–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benmakrouha:1997:CFM

[707] F. Benmakrouha. Construction of a fuzzy model with few examples. Lecture Notes in Computer Science, 1226: 285–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBaets:1997:FZI

[708] B. De Baets, M. Mares, and R. Mesiar. Fuzzy zeroes and indistinguishability of real numbers. Lecture Notes in Computer Science, 1226:299–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berger:1997:DSN

[709] A. Berger, D. P. F. Moeller, and M. Reuter. Detection of sleep with new preprocessing methods for EEG analyzing. Lecture Notes in Computer Science, 1226:304–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gedeon:1997:WCC

[710] T. D. Gedeon, R. A. Bustos, B. J. Briedis, and G. Greenleaf. Wordconcept clusters in a legal document collection. Lecture Notes in Computer Science, 1226:329–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fogel:1997:RVO

[711] L. J. Fogel. A retrospective view and outlook on evolutionary algorithms. Lecture Notes in Computer Science, 1226:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:NTF

[712] J. Zhang, A. Knoll, and K. V. Le. A new type of fuzzy logic system for adaptive modeling and control. Lecture Notes in Computer Science, 1226: 363–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Binh:1997:SOL

[713] T. T. Binh and U. Korn. Scalar optimization with linear and nonlinear constraints using evolution strategies. Lecture Notes in Computer Science, 1226:381–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karakas:1997:DAF

[714] E. Karakas, H. Kueluenk, E. Oezdemir, and E. Oelcer. Design of adaptive fuzzy controller for Bolu highway tunnel ventilation system. Lecture Notes in Computer Science, 1226:403–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciuca:1997:LNN

[715] I. Ciuca and J. A. Ware. Layered neural networks as universal approximators. Lecture Notes in Computer Science, 1226:411–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerke:1997:FBT

[716] M. Gerke and H. Hoyer. Fuzzy backpropagation training of neural networks. Lecture Notes in Computer Science, 1226:416–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voutchkov:1997:IWW

[717] I. I. Voutchkov and K. D. Velev. Identification of waste water treatment plant using neural networks. Lecture Notes in Computer Science, 1226:478–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derdiyok:1997:FLB

[718] A. Derdiyok, N. Inanc, V. Oebulur, and H. Pastaci. Fuzzy logic based control of switched reluctance motor to reduce torque ripple. Lecture Notes in Computer Science, 1226:484–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1997:AEA

[719] S. Wagner and H.-D. Kochs. An adjusted evolutionary algorithm for the optimization of fuzzy controllers. Lecture Notes in Computer Science, 1226: 517–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agapie:1997:EFU

[720] Adriana Agapie and Alexandru Agapie. Economic forecasting using genetic algorithms. Lecture Notes in Computer Science, 1226:543–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albert:1997:IGA

[721] J. Albert and J. Schoof. Interactive genetic algorithms and evolution based cooperative problem-solving. Lecture Notes in Computer Science, 1226:545– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alpbaz:1997:FMB

[722] M. Alpbaz, Z. Zeybek, L. Oezbek, and F. Aliev. A fuzzy model based control of well mixed reactor with cooling jacket. Lecture Notes in Computer Science, 1226:546–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blume:1997:OVS

[723] H. Blume, O. Franzen, and M. Schmidt. Optimizing video signal processing algorithms by evolution strategies. Lecture Notes in Computer Science, 1226: 547–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borgulya:1997:EDS

[724] I. Borgulya. An example for decision support in insurance. Lecture Notes in Computer Science, 1226:549–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damiani:1997:FID

[725] E. Damiani and M. G. Fugini. Fuzzy identification of distributed components. Lecture Notes in Computer Science, 1226:550–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Englert:1997:NCN

[726] M. Englert, R. Trapp, and R. De Klerk. Neural control of a nonlinear elastic two-mass system. Lecture Notes in Computer Science, 1226:553–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freisleben:1997:LFS

[727] B. Freisleben and St. Strelen. A learning fuzzy system for predicting overshootings in process control. Lecture Notes in Computer Science, 1226:554– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halinka:1997:ABP

[728] A. Halinka, M. Szewczyk, and B. Witek. ANN-based power unit protective system. Lecture Notes in Computer Science, 1226:556–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamburg:1997:SPM

[729] I. Hamburg and L. Marin. A scheduling problem with multicriteria evaluation. Lecture Notes in Computer Science, 1226:558–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Izak:1997:HIM

[730] R. Izak, K. Trott, T. Zahn, and U. Markl. Hardware implementation of a mixed analog-digital neural network. Lecture Notes in Computer Science, 1226:560–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaulent:1997:FSM

[731] M. C. Jaulent, V. Bombardiet, A. Bubel, and J. Bremont. A fuzzy segmentation method for digital substract angiograms analysis. Lecture Notes in Computer Science, 1226:561–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wasiewicz:1997:GPO

[732] P. Wasiewicz and J. Mulawka. Genetic programming in optimization of algorithms. Lecture Notes in Computer Science, 1226:581–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zivaljevic:1997:AFL

[733] M. Zivaljevic and B. Rasuo. Application of fuzzy logic in solving problems from the area of reliability and maintenance of aircraft systems. Lecture Notes in Computer Science, 1226: 583–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgieva:1997:FCD

[734] O. Georgieva. Fuzzy control design of an activated sludge plant. Lecture Notes in Computer Science, 1226:593– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barabash:1997:UFO

[735] S. Barabash. Use of fuzzy objects for statistical description of discrete systems with complex behavior. Lecture Notes in Computer Science, 1226:595– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cistelecan:1997:AGA

[736] M. Cistelecan. On applying genetic algorithms to an intelligent slidingmode control system training. Lecture Notes in Computer Science, 1226:598– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heirman:1997:NNE

[737] P. Heirman and R. Serneels. Neural networks and edge detection. Lecture Notes in Computer Science, 1226:601– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bookstein:1997:QVD

[738] F. L. Bookstein. Quadratic variation of deformations. Lecture Notes in Computer Science, 1230:15–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaturvedi:1997:AEI

[739] P. Chaturvedi, M. F. Insana, and T. J. Hall. Acoustic and elastic imaging to model disease-induced changes in soft tissue structure. Lecture Notes in Computer Science, 1230:43–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buvat:1997:UST

[740] I. Buvat, M. L. Bartlett, A. N. Kitsiou, and J. M. Carson. Using spatialtemporal images for analysis of gated cardiac tomographic data: The “Mmode” revisited. Lecture Notes in Computer Science, 1230:85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Worth:1997:ASB

[741] A. J. Worth, N. Makris, J. W. Meyer, and V. S. Caviness. Automated segmentation of brain exterior in MR images driven by empirical procedures and anatomical knowledge. Lecture Notes in Computer Science, 1230:99– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1997:ASU

[742] T. N. Jones and D. N. Metaxas. Automated 3D segmentation using deformable models and fuzzy affinity. Lecture Notes in Computer Science, 1230: 113–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fritsch:1997:SMI

[743] D. Fritsch, S. M. Pizer, L. Yu, and V. Johnson. Segmentation of medical image objects using deformable shape loci. Lecture Notes in Computer Science, 1230:127–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrett:1997:BDR

[744] H. H. Barrett and C. K. Abbey. Bayesian detection of random signals on random backgrounds. Lecture Notes in Computer Science, 1230:155–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abbey:1997:ADM

[745] C. K. Abbey, E. Clarkson, H. H. Barrett, and S. P. Mueller. Approximate distributions for maximum likelihood and maximum a posteriori estimates under a Gaussian noise model. Lecture Notes in Computer Science, 1230:167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aylward:1997:CGM

[746] S. Aylward and S. M. Pizer. Continuous Gaussian mixture modeling. Lecture Notes in Computer Science, 1230: 176–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yavuz:1997:NSM

[747] M. Yavuz and J. A. Fessler. New statistical models for randoms-precorrected PET scans. Lecture Notes in Computer Science, 1230:190–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edwards:1997:DIG

[748] P. J. Edwards, D. L. G. Hill, J. A. Little, and D. J. Hawkes. Deformation for image-guided interventions using a three component tissue model. Lecture Notes in Computer Science, 1230:218– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benali:1997:STS

[749] H. Benali, I. Buvat, J. L. Anton, and M. Pelegrini. Space-time statistical model for functional MRI image sequences. Lecture Notes in Computer Science, 1230:285–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Houston:1997:JND

[750] A. S. Houston. Jackknifing a normal database for regional cerebral blood flow SPECT obtained using principal components. Lecture Notes in Computer Science, 1230:299–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerwin:1997:GCM

[751] W. S. Kerwin and J. L. Prince. Generating 3D cardiac material markers using tagged MRI. Lecture Notes in Computer Science, 1230:313–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denney:1997:IMT

[752] T. S. Denney. Identification of myocardial tags in tagged MR images without prior knowledge of myocardial contours. Lecture Notes in Computer Science, 1230:327–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atkinson:1997:AAA

[753] D. Atkinson, D. L. G. Hill, P. N. R. Stoyle, and P. E. Summers. An autofocus algorithm for the automatic correction of motion artifacts in MR images. Lecture Notes in Computer Science, 1230:341–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freeborough:1997:MGR

[754] P. A. Freeborough and N. C. Fox. Measurement of global and regional cerebral volume changes by integrating boundary shifts between registered serial 3D MR scans. Lecture Notes in Computer Science, 1230:355–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duta:1997:SIM

[755] N. Duta and M. Sonka. Segmentation and interpretation of MR brain images using an improved knowledge-based active shape model. Lecture Notes in Computer Science, 1230:375–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joshi:1997:GRF

[756] S. C. Joshi, A. Banerjee, G. E. Christensen, and J. G. Csernansky. Gaussian random fields on sub-manifolds for characterizing brain surfaces. Lecture Notes in Computer Science, 1230:381– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hata:1997:FLA

[757] Y. Hata, S. Kobashi, N. Kamiura, and M. Ishikawa. Fuzzy logic approach to 3D magnetic resonance image segmentation. Lecture Notes in Computer Science, 1230:387–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dean:1997:CTB

[758] D. Dean, K. Subramanyan, J. Kamath, and F. L. Bookstein. Comparison of traditional brain segmentation tools with 3D self-organizing maps. Lecture Notes in Computer Science, 1230:393–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1997:FBC

[759] C. Xu, D. Pham, and J. Prince. Finding the brain cortex using fuzzy segmentation, isosurfaces, and deformable surface models. Lecture Notes in Computer Science, 1230:399–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zwiggelaar:1997:SML

[760] R. Zwiggelaar, T. C. Parr, C. R. M. Boggis, and S. M. Astley. Statistical modelling of lines and structures in mammograms. Lecture Notes in Computer Science, 1230:405–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1997:SCV

[761] D. L. Wilson and J. A. Noble. Segmentation of cerebral vessels and aneurysms from MR angiography data. Lecture Notes in Computer Science, 1230:423–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aharon:1997:CEP

[762] S. Aharon, B. M. Cameron, and R. A. Robb. Computation of efficient patientspecific models from 3D medical images: Use in virtual endoscopy and surgery rehearsal. Lecture Notes in Computer Science, 1230:429–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:HAD

[763] G.-Z. Yang and D. M. Hansell. A hybrid approach for the detection of small airways disease from computed tomographic images. Lecture Notes in Computer Science, 1230:447–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1997:CSC

[764] S. P. Huang, B. E. Chapman, J. E. Muhlestein, and D. D. Blatter. Computer simulation of convection and diffusion effects on velocity estimations from X-ray contrast density time curves. Lecture Notes in Computer Science, 1230:453–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hill:1997:ALI

[765] A. Hill, A. D. Brett, and C. J. Taylor. Automatic landmark identification using a new method of non-rigid correspondence. Lecture Notes in Computer Science, 1230:483–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cotton:1997:NSI

[766] S. Cotton, E. Claridge, and P. Hall. Noninvasive skin imaging. Lecture Notes in Computer Science, 1230:501– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolfe:1997:MIS

[767] R. Wolfe and D. Laurendeau. Modelling interreflections in a system based on the differential absorption of light (DAL) for the measurement of range images. Lecture Notes in Computer Science, 1230:507–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bullitt:1997:RIV

[768] E. Bullitt, A. Liu, S. Aylward, and S. M. Pizer. Reconstruction of the intracerebral vasculature from MRA and a pair of projection views. Lecture Notes in Computer Science, 1230:537– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapman:1997:TAF

[769] B. E. Chapman, K. C. Goodrich, A. L. Alexander, and D. D. Blatter. Two alternative forced-choice evaluation of vessel visibility increases due to zerofilled interpolation in MR angiography. Lecture Notes in Computer Science, 1230:543–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarkson:1997:BDA

[770] E. Clarkson and H. Barrett. Bayesian detection with amplitude, scale, orientation and position uncertainty. Lecture Notes in Computer Science, 1230: 549–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonner:1997:WWL

[771] A. Bonner. What was Llull up to? Lecture Notes in Computer Science, 1231:

1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjoerner:1997:DVR

[772] N. S. Bjoerner, Z. Manna, H. B. Sipma, and T. E. Uribe. Deductive verification of real-time systems using STeP. Lecture Notes in Computer Science, 1231:22–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1997:RT

[773] M. Broy. Refinement of time. Lecture Notes in Computer Science, 1231:44– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campos:1997:VLR

[774] S. Campos and E. Clarke. The Verus language: Representing time efficiently with BDDs. Lecture Notes in Computer Science, 1231:64–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cau:1997:RIT

[775] A. Cau and H. Zedan. Refining interval temporal logic specifications. Lecture Notes in Computer Science, 1231: 79–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dierks:1997:PAN

[776] H. Dierks. PLC-Automata: A new class of implementable real-time automata. Lecture Notes in Computer Science, 1231:111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DArgenio:1997:RPT

[777] P. R. D’Argenio. Regular processes and timed automata. Lecture Notes in Computer Science, 1231:141–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feyerabend:1997:VFR

[778] K. Feyerabend and B. Josko. A visual formalism for real-time requirement specifications. Lecture Notes in Computer Science, 1231:156–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamane:1997:FSV

[779] S. Yamane. Formal specification and verification method of concurrent and distributed systems by restricted timed automata. Lecture Notes in Computer Science, 1231:169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertran:1997:TMC

[780] M. Bertran. A transformation of monitor into communication synchronized parallel processes: A systematic refinement step in design. Lecture Notes in Computer Science, 1231:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fevrier:1997:CO

[781] A. Fevrier, E. Najm, and J.-B. Stefani. Contracts for ODP. Lecture Notes in Computer Science, 1231:216–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Back:1997:ABC

[782] R. Back, M. Buechi, and E. Sekerinski. Action-based concurrency and synchronization for objects. Lecture Notes in Computer Science, 1231:248–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertran:1997:CEA

[783] M. Bertran, F. Alvarez-Cuevas, and A. Duran. Communication extended abstract types in the refinement of parallel communicating processes. Lecture Notes in Computer Science, 1231:263– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galan:1997:VRD

[784] L. A. Galán and R. Peña. Verification and refinement of distributed programs in a fair framework. Lecture Notes in Computer Science, 1231:280–294, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldsack:1997:SRC

[785] S. Goldsack, K. Lano, and E. Durr. Specification and refinement of continuous real-time systems. Lecture Notes in Computer Science, 1231:310– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:HLE

[786] F. Wang. High-level execution time analysis. Lecture Notes in Computer Science, 1231:325–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomez:1997:SCP

[787] F. C. Gomez, D. De Frutos Escrig, and V. V. Ruiz. A sound and complete proof system for probabilistic processes. Lecture Notes in Computer Science, 1231:340–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gregorio-Rodriguez:1997:TSP

[788] C. Gregorio-Rodriguez, L. Llana-Diaz, M. Nuthez, and Pedro Palao-Gostanza. Testing semantics for a probabilistictimed process algebra. Lecture Notes in Computer Science, 1231:353–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bowman:1997:ELT

[789] H. Bowman and J. Derrick. Extending LOTOS with time: A true concurrency perspective. Lecture Notes in Computer Science, 1231:383–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hernalsteen:1997:ISR

[790] C. Hernalsteen and A. Fevrier. Introduction of a suspend/resume operator in ET-LOTOS. Lecture Notes in Computer Science, 1231:400–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCune:1997:WBS

[791] W. McCune. Well behaved search and the Robbins problem. Lecture Notes in Computer Science, 1232:1–7, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lynch:1997:GDC

[792] C. Lynch. Goal-directed completion using SOUR graphs. Lecture Notes in Computer Science, 1232:8–22, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1997:SCC

[793] D. Kapur. Shostak’s congruence closure as completion. Lecture Notes in Computer Science, 1232:23–37, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuehler:1997:CES

[794] U. Kuehler and C.-P. Wirth. Conditional equational specifications of data types with partial operations for inductive theorem proving. Lecture Notes in Computer Science, 1232:38–52, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otto:1997:CSF

[795] F. Otto, M. Katsura, and Y. Kobayashi. Cross-sections for finitely presented monoids with decidable word problems. Lecture Notes in Computer Science, 1232:53–67, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sattler-Klein:1997:NUR

[796] A. Sattler-Klein. New undecidability results for finitely presented monoids. Lecture Notes in Computer Science, 1232:68–82, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otto:1997:PPR

[797] F. Otto. On the property of preserving regularity for string-rewriting systems. Lecture Notes in Computer Science, 1232:83–97, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Contejean:1997:RSN

[798] E. Contejean, C. Marche, and L. Rabehasaina. Rewrite systems for natural, integral, and rational arithmetic. Lecture Notes in Computer Science, 1232:98–112, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1997:BPI

[799] L. Bachmair and A. Tiwari. D-bases for polynomial ideals over commutative Noetherian rings. Lecture Notes in Computer Science, 1232:113–127, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Struth:1997:WPF

[800] G. Struth. On the word problem for free lattices. Lecture Notes in Computer Science, 1232:128–141, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1997:TGP

[801] D. Kapur and G. Sivakumar. A total ground path ordering for proving termination of AC rewrite systems. Lecture Notes in Computer Science, 1232:142–156, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arts:1997:PIN

[802] T. Arts and J. Giesl. Proving innermost normalisation automatically. Lecture Notes in Computer Science, 1232:157– 171, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zantema:1997:TCS

[803] H. Zantema. Termination of contextsensitive rewriting. Lecture Notes in Computer Science, 1232:172–186, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oyamaguchi:1997:NPC

[804] M. Oyamaguchi and Y. Ohta. A new parallel closed condition for Church– Rosser of left-linear term rewriting systems. Lecture Notes in Computer Science, 1232:187–201, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1997:ICS

[805] N. Dershowitz. Innocuous constructorsharing combinations. Lecture Notes in Computer Science, 1232:202–216, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perlo-Freeman:1997:SCT

[806] S. Perlo-Freeman and P. Proehle. Scott’s conjecture is true, position sensitive weights. Lecture Notes in Computer Science, 1232:217–227, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lafont:1997:TDR

[807] Y. Lafont. Two dimensional rewriting. Lecture Notes in Computer Science, 1232:228–229, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beeher:1997:CAI

[808] D. Beeher, P. De Groote, and C. Retore. A complete axiomatisation for the inclusion of series-parallel partial orders. Lecture Notes in Computer Science, 1232:230–240, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marcinkowski:1997:UFO

[809] J. Marcinkowski. Undecidability of the first order theory of one-step right ground rewriting. Lecture Notes in Computer Science, 1232:241–253, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vorobyov:1997:FOT

[810] S. Vorobyov. The first-order theory of one step rewriting in linear Noetherian systems is undecidable. Lecture Notes in Computer Science, 1232:254–268, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomas:1997:SLD

[811] A. P. Tomas and M. Filgueiras. Solving linear Diophantine equations using the geometric structure of the solution space. Lecture Notes in Computer Science, 1232:269–283, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1997:CIU

[812] K. Schulz. A criterion for intractability of E-unification with free function symbols and its relevance for combination of unification algorithms. Lecture Notes in Computer Science, 1232:284–307, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Statman:1997:ERC

[813] R. Statman. Effective reduction and conversion strategies for combinators. Lecture Notes in Computer Science, 1232:299–307, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanOostrom:1997:FFD

[814] V. Van Oostrom. Finite family developments. Lecture Notes in Computer Science, 1232:308–322, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ringeissen:1997:PCU

[815] C. Ringeissen. Prototyping combination of unification algorithms with the ELAN rule-based programming language. Lecture Notes in Computer Science, 1232:323–326, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goebel:1997:IPM

[816] M. Goebel. The invariant package of MAS. Lecture Notes in Computer Science, 1232:327–334, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Green:1997:OSC

[817] E. Green, L. Heath, and B. Keller. Opal: A system for computing noncommutative Gröbner bases. Lecture Notes in Computer Science, 1232:331–334, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ogata:1997:TAM

[818] K. Ogata, K. Ohhara, and K. Futatsugi. TRAM: an abstract machine for order-sorted conditional term rewriting systems. Lecture Notes in Computer Science, 1232:335–338, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borst:1997:TAR

[819] Johan Borst, Lars R. Knudsen, and Vincent Rijmen. Two attacks on reduced IDEA. Lecture Notes in Computer Science, 1233:1– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330001.
pdf.

Zieschang:1997:CPB

[820] Thilo Zieschang. Combinatorial properties of basic encryption operations. Lecture Notes in Computer Science, 1233:14–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1233/12330014.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1233/12330014.
pdf.

Naccache:1997:NPK

[821] David Naccache and Jacques Stern. A new public-key cryptosystem. Lecture Notes in Computer Science, 1233:27– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330027.
pdf.

Boneh:1997:ICC

[822] Dan Boneh, Richard A. DeMillo, and Richard J. Lipton. On the importance of checking cryptographic protocols for faults. Lecture Notes in Computer Science, 1233:37–51, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330037.
pdf;   http://theory.stanford.edu/
~dabo/papers/faults.ps.gz.

Coppersmith:1997:LAN

[823] Don Coppersmith and Adi Shamir. Lattice attacks on NTRU. Lecture Notes in Computer Science, 1233:52– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330052.
pdf.

Young:1997:KUC

[824] Adam Young and Moti Yung. Kleptography: Using cryptography against cryptography. Lecture Notes in Computer Science, 1233:62–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/series/0558/bibs/1233/12330062.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1233/12330062.pdf.

Cramer:1997:FSI

[825] Ronald Cramer and Ivan Bjerre Damgård. Fast and secure immunization against adaptive man-in-the-middle impersonation. Lecture Notes in Computer Science, 1233:75– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1233/12330075.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1233/12330075.
                                        pdf.

Pfitzmann:1997:AF

[826] Birgit Pfitzmann and Michael Waidner. Anonymous fingerprinting. Lecture Notes in Computer Science, 1233: 88–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1233/12330088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1233/12330088.
                                        pdf.

Cramer:1997:SOE

[827] Ronald Cramer, Rosario Gennaro, and Berry Schoenmakers. A secure and optimally efficient multiauthority election scheme. Lecture Notes in Computer Science, 1233:103– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1233/12330103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330103.
pdf.

Verheul:1997:BEF

[828] Eric R. Verheul and Henk C. A. van Tilborg. Binding ElGamal: A fraud-detectable alternative to keyescrow proposals. Lecture Notes in Computer Science, 1233:119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330119.
pdf.

Anderson:1997:GPP

[829] Ross Anderson and Michael Roe. The GCHQ protocol and its problems. Lecture Notes in Computer Science, 1233: 134–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330134.
pdf.

Johansson:1997:BHS

[830] Thomas Johansson. Bucket hashing with a small key size. Lecture Notes in Computer Science, 1233:149– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330149.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1233/12330149.
                                        pdf.

Bellare:1997:NPC

[831] Mihir Bellare and Daniele Micciancio. A new paradigm for collision-free hashing: Incrementality at reduced cost. Lecture Notes in Computer Science, 1233:163–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1233/12330163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1233/12330163.
                                        pdf.

Cachin:1997:SER

[832] Christian Cachin. Smooth entropy and Rényi entropy. Lecture Notes in Computer Science, 1233:193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1233/12330193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1233/12330193.
                                        pdf.

Maurer:1997:ITS

[833] Ueli M. Maurer. Informationtheoretically secure secret-key agreement by NOT authenticated public discussion. Lecture Notes in Computer Science, 1233:209–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1233/12330209.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
1233/12330209.pdf.

Golic:1997:LSW

[834] Jovan Dj. Golic. Linear statistical weakness of alleged RC4 keystream generator. Lecture Notes in Computer Science, 1233:226–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330226.
pdf.

Golic:1997:CAA

[835] Jovan Dj. Golic. Cryptanalysis of alleged A5 stream cipher. Lecture Notes in Computer Science, 1233:239– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330239.
pdf.

Shoup:1997:LBD

[836] Victor Shoup. Lower bounds for discrete logarithms and related problems. Lecture Notes in Computer Science, 1233:256–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330256.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330256.
pdf.

Fischlin:1997:SSP

[837] Roger Fischlin and Claus P. Schnorr. Stronger security proofs for RSA and Rabin bits. Lecture Notes in Computer Science, 1233:267–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1233/12330267.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1233/12330267.
                                        pdf.

Bellare:1997:ROZ

[838] Mihir Bellare, Markus Jakobsson, and Moti Yung. Round-optimal zeroknowledge arguments based on any one-way function. Lecture Notes in Computer Science, 1233:280–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1233/12330280.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1233/12330280.
                                        pdf.

Crepeau:1997:ECP

[839] Claude Crépeau. Efficient cryptographic protocols based on noisy channels. Lecture Notes in Computer Science, 1233:306–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1233/12330306.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1233/12330306.pdf.

Brands:1997:RDL

[840] Stefan Brands. Rapid demonstration of linear relations connected by Boolean operators. Lecture Notes in Computer Science, 1233:318–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330318.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330318.
pdf.

Brassard:1997:OTP

[841] Gilles Brassard and Claude Crépeau. Oblivious transfers and privacy amplification. Lecture Notes in Computer Science, 1233:334–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1233/12330334.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1233/12330334.pdf.

Bosselaers:1997:SDP

[842] Antoon Bosselaers, René Govaerts, and Joos Vandewalle. SHA: A design for parallel architectures? Lecture Notes in Computer Science, 1233: 348–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330348.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330348.
pdf.

Paar:1997:FAA

[843] Christof Paar and Pedro Soria-Rodriguez. Fast arithmetic architectures for public-key algorithms over Galois fields GF ((2^{n})^{m}). Lecture Notes in Computer Science, 1233:363– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330363.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330363.
pdf.

Lercier:1997:FGR

[844] Reynald Lercier. Finding good random elliptic curves for cryptosystems defined over F_{2^{n}}. Lecture Notes in Computer Science, 1233:379–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330379.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330379.
pdf.

Fischlin:1997:ICM

[845] Marc Fischlin. Incremental cryptography and memory checkers. Lecture Notes in Computer Science, 1233:393– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330393.
pdf.

Kurosawa:1997:AWI

[846] Kaoru Kurosawa, Thomas Johansson, and Douglas R. Stinson. Almost kwise independent sample spaces and their cryptologic applications. Lecture Notes in Computer Science, 1233:409– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330409.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330409.
pdf.

Carlet:1997:MCI

[847] Claude Carlet. More correlationimmune and resilient functions over Galois fields and Galois rings. Lecture Notes in Computer Science, 1233:422– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330422.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330422.
pdf.

Kurosawa:1997:DSP

[848] Kaoru Kurosawa and Takashi Satoh. Design of SAC/PC(l ) of order k Boolean functions and three other cryptographic criteria. Lecture Notes in Computer Science, 1233:434–449, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330434.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330434.
pdf.

Jakobsson:1997:DMI

[849] Markus Jakobsson and Moti Yung. Distributed “magic ink” signatures. Lecture Notes in Computer Science, 1233: 450–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1233/12330450.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1233/12330450.
                                        pdf.

Camenisch:1997:EGG

[850] Jan L. Camenisch. Efficient and generalized group signatures. Lecture Notes in Computer Science, 1233:465– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1233/12330465.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1233/12330465.
                                        pdf.

Baric:1997:CFA

[851] Niko Baric and Birgit Pfitzmann. Collision-free accumulators and failstop signature schemes without trees. Lecture Notes in Computer Science, 1233:480–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1233/12330480.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1233/12330480.
                                        pdf.

Girault:1997:SFR

[852] Marc Girault and Jean-François Misarsky. Selective forgery of RSA signatures using redundancy. Lecture Notes in Computer Science, 1233:495– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1233/12330495.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1233/12330495.
pdf.

Anonymous:1997:AIa

[853] Anonymous. Author index. Lecture Notes in Computer Science, 1233:509– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1233/1233indx.pdf.

Artemov:1997:TSH

[854] S. Artemov, J. Davoren, and A. Nerode. Topological semantics for hybrid systems. Lecture Notes in Computer Science, 1234:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1997:DFP

[855] G. Barthe and M. H. Soerensen. Domain-free pure type systems. Lecture Notes in Computer Science, 1234:9–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belegradek:1997:GQQ

[856] O. V. Belegradek, A. P. Stolboushkin, and M. A. Taitslin. Generic queries over quasi-o-minimal domains. Lecture Notes in Computer Science, 1234:21– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bukatin:1997:TCD

[857] M. A. Bukatin and J. S. Scott. Towards computing distances between programs via Scott domains. Lecture Notes in Computer Science, 1234:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clote:1997:SRS

[858] P. Clote. A safe recursion scheme for exponential time. Lecture Notes in Computer Science, 1234:44–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcelle:1997:FMT

[859] B. Courcelle. Finite model theory, universal algebra and graph grammars. Lecture Notes in Computer Science, 1234:53–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dantsin:1997:CQA

[860] E. Dantsin and A. Voronkov. Complexity of query answering in logic databases with complex values. Lecture Notes in Computer Science, 1234: 56–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dekhtyar:1997:RDD

[861] M. I. Dekhtyar and A. J. Dikovsky. Recognition of deductive data base stability. Lecture Notes in Computer Science, 1234:67–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dudakov:1997:CCH

[862] S. M. Dudakov. The concurrency complexity for the Horn fragment of linear logic. Lecture Notes in Computer Science, 1234:78–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durnev:1997:SAP

[863] V. Durnev. Studying algorithmic problems for free semi-groups and groups. Lecture Notes in Computer Science, 1234:88–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1997:LSP

[864] R. Freivalds, G. Tervits, R. Wiehagen, and C. Smith. Learning small programs with additional information. Lecture Notes in Computer Science, 1234:102– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gavrilovich:1997:CES

[865] M. Gavrilovich. Cut elimination for the second order propositional logic with Hilbert’s symbol, extensionality, and full comprehension. Lecture Notes in Computer Science, 1234:113–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golovanov:1997:FBA

[866] M. Golovanov. Finite bases of admissible rules for the logic S52C. Lecture Notes in Computer Science, 1234:119– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGroote:1997:ACC

[867] P. De Groote. An algebraic correctness criterion for intuitionistic proofnets. Lecture Notes in Computer Science, 1234:130–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harel:1997:TTR

[868] D. Harel. Towards a theory of recursive structures. Lecture Notes in Computer Science, 1234:141–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heinemann:1997:CPF

[869] B. Heinemann. On the complexity of prefix formulas in modal logic of subset spaces. Lecture Notes in Computer Science, 1234:145–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xi:1997:SEE

[870] H. Xi. Simulating eta-expansions with β-reductions in the second-order polymorphic lambda-calculus. Lecture Notes in Computer Science, 1234:399– 409, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yavorsky:1997:LSF

[871] R. E. Yavorsky. Logical schemes for first order theories. Lecture Notes in Computer Science, 1234:410–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zakharov:1997:VPF

[872] V. Zakharov. Verification of PLTL formulae by means of monotone disjunctive normal forms. Lecture Notes in Computer Science, 1234:419–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conradi:1997:TUV

[873] R. Conradi and B. Westfechtel. Towards a uniform version model for software configuration management. Lecture Notes in Computer Science, 1235: 1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buffenbarger:1997:WYD

[874] J. Buffenbarger and K. Gruell. What have you done for me lately? (branches, merges, and change logs). Lecture Notes in Computer Science, 1235:18–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1997:CSV

[875] D. W. Weber. Change sets versus change packages: Comparing implementations of change-based SCM. Lecture Notes in Computer Science, 1235: 25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wingerd:1997:CLP

[876] L. Wingerd and C. Seiwald. Constructing a large product with jam. Lecture Notes in Computer Science, 1235:36– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estublier:1997:HLP

[877] J. Estublier, S. Dami, and M. Amiour. High level process modeling for SCM systems. Lecture Notes in Computer Science, 1235:81–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ci:1997:SDS

[878] J. X. Ci, M. Poonawala, W. Tsai, and A. K. Onoma. ScmEngine: A distributed software configuration management environment on X.500. Lecture Notes in Computer Science, 1235: 108–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunt:1997:DCM

[879] J. J. Hunt, F. Lamers, J. Reuter, and W. F. Tichy. Distributed configuration management via Java and the World Wide Web. Lecture Notes in Computer Science, 1235:161–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Froehlich:1997:WCM

[880] P. Froehlich and W. Nejdl. WebRC: Configuration management for a cooperation tool. Lecture Notes in Computer Science, 1235:175–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kilpi:1997:PMR

[881] T. Kilpi. Product management requirements for SCM discipline. Lecture Notes in Computer Science, 1235: 186–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asklund:1997:CSC

[882] U. Asklund and B. Magnusson. A case-study of configuration management with ClearCase in an industrial environment. Lecture Notes in Computer Science, 1235:201–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crnkovic:1997:ECO

[883] I. Crnkovic. Experience with changeoriented SCM tools. Lecture Notes in Computer Science, 1235:222–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernsen:1997:UES

[884] N. O. Bernsen, L. Dybkjaer, and H. Dybkjaer. User errors in spoken humanmachine dialogue. Lecture Notes in Computer Science, 1236:14–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahlbaeck:1997:TDT

[885] N. Dahlbaeck. Towards a dialogue taxonomy. Lecture Notes in Computer Science, 1236:29–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grote:1997:SPH

[886] B. Grote, E. Hagen, A. Stein, and E. Teich. Speech production in humanmachine dialogue: A natural language generation perspective. Lecture Notes in Computer Science, 1236:70– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abella:1997:DPD

[887] A. Abella, M. K. Brown, and B. Buntschuh. Development principles for dialog-based interfaces. Lecture Notes in Computer Science, 1236:

141–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnett:1997:DPS

[888] J. Barnett and M. Singh. Designing a portable spoken dialogue system. Lecture Notes in Computer Science, 1236:156–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Araki:1997:AEE

[889] M. Araki and S. Doshita. Automatic evaluation environment for spoken dialogue systems. Lecture Notes in Computer Science, 1236:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gates:1997:EEE

[890] D. Gates, A. Lavie, L. Levin, and A. Waibel. End-to-end evaluation in JANUS: A speech-to-speech translation system. Lecture Notes in Computer Science, 1236:195–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wellman:1997:MAA

[891] M. P. Wellman. Market-aware agents for a multiagent world. Lecture Notes in Computer Science, 1237:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ermoliev:1997:LAV

[892] Y. M. Ermoliev. Learning and adoption versus optimization. Lecture Notes in Computer Science, 1237:2–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castelfranchi:1997:LSR

[893] C. Castelfranchi. Limits of strategic rationality for agents and M-A systems. Lecture Notes in Computer Science, 1237:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boman:1997:NCR

[894] M. Boman. Norms as constraints on real-time autonomous agent action. Lecture Notes in Computer Science, 1237:36–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dragoni:1997:DBR

[895] A. F. Dragoni, P. Giorgini, and M. Baffetti. Distributed belief revision vs. belief revision in a multi-agent environment: First results of a simulation experiment. Lecture Notes in Computer Science, 1237:45–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeJong:1997:MAC

[896] E. De Jong. Multi-agent coordination by communication of evaluations. Lecture Notes in Computer Science, 1237:63–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaib-draa:1997:CRM

[897] B. Chaib-draa. Causal reasoning in multi-agent systems. Lecture Notes in Computer Science, 1237:79–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glaser:1997:RSA

[898] N. Glaser and P. Morignot. The reorganization of societies of autonomous agents. Lecture Notes in Computer Science, 1237:98–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1997:DPS

[899] M. Fisher and M. Wooldridge. Distributed problem-solving as concurrent theorem proving. Lecture Notes in Computer Science, 1237:128–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ygge:1997:MCM

[900] F. Ygge and H. Akkermans. Making a case for multi-agent systems. Lecture Notes in Computer Science, 1237: 156–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baerentzen:1997:LND

[901] L. Baerentzen, P. Avila, and S. N. Talukdar. Learning network designs for asynchronous teams. Lecture Notes in Computer Science, 1237:177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joosen:1997:BMA

[902] W. Joosen, S. Bijnens, F. Matthijs, and B. Robben. Building multi-agent systems with CORRELATE. Lecture Notes in Computer Science, 1237:197– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonker:1997:MAM

[903] C. M. Jonker and J. Treur. Modelling an agent’s mind and matter. Lecture Notes in Computer Science, 1237:210– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castelfranchi:1997:DC

[904] C. Castelfranchi and R. Falcone. Delegation conflicts. Lecture Notes in Computer Science, 1237:234–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeZen:1997:VDI

[905] G. De Zen and L. Faglia. Video on demand in an integrated IN/B-ISDN scenario. Lecture Notes in Computer Science, 1238:5–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evans:1997:NRA

[906] M. P. Evans, K. T. Kettunen, G. K. Blackwell, and S. M. Furnell. Network resource adaptation in the DOL-MEN service machine. Lecture Notes in Computer Science, 1238:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huelamo:1997:TBP

[907] J. Huelamo, H. Vanderstraeten, J. C. Garcia, and P. Coppo. A TINA based prototype for a multimedia multiparty mobility service. Lecture Notes in Computer Science, 1238:35–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1997:DIC

[908] T. Choi, E. Son, and K.-Y. Yu. DAVIC and the Internet convergence: A convergence architecture using WWW, JAVA, and CORBA. Lecture Notes in Computer Science, 1238:77–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hussmann:1997:IAN

[909] H. Hussmann. Intelligent ATM networks: Services and realisation alternatives. Lecture Notes in Computer Science, 1238:87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dede:1997:OCM

[910] A. Dede, S. Arsenis, A. Tosti, and F. Lucidi. OSAM component model, A key concept for the efficient design of future telecommunication systems. Lecture Notes in Computer Science, 1238:127–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeZen:1997:PSS

[911] G. De Zen, L. Faglia, H. Hussmann, and A. Van der Vekens. Proposal for an IN switching state model in an integrated IN/B-ISDN scenario. Lecture Notes in Computer Science, 1238:179– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johansen:1997:HIB

[912] J. Johansen. Harmonisation/integration of B-ISDN and IN (EURESCOM project P506). Lecture Notes in Computer Science, 1238:189–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1997:GSO

[913] Y. B. Choi and A. Tang. A generic service order handling interface for the cooperative service providers in the deregulating and competitive telecommunications environment. Lecture Notes in Computer Science, 1238:211– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herzog:1997:TTP

[914] U. Herzog and T. Magedanz. From IN toward TINA — potential migration steps. Lecture Notes in Computer Science, 1238:219–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wade:1997:MDI

[915] V. Wade, D. Lewis, M. Sheppard, and M. Tschichholz. A methodology for developing integrated multi-domain service management systems. Lecture Notes in Computer Science, 1238:245– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Covaci:1997:THP

[916] S. Covaci and D. Dragan. Towards harmonised pan-European TMN customer care solutions: Interoperable trouble ticketing management service. Lecture Notes in Computer Science, 1238:255– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griffin:1997:ITL

[917] D. Griffin, G. Pavlou, and T. Tin. Implementing TMN-like management services in a TINA compliant architecture: A case study on resource configuration management. Lecture Notes in Computer Science, 1238:263–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bleakley:1997:TSS

[918] C. Bleakley, W. Donnelly, A. Lindgren, and H. Vuorela. TMN specifications to support inter-domain exchange of accounting, billing and charging information. Lecture Notes in Computer Science, 1238:275–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjerring:1997:DIF

[919] L. H. Bjerring and P. Vorm. Domain interoperability for federated connectivity management. Lecture Notes in Computer Science, 1238:293–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galis:1997:TIN

[920] A. Galis, C. Brianza, C. Leone, and C. Salvatori. Towards integrated network management for ATM and SDH networks supporting a global broadband connectivity management service. Lecture Notes in Computer Science, 1238:303–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deri:1997:SVD

[921] L. Deri and B. Ban. Static vs. dynamic CMIP/SNMP network management using CORBA. Lecture Notes in Computer Science, 1238:329–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dassow:1997:STA

[922] H. Dassow and G. Lehr. SNMP and TMN: Aspects of migration and integration. Lecture Notes in Computer Science, 1238:339–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bogler:1997:INI

[923] G. Bogler. Internet — new inspiration for telecommunications management network. Lecture Notes in Computer Science, 1238:359–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Francis:1997:TGM

[924] J. C. Francis and B. Diem. Towards global multimedia mobility. Lecture Notes in Computer Science, 1238:371– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adamidis:1997:CIU

[925] E. P. Adamidis, G. Fleming, and E. D. Sykas. Consistency issues in the UMTS distributed database. Lecture Notes in Computer Science, 1238:403–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gagnon:1997:SAT

[926] F. Gagnon, D. Maillot, J. Oelnes, and L. Hofseth. A security architecture for TMN inter-domain management. Lecture Notes in Computer Science, 1238:417–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whinnett:1997:EUA

[927] D. Whinnett. End user acceptance of security technology for electronic commerce. Lecture Notes in Computer Science, 1238:447–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fink:1997:PHI

[928] J. Fink, A. Kobsa, and J. Schreck. Personalized hypermedia information provision through adaptive and adaptable system features: User modelling, privacy and security issues. Lecture Notes in Computer Science, 1238:459– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deri:1997:VAN

[929] L. Deri and D. Manikis. VRML: Adding 3D to network management. Lecture Notes in Computer Science, 1238:469–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooper:1997:CLR

[930] K. Cooper, K. Kennedy, and N. McIntosh. Cross-loop reuse analysis and its application to cache optimizations. Lecture Notes in Computer Science, 1239:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayguade:1997:DDL

[931] E. Ayguade, J. Garcia, M. L. Grande, and J. Labarta. Data distribution and loop parallelization for shared-memory multiprocessors. Lecture Notes in Computer Science, 1239:41–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoshida:1997:DLU

[932] A. Yoshida and H. Kasahara. Data localization using loop aligned decomposition for macro-dataflow processing. Lecture Notes in Computer Science, 1239:56–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Creusillet:1997:EVA

[933] B. Creusillet and F. Irigoin. Exact versus approximate array region analyses. Lecture Notes in Computer Science, 1239:86–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Autrey:1997:IRG

[934] T. Autrey and M. Wolfe. Initial results for glacial variable analysis. Lecture Notes in Computer Science, 1239: 120–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fang:1997:CAI

[935] J. Fang. Compiler algorithms on ifconversion, speculative predicate assignment and predicated code optimizations. Lecture Notes in Computer Science, 1239:135–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donaldson:1997:DAP

[936] V. Donaldson and J. Ferrante. Determining asynchronous pipeline execution times. Lecture Notes in Computer Science, 1239:154–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berson:1997:IPO

[937] D. A. Berson, P. Chang, R. Gupta, and M. L. Soffa. Integrating program optimizations and transformations with the scheduling of instruction level parallelism. Lecture Notes in Computer Science, 1239:207–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1997:BSN

[938] P. Chang, D.-Y. Chen, Y.-F. Lee, and Y. Wu. Bidirectional scheduling: A new global code scheduling approach. Lecture Notes in Computer Science, 1239:222–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeDinechin:1997:PCM

[939] B. D. De Dinechin. Parametric computation of margins and of minimum cumulative register lifetime dates. Lecture Notes in Computer Science, 1239: 231–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diniz:1997:LCE

[940] P. Diniz and M. Rinard. Lock coarsening: Eliminating lock overhead in automatically parallelized object-based programs. Lecture Notes in Computer Science, 1239:285–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eigenmann:1997:PWR

[941] R. Eigenmann, I. Park, and M. J. Voss. Are parallel workstations the right target for parallelizing compilers? Lecture Notes in Computer Science, 1239:300–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xue:1997:CMT

[942] J. Xue. Communication-minimal tiling of uniform dependence loops. Lecture Notes in Computer Science, 1239:330– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barua:1997:CMP

[943] R. Barua, D. Kranz, and A. Agarwal. Communication-minimal partitioning of parallel loops and data arrays for cache-coherent distributed-memory multiprocessors. Lecture Notes in Computer Science, 1239:350–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennedy:1997:RBC

[944] K. Kennedy and A. Sethi. Resourcebased communication placement analysis. Lecture Notes in Computer Science, 1239:369–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chandra:1997:HFG

[945] S. Chandra and J. R. Larus. HPF on fine-grain distributed shared memory: Early experience. Lecture Notes in Computer Science, 1239:450–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bik:1997:SQE

[946] A. J. C. Bik and H. A. G. Wijshoff. Simple qualitative experiments with a sparse compiler. Lecture Notes in Computer Science, 1239:466–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chamberlain:1997:FJU

[947] B. L. Chamberlain, S.-E. Choi, E. C. Lewis, and L. Snyder. Factor-join: A unique approach to compiling array languages for parallel machines. Lecture Notes in Computer Science, 1239: 481–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

John:1997:CCS

[948] A. John and J. C. Browne. Compilation of constraint systems to procedural parallel programs. Lecture Notes in Computer Science, 1239:501–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1997:MSC

[949] Arvind, A. Caro, J.-W. Maessen, and S. Aditya. A multithreaded substrate and compilation model for the implicitly parallel language pH. Lecture Notes in Computer Science, 1239:519– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kale:1997:TIP

[950] L. V. Kale, J. Yelon, and T. Knauff. Threads for interoperable parallel programming. Lecture Notes in Computer Science, 1239:534–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1997:SEA

[951] Y.-S. Hwang and J. Saltz. Side effect analysis on user-defined reduction functions with dynamic pointer-linked data structures. Lecture Notes in Computer Science, 1239:594–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griebl:1997:LPL

[952] M. Griebl and C. Lengauer. The loop parallelizer LooPo — announcement. Lecture Notes in Computer Science, 1239:603–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dechering:1997:GFC

[953] P. F. G. Dechering, L. C. Breebaart, F. Kuijlman, and C. Van Reeuwijk. A generalized forall concept for parallel languages. Lecture Notes in Computer Science, 1239:605–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFelipe:1997:MB

[954] J. De Felipe. Microcircuits in the brain. Lecture Notes in Computer Science, 1240:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alonso:1997:IBE

[955] R. Alonso, I. Lopez-Coviella, F. Hernandez-Diaz, and P. Abreu. Interactions between environmental and hormonal oscillations induce plastic changes in a simple neuroendocrine transducer. Lecture Notes in Computer Science, 1240: 72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aleksandrovsky:1997:NNM

[956] B. Aleksandrovsky, F. Bruecher, G. Lynch, and R. Granger. Neural network model of striatal complex. Lecture Notes in Computer Science, 1240: 103–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anastasio:1997:SSO

[957] T. J. Anastasio. Symmetry and selforganization of the oculo-motor neural integrator. Lecture Notes in Computer Science, 1240:116–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delgado-Garcia:1997:QNM

[958] J. M. Delgado-Garcia, A. Gruart, J. A. Domingo, and J. A. Trigo. Quantal neural mechanisms underlying movement execution and motor learning. Lecture Notes in Computer Science, 1240:124–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreu:1997:BBI

[959] E. Andreu, R. Pomares, B. Soria, and J. V. Sanchez-Andres. Balance between intercellular coupling and input resistence as a necessary requirement for oscillatory electrical activity in pancreatic β-cells. Lecture Notes in Computer Science, 1240:146–153, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalo:1997:AJG

[960] I. Gonzalo. Allometry in the Justo Gonzalo’s model of the sensorial cortex. Lecture Notes in Computer Science, 1240:169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gimenez:1997:CPF

[961] V. Gimenez, M. Perez-Castellanos, J. Rios Carrion, and F. De Mingo. Capacity and parasitic fixed points control in a recursive neural network. Lecture Notes in Computer Science, 1240: 217–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hilario:1997:UPK

[962] M. Hilario and A. Rida. The use of prior knowledge in neural network configuration and training. Lecture Notes in Computer Science, 1240:227– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coolen:1997:CBF

[963] A. C. C. Coolen and L. Viana. Competition between feed-forward and lateral information processing in layered neural networks. Lecture Notes in Computer Science, 1240:257–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amari:1997:SAR

[964] S.-I. Amari and N. Murata. Statistical analysis of regularization constant — from Bayes, MDL and NIC points of view. Lecture Notes in Computer Science, 1240:284–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GarciaDelValle:1997:GNN

[965] M. Garcia Del Valle, C. Garcia-Orellana, F. J. Lopez-Aligue, and I. Acevedo-Sotoca. Generic neural network model and simulation toolkit. Lecture Notes in Computer Science, 1240:313–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feng:1997:VCN

[966] J. Feng and D. Brown. Viewing a class of neurodynamics on parameter space. Lecture Notes in Computer Science, 1240:546–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joya:1997:HNN

[967] G. Joya, M. A. Atencia, and F. Sandoval. Hopfield neural network applied to optimization problems: Some theoretical and simulation results. Lecture Notes in Computer Science, 1240:556– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gagne:1997:PTC

[968] F. Gagne and C. Blaise. Predicting toxicity of complex mixtures by artificial neural networks. Lecture Notes in Computer Science, 1240:576–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grabec:1997:CCF

[969] I. Grabec and S. Mandelj. Continuation of chaotic fields by RBFNN. Lecture Notes in Computer Science, 1240: 597–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chinea:1997:IPP

[970] A. Chinea, J. M. Moreno, J. Madrenas, and J. Cabestany. Improving the performance of piecewise linear separation incremental algorithms for practical hardware implementations. Lecture Notes in Computer Science, 1240: 607–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bologna:1997:ADS

[971] G. Bologna and C. Pellegrini. Accurate decomposition of standard MLP classification responses into symbolic rules. Lecture Notes in Computer Science, 1240:617–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guijarro-Berdinas:1997:HIS

[972] B. Guijarro-Berdiñas, A. Alonso-Betanzos, S. Prados-Mendez, and O. Fernandez-Chaves. A hybrid intelligent system for the preprocessing of fetal heart rate signals in antenatal testing. Lecture Notes in Computer Science, 1240:628–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coward:1997:PEA

[973] L. A. Coward. The pattern extraction architecture: A connectionist alternative to the Von Neumann architecture. Lecture Notes in Computer Science, 1240:634–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BarahonaDaFonseca:1997:CPS

[974] J. Barahona Da Fonseca, I. Barahona Da Fonseca, and J. Simoes Da Fonseca. Cognitive processes in social interaction — A neural networks approach. Lecture Notes in Computer Science, 1240:660–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corbacho:1997:SBL

[975] F. J. Corbacho and M. A. Arbib. Schema-based learning: Biologically inspired principles of dynamic organization. Lecture Notes in Computer Science, 1240:678–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ienne:1997:DCH

[976] P. Ienne. Digital connectionist hardware: Current problems and future challenges. Lecture Notes in Computer Science, 1240:688–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beiu:1997:TBS

[977] V. Beiu and T. De Pauw. Tight bounds on the size of neural networks for classification problems. Lecture Notes in Computer Science, 1240:743–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Draghici:1997:PLP

[978] S. Draghici and I. K. Sethi. On the possibilities of the limited precision weights neural networks in classification problems. Lecture Notes in Computer Science, 1240:753–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernier:1997:MBA

[979] J. L. Bernier, J. Ortega, and A. Prieto. A modified backpropagation algorithm to tolerate weight errors. Lecture Notes in Computer Science, 1240: 763–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asanovic:1997:FKN

[980] K. Asanovic. A fast Kohonen net implementation for Spert-II. Lecture Notes in Computer Science, 1240:792– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colodro:1997:HIC

[981] F. Colodro, A. Torralba, R. Gonzalez, and L. G. Franquelo. A hardware implementation of CNNs based on pulse stream techniques. Lecture Notes in Computer Science, 1240:811–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Granado:1997:ECN

[982] B. Granado and P. Garda. Evaluation of the CNAPS neuro-computer for the simulation of MLPs with receptive fields. Lecture Notes in Computer Science, 1240:817–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carvajal:1997:ATA

[983] R. G. Carvajal, A. Torralba, F. Colodro, and L. G. Franquelo. AFAN, a tool for the automatic design of fuzzy and neural controllers. Lecture Notes in Computer Science, 1240:825–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colodro:1997:FSF

[984] F. Colodro, A. Torralba, R. Gonzalez, and L. G. Franquelo. A fully stochastic fuzzy logic controller. Lecture Notes in Computer Science, 1240:834–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chebira:1997:MNN

[985] A. Chebira, K. Madani, and G. Mercier. Multi-neural networks hardware and software architecture: Application of the divide to simplify paradigm DTS. Lecture Notes in Computer Science, 1240:841–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cottrell:1997:KAP

[986] M. Cottrell and P. Rousset. The Kohonen algorithm: A powerful tool for analyzing and representing multidimensional quantitative and qualitative data. Lecture Notes in Computer Science, 1240:861–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerin-Dugue:1997:CNN

[987] A. Guerin-Dugue, P. Teissier, J.-L. Schwartz, and J. Herault. Constrained neural network for estimating sensor reliability in sensors fusion. Lecture Notes in Computer Science, 1240:872– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alvarez:1997:SLC

[988] J. R. Alvarez. Structural level comparison of two basic paradigms in neural computation. Lecture Notes in Computer Science, 1240:892–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barakova:1997:SBI

[989] E. L. Barakova and I. Spaanenburg. Symmetry: Between indecision and equality of choice. Lecture Notes in Computer Science, 1240:903–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dopazo:1997:NTU

[990] J. Dopazo, H. Wang, and J. M. Carazo. A new type of unsupervised growing neural network for biological sequence classification that adopts the topology of a phylogenetic tree. Lecture Notes in Computer Science, 1240:932–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braithwaite:1997:COR

[991] E. Braithwaite, J. Dripps, A. Lyon, and A. F. Murray. Classification of the onset of respiratory difficulties in ventilation assisted neonates. Lecture Notes in Computer Science, 1240:942–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duro:1997:EBC

[992] R. J. Duro and J. Santos. ECG beat classification with synaptic delay based artificial neural networks. Lecture Notes in Computer Science, 1240: 962–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrianasy:1997:NNB

[993] F. Andrianasy and M. Milgram. Neural network-based insulin infusion control for an insulin-pump, using discontinuous blood glucose measurements. Lecture Notes in Computer Science, 1240: 971–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alonso-Betanzos:1997:CAN

[994] A. Alonso-Betanzos, E. Mosquera-Rey, and B. Baldonedo Del Rio. A comparative analysis of the neonatal prognosis problem using artificial neural networks, statistical techniques and certainty management techniques. Lecture Notes in Computer Science, 1240:995– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eckhorn:1997:MVP

[995] R. Eckhorn. Models of visual processing derived from cortical microelectrode recordings. Lecture Notes in Computer Science, 1240:1005–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joya:1997:NNB

[996] G. Joya and F. Sandoval. Neural networks based projectivity invariant recognition of flat patterns. Lecture Notes in Computer Science, 1240:1038– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrandez:1997:BFE

[997] J. M. Ferrandez, D. Del Valle, V. Rodellar, and P. Gomez. A biological front-end processing for speech recognition. Lecture Notes in Computer Science, 1240:1058–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

HernandezAbrego:1997:NPC

[998] G. Hernandez Abrego, E. Monte, and J. B. Maritho. Non parametric coding of speech by means of a MLP with hints. Lecture Notes in Computer Science, 1240:1068–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Croll:1997:MFI

[999] C. Croll, D. Pellerin, and J. Herault. Multidimensional filtering inspired by retino-cortical projection: Application to texture segmentation. Lecture Notes in Computer Science, 1240:1085–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yi:1997:NBI

[1000] C.-H. Yi, R. Schlabbach, H. Kroth, and H. Klar. A new bio-inspired algorithm for early vision edge detection and image segmentation. Lecture Notes in Computer Science, 1240:1107–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishii:1997:FBA

[1001] N. Ishii and K.-I Naka. Function of biological asymetrical neural networks. Lecture Notes in Computer Science, 1240:1115–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofman:1997:ISF

[1002] P. Hofman and J. Van Opstal. Identification of spectral features as sound localization cues in the external ear acoustics. Lecture Notes in Computer Science, 1240:1126–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freitag:1997:PRM

[1003] F. Freitag and E. Monte. Phoneme recognition by means of predictive neural networks. Lecture Notes in Computer Science, 1240:1136–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Won:1997:SIF

[1004] Y. Won, B.-H. Lee, Y.-C. Baek, and J.-S. Lee. Shift-invariant fuzzymorphology neural network for occluded target recognition. Lecture Notes in Computer Science, 1240:1144– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faundez:1997:CSB

[1005] M. Faundez, E. Monte, and F. Vallverdu. A comparative study between linear and nonlinear speech prediction. Lecture Notes in Computer Science, 1240:1154–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalocsai:1997:RME

[1006] P. Kalocsai. Recognition model with extension fields. Lecture Notes in Computer Science, 1240:1164–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brezmes:1997:AAN

[1007] J. Brezmes, N. Canyellas, E. Llobet, and X. Vilanova. Application of artificial neural networks to the design and implementation of electronic olfactory systems. Lecture Notes in Computer Science, 1240:1183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz:1997:SDR

[1008] F. Diaz, J. M. Ferrandez, P. Gomez, and V. Rodellar. Spoken-digit recognition using self-organizing maps with perceptual pre-processing. Lecture Notes in Computer Science, 1240:1203– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia-Berdones:1997:NQD

[1009] C. Garcia-Berdones, J. Narvaez, U. Fernandez, and F. Sandoval. A new QRS detector based on neural network. Lecture Notes in Computer Science, 1240:1260–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bella:1997:ANN

[1010] O. Bella, C. Cortes, and A. Tomas. Application of neural networks for local modelization of the boiler furnace in thermal power plants. Lecture Notes in Computer Science, 1240:1270–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hernandez-Arauzo:1997:ANN

[1011] P. Hernandez-Arauzo, S. Lopez-Garcia, and A. Bahamonde. Artificial neural networks for the computation of traffic queues. Lecture Notes in Computer Science, 1240:1288–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Isasi-Vithuela:1997:UNN

[1012] P. Isasi-Vithuela, J. M. Molina-Lopez, and A. Sanchis De Miguel. Unsupervised neural network for forecasting alarms in hydroelectric power plant. Lecture Notes in Computer Science, 1240:1298–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felzer:1997:MSP

[1013] T. Felzer, P. Hartmann, K. Hohm, and P. Marenbach. Motor sequence processing with an artificial learning system. Lecture Notes in Computer Science, 1240:1307–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cervera:1997:ISO

[1014] E. Cervera and A. P. Del Poblil. Integration of self-organizing feature maps and reinforcement learning in robotics. Lecture Notes in Computer Science, 1240:1344–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iglesias:1997:IBR

[1015] R. Iglesias, C. V. Regueiro, D. Correa, and S. Barro. Implementation of a basic reactive behavior in mobile robotics through artificial neural networks. Lecture Notes in Computer Science, 1240:1364–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeTremiolles:1997:VPM

[1016] G. De Tremiolles, P. Tannhof, B. Plougonven, and C. Demarigny. Visual probe mark inspection, using hardware implementation of artificial neural networks, in VLSI production. Lecture Notes in Computer Science, 1240:1374– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nygaard:1997:GAS

[1017] Kristen Nygaard. GOODS to appear on the stage. Lecture Notes in Computer Science, 1241:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1241/12410001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1241/12410001.
pdf.

Almeida:1997:BTC

[1018] Paulo Sérgio Almeida. Balloon types: Controlling sharing of state in data types. Lecture Notes in Computer Science, 1241:32–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1241/12410032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1241/12410032.
pdf.

Benzaken:1997:SIC

[1019] Véronique Benzaken and Xavier Schaefer. Static integrity constraint management in object-oriented database programming languages via predicate transformers. Lecture Notes in Computer Science, 1241:60–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1241/12410060.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1241/12410060.
                                        pdf.

Miller:1997:IEH

[1020] Robert Miller and Anand Tripathi. Issues with exception handling in object-oriented systems. Lecture Notes in Computer Science, 1241:85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1241/12410085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1241/12410085.
                                        pdf.

Bruce:1997:SGM

[1021] Kim B. Bruce, Leaf Petersen, and Adrian Fiech. Subtyping is not a good “match” for object-oriented languages. Lecture Notes in Computer Science, 1241:104–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1241/12410104.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1241/12410104.
                                        pdf.

Krall:1997:NOH

[1022] Andreas Krall, Jan Vitek, and R. Nigel Horspool. Near optimal hierarchical encoding of types. Lecture Notes in Computer Science, 1241:128–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1241/12410128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1241/12410128.
pdf.

Liquori:1997:ETP

[1023] Luigi Liquori. An extended theory of primitive objects: First order system. Lecture Notes in Computer Science, 1241:146–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1241/12410146.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1241/12410146.
pdf.

Lunau:1997:RAP

[1024] Charlotte Pii Lunau. A reflective architecture for process control applications. Lecture Notes in Computer Science, 1241:170–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1241/12410170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1241/12410170.
pdf.

Mezini:1997:DOE

[1025] Mira Mezini. Dynamic object evolution without name collisions. Lecture Notes in Computer Science, 1241:190– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1241/12410190.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1241/12410190.
pdf.

Kiczales:1997:AOP

[1026] Gregor Kiczales, John Lamping, Anurag Mendhekar, Chris Maeda, Cristina Lopes, Jean-Marc Loingtier, and John Irwin. Aspect-oriented programming. Lecture Notes in Computer Science, 1241:220–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1241/12410220.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1241/12410220.pdf.

Evans:1997:DRT

[1027] Huw Evans and Peter Dickman. DRASTIC: A run-time architecture for evolving, distributed, persistent systems. Lecture Notes in Computer Science, 1241:243–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1241/12410243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1241/12410243.
pdf.

Holst:1997:GFI

[1028] Wade Holst and Duane Szafron. A general framework for inheritance management and method dispatch in objectoriented languages. Lecture Notes in Computer Science, 1241:276–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1241/12410276.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1241/12410276.
pdf.

Onodera:1997:OSS

[1029] Tamiya Onodera and Hiroaki Nakamura. Optimizing Smalltalk by selector code indexing can be practical. Lecture Notes in Computer Science, 1241: 302–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1241/12410302.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1241/12410302.
pdf.

Bicarregui:1997:OAS

[1030] J. C. Bicarregui, K. C. Lano, and T. S. E. Maibaum. Objects, associations and subsystems: A hierarchical approach to encapsulation. Lecture Notes in Computer Science, 1241:324– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1241/12410324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1241/12410324.
pdf.

Breu:1997:TFU

[1031] Ruth Breu, Ursula Hinkel, Christoph Hofmann, Cornel Klein, Barbara Paech, Bernhard Rumpe, and Veronika Thurner. Towards a formalization of the unified modeling language. Lecture Notes in Computer Science, 1241: 344–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1241/12410344.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1241/12410344.
                                        pdf.

Puntigam:1997:CRE

[1032] Franz Puntigam. Coordination requirements expressed in types for active objects. Lecture Notes in Computer Science, 1241:367–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1241/12410367.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1241/12410367.
                                        pdf.

Drossopoulou:1997:JTS

[1033] Sophia Drossopoulou and Susan Eisenbach. Java is type safe — probably. Lecture Notes in Computer Science, 1241:389–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1241/12410389.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1241/12410389.
                                        pdf.

Prehofer:1997:FOP

[1034] Christian Prehofer. Feature-oriented programming: A fresh look at objects. Lecture Notes in Computer Science, 1241:419–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1241/12410419.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/1241/12410419.
pdf.

Thorup:1997:GJV

[1035] Kresten Krab Thorup. Genericity in Java with virtual types. Lecture Notes in Computer Science, 1241:444– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1241/12410444.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1241/12410444.
pdf.

Florijn:1997:TSO

[1036] Gert Florijn, Marco Meijers, and Pieter van Winsen. Tool support for object-oriented patterns. Lecture Notes in Computer Science, 1241:472– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1241/12410472.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1241/12410472.
pdf.

Meusel:1997:MSU

[1037] Matthias Meusel, Krzysztof Czarnecki, and Wolfgang Köpf. A model for structuring user documentation of object- oriented frameworks using patterns and hypertext. Lecture Notes in Computer Science, 1241:496– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1241/12410496.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1241/12410496.
                                        pdf.

Odenthal:1997:UPD

[1038] Georg Odenthal and Klaus Quibeldey-Cirkel. Using patterns for design and documentation. Lecture Notes in Computer Science, 1241:511–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1241/12410511.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1241/12410511.
                                        pdf.

Gamma:1997:GBO

[1039] Erich Gamma. Going beyond objects with design patterns. Lecture Notes in Computer Science, 1241:530– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1241/12410530.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1241/12410530.
                                        pdf.

Anonymous:1997:AIb

[1040] Anonymous. Author index. Lecture Notes in Computer Science, 1241:531– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1241/1241ind.pdf.

Achtmann:1997:ABD

[1041] K. Achtmann, K.-H. Doering, R. Herher, and G. Komp. An ATM-based demonstration model for multimedia services using different access networks. Lecture Notes in Computer Science, 1242:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hussmann:1997:TBI

[1042] H. Hussmann, F. J. Herrera, R. Pasquali, and P. Todorova. A transnational IN/ B-ISDN integrated network for the provision of multimedia services. Lecture Notes in Computer Science, 1242:19– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johansson:1997:DYS

[1043] B.-E. Johansson, M. Andreasson, H. Flinck, and M. Kalervo. Divide your systems to conquer the market. Lecture Notes in Computer Science, 1242: 37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cole:1997:DIM

[1044] A. Cole, J. Robinson, and C. Athwal. The design and implementation of a multimedia system for the interactive and time-synchronised analysis of vehicle crash tests. Lecture Notes in Computer Science, 1242:79–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waddington:1997:RMD

[1045] D. Waddington, C. Edwards, and D. Hutchison. Resource management for distributed multimedia applications. Lecture Notes in Computer Science, 1242:95–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenhalgh:1997:MNA

[1046] C. Greenhalgh and S. Benford. A multicast network architecture for large scale collaborative virtual environments. Lecture Notes in Computer Science, 1242:113–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeHert:1997:CMA

[1047] E. De Hert, R. Guarneri, S. Malmros, and K. Start. Consensus making activities in Europe: The example of multimedia in ACTS. Lecture Notes in Computer Science, 1242:129–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delaigle:1997:LCW

[1048] J. F. Delaigle, C. De Vleeschouwer, F. Goffin, and B. Macq. Low cost watermarking based on a human visual model. Lecture Notes in Computer Science, 1242:153–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bessler:1997:DMA

[1049] S. Bessler, M. Hager, H. Benz, and R. Mecklenburg. DIANE: A multimedia annotation system. Lecture Notes in Computer Science, 1242:183– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniazzi:1997:ART

[1050] S. Antoniazzi, C. Ievola, E. Marcozzi, and R. Penzo. An authoring and runtime environment for interactive multimedia services. Lecture Notes in Computer Science, 1242:245–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badique:1997:TMC

[1051] E. Badique and A. Barbagli. Tools for multimedia content creation, manipulation and management: The ACTS perspective. Lecture Notes in Computer Science, 1242:261–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allan:1997:IDB

[1052] R. Allan and C. Taylor. Interaction with digital broadcast services through a terrestrial return channel. Lecture Notes in Computer Science, 1242:277– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guirao:1997:WAI

[1053] F. Guirao and L. Van Noorden. The work of ACTS in interactive distribution and transmission: Towards interoperable interactive multimedia services. Lecture Notes in Computer Science, 1242:309–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brelot:1997:OEM

[1054] M. Brelot and G. Privat. Object encapsulation of multiscale image region representations. Lecture Notes in Computer Science, 1242:375–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartung:1997:CPV

[1055] F. Hartung and B. Girod. Copyright protection in video delivery networks by watermarking of pre-compressed video. Lecture Notes in Computer Science, 1242:423–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djeraba:1997:TIR

[1056] C. Djeraba and H. Briand. Temporal and interactive relations in a multimedia database system. Lecture Notes in Computer Science, 1242:457–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christodoulakis:1997:KMI

[1057] S. Christodoulakis, N. Pappas, G. Kyriakaki, and Y. Maragoudakis. The KY-DONIA multimedia information server. Lecture Notes in Computer Science, 1242:491–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hutchison:1997:CMI

[1058] D. Hutchison, R. El-Marakby, and L. Mathy. A critique of modern Internet protocols: The issue of support for multimedia. Lecture Notes in Computer Science, 1242:507–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolla:1997:CCS

[1059] R. Bolla, M. Marchese, and S. Zappatore. A congestion control scheme for multimedia traffic in packet switching best-effort networks. Lecture Notes in Computer Science, 1242:523–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alfano:1997:URR

[1060] M. Alfano. User requirements and resource control for cooperative multimedia applications. Lecture Notes in Computer Science, 1242:537–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeLang:1997:STT

[1061] D. De Lang, P. Balsiger, H. Eggenhuisen, and P. Gardien. Swiss TopVision, the trial with interactive content. Lecture Notes in Computer Science, 1242:571–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harper:1997:CIC

[1062] N. Harper. Consumer interactivity in cambridge. Lecture Notes in Computer Science, 1242:581–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carducci:1997:IPD

[1063] F. Carducci and R. Giubilei. ISIS project — demonstration system and pilot trials. Lecture Notes in Computer Science, 1242:609–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniazzi:1997:OSA

[1064] S. Antoniazzi and G. Schapeler. An open software architecture for multimedia consumer terminals. Lecture Notes in Computer Science, 1242:621– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charlton:1997:UAM

[1065] P. Charlton, F. Espinoza, E. Mamdani, and O. Olsson. Using an asset model for integration of agents and multimedia to provide an open service architecture. Lecture Notes in Computer Science, 1242:635–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zahariadis:1997:TNT

[1066] T. Zahariadis, A. Dede, I. Maffett, and G. Stassinopoulos. Telelearning: New trends in education through global multimedia networks. Lecture Notes in Computer Science, 1242:681–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomez:1997:BTA

[1067] E. J. Gomez, P. Cano, F. Del Pozo, and H. Rahms. The BONAPARTE telemedicine ATM multimedia applications. Lecture Notes in Computer Science, 1242:693–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elbert:1997:MSP

[1068] H. Elbert, J. Enssle, M. Frank, and H. Hessenmueller. Multimedia service provisioning in the Munich trial of the AMUSE project. Lecture Notes in Computer Science, 1242:739–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunawardena:1997:RDM

[1069] J. Gunawardena. Recent developments in the mathematics of reactive systems. Lecture Notes in Computer Science, 1243:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henriksen:1997:PVD

[1070] J. G. Henriksen and P. S. Thiagarajan.

A product version of dynamic linear time temporal logic. Lecture Notes in Computer Science, 1243:45–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:1997:RAC

[1071] M. Abadi and A. D. Gordon. Reasoning about cryptographic protocols in the Spi calculus. Lecture Notes in Computer Science, 1243:59–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1997:MTH

[1072] R. Alur and T. A. Henzinger. Modularity for timed and hybrid systems. Lecture Notes in Computer Science, 1243: 74–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1997:URM

[1073] H. R. Andersen, S. Moerk, and M. U.

Soerensen. A universal reactive machine. Lecture Notes in Computer Science, 1243:89–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeten:1997:BSB

[1074] J. C. M. Baeten and J. A. Bergstra.

Bounded stacks, bags and queues. Lecture Notes in Computer Science, 1243: 104–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhat:1997:APO

[1075] G. Bhat and D. Peled. Adding partial orders to linear temporal logic. Lecture Notes in Computer Science, 1243: 119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1997:RAP

[1076] A. Bouajjani, J. Esparza, and O. Maler. Reachability analysis of pushdown automata: application to model-checking. Lecture Notes in Computer Science, 1243:135–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busi:1997:SNI

[1077] N. Busi and G. M. Pinna. Synthesis of nets with inhibitor arcs. Lecture Notes in Computer Science, 1243: 151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleaveland:1997:ATM

[1078] R. Cleaveland, G. Luettgen, and M. Mendler. An algebraic theory of multiple clocks. Lecture Notes in Computer Science, 1243:166–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1997:HVS

[1079] H. Ehrig, R. Geisler, M. Klar, and J. Padberg. Horizontal and vertical structuring techniques for statecharts. Lecture Notes in Computer Science, 1243:181–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fournet:1997:ITM

[1080] C. Fournet, C. Laneve, L. Maranget, and D. Remy. Implicit typing a la ML for the join-calculus. Lecture Notes in Computer Science, 1243:196–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fribourg:1997:PSP

[1081] L. Fribourg and H. Olsen. Proving safety properties of infinite state systems by compilation into Presburger arithmetic. Lecture Notes in Computer Science, 1243:213–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1997:PCC

[1082] V. Gupta, R. Jagadeesan, and V. Saraswat. Probabilistic concurrent constraint programming. Lecture Notes in Computer Science, 1243:243–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harel:1997:CVC

[1083] D. Harel, O. Kupferman, and M. Y.

Vardi. On the complexity of verifying concurrent transition systems. Lecture Notes in Computer Science, 1243: 258–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1997:FS

[1084] T. A. Henzinger, O. Kupferman, and S. K. Rajamani. Fair simulation. Lecture Notes in Computer Science, 1243: 273–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wimmel:1997:ACP

[1085] H. Wimmel and L. Priese. Algebraic characterization of Petri net pomset semantics. Lecture Notes in Computer Science, 1243:406–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Appriou:1997:MDF

[1086] A. Appriou. Multisensor data fusion in situation assessment processes. Lecture Notes in Computer Science, 1244:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bell:1997:DMR

[1087] S. Bell. Dependency mining in relational databases. Lecture Notes in Computer Science, 1244:16–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benferhat:1997:SCU

[1088] S. Benferhat, D. Dubois, and H. Prade. Syntactic combination of uncertain information: A possibilistic approach. Lecture Notes in Computer Science, 1244:30–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benferhat:1997:CBA

[1089] S. Benferhat and L. Garcia. A coherence-based approach to default reasoning. Lecture Notes in Computer Science, 1244:43–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bison:1997:SAD

[1090] P. Bison, G. Chemello, C. Sossai, and G. Trainito. A syntactical approach to data fusion. Lecture Notes in Computer Science, 1244:58–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borgelt:1997:SER

[1091] C. Borgelt and R. Kruse. Some experimental results on learning probabilistic and possibilistic networks with different evaluation measures. Lecture Notes in Computer Science, 1244:71– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cholvy:1997:IFL

[1092] L. Cholvy and A. Hunter. Information fusion in logic: A brief overview. Lecture Notes in Computer Science, 1244: 86–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1997:FVB

[1093] D. Dubois and H. Prade. Focusing vs. belief revision: A fundamental distinction when dealing with generic knowledge. Lecture Notes in Computer Science, 1244:96–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gebhardt:1997:BPP

[1094] J. Gebhardt and R. Kruse. Background and perspectives of possibilistic graphical models. Lecture Notes in Computer Science, 1244:108–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Woods:1997:JHS

[1095] J. Woods. Just how stupid is postmodernism? Lecture Notes in Computer Science, 1244:154–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amgoud:1997:IPO

[1096] L. Amgoud and C. Cayrol. Integrating preference orderings into argumentbased reasoning. Lecture Notes in Computer Science, 1244:159–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anrig:1997:ABM

[1097] B. Anrig, R. Haenni, J. Kohlas, and N. Lehmann. Assumption-based modeling using ABEL. Lecture Notes in Computer Science, 1244:171–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besnard:1997:PQC

[1098] P. Besnard, J.-M. Guinnebault, and E. Mayer. Propositional quantification for conditional logic. Lecture Notes in Computer Science, 1244:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bissig:1997:FDA

[1099] R. Bissig, J. Kohlas, and N. Lehmann. Fast-division architecture for Dempster-Shafer belief functions. Lecture Notes in Computer Science, 1244:198–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delechamp:1997:GMA

[1100] J. Delechamp and B. Bouchon-Meunier. Graduality by means of analogical reasoning. Lecture Notes in Computer Science, 1244:210–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DupinDeSaint-Cyr:1997:RAU

[1101] F. Dupin De Saint-Cyr and J. Lang.

Reasoning about unpredicted change and explicit time. Lecture Notes in Computer Science, 1244:223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egly:1997:NES

[1102] U. Egly and H. Tompits. Nonelementary speed-ups in default reasoning. Lecture Notes in Computer Science, 1244:237–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1997:CRS

[1103] J. Engelfriet and J. Treur. A compositional reasoning system for executing nonmonotonic theories of reasoning. Lecture Notes in Computer Science, 1244:252–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabbay:1997:SBB

[1104] D. Gabbay and O. Rodrigues. Structured belief bases: A practical approach to prioritised base revision. Lecture Notes in Computer Science, 1244: 267–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgatos:1997:ERU

[1105] K. Georgatos. Entrenchment relations:

A uniform approach to nonmonotonicity. Lecture Notes in Computer Science, 1244:282–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heinemann:1997:MLR

[1106] B. Heinemann. A modal logic for reasoning about knowledge and time on binary subset trees. Lecture Notes in Computer Science, 1244:298–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herzig:1997:HCF

[1107] A. Herzig. How to change factual beliefs using laws and dependence information. Lecture Notes in Computer Science, 1244:311–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunter:1997:UDL

[1108] A. Hunter. Using default logic for lexical knowledge. Lecture Notes in Computer Science, 1244:322–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibargueengoytia:1997:LAA

[1109] P. H. Ibargueengoytia, S. Vadera, and L. E. Sucar. A layered, anytime approach to sensor validation. Lecture Notes in Computer Science, 1244:336– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jitnah:1997:TFA

[1110] N. Jitnah and A. Nicholson. treeNets: A framework for anytime evaluation of belief networks. Lecture Notes in Computer Science, 1244:350–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kern-Isberner:1997:LSM

[1111] G. Kern-Isberner. A logically sound method for uncertain reasoning with quantified conditionals. Lecture Notes in Computer Science, 1244:365–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voorbraak:1997:CEU

[1112] F. Voorbraak. Combining evidence under partial ignorance. Lecture Notes in Computer Science, 1244:574–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weydert:1997:RDQ

[1113] E. Weydert. Rational default quantifier logic. Lecture Notes in Computer Science, 1244:589–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:DUM

[1114] Y. Zhang. Disjunctive update, minimal change, and default reasoning. Lecture Notes in Computer Science, 1244: 600–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demolombe:1997:TUL

[1115] R. Demolombe, A. J. I. Jones, and J. Carmo. Toward a uniform logical representation of different kinds of integrity constraints. Lecture Notes in Computer Science, 1244:614–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carmo:1997:TIP

[1116] R. M. L. R. Carmo, L. R. De Carvalho, E. De Souza e. Silva, and M. C.

Diniz. TANGRAM-II: A performability modeling environment tool. Lecture Notes in Computer Science, 1245:6–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMeer:1997:PDE

[1117] H. De Meer and H. Sevcikova. PENE-LOPE: Dependability evaluation and the optimization of performability. Lecture Notes in Computer Science, 1245: 19–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciardo:1997:SAL

[1118] G. Ciardo and A. S. Miner. Storage alternatives for large structured state spaces. Lecture Notes in Computer Science, 1245:44–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deavours:1997:EDB

[1119] D. D. Deavours and W. H. Sanders. An efficient disk-based tool for solving very large Markov models. Lecture Notes in Computer Science, 1245:58–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernat:1997:ETO

[1120] G. Bernat, A. Burns, and A. Llamosi. Efficient transient overload tests for real-time systems. Lecture Notes in Computer Science, 1245:72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasslinger:1997:TAT

[1121] G. Hasslinger. Towards an analytical tool for performance modeling of ATM networks by decomposition. Lecture Notes in Computer Science, 1245:83– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conway:1997:STW

[1122] A. E. Conway, S. B. Moon, and P. Skelly. Synchronized two-way voice simulation for Internet phone performance analysis and evaluation. Lecture Notes in Computer Science, 1245: 108–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

German:1997:SPL

[1123] R. German. SPNL: Processes as language-oriented building blocks of stochastic Petri nets. Lecture Notes in Computer Science, 1245:123–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dilley:1997:MTM

[1124] J. Dilley, R. Friedrich, T. Jin, and J. Rolia. Measurement tools and modeling techniques for evaluating WEB server performance. Lecture Notes in Computer Science, 1245:155–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braun:1997:IBW

[1125] M. Braun and G. Kotsis. Interval based workload characterization for distributed systems. Lecture Notes in Computer Science, 1245:181–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fourneau:1997:BLR

[1126] J.-M. Fourneau, L. Mokdad, and N. Pekergin. Bounding the loss rates in a multistage ATM switch. Lecture Notes in Computer Science, 1245:193– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balsamo:1997:QLM

[1127] S. Balsamo and I. Mura. On queue length moments in fork and join queuing networks with general service times. Lecture Notes in Computer Science, 1245:218–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnes:1997:ARL

[1128] John G. P. Barnes. Ada 95 rationale: the language, the standard libraries. Lecture Notes in Computer Science, 1247:xii + 458, 1997. CO-DEN LNCSD9. ISBN 3-540-63143-7. ISSN 0302-9743 (print), 1611-3349 (electronic). Document to accompany the new Ada standard, officially ISO/IEC 8652: 1995.

Esparza:1997:MCL

[1129] J. Esparza and S. Melzer. Model checking LTL using constraint programming. Lecture Notes in Computer Science, 1248:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habermehl:1997:CLT

[1130] P. Habermehl. On the complexity of the linear-time μ-calculus for Petri nets. Lecture Notes in Computer Science, 1248:102–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraczak:1997:IIN

[1131] W. Fraczak and E. Pelz. Improved implementations via a new structural equivalence on labelled nets. Lecture Notes in Computer Science, 1248:117– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aura:1997:TPT

[1132] T. Aura and J. Lilius. Time processes for time Petri nets. Lecture Notes in Computer Science, 1248:136–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haddad:1997:EHP

[1133] S. Haddad, P. Moreaux, and G. Chiola. Efficient handling of phase-type distributions in generalized stochastic Petri nets. Lecture Notes in Computer Science, 1248:175–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bause:1997:APN

[1134] F. Bause. Analysis of Petri nets with a dynamic priority method. Lecture Notes in Computer Science, 1248:215– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:1997:TCP

[1135] S. Christensen and K. H. Mortensen.

Teaching coloured Petri nets — A gentle introduction to formal methods in a distributed systems course. Lecture Notes in Computer Science, 1248:290– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boubour:1997:FDT

[1136] R. Boubour and C. Jard. Fault detection in telecommunication networks based on Petri net representation of alarm propagation. Lecture Notes in Computer Science, 1248:367–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:1997:ABO

[1137] S. Christensen and J. B. Jørgensen. Analysing Bang and Olufsen’s BeoLink[R] audio/video system using coloured Petri nets. Lecture Notes in Computer Science, 1248:387–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1997:OTC

[1138] E. Best and T. Thielke. Orthogonal transformations for coloured Petri nets. Lecture Notes in Computer Science, 1248:447–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1997:CSM

[1139] W.-T. Wu. The char-set method and its applications to automated reasoning. Lecture Notes in Computer Science, 1249:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durand:1997:DCN

[1140] I. Durand and A. Middeldorp. Decidable call by need computations in term rewriting. Lecture Notes in Computer Science, 1249:4–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1997:NAC

[1141] F. Baader and C. Tinelli. A new approach for combining decision procedures for the word problem, and its connection to the Nelson-Oppen combination method. Lecture Notes in Computer Science, 1249:19–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonacina:1997:CDT

[1142] M. P. Bonacina. The clause-diffusion theorem prover peers-mcd. Lecture Notes in Computer Science, 1249:53– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahn:1997:IAI

[1143] B. I. Dahn, J. Gehne, T. Honigmann, and A. Wolf. Integration of automated and interactive theorem proving in ILF. Lecture Notes in Computer Science, 1249:57–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1997:ISP

[1144] A. Wolf and J. Schumann. ILF-SETHEO: Processing model elimination proofs for natural language output. Lecture Notes in Computer Science, 1249:61–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1997:SGS

[1145] B. Fischer and J. Schumann. SETHEO goes software engineering: Application of ATP to software reuse. Lecture Notes in Computer Science, 1249:65– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:PSA

[1146] L. Yang, H. Fu, and Z. Zeng. A practical symbolic algorithm for the inverse kinematics of 6R manipulators with simple geometry. Lecture Notes in Computer Science, 1249:73–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjoerner:1997:PIF

[1147] N. S. Bjoerner, M. E. Stickel, and T. E. Uribe. A practical integration of first-order reasoning and decision procedures. Lecture Notes in Computer Science, 1249:101–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egly:1997:SPL

[1148] U. Egly. Some pitfalls of LK-to-LJ translations and how to avoid them. Lecture Notes in Computer Science, 1249:116–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwanuma:1997:LMP

[1149] K. Iwanuma. Lemma matching for a PTTP-based top-down theorem prover. Lecture Notes in Computer Science, 1249:146–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasegawa:1997:NHM

[1150] R. Hasegawa, K. Inoue, Y. Ohta, and M. Koshimura. Non-Horn magic sets to incorporate top-down inference into bottom-up theorem proving. Lecture Notes in Computer Science, 1249:176– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harland:1997:RDB

[1151] J. Harland and D. Pym. Resourcedistribution via Boolean constraints. Lecture Notes in Computer Science, 1249:222–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cryan:1997:CNF

[1152] M. Cryan and A. Ramsay. Constructing a normal form for property theory. Lecture Notes in Computer Science, 1249:237–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benzmueller:1997:OTM

[1153] C. Benzmueller, L. Cheikhrouhou, D. Fehrer, and A. Fiedler. omegaMEGA: Towards a mathematical assistant. Lecture Notes in Computer Science, 1249: 252–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1997:CPP

[1154] D. Fuchs and M. Fuchs. CODE: A powerful prover for problems of condensed detachment. Lecture Notes in Computer Science, 1249:260–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giunchiglia:1997:NMT

[1155] F. Giunchiglia, M. Roveri, and R. Sebastiani. A new method for testing decision procedures in modal logics. Lecture Notes in Computer Science, 1249:264–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:SEP

[1156] H. Zhang. SATO: An efficient propositional prover. Lecture Notes in Computer Science, 1249:272–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dennis:1997:UGC

[1157] L. Dennis, A. Bundy, and I. Green. Using a generalisation critic to find bisimulations for coinductive proofs. Lecture Notes in Computer Science, 1249:276– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hutter:1997:CVL

[1158] D. Hutter and M. Kohlhase. A colored version of the lambda-calculus. Lecture Notes in Computer Science, 1249: 291–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganzinger:1997:STO

[1159] H. Ganzinger, C. Meyer, and C. Weidenbach. Soft typing for ordered resolution. Lecture Notes in Computer Science, 1249:321–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNivelle:1997:CNL

[1160] H. De Nivelle. A classification of nonliftable orders for resolution. Lecture Notes in Computer Science, 1249:336– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felty:1997:HIT

[1161] A. P. Felty and D. J. Howe. Hybrid interactive theorem proving using Nuprl and HOL. Lecture Notes in Computer Science, 1249:351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eastaughffe:1997:PTT

[1162] K. A. Eastaughffe, M. A. Ozols, and A. Cant. Proof tactics for a theory of state machines in a graphical environment. Lecture Notes in Computer Science, 1249:366–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonOheimb:1997:RMS

[1163] D. Von Oheimb and T. F. Gritzner.

RALL: Machine-supported proofs for relation algebra. Lecture Notes in Computer Science, 1249:380–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hickey:1997:NLI

[1164] J. J. Hickey. Nuprl-Light: An implementation framework for higher-order logics. Lecture Notes in Computer Science, 1249:395–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kettner:1997:TBS

[1165] M. Kettner and N. Eisinger. The tableau browser SNARKS. Lecture Notes in Computer Science, 1249:408– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bornat:1997:JCA

[1166] R. Bornat and B. Sufrin. Jape: A calculator for animating proof-on-paper. Lecture Notes in Computer Science, 1249:412–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1997:EC

[1167] M. Fuchs. Evolving combinators. Lecture Notes in Computer Science, 1249: 416–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Defourneaux:1997:PMA

[1168] G. Defourneaux and N. Peltier. Partial matching for analogy discovery in proofs and counter-examples. Lecture Notes in Computer Science, 1249:431– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrensberger:1997:DSD

[1169] J. Ehrensberger and C. Zinn. DIA-LOG: A system for dialogue logic. Lecture Notes in Computer Science, 1249: 446–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ceri:1997:SIM

[1170] S. Ceri and P. Fraternali. The story of the IDEA methodology. Lecture Notes in Computer Science, 1250:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wieringa:1997:ISF

[1171] R. Wieringa, E. Dubois, and S. Huyts. Integrating semi-formal and formal requirements. Lecture Notes in Computer Science, 1250:19–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz:1997:SBP

[1172] O. Diaz and N. W. Paton. Stimuli and business policies as modelling constructs: Their definition and validation through the event calculus. Lecture Notes in Computer Science, 1250: 33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filippidou:1997:USV

[1173] D. Filippidou and P. Loucopoulos. Using scenarios to validate requirements in a plausibility-centred approach. Lecture Notes in Computer Science, 1250: 47–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jajodia:1997:ISB

[1174] S. Jajodia, I. Ray, and P. Ammann. Implementing semantic-based decomposition of transactions. Lecture Notes in Computer Science, 1250:75–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castano:1997:MPF

[1175] S. Castano and V. De Antonellis. A multi-perspective framework for the analysis of legacy information systems. Lecture Notes in Computer Science, 1250:117–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carrico:1997:FAD

[1176] L. Carrico and N. Guimaraes. Facilitating analysis and diagnosis in organisations. Lecture Notes in Computer Science, 1250:131–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altmeyer:1997:AGB

[1177] J. Altmeyer, J. P. Riegel, B. Schuermann, and M. Schuetze. Application of a generator-based software development method supporting model reuse. Lecture Notes in Computer Science, 1250:159–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brataas:1997:PEH

[1178] G. Brataas, P. H. Hughes, and A. Sølvberg. Performance engineering of human and computerised workflows. Lecture Notes in Computer Science, 1250:187–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ebert:1997:MCP

[1179] J. Ebert, R. Suettenbach, and I. Uhe. Meta-CASE in practice: A case for KOGGE. Lecture Notes in Computer Science, 1250:203–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaipala:1997:ACT

[1180] J. Kaipala. Augmenting CASE tools with hypertext: Desired functionality and implementation issues. Lecture Notes in Computer Science, 1250:217– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

FalcaoGoncalves:1997:DNP

[1181] P. Falcao Goncalves, A. C. Salgado, and S. Lemos Meira. Digital neighbourhoods: Partitioning the Web for information indexing and searching. Lecture Notes in Computer Science, 1250: 289–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dominguez:1997:CAM

[1182] E. Dominguez, M. A. Zapata, and J. Rubio. A conceptual approach to meta-modelling. Lecture Notes in Computer Science, 1250:319–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cathala:1997:PBW

[1183] F. Cathala and P. Poncelet. Preserving behaviour: Why and how. Lecture Notes in Computer Science, 1250:333– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bichler:1997:WT

[1184] P. Bichler, G. Preuner, and M. Schrefl.

Workflow transparency. Lecture Notes in Computer Science, 1250:423–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barros:1997:TRS

[1185] A. P. Barros, A. H. M. Ter Hofstede, and H. A. Proper. Towards real-scale business transaction workflow modelling. Lecture Notes in Computer Science, 1250:437–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzmueller:1997:FUA

[1186] B. Holzmueller and E. Ploeedereder. Finite unions for Ada 95. Lecture Notes in Computer Science, 1251:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:1997:ABC

[1187] T. P. Baker and D. I. Oh. Ada bindings for C interfaces: Lessons learned from the Florist Implementation. Lecture Notes in Computer Science, 1251:13–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldman:1997:ASR

[1188] M. B. Feldman. An Ada 95 sort race construction set. Lecture Notes in Computer Science, 1251:23–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooling:1997:CAT

[1189] J. E. Cooling, N. Duff, and J. Cooling. Computer aided teaching of programming languages: An Ada-specific development. Lecture Notes in Computer Science, 1251:35–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dupouy:1997:SP

[1190] B. Dupouy, O. Hainque, L. Pautet, and S. Tardieu. The SPIF project. Lecture Notes in Computer Science, 1251: 46–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jovanovic:1997:DSC

[1191] S. S. Jovanovic, D. M. Sotirovski, and P. Van Aswegen. Developing scripting capabilities for an Ada-based simulator. Lecture Notes in Computer Science, 1251:56–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wegener:1997:SUT

[1192] J. Wegener and I. Fey. Systematic unit-testing of Ada programs. Lecture Notes in Computer Science, 1251:64– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franch:1997:INF

[1193] X. Franch. Including non-functional issues in Anna/Ada programs for automatic implementation selection. Lecture Notes in Computer Science, 1251: 88–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yeung:1997:SBS

[1194] W. L. Yeung. Semantics-based support tools for high integrity Ada software. Lecture Notes in Computer Science, 1251:100–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bell:1997:ATA

[1195] A. E. Bell. An alternative toolset for analysis of Ada programs. Lecture Notes in Computer Science, 1251:112– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GonzalezHarbour:1997:IAL

[1196] M. Gonzalez Harbour, J. J. Gutierrez Garcia, and J. C. Palencia Gutierrez. Implementing application-level sporadic server schedulers in Ada 95. Lecture Notes in Computer Science, 1251:125–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wellings:1997:TTA

[1197] A. J. Wellings, A. Burns, and O. Pazy. Task termination in Ada 95. Lecture Notes in Computer Science, 1251:149– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gellerich:1997:PIA

[1198] W. Gellerich and E. Ploeedereder. Parameter-induced aliasing and related problems can be avoided. Lecture Notes in Computer Science, 1251:161– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheung:1997:TIS

[1199] K. S. Cheung, K. O. Chow, and T. Y. Cheung. Towards an integration of syntactic constructs and structural features for formalised objectoriented methods. Lecture Notes in Computer Science, 1251:173–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

English:1997:DI

[1200] J. English. The dangers of inheritance. Lecture Notes in Computer Science, 1251:185–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Celier:1997:CDC

[1201] V. M. Celier, D. Sotirovski, and C. J.

Thompson. Code-data consistency in Ada. Lecture Notes in Computer Science, 1251:209–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerra:1997:ALP

[1202] F. Guerra, J. Miranda, A. Alvarez, and S. Arevalo. An Ada library to program fault-tolerant distributed applications. Lecture Notes in Computer Science, 1251:230–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobbing:1997:STS

[1203] B. Dobbing and M. Richard-Foy. Tsmart — task-safe, minimal Ada realtime toolset. Lecture Notes in Computer Science, 1251:244–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waterman:1997:TTA

[1204] S. R. Waterman. Techniques for testing Ada 95. Lecture Notes in Computer Science, 1251:278–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weickert:1997:RND

[1205] J. Weickert. A review of nonlinear diffusion filtering. Lecture Notes in Computer Science, 1252:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caselles:1997:SSV

[1206] V. Caselles, B. Coll, and J.-M Morel.

Scale-space versus topographic map for natural images. Lecture Notes in Computer Science, 1252:29–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalitzin:1997:IOB

[1207] S. N. Kalitzin, B. M. Ter Haar Romeny, and M. A. Viergever. Invertible orientation bundles on 2D scalar images. Lecture Notes in Computer Science, 1252:77–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bretzner:1997:HST

[1208] L. Bretzner and T. Lindeberg. On the handling of spatial and temporal scales in feature tracking. Lecture Notes in Computer Science, 1252:128–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fidrich:1997:FFL

[1209] M. Fidrich. Following feature lines across scale. Lecture Notes in Computer Science, 1252:140–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckman:1997:SDP

[1210] F. J. Beckman, E. T. P. Slijpen, and W. J. Niessen. Supervised diffusion parameter selection for filtering SPECT brain images. Lecture Notes in Computer Science, 1252:164–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furst:1997:ILR

[1211] J. D. Furst, R. S. Keller, J. E. Miller, and S. M. Pizer. Image loci are ridges in geometric spaces. Lecture Notes in Computer Science, 1252:176–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coulon:1997:MML

[1212] O. Coulon, I. Bloch, V. Frouin, and J.- F. Mangin. Multiscale measures in linear scale-space for characterizing cerebral functional activations in 3D PET difference images. Lecture Notes in Computer Science, 1252:188–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arridge:1997:MSP

[1213] S. R. Arridge and A. Simmons. Multispectral probabilistic diffusion using Bayesian classification. Lecture Notes in Computer Science, 1252:224–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weickert:1997:RSS

[1214] J. Weickert. Recursive separable schemes for nonlinear diffusion filters. Lecture Notes in Computer Science, 1252:260–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faugeras:1997:LSM

[1215] O. Faugeras and R. Keriven. Level set methods and the stereo problem. Lecture Notes in Computer Science, 1252:272–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abbasi:1997:RCC

[1216] S. Abbasi, F. Mokhtarian, and J. Kittler. Reliable classification of chrysanthemum leaves through curvature scale space. Lecture Notes in Computer Science, 1252:284–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berkner:1997:RSS

[1217] K. Berkner. Reconstruction of selfsimilar functions from scale-space. Lecture Notes in Computer Science, 1252: 311–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hontani:1997:MSD

[1218] H. Hontani and K. Deguchi. Multiscale detection of characteristic figure structures using principal curvatures of image gray-level profile. Lecture Notes in Computer Science, 1252:315– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bajcsy:1997:NFH

[1219] P. Bajcsy and N. Ahuja. A new framework for hierarchical segmentation using similarity analysis. Lecture Notes in Computer Science, 1252:319–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Black:1997:RAD

[1220] M. J. Black, G. Sapiro, D. Marimont, and D. Heeger. Robust anisotropic diffusion: connections between robust statistics, line processing, and anisotropic diffusion. Lecture Notes in Computer Science, 1252:323–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischl:1997:FAA

[1221] B. Fischl and E. L. Schwartz. Fast adaptive alternatives to nonlinear diffusion in image enhancement: Green’s function approximators and nonlocal filters. Lecture Notes in Computer Science, 1252:327–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forte:1997:SSA

[1222] P. Forte and D. Greenhill. A scalespace approach to shape similarity. Lecture Notes in Computer Science, 1252:333–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harvey:1997:SSF

[1223] R. Harvey, J. A. Bangham, and A. Bosson. Scale-space filters and their robustness. Lecture Notes in Computer Science, 1252:341–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imiya:1997:ESF

[1224] A. Imiya and R. Katsuta. Extraction of a structure feature from threedimensional objects by scale-space analysis. Lecture Notes in Computer Science, 1252:353–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kacur:1997:SAD

[1225] J. Kacur and K. Mikula. Slowed anisotropic diffusion. Lecture Notes in Computer Science, 1252:357–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armande:1997:TNE

[1226] N. Armande, P. Montesinos, and O. Monga. Thin nets extraction using a multi-scale approach. Lecture Notes in Computer Science, 1252:361–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Argueello:1997:PMB

[1227] F. Argueello, N. Guil, J. Lopez, and M. Amor. A probabilistic model for best-first search B and B algorithms. Lecture Notes in Computer Science, 1253:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buecker:1997:VBG

[1228] H. M. Buecker and M. Sauren. A variant of the biconjugate gradient method suitable for massively parallel computing. Lecture Notes in Computer Science, 1253:72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:EIIa

[1229] T. Yang and H.-X. Lin. Efficient implementation of the improved quasiminimal residual method on massively distributed memory computers. Lecture Notes in Computer Science, 1253: 80–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobson:1997:PSD

[1230] S. Dobson and D. Goodeve. Programming with shared data abstractions. Lecture Notes in Computer Science, 1253:93–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diekmann:1997:EDL

[1231] R. Diekmann, S. Muthukrishnan, and M. V. Nayakkankuppam. Engineering diffusive load balancing algorithms using experiments. Lecture Notes in Computer Science, 1253:111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1997:NAA

[1232] K. Jansen and J. Reiter. A new approximation algorithm for the register allocation problem. Lecture Notes in Computer Science, 1253:135–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Decker:1997:VDS

[1233] T. Decker. Virtual data space — A universal load balancing scheme. Lecture Notes in Computer Science, 1253: 159–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allmaier:1997:PSM

[1234] S. C. Allmaier and G. Horton. Parallel shared-memory state-space exploration in stochastic modeling. Lecture Notes in Computer Science, 1253:207– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Formella:1997:PSC

[1235] A. Formella and J. Keller. Parallel software caches. Lecture Notes in Computer Science, 1253:219–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeumker:1997:CEP

[1236] A. Baeumker and F. Meyer auf der Heide. Communication efficient parallel searching. Lecture Notes in Computer Science, 1253:233–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hughes:1997:FVD

[1237] R. B. Hughes. Formal verification of digital systems, from ASICs to HW/SW codesign — a pragmatic approach. Lecture Notes in Computer Science, 1254:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boraelv:1997:ISV

[1238] A. Boraelv. The industrial success of verification tools based on Stalmarck’s method. Lecture Notes in Computer Science, 1254:7–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaivola:1997:UCP

[1239] R. Kaivola. Using compositional preorders in the verification of sliding window protocol. Lecture Notes in Computer Science, 1254:48–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cyrluk:1997:EDP

[1240] D. Cyrluk, O. Moeller, and H. Ruess. An efficient decision procedure for the theory of fixed-sized bit-vectors. Lecture Notes in Computer Science, 1254: 60–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1997:CAS

[1241] S. Graf and H. Saidi. Construction of abstract state graphs with PVS. Lecture Notes in Computer Science, 1254:72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamhi:1997:ADE

[1242] G. Kamhi, O. Weissberg, L. Fix, and Z. Binyamini. Automatic datapath extraction for efficient usage of HDD. Lecture Notes in Computer Science, 1254:95–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baier:1997:WBF

[1243] C. Baier and H. Hermanns. Weak bisimulation for fully probabilistic processes. Lecture Notes in Computer Science, 1254:119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolignano:1997:TMC

[1244] D. Bolignano. Towards a mechanization of cryptographic protocol verification. Lecture Notes in Computer Science, 1254:131–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisler:1997:CRL

[1245] K. Fisler. Containment of regular languages in non-regular timing diagram languages is decidable. Lecture Notes in Computer Science, 1254:155–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boigelot:1997:IRA

[1246] B. Boigelot, L. Bronne, and S. Rassart. An improved reachability analysis method for strongly linear hybrid systems. Lecture Notes in Computer Science, 1254:167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bozga:1997:SPS

[1247] M. Bozga, O. Maler, A. Pnueli, and S. Yovine. Some progress in the symbolic verification of timed automata. Lecture Notes in Computer Science, 1254:179–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cimatti:1997:PCE

[1248] A. Cimatti, F. Giunchiglia, P. Pecchiari, and B. Pietra. A provably correct embedded verifier for the certification of safety critical software. Lecture Notes in Computer Science, 1254: 202–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrett:1997:MCM

[1249] G. Barrett and A. McIssac. Model checking in a microprocessor design project. Lecture Notes in Computer Science, 1254:214–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harel:1997:STS

[1250] D. Harel. Some thoughts on statecharts, 13 years later. Lecture Notes in Computer Science, 1254:226–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gyuris:1997:FMC

[1251] V. Gyuris and A. P. Sistla. On-the-fly model checking under fairness that exploits symmetry. Lecture Notes in Computer Science, 1254:232–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hardin:1997:NHB

[1252] R. H. Hardin, R. P. Kurshan, S. K.

Shukla, and M. Y. Vardi. A new heuristic for bad cycle detection using BDDs. Lecture Notes in Computer Science, 1254:268–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beer:1997:EDV

[1253] I. Beer, S. Ben-David, U. Eisner, and Y. Rodeh. Efficient detection of vacuity in ACTL formulas. Lecture Notes in Computer Science, 1254:279–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Immerman:1997:MCT

[1254] N. Immerman and M. Y. Vardi. Model checking and transitive-closure logic. Lecture Notes in Computer Science, 1254:291–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berry:1997:BAN

[1255] G. Berry. Boolean and 2-adic numbers based techniques for verifying synchronous designs. Lecture Notes in Computer Science, 1254:303–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cece:1997:PQS

[1256] G. Cece and A. Finkel. Programs with quasi-stable channels are effectively recognizable. Lecture Notes in Computer Science, 1254:304–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1997:CCS

[1257] W. Chan, R. Anderson, P. Beame, and D. Notkin. Combining constraint solving and symbolic model checking for a class of systems with non-linear constraints. Lecture Notes in Computer Science, 1254:316–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1997:POR

[1258] R. Alur, R. K. Brayton, T. A. Henzinger, and S. Qadeer. Partial-order reduction in symbolic state space exploration. Lecture Notes in Computer Science, 1254:340–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuan:1997:CFI

[1259] J. Yuan, J. Shen, J. Abraham, and A. Aziz. On combining formal and informal verification. Lecture Notes in Computer Science, 1254:376–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bultan:1997:SMC

[1260] T. Bultan, R. Gerber, and W. Pugh. Symbolic model checking of infinite state systems using Presburger arithmetic. Lecture Notes in Computer Science, 1254:400–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kesten:1997:SMC

[1261] Y. Kesten, O. Maler, M. Marcus, and A. Pnueli. Symbolic model checking with rich assertional languages. Lecture Notes in Computer Science, 1254: 424–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grahlmann:1997:PT

[1262] B. Grahlmann. The PEP tool. Lecture Notes in Computer Science, 1254: 440–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelb:1997:MSE

[1263] P. Kelb, T. Margaria, M. Mendler, and C. Gsottberger. MOSEL: A sound and efficient tool for M2L(Str). Lecture Notes in Computer Science, 1254:448– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campos:1997:VTQ

[1264] S. Campos, E. Clarke, and M. Minea.

The Verus tool: A quantitative approach to the formal verification of real-time systems. Lecture Notes in Computer Science, 1254:452–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1997:HMC

[1265] T. A. Henzinger, P.-H. Ho, and H. Wong-Toi. HYTECH: A model checker for hybrid systems. Lecture Notes in Computer Science, 1254:460– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biere:1997:MEM

[1266] A. Biere. mucke — efficient mucalculus model checking. Lecture Notes in Computer Science, 1254:468–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godefroid:1997:VTA

[1267] P. Godefroid. VeriSoft: A tool for the automatic analysis of concurrent reactive software. Lecture Notes in Computer Science, 1254:476–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beer:1997:RMC

[1268] I. Beer, S. Ben-David, C. Eisner, and D. Geist. RuleBase: Model checking at IBM. Lecture Notes in Computer Science, 1254:480–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barg:1997:MDD

[1269] A. Barg. Minimum distance decoding algorithms for linear codes. Lecture Notes in Computer Science, 1255:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernardin:1997:EMF

[1270] L. Bernardin and M. B. Monagan. Efficient multivariate factorization over finite fields. Lecture Notes in Computer Science, 1255:15–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyvalenkov:1997:MCP

[1271] P. Boyvalenkov and D. Danev. On maximal codes in polynomial metric spaces. Lecture Notes in Computer Science, 1255:29–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caboara:1997:YAI

[1272] M. Caboara, P. Conti, and C. Traverso. Yet another ideal decomposition algorithm. Lecture Notes in Computer Science, 1255:39–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

CarraFerro:1997:RTO

[1273] G. Carra Ferro. A resultant theory for ordinary algebraic differential equations. Lecture Notes in Computer Science, 1255:55–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costa:1997:SGZ

[1274] S. R. Costa, J. R. Geronimo, R. Palazzo, and J. C. Interlando. The symmetry group of Z0nq in the Lee space and the Zqn-linear codes. Lecture Notes in Computer Science, 1255: 66–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Encheva:1997:RRC

[1275] S. Encheva. On repeated-root cyclic codes and the two-way chain condition. Lecture Notes in Computer Science, 1255:78–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonzurGathen:1997:EFF

[1276] J. Von zur Gathen and M. Noecker.

Exponentiation in finite fields: Theory and practice. Lecture Notes in Computer Science, 1255:88–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grigoriev:1997:CML

[1277] D. Grigoriev and A. Slissenko. Computing minimum-link path in a homotopy class amidst semi-algebraic obstacles in the plane. Lecture Notes in Computer Science, 1255:114–129, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gulliver:1997:CSD

[1278] T. A. Gulliver and M. Harada. Certain self-dual codes over Z4 and the odd leech lattice. Lecture Notes in Computer Science, 1255:130–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoeholdt:1997:OFE

[1279] T. Hoeholdt, J. H. Van Lint, and R. Pellikaan. Order functions and evaluation codes. Lecture Notes in Computer Science, 1255:138–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horadam:1997:CC

[1280] K. J. Horadam and A. A. I. Perera. Codes from cocycles. Lecture Notes in Computer Science, 1255:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaffe:1997:BTS

[1281] D. B. Jaffe. A brief tour of split linear programming. Lecture Notes in Computer Science, 1255:164–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jean:1997:EAE

[1282] F. Jean and P.-V. Koseleff. Elementary approximation of exponentials of Lie polynomials. Lecture Notes in Computer Science, 1255:174–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jensen:1997:AIC

[1283] J. M. Jensen. On the q-ary image of cyclic codes. Lecture Notes in Computer Science, 1255:189–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasami:1997:SWE

[1284] T. Kasami, T. Sugita, and T. Fujiwara. The split weight (wL, wR) enumeration of Reed-Müller codes for wL + wR, 2dmin. Lecture Notes in Computer Science, 1255:197–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wood:1997:ETL

[1285] J. A. Wood. Extension theorems for linear codes over finite rings. Lecture Notes in Computer Science, 1255:329– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xia:1997:NCE

[1286] S. Xia and F. Fu. Nonperiodic cyclic equivalence classes of cyclic codes and algebraic constructions of cyclically permutable codes. Lecture Notes in Computer Science, 1255:341–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carton:1997:WWH

[1287] O. Carton and D. Perrin. The Wadge-Wegner hierarchy of omega-rational sets. Lecture Notes in Computer Science, 1256:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apt:1997:CIC

[1288] K. R. Apt. From chaotic iteration to constraint propagation. Lecture Notes in Computer Science, 1256:36– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durand:1997:TQ

[1289] B. Durand. Tilings and quasiperiodicity. Lecture Notes in Computer Science, 1256:65–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bassino:1997:ESL

[1290] F. Bassino, M.-P. Beal, and D. Perrin. Enumerative sequences of leaves in rational trees. Lecture Notes in Computer Science, 1256:76–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruyere:1997:CAC

[1291] V. Bruyere. A completion algorithm for codes with bounded synchronization delay. Lecture Notes in Computer Science, 1256:87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karhumaeki:1997:ELR

[1292] J. Karhumaeki, W. Plandowski, and F. Mignosi. The expressibility of languages and relations by word equations. Lecture Notes in Computer Science, 1256:98–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beaudry:1997:FLR

[1293] M. Beaudry, F. Lemieux, and D. Therien. Finite loops recognize exactly the regular open languages. Lecture Notes in Computer Science, 1256:110– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiGianantonio:1997:ADT

[1294] P. Di Gianantonio. An abstract data type for real numbers. Lecture Notes in Computer Science, 1256:121–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bournez:1997:SBC

[1295] O. Bournez. Some bounds on the computational power of piecewise constant derivative systems. Lecture Notes in Computer Science, 1256:143–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gurevich:1997:MSR

[1296] Y. Gurevich and A. Voronkov. Monadic simultaneous rigid E-unification and related problems. Lecture Notes in Computer Science, 1256:154–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weihrauch:1997:CPM

[1297] K. Weihrauch. Computability on the probability measures on the Borel sets of the unit interval. Lecture Notes in Computer Science, 1256:166–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreev:1997:WCH

[1298] A. A. Andreev, A. E. F. Clementi, and J. D. P. Rolim. Worst-case hardness suffices for derandomization: A new method for hardness-randomness trade-offs. Lecture Notes in Computer Science, 1256:177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buhrman:1997:RRB

[1299] H. Buhrman, S. Fenner, and L. Fortnow. Results on resource-bounded measure. Lecture Notes in Computer Science, 1256:188–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ablayev:1997:RNI

[1300] F. Ablayev. Randomization and nondeterminism are incomparable for ordered read-once branching programs. Lecture Notes in Computer Science, 1256:195– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codenotti:1997:CPP

[1301] B. Codenotti, F. Erguen, P. S. Gemmell, and S. Ravi Kumar. Checking properties of polynomials. Lecture Notes in Computer Science, 1256:203– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemaspaandra:1997:EAD

[1302] E. Hemaspaandra, L. A. Hemaspaandra, and J. Rothe. Exact analysis of Dodgson elections: Lewis Carroll’s 1876 voting system is complete for parallel access to NP. Lecture Notes in Computer Science, 1256:214–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honda:1997:GTA

[1303] K. Honda and N. Yoshida. Game theoretic analysis of call-by-value computation. Lecture Notes in Computer Science, 1256:225–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiCosmo:1997:MPH

[1304] R. Di Cosmo and N. Ghani. On modular properties of higher order extensional lambda-calculi. Lecture Notes in Computer Science, 1256:237–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asperti:1997:DSG

[1305] A. Asperti and C. Laneve. On the dynamics of sharing graphs. Lecture Notes in Computer Science, 1256:259– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alstrup:1997:MDD

[1306] S. Alstrup, J. Holm, K. De Lichtenberg, and M. Thorup. Minimizing diameters of dynamic trees. Lecture Notes in Computer Science, 1256:270–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagerup:1997:DAG

[1307] T. Hagerup. Dynamic algorithms for graphs of bounded treewidth. Lecture Notes in Computer Science, 1256:292– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fu:1997:PTA

[1308] Y. Fu. A proof theoretical approach to communication. Lecture Notes in Computer Science, 1256:325–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diekert:1997:STE

[1309] V. Diekert, Y. Matiyasevich, and A. Muscholl. Solving trace equations using lexicographical normal forms. Lecture Notes in Computer Science, 1256:336–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilke:1997:SFP

[1310] T. Wilke. Star-free picture expressions are strictly weaker than firstorder logic. Lecture Notes in Computer Science, 1256:347–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernardo:1997:ABM

[1311] M. Bernardo. An algebra-based method to associate rewards with EMPA terms. Lecture Notes in Computer Science, 1256:358–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caceres:1997:EPG

[1312] E. Caceres, F. Dehne, A. Ferreira, and P. Flocchini. Efficient parallel graph algorithms for coarse grained multicomputers and BSP. Lecture Notes in Computer Science, 1256:390–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambainis:1997:UBC

[1313] A. Ambainis. Upper bound on the communication complexity of private information retrieval. Lecture Notes in Computer Science, 1256:401–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harel:1997:CPL

[1314] D. Harel and E. Singerman. Computation paths logic: An expressive, yet elementary, process logic. Lecture Notes in Computer Science, 1256:408– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkart:1997:MCF

[1315] O. Burkart and B. Steffen. Model checking the full modal mu-calculus for infinite sequential processes. Lecture Notes in Computer Science, 1256:419– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baier:1997:SMC

[1316] C. Baier, E. M. Clarke, V. Hartonas-Garmhausen, and M. Kwiatkowska. Symbolic model checking for probabilistic processes. Lecture Notes in Computer Science, 1256:430–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeVink:1997:BPT

[1317] E. P. De Vink and J. J. M. M. Rutten. Bisimulation for probabilistic transition systems: A coalgebraic approach. Lecture Notes in Computer Science, 1256:460–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boreale:1997:BOP

[1318] M. Boreale, R. De Nicola, and R. Pugliese. Basic observables for processes. Lecture Notes in Computer Science, 1256:482–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaklamanis:1997:CBE

[1319] C. Kaklamanis, P. Persiano, T. Erlebach, and K. Jansen. Constrained bipartite edge coloring with applications to wavelength routing. Lecture Notes in Computer Science, 1256:493–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gargano:1997:CPD

[1320] L. Gargano, P. Hell, and S. Perennes. Colouring paths in directed symmetric trees with applications to WDM routing. Lecture Notes in Computer Science, 1256:505–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bartal:1997:LRA

[1321] Y. Bartal and S. Leonardi. On-line routing in all-optical networks. Lecture Notes in Computer Science, 1256: 516–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eilam:1997:CCP

[1322] T. Eilam, M. Flammini, and S. Zaks. A complete characterization of the path layout construction problem for ATM networks with given hop count and load. Lecture Notes in Computer Science, 1256:527–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jancar:1997:BED

[1323] P. Jancar. Bisimulation equivalence is decidable for one-counter processes. Lecture Notes in Computer Science, 1256:549–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1997:SRA

[1324] A. Bouajjani and P. Habermehl. Symbolic reachability analysis of FIFO-Channel systems with nonregular sets of configurations. Lecture Notes in Computer Science, 1256:560–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fokkink:1997:APL

[1325] W. Fokkink. Axiomatizations for the perpetual loop in process algebra. Lecture Notes in Computer Science, 1256:571–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1997:DTC

[1326] T. A. Henzinger and P. W. Kopke.

Discrete-time control for rectangular hybrid automata. Lecture Notes in Computer Science, 1256:582–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grossi:1997:ESM

[1327] R. Grossi and G. F. Italiano. Efficient splitting and merging algorithms for order decomposable problems. Lecture Notes in Computer Science, 1256:605– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khanna:1997:EAP

[1328] S. Khanna, S. Muthukrishnan, and S. Skiena. Efficient array partitioning. Lecture Notes in Computer Science, 1256:616–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1997:CLT

[1329] H. L. Bodlaender and D. M. Thilikos. Constructive linear time algorithms for branchwidth. Lecture Notes in Computer Science, 1256:627–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khasidashvili:1997:GOR

[1330] Z. Khasidashvili and J. Glauert. The geometry of orthogonal reduction spaces. Lecture Notes in Computer Science, 1256:649–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Droste:1997:RRF

[1331] M. Droste and P. Gastin. On recognizable and rational formal power series in partially commuting variables. Lecture Notes in Computer Science, 1256:682–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cassaigne:1997:CJS

[1332] J. Cassaigne. On a conjecture of J. shallit. Lecture Notes in Computer Science, 1256:693–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frankel:1997:CEE

[1333] Y. Frankel and M. Yung. On characterizations of escrow encryption schemes. Lecture Notes in Computer Science, 1256:705–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSantis:1997:REN

[1334] A. De Santis, G. Di Crescenzo, and P. Persiano. Randomness-efficient noninteractive zero knowledge. Lecture Notes in Computer Science, 1256:716– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1997:ARO

[1335] K. Jansen. Approximation results for the optimum cost chromatic partition problem. Lecture Notes in Computer Science, 1256:727–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bar-Noy:1997:MCS

[1336] A. Bar-Noy and G. Kortsarz. The minimum color sum of bipartite graphs. Lecture Notes in Computer Science, 1256:738–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujito:1997:PDA

[1337] T. Fujito. A primal-dual approach to approximation of node-deletion problems for matroidal properties. Lecture Notes in Computer Science, 1256:749– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broersma:1997:ISA

[1338] H. Broersma, T. Kloks, D. Kratsch, and H. Mueller. Independent sets in asteroidal triple-free graphs. Lecture Notes in Computer Science, 1256:760– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giacobazzi:1997:RCA

[1339] R. Giacobazzi and F. Ranzato. Refining and compressing abstract domains. Lecture Notes in Computer Science, 1256:771–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dami:1997:LRR

[1340] L. Dami. Labelled reductions, runtime errors, and operational subsumption. Lecture Notes in Computer Science, 1256:782–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kabanets:1997:RED

[1341] V. Kabanets. Recognizability equals definability for partial k-paths. Lecture Notes in Computer Science, 1256: 805–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beigel:1997:MCB

[1342] R. Beigel and B. Fu. Molecular computing, bounded nondeterminism, and efficient recursion. Lecture Notes in Computer Science, 1256:816–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erdos:1997:CBT

[1343] P. L. Erdos, M. A. Steel, L. A. Szekely, and T. J. Warnow. Constructing big trees from short sequences. Lecture Notes in Computer Science, 1256:827– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buccafurri:1997:EPU

[1344] F. Buccafurri, S. Greco, and D. Sacca.

The expressive power of unique total stable model semantics. Lecture Notes in Computer Science, 1256:849– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeman:1997:PG

[1345] J. Zeman. Peirce’s graphs. Lecture Notes in Computer Science, 1257:12– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chein:1997:CPC

[1346] M. Chein. The CORALI project: From conceptual graphs to conceptual graphs via labelled graphs. Lecture Notes in Computer Science, 1257:65–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chein:1997:PNC

[1347] M. Chein and M.-L. Mugnier. Positive nested conceptual graphs. Lecture Notes in Computer Science, 1257:95– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wermelinger:1997:DPC

[1348] M. Wermelinger. A different perspective on canonicity. Lecture Notes in Computer Science, 1257:110–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faron:1997:RDE

[1349] C. Faron and J.-B. Ganascia. Representation of defaults and exceptions in conceptual graphs formalism. Lecture Notes in Computer Science, 1257:153– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Angelova:1997:TDA

[1350] G. Angelova and K. Bontcheva. Taskdependent aspects of knowledge acquisition: A case study in a technical domain. Lecture Notes in Computer Science, 1257:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dick:1997:MCE

[1351] J. P. Dick. Modeling cause and effect in legal text. Lecture Notes in Computer Science, 1257:244–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bos:1997:MSH

[1352] C. Bos, B. Botella, and P. Vanheeghe. Modeling and simulating human behaviors with conceptual graphs. Lecture Notes in Computer Science, 1257: 275–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wille:1997:CGF

[1353] R. Wille. Conceptual graphs and formal concept analysis. Lecture Notes in Computer Science, 1257:290–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biedermann:1997:HTD

[1354] K. Biedermann. How triadic diagrams represent conceptual structures. Lecture Notes in Computer Science, 1257: 304–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellis:1997:OKU

[1355] G. Ellis and S. Callaghan. Organization of knowledge using order factors. Lecture Notes in Computer Science, 1257:342–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerdiles:1997:SCC

[1356] G. Kerdiles and E. Salvat. A sound and complete CG proof procedure combining projections with analytic tableaux. Lecture Notes in Computer Science, 1257:371–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cao:1997:FUR

[1357] T. H. Cao, P. N. Creasy, and V. Wuwongse. Fuzzy unification and resolution proof procedure for fuzzy conceptual graph programs. Lecture Notes in Computer Science, 1257:386–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cao:1997:UMF

[1358] T. H. Cao and P. N. Creasy. Universal marker and functional relation: Semantics and operations. Lecture Notes in Computer Science, 1257:416–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bournaud:1997:ADK

[1359] I. Bournaud and J.-G. Ganascia. Accounting for domain knowledge in the construction of a generalization space. Lecture Notes in Computer Science, 1257:446–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerbe:1997:CGC

[1360] O. Gerbe. Conceptual graphs for corporate knowledge repositories. Lecture Notes in Computer Science, 1257:474– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Genest:1997:EDR

[1361] D. Genest and M. Chein. An experiment in document retrieval using conceptual graphs. Lecture Notes in Computer Science, 1257:489–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keeler:1997:PTP

[1362] M. A. Keeler, L. F. Searle, and C. Kloesel. PORT: A testbed paradigm for knowledge processing in the humanities. Lecture Notes in Computer Science, 1257:505–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clark:1997:UAP

[1363] P. Clark and B. Porter. Using access paths to guide inference with conceptual graphs. Lecture Notes in Computer Science, 1257:521–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMoor:1997:ACG

[1364] A. De Moor. Applying conceptual graph theory to the user-driven specification of network information systems. Lecture Notes in Computer Science, 1257:536–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garner:1997:DTC

[1365] B. Garner, E. Tsui, and D. Lukose.

Deakin toolset: Conceptual graphs based knowledge acquisition, management, and processing tools. Lecture Notes in Computer Science, 1257:589– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Angelova:1997:MBI

[1366] G. Angelova, S. Damyanova, K. Toutanova, and K. Bontcheva. Menu-based interfaces to conceptual graphs: The CGLex approach. Lecture Notes in Computer Science, 1257:603–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cyre:1997:KET

[1367] W. R. Cyre. Knowledge extractor:

A tool for extracting knowledge from text. Lecture Notes in Computer Science, 1257:607–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1997:CML

[1368] G. Fuchs and R. Levinson. The CG Mars lander. Lecture Notes in Computer Science, 1257:611–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1997:POT

[1369] R. P. Wolf and H. S. Delugach. PCCG: An operational tracked grid for creating conceptual graphs. Lecture Notes in Computer Science, 1257:617–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arieli:1997:FVD

[1370] O. Arieli and A. Avron. Four-valued diagnoses for stratified knowledge-bases. Lecture Notes in Computer Science, 1258:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1997:FCE

[1371] M. Baaz and A. Leitsch. Fast cutelimination by projection. Lecture Notes in Computer Science, 1258:18– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1997:SRP

[1372] G. Barthe and P.-A. Mellies. On the subject reduction property for algebraic type systems. Lecture Notes in Computer Science, 1258:34–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattani:1997:PMCa

[1373] G. L. Cattani and G. Winskel. Presheaf models for concurrency. Lecture Notes in Computer Science, 1258:58– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Danos:1997:DVR

[1374] V. Danos, M. Pedicini, and L. Regnier. Directed virtual reductions. Lecture Notes in Computer Science, 1258: 76–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dziembowski:1997:BVF

[1375] S. Dziembowski. Bounded-variable fixpoint queries are PSPACE-complete. Lecture Notes in Computer Science, 1258:89–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiore:1997:EMA

[1376] M. P. Fiore and G. D. Plotkin. An extension of models of axiomatic domain theory to models of synthetic domain theory. Lecture Notes in Computer Science, 1258:129–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geser:1997:RUTb

[1377] A. Geser, A. Middeldorp, E. Ohlebusch, and H. Zantema. Relative undecidability in term rewriting. Lecture Notes in Computer Science, 1258:150– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geuvers:1997:EMS

[1378] H. Geuvers. Extending models of second order predicate logic to models of second order dependent type theory. Lecture Notes in Computer Science, 1258:167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghani:1997:EEF

[1379] N. Ghani. Eta-expansions in F0omega. Lecture Notes in Computer Science, 1258:182–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gore:1997:CFD

[1380] R. Gore. Cut-free display calculi for relation algebras. Lecture Notes in Computer Science, 1258:198–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imhof:1997:CAA

[1381] H. Imhof. Computational aspects of arity hierarchies. Lecture Notes in Computer Science, 1258:211–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ingolfsdottir:1997:WSB

[1382] A. Ingolfsdottir. Weak semantics based on lighted button pressing experiments. Lecture Notes in Computer Science, 1258:226–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yashin:1997:NIL

[1383] A. D. Yashin. New intuitionistic logical constants: Undecidability of the conservativeness problem. Lecture Notes in Computer Science, 1258:460–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aleksander:1997:ILN

[1384] I. Aleksander. Iconic learning in networks of logical neurons. Lecture Notes in Computer Science, 1259:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armstrong:1997:HRF

[1385] W. W. Armstrong. Hardware requirements for fast evaluation of functions learned by adaptive logic networks. Lecture Notes in Computer Science, 1259:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1997:PCE

[1386] X. Yao and T. Higuchi. Promises and challenges of evolvable hardware. Lecture Notes in Computer Science, 1259: 55–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goeke:1997:OAE

[1387] M. Goeke, M. Sipper, D. Mange, and A. Stauffer. Online autonomous evolware. Lecture Notes in Computer Science, 1259:96–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keymeulen:1997:ERN

[1388] D. Keymeulen, M. Durantez, K. Konaka, and Y. Kuniyoshi. An evolutionary robot navigation system using a gate-level evolvable hardware. Lecture Notes in Computer Science, 1259:195– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamamoto:1997:ARE

[1389] J. Yamamoto and Y. Anzai. Autonomous robot with evolving algorithm based on biological systems. Lecture Notes in Computer Science, 1259: 220–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1997:MBN

[1390] H. Ito and T. Furuya. Memory-based neural network and its application to a mobile robot with evolutionary and experience learning. Lecture Notes in Computer Science, 1259:234–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ae:1997:SPB

[1391] T. Ae, H. Fukumoto, and S. Hiwatashi. Special-purpose brainware architecture for data processing. Lecture Notes in Computer Science, 1259:289–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iba:1997:MLA

[1392] H. Iba, M. Iwata, and T. Higuchi. Machine learning approach to gate-level evolvable hardware. Lecture Notes in Computer Science, 1259:327–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zebulum:1997:ESH

[1393] R. S. Zebulum, M. A. Pacheco, and M. Vellasco. Evolvable systems in hardware design: Taxonomy, survey and applications. Lecture Notes in Computer Science, 1259:344–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kazimierczak:1997:STB

[1394] J. Kazimierczak. From some tasks to biology and then to hardware. Lecture Notes in Computer Science, 1259: 359–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harvey:1997:TLE

[1395] L. Harvey and A. Thompson. Through the labyrinth evolution finds a way: A silicon ridge. Lecture Notes in Computer Science, 1259:406–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hikage:1997:HES

[1396] T. Hikage, H. Hemmi, and K. Shimohara. Hardware evolution system introducing dominant and recessive heredity. Lecture Notes in Computer Science, 1259:423–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gers:1997:CBN

[1397] F. Gers and H. De Garis. CAM-Brain: A new model for ATR’s cellular automata based artificial brain project. Lecture Notes in Computer Science, 1259:437–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bennett:1997:EDO

[1398] F. H. Bennett, J. R. Koza, D. Andre, and M. A. Keane. Evolution of a 60 decibel op amp using genetic programming. Lecture Notes in Computer Science, 1259:455–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biehl:1997:AGT

[1399] M. Biehl, N. Klarlund, and T. Rauhe. Algorithms for guided tree automata. Lecture Notes in Computer Science, 1260:6–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bone:1997:TSF

[1400] R. Bone, C. Daguin, A. Georgevail, and D. Maurel. Time series forecasting by finite-state automata. Lecture Notes in Computer Science, 1260:26–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garzon:1997:DIN

[1401] M. Garzon and E. Eberbach. Dynamical implementation of nondeterministic automata and concurrent systems. Lecture Notes in Computer Science, 1260:35–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glenn:1997:IWF

[1402] J. Glenn and W. Gasarch. Implementing WSIS via finite automata. Lecture Notes in Computer Science, 1260:50– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnson:1997:ICS

[1403] J. H. Johnson and D. Wood. Instruction computation in subset construction. Lecture Notes in Computer Science, 1260:64–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1997:FLF

[1404] B. W. Watson. FIRE lite: FAs and REs in C++. Lecture Notes in Computer Science, 1260:167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abel:1997:SIM

[1405] D. J. Abel. Spatial Internet marketplaces: A grand challenge? Lecture Notes in Computer Science, 1262:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ester:1997:SDM

[1406] M. Ester, H.-P. Kriegel, and J. Sander.

Spatial data mining: A database approach. Lecture Notes in Computer Science, 1262:47–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1997:GEI

[1407] E. P. F. Chan and J. N. H. Ng.

A general and efficient implementation of geometric operators and predicates. Lecture Notes in Computer Science, 1262:69–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belussi:1997:MSD

[1408] A. Belussi, E. Bertino, and B. Catania. Manipulating spatial data in constraint databases. Lecture Notes in Computer Science, 1262:115–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chomicki:1997:CBI

[1409] J. Chomicki and P. Z. Revesz.

Constraint-based interoperability of spatiotemporal databases. Lecture Notes in Computer Science, 1262:142– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1997:ISI

[1410] Y.-W. Huang, M. Jones, and E. A. Rundensteiner. Improving spatial intersect joins using symbolic intersect detection. Lecture Notes in Computer Science, 1262:165–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1997:DPP

[1411] X. Zhou, D. J. Abel, and D. Truffet. Data partitioning for parallel spatial join processing. Lecture Notes in Computer Science, 1262:178–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esperanca:1997:OPB

[1412] C. Esperanca and H. Samet. Orthogonal polygons as bounding structures in filter-refine query processing strategies. Lecture Notes in Computer Science, 1262:197–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balovnev:1997:GGS

[1413] O. Balovnev, M. Breunig, and A. B. Cremers. From GeoStore to GeoToolKit: The second step. Lecture Notes in Computer Science, 1262: 223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alonso:1997:GOW

[1414] G. Alonso and C. Hagen. Geo-Opera:

Workflow concepts for spatial processes. Lecture Notes in Computer Science, 1262:238–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erwig:1997:VR

[1415] M. Erwig and M. Schneider. Vague regions. Lecture Notes in Computer Science, 1262:298–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delis:1997:AGC

[1416] V. Delis and T. Hadzilacos. On the assessment of generalisation consistency. Lecture Notes in Computer Science, 1262:321–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ang:1997:NLN

[1417] C. H. Ang and T. C. Tan. New linear node splitting algorithm for Rtrees. Lecture Notes in Computer Science, 1262:339–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aggarwal:1997:TEI

[1418] C. Aggarwal, J. Wolf, P. Yu, and M. Epelman. The S-tree: An efficient, index for multidimensional objects. Lecture Notes in Computer Science, 1262:350–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foss:1997:KDC

[1419] B. A. Foss. Knowledge discovery — A control theory perspective. Lecture Notes in Computer Science, 1263:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carter:1997:SBM

[1420] C. L. Carter, H. J. Hamilton, and N. Cercone. Share based measures for itemsets. Lecture Notes in Computer Science, 1263:14–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hilderman:1997:PKD

[1421] R. J. Hilderman, H. J. Hamilton, R. J. Kowalchuk, and N. Cercone. Parallel knowledge discovery using domain generalization graphs. Lecture Notes in Computer Science, 1263:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bull:1997:DHR

[1422] M. Bull, G. Kundt, and L. Gierl. Discovering of health risks and case-based forecasting of epidemics in a health surveillance system. Lecture Notes in Computer Science, 1263:68–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wrobel:1997:AMR

[1423] S. Wrobel. An algorithm for multirelational discovery of subgroups. Lecture Notes in Computer Science, 1263: 78–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1997:FST

[1424] G. Das, D. Gunopulos, and H. Mannila. Finding similar time series. Lecture Notes in Computer Science, 1263: 88–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldman:1997:PBB

[1425] R. Feldman, W. Kloesgen, Y. Ben-Yehuda, and G. Kedar. Pattern based browsing in document collections. Lecture Notes in Computer Science, 1263:112–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elomaa:1997:EMN

[1426] T. Elomaa and J. Rousu. Efficient multisplitting on numerical data. Lecture Notes in Computer Science, 1263:178– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andre-Joensson:1997:USF

[1427] H. Andre-Joensson and D. Z. Badal.

Using signature files for querying timeseries data. Lecture Notes in Computer Science, 1263:211–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amir:1997:NVM

[1428] A. Amir, R. Feldman, and R. Kashi.

A new and versatile method for association generation. Lecture Notes in Computer Science, 1263:221–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bioch:1997:BDT

[1429] J. C. Bioch, O. Van der Meer, and R. Potharst. Bivariate decision trees. Lecture Notes in Computer Science, 1263:232–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wirth:1997:TPO

[1430] R. Wirth, C. Shearer, U. Grimmer, and T. Reinartz. Towards process-oriented tool support for KDD. Lecture Notes in Computer Science, 1263:243–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gebhardt:1997:FSC

[1431] F. Gebhardt. Finding spatial clusters. Lecture Notes in Computer Science, 1263:277–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudaillier:1997:IIH

[1432] E. Boudaillier and G. Hebrail. Interactive interpretation of hierarchical clustering. Lecture Notes in Computer Science, 1263:288–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freitas:1997:PTB

[1433] A. A. Freitas. The principle of transformation between efficiency and effectiveness: Towards a fair evaluation of the cost-effectiveness of KDD techniques. Lecture Notes in Computer Science, 1263:299–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berka:1997:RRD

[1434] P. Berka. Recognizing reliability of discovered knowledge. Lecture Notes in Computer Science, 1263:307–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watts:1997:TIM

[1435] R. J. Watts, A. L. Porter, S. Cunningham, and D. Zhu. TOAS intelligence mining; analysis of NLP and computational linguistics. Lecture Notes in Computer Science, 1263:323–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahonen:1997:MPF

[1436] H. Ahonen, O. Heinonen, M. Klemettinen, and A. I. Verkamo. Mining in the phrasal frontier. Lecture Notes in Computer Science, 1263:343–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjorvand:1997:MTS

[1437] A. T. Bjorvand. Mining time series using rough sets. Lecture Notes in Computer Science, 1263:351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhong:1997:MLO

[1438] N. Zhong, S. Ohsuga, C. Liu, and Y. Kakemoto. On meta levels of an organized society of KDD agents. Lecture Notes in Computer Science, 1263: 367–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1997:UNN

[1439] Y. Zhou, Y. Lu, and C. Shi. Using neural network to extract knowledge from database. Lecture Notes in Computer Science, 1263:376–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grossman:1997:TIH

[1440] R. Grossman. A tutorial introduction to high performance data mining. Lecture Notes in Computer Science, 1263:395–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carbonara:1997:DMT

[1441] L. Carbonara, H. Roberts, and B. Egan. Data mining in the telecommunications industry. Lecture Notes in Computer Science, 1263:396–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1997:EM

[1442] G. Das, R. Fleischer, L. Gasieniec, and D. Gunopulos. Episode matching. Lecture Notes in Computer Science, 1264:12–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berman:1997:CPM

[1443] P. Berman, M. Karpinski, L. L. Larmore, and W. Plandowski. On the complexity of pattern matching for highly compressed two-dimensional texts. Lecture Notes in Computer Science, 1264:40–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraenkel:1997:WCS

[1444] A. S. Fraenkel, J. Simpson, and M. Paterson. On weak circular squares in binary words. Lecture Notes in Computer Science, 1264:76–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasieniec:1997:EIP

[1445] L. Gasieniec, P. Indyk, and P. Krysta.

External inverse pattern matching. Lecture Notes in Computer Science, 1264:90–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crochemore:1997:DCC

[1446] M. Crochemore and R. Verin. Direct construction of compact directed acyclic word graphs. Lecture Notes in Computer Science, 1264:116–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cryan:1997:AAF

[1447] M. Cryan, L. A. Goldberg, and C. Phillips. Approximation algorithms for the fixed-topology phylogenetic number problem. Lecture Notes in Computer Science, 1264:130–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdeddaim:1997:ICT

[1448] S. Abdeddaim. On incremental computation of transitive closure and greedy alignment. Lecture Notes in Computer Science, 1264:167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvestad:1997:ACD

[1449] L. Arvestad. Aligning coding DNA in the presence of frame-shift errors. Lecture Notes in Computer Science, 1264: 180–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:BDS

[1450] T. Chen and S. S. Skiena. Thebased data structures for sequence assembly. Lecture Notes in Computer Science, 1264:206–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-Dor:1997:BBC

[1451] A. Ben-Dor, G. Lancia, J. Perone, and R. Ravi. Banishing bias from consensus sequences. Lecture Notes in Computer Science, 1264:247–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dress:1997:IVS

[1452] A. Dress. Iterative versus simultaneous multiple sequence alignment. Lecture Notes in Computer Science, 1264: 275–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaffar:1997:FBC

[1453] J. Jaffar, B. Liu, and R. H. C. Yap. Forward and backward chaining in constraint programming. Lecture Notes in Computer Science, 1265:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buccafurri:1997:SWC

[1454] F. Buccafurri, N. Leone, and P. Rullo.

Strong and weak constraints in disjunctive datalog. Lecture Notes in Computer Science, 1265:2–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zukowski:1997:IAF

[1455] U. Zukowski, S. Brass, and B. Freitag. Improving the alternating fixpoint: The transformation approach. Lecture Notes in Computer Science, 1265:40–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egly:1997:NMR

[1456] U. Egly and H. Tompits. Is nonmonotonic reasoning always harder? Lecture Notes in Computer Science, 1265:60–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blair:1997:SBP

[1457] H. A. Blair, F. Dushin, and P. Humenn. Simulations between programs as cellular automata. Lecture Notes in Computer Science, 1265:115–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janhunen:1997:SDB

[1458] T. Janhunen. Separating disbeliefs from beliefs in autoepistemic reasoning. Lecture Notes in Computer Science, 1265:132–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:PD

[1459] G.-Q. Zhang and W. C. Rounds. Power defaults. Lecture Notes in Computer Science, 1265:152–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bochman:1997:SPS

[1460] A. Bochman. A study of Przymusinski’s static semantics. Lecture Notes in Computer Science, 1265:170–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonatti:1997:RSS

[1461] P. A. Bonatti. Resolution for skeptical stable semantics. Lecture Notes in Computer Science, 1265:185–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiter:1997:CNG

[1462] T. Eiter, J. Lu, and V. S. Subrahmanian. Computing non-ground representations of stable models. Lecture Notes in Computer Science, 1265:198– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karig:1997:INI

[1463] M. R. Karig. Industry needs for integrated information services. Lecture Notes in Computer Science, 1265:218– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchberger:1997:CSP

[1464] B. Buchberger. Computing, solving, proving: A report of the theorema project. Lecture Notes in Computer Science, 1265:220–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gelfond:1997:TSA

[1465] M. Gelfond. Towards a systematic approach to representing knowledge in declarative logic programming. Lecture Notes in Computer Science, 1265:222– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damasio:1997:PSC

[1466] C. V. Damasio and L. M. Pereira.

A paxaconsistent semantics with contradiction support detection. Lecture Notes in Computer Science, 1265:224– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dekhtyar:1997:CEU

[1467] M. Dekhtyar, A. Dikorsky, and N. Spyratos. On conservative enforced updates. Lecture Notes in Computer Science, 1265:244–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Witteveen:1997:GFR

[1468] C. Witteveen and W. Van Der Hoek.

A general framework for revising nonmonotonic theories. Lecture Notes in Computer Science, 1265:258–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1997:CGL

[1469] A. Brogi, S. Contiero, and F. Turini.

Composing general logic programs. Lecture Notes in Computer Science, 1265:273–??, 1997. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiter:1997:MLP

[1470] T. Eiter, G. Gottlob, and H. Veith. Modular logic programming and generalized quantifiers. Lecture Notes in Computer Science, 1265:289–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yahya:1997:GQA

[1471] A. H. Yahya. Generalized query answering in disjunctive deductive databases: Procedural and nonmonotonic aspects. Lecture Notes in Computer Science, 1265:324–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aravindan:1997:DTD

[1472] C. Aravindan, J. Dix, and I. Niemelae. DisLoP: Towards a disjunctive logic programming system. Lecture Notes in Computer Science, 1265:341–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damasio:1997:RLP

[1473] C. V. Damasio, L. M. Pereira, and M. Schroeder. REVISE: Logic programming and diagnosis. Lecture Notes in Computer Science, 1265:353–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiter:1997:DSN

[1474] T. Eiter, N. Leone, C. Mateis, and G. Pfeifer. A deductive system for nonmonotonic reasoning. Lecture Notes in Computer Science, 1265:363–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zukowski:1997:DDS

[1475] U. Zukowski and B. Freitag. The deductive database system LOLA. Lecture Notes in Computer Science, 1265: 375–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kakas:1997:AFS

[1476] A. C. Kakas and C. Mourlas. ACLP:

Flexible solutions to complex problems. Lecture Notes in Computer Science, 1265:387–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kandzia:1997:NMR

[1477] P.-T. Kandzia. Non-monotonic reasoning in FLORID. Lecture Notes in Computer Science, 1265:399–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheetham:1997:CBR

[1478] W. Cheetham and J. Graf. Case-based reasoning in color matching. Lecture Notes in Computer Science, 1266:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finnie:1997:ESD

[1479] G. R. Finnie, G. W. Wittig, and J.- M. Desharnais. Estimating software development effort with case-based reasoning. Lecture Notes in Computer Science, 1266:13–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1997:ACB

[1480] Marc Fuchs and Matt Fuchs. Applying case-based reasoning to automated deduction. Lecture Notes in Computer Science, 1266:23–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomes:1997:CBA

[1481] P. Gomes and C. Bento. A casebased approach for elaboration of design requirements. Lecture Notes in Computer Science, 1266:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarmulak:1997:CBR

[1482] J. Jarmulak, E. J. H. Kerckhoffs, and P. P. Van’t Veen. Case-based reasoning in an ultrasonic rail-inspection system. Lecture Notes in Computer Science, 1266:43–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joh:1997:CCE

[1483] D. Joh. CBR in a changing environment. Lecture Notes in Computer Science, 1266:53–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber-Lee:1997:LCB

[1484] R. Weber-Lee, R. M. Barcia, M. C. Da Costa, and I. W. Rodrigues Filho. A large case-based reasoner for legal cases. Lecture Notes in Computer Science, 1266:190–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wisdo:1997:SAQ

[1485] C. Wisdo. A scalable approach for question-based indexing of encyclopedic texts. Lecture Notes in Computer Science, 1266:200–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cox:1997:ERR

[1486] M. T. Cox. An explicit representation of reasoning failures. Lecture Notes in Computer Science, 1266:211–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamp:1997:ACD

[1487] G. Kamp. On the admissibility of concrete domains for CBR based on description logics. Lecture Notes in Computer Science, 1266:223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1997:EHC

[1488] I. Watson and S. Perera. The evaluation of a hierarchical case representation using context guided retrieval. Lecture Notes in Computer Science, 1266: 255–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aha:1997:RCC

[1489] D. W. Aha and L. A. Breslow. Refining conversational case libraries. Lecture Notes in Computer Science, 1266: 267–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arcos:1997:PDB

[1490] J. L. Arcos and R. Lopez de Mantaras. Perspectives: A declarative bias mechanism for case retrieval. Lecture Notes in Computer Science, 1266:279– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonzano:1997:UIL

[1491] A. Bonzano, P. Cunningham, and B. Smyth. Using introspective learning to improve retrieval in CBR: A case study in air traffic control. Lecture Notes in Computer Science, 1266: 291–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brueninghaus:1997:UML

[1492] S. Brueninghaus and K. D. Ashley. Using machine learning for assigning indices to textual cases. Lecture Notes in Computer Science, 1266:303–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daengdej:1997:HCB

[1493] J. Daengdej and D. Lukose. How case-based reasoning and cooperative query answering techniques support RICAD? Lecture Notes in Computer Science, 1266:315–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daniels:1997:WYS

[1494] J. J. Daniels and E. L. Rissland. What you saw is what you want: Using cases to seed information retrieval. Lecture Notes in Computer Science, 1266:325–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanney:1997:AKB

[1495] K. Hanney and M. T. Keane. The adaptation knowledge bottleneck: How to ease it by learning from cases. Lecture Notes in Computer Science, 1266: 359–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cox:1997:LCF

[1496] M. T. Cox. Loose coupling of failure explanation and repair: Using learning goals to sequence learning methods. Lecture Notes in Computer Science, 1266:425–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drummond:1997:UCB

[1497] C. Drummond. Using a case-base of surfaces to speed-up reinforcement learning. Lecture Notes in Computer Science, 1266:435–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griffiths:1997:PAS

[1498] A. D. Griffiths and D. G. Bridge. PAC analyses of a ‘similarity learning’ IBL algorithm. Lecture Notes in Computer Science, 1266:445–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Howe:1997:ELV

[1499] N. Howe and C. Cardie. Examining locally varying weights for nearest neighbor algorithms. Lecture Notes in Computer Science, 1266:455–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aarts:1997:QKS

[1500] R. J. Aarts and J. Rousu. Qualitative knowledge to support reasoning about cases. Lecture Notes in Computer Science, 1266:489–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

An:1997:IRI

[1501] A. An, N. Cercone, and C. Chan. Integrating rule induction and casebased reasoning to enhance problem solving. Lecture Notes in Computer Science, 1266:499–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergmann:1997:USP

[1502] R. Bergmann, W. Wilke, and J. Schumacher. Using software process modeling for building a case-based reasoning methodology: Basic approach and case study. Lecture Notes in Computer Science, 1266:509–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branting:1997:SCB

[1503] L. K. Branting. Stratified case-based reasoning in non-refinable abstraction hierarchies. Lecture Notes in Computer Science, 1266:519–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cox:1997:SCH

[1504] M. T. Cox and M. M. Veloso. Supporting combined human and machine planning: An interface for planning by analogical reasoning. Lecture Notes in Computer Science, 1266:531–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karacapilidis:1997:UCB

[1505] N. Karacapilidis, B. Trousse, and D. Papadias. Using case-based reasoning for argumentation with multiple viewpoints. Lecture Notes in Computer Science, 1266:541–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhatta:1997:ATC

[1506] S. R. Bhatta and A. K. Goel. An analogical theory of creativity in design. Lecture Notes in Computer Science, 1266:565–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1997:FMC

[1507] D. Dubois, F. Esteva, P. Garcia, and L. Godo. Fuzzy modeling of casebased reasoning and decision. Lecture Notes in Computer Science, 1266:599– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faltings:1997:PIC

[1508] B. Faltings. Probabilistic indexing for case-based prediction. Lecture Notes in Computer Science, 1266:611–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yager:1997:CBR

[1509] R. R. Yager. Case-based reasoning, fuzzy systems modeling and solution composition. Lecture Notes in Computer Science, 1266:633–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Handschuh:1997:CSE

[1510] Helena Handschuh and Henri Gilbert.

χ^{2} cryptanalysis of the SEAL encryption algorithm. Lecture Notes in Computer Science, 1267:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1267/12670001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670001.
pdf.

Harpes:1997:PC

[1511] Carlo Harpes and James L. Massey. Partitioning cryptanalysis. Lecture Notes in Computer Science, 1267:13– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670013.
pdf.

Jakobsen:1997:IAB

[1512] Thomas Jakobsen and Lars R. Knudsen. The interpolation attack on block ciphers. Lecture Notes in Computer Science, 1267:28–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670028.
pdf.

Aoki:1997:BDC

[1513] Kazumaro Aoki, Kunio Kobayashi, and Shiho Moriai. Best differential characteristic search of FEAL. Lecture Notes in Computer Science, 1267:41– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670041.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1267/12670041.
pdf.

Matsui:1997:NBE

[1514] Mitsuru Matsui. New block encryption algorithm MISTY. Lecture Notes in Computer Science, 1267:54– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670054.
pdf.

Kwan:1997:DIE

[1515] Matthew Kwan. The design of the ICE encryption algorithm. Lecture Notes in Computer Science, 1267:69– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670069.
pdf.

Anonymous:1997:AES

[1516] Anonymous. Advanced Encryption Standard, draft minimum requirements and evaluation criteria. Lecture Notes in Computer Science, 1267:83– 87, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670083.
pdf.

Ding:1997:TFS

[1517] Cunsheng Ding, Valtteri Niemi, Ari Renvall, and Arto Salomaa. TWOPRIME: A fast stream ciphering algorithm. Lecture Notes in Computer Science, 1267:88–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1267/12670088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1267/12670088.
                                        pdf.

Dichtl:1997:NFG

[1518] Markus Dichtl. On nonlinear filter generators. Lecture Notes in Computer Science, 1267:103–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1267/12670103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1267/12670103.
                                        pdf.

Anderson:1997:CNK

[1519] Ross Anderson and Charalampos Manifavas. Chameleon — A new kind of stream cipher. Lecture Notes in Computer Science, 1267:107–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1267/12670107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1267/12670107.
                                        pdf.

Sakurai:1997:ILC

[1520] Kouichi Sakurai and Souichi Furuya. Improving linear cryptanalysis of LOKI91 by Probabilistic Counting method. Lecture Notes in Computer Science, 1267:114–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670114.
pdf.

Biham:1997:CL

[1521] Eli Biham. Cryptanalysis of Ladder-DES. Lecture Notes in Computer Science, 1267:134–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1267/12670134.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1267/12670134.pdf.

Rijmen:1997:FTC

[1522] Vincent Rijmen and Bart Preneel. A family of trapdoor ciphers. Lecture Notes in Computer Science, 1267:139– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670139.
pdf.

Daemen:1997:BCS

[1523] Joan Daemen, Lars Knudsen, and Vincent Rijmen. The block cipher SQUARE. Lecture Notes in Computer Science, 1267:149–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670149.
pdf.

MRaihi:1997:XFO

[1524] David M’Raïhi, David Naccache, Jacques Stern, and Serge Vaudenay. XMX: A firmware-oriented block cipher based on modular multiplications. Lecture Notes in Computer Science, 1267:166–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1267/12670166.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1267/12670166.
pdf.

Halevi:1997:MSM

[1525] Shai Halevi and Hugo Krawczyk. MMH: Software message authentication in the Gbit/second rates. Lecture Notes in Computer Science, 1267:172– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670172.
pdf.

Afanassiev:1997:FMA

[1526] Valentine Afanassiev, Christian Gehrmann, and Ben Smeets. Fast message authentication using efficient polynomial evaluation. Lecture Notes in Computer Science, 1267:190–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670190.
pdf.

Yuval:1997:RTE

[1527] Gideon Yuval. Reinventing the travois:

Encryption/MAC in 30 ROM bytes. Lecture Notes in Computer Science, 1267:205–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1267/12670205.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1267/12670205.
pdf.

Rivest:1997:ANE

[1528] Ronald L. Rivest. All-or-nothing encryption and the package transform. Lecture Notes in Computer Science, 1267:210–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1267/12670210.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1267/12670210.
pdf.

Lucks:1997:SRK

[1529] Stefan Lucks. On the security of remotely keyed encryption. Lecture Notes in Computer Science, 1267:219– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1267/12670219.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1267/12670219.
                                        pdf.

Young:1997:SEC

[1530] Adam Young and Moti Yung. Sliding encryption: A cryptographic tool for mobile agents. Lecture Notes in Computer Science, 1267:230–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1267/12670230.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1267/12670230.
                                        pdf.

Schneier:1997:FSE

[1531] Bruce Schneier and Doug Whiting. Fast software encryption: Designing encryption algorithms for optimal software speed on the Intel Pentium processor. Lecture Notes in Computer Science, 1267:242–259, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1267/12670242.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1267/12670242.
                                        pdf;   http://www.counterpane.com/
                                        fast_software_encryption.html.

Biham:1997:FNI

[1532] Eli Biham. A fast new DES implementation in software. Lecture Notes in Computer Science, 1267:260– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670260.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670260.
pdf.

Clapp:1997:OFS

[1533] Craig S. K. Clapp. Optimizing a fast stream cipher for VLIW, SIMD, and superscalar processors. Lecture Notes in Computer Science, 1267:273– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1267/12670273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1267/12670273.
pdf.

Braine:1997:ICO

[1534] L. Braine and C. Clack. Introducing clover: An object-oriented functional language. Lecture Notes in Computer Science, 1268:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erwig:1997:AP

[1535] M. Erwig. Active patterns. Lecture Notes in Computer Science, 1268:21– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boquist:1997:GPH

[1536] U. Boquist and T. Johnsson. The Grin Project: A highly optimising back end for lazy functional languages. Lecture Notes in Computer Science, 1268:58– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breitinger:1997:IET

[1537] S. Breitinger, U. Klusik, and R. Loogen. An implementation of Eden on top of Concurrent Haskell. Lecture Notes in Computer Science, 1268:141–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Assmann:1997:CFP

[1538] C. Assmann. Coordinating functional processes using Petri nets. Lecture Notes in Computer Science, 1268:162– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarvis:1997:RPL

[1539] S. A. Jarvis and R. G. Morgan. The results of: Profiling large-scale lazy functional programs. Lecture Notes in Computer Science, 1268:200–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debbabi:1997:TBA

[1540] M. Debbabi, A. Faour, and N. Tawbi. A type-based algorithm for the controlflow analysis of higher-order concurrent programs. Lecture Notes in Computer Science, 1268:247–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dosch:1997:CFM

[1541] W. Dosch and B. Moeller. Calculating a functional module for binary search trees. Lecture Notes in Computer Science, 1268:267–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joesang:1997:PMT

[1542] A. Joesang. Prospectives for modelling trust in information security. Lecture Notes in Computer Science, 1270:2–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bai:1997:AIF

[1543] Y. Bai and V. Varadharajan. Analysis and implementation of a formal authorisation policy design approach. Lecture Notes in Computer Science, 1270:14–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holford:1997:POS

[1544] J. Holford and G. Mohay. ProtectOS:

Operating system and hardware support for small objects. Lecture Notes in Computer Science, 1270:102–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischlin:1997:PMC

[1545] M. Fischlin. Practical memory checkers for stacks, queues and deques. Lecture Notes in Computer Science, 1270: 114–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charney:1997:PSC

[1546] S. Charney, S. Perrin, S. Orlowski, and N. Reaburn. Panel session: Cryptographic policy guidelines. Lecture Notes in Computer Science, 1270:126– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yi:1997:NHF

[1547] X. Yi and K.-Y. Lam. A new hash function based on block cipher. Lecture Notes in Computer Science, 1270: 139–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:NLB

[1548] X.-M. Zhang and Y. Zheng. New lower bounds on nonlinearity and a class of highly nonlinear functions. Lecture Notes in Computer Science, 1270:147– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:SSR

[1549] L. Chen, D. Gollmann, C. J. Mitchell, and P. Wild. Secret sharing with reusable polynomials. Lecture Notes in Computer Science, 1270:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bakhtiari:1997:MAC

[1550] S. Bakhtiari, R. Safavi-Naini, and J. Pieprzyk. A message authentication code based on Latin squares. Lecture Notes in Computer Science, 1270:194– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bao:1997:DAT

[1551] F. Bao, R. Deng, Y. Han, and A. Jeng. Design and analyses of two basic protocols for use in TTP-based key escrow. Lecture Notes in Computer Science, 1270:261–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyd:1997:KAC

[1552] C. Boyd. On keg agreement and conference keg agreement. Lecture Notes in Computer Science, 1270:294–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dittmann:1997:ETT

[1553] J. Dittmann and A. Steinmetz. Enabling technology for the trading of MPEG-encoded video. Lecture Notes in Computer Science, 1270:314–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abe:1997:IDM

[1554] T. Abe, H. Fujii, and K. Kushima. Image distribution method with embedded identifier scheme for copyright protection. Lecture Notes in Computer Science, 1270:325–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerschberg:1997:RIS

[1555] L. Kerschberg. The role of intelligent software agents in advanced information systems. Lecture Notes in Computer Science, 1271:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bratsberg:1997:LRI

[1556] S. E. Bratsberg, S.-O. Hvasshovd, and O. Torbjoernsen. Location and replication independent recovery in a highly available database. Lecture Notes in Computer Science, 1271:23–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groevlen:1997:CBQ

[1557] O. Groevlen, O. Torbjoernsen, and S.- O. Hvasshovd. Compensation-based query processing in on-line transaction processing systems. Lecture Notes in Computer Science, 1271:38–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFerreiraRezende:1997:DAD

[1558] F. De Ferreira Rezende, T. Haerder, A. Gloeckner, and J. Lutze. Detection arcs for deadlock management in nested transactions and their performance. Lecture Notes in Computer Science, 1271:54–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bratbergsengen:1997:IOP

[1559] K. Bratbergsengen and K. Noervaag. Improved and optimized partitioning techniques in database query processing. Lecture Notes in Computer Science, 1271:69–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grust:1997:QEC

[1560] T. Grust, J. Kroeger, D. Gluche, and A. Heuer. Query evaluation in CROQUE — calculus and algebra coincide. Lecture Notes in Computer Science, 1271:84–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zurek:1997:OPT

[1561] T. Zurek. Optimisation of partitioned temporal joins. Lecture Notes in Computer Science, 1271:101–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balownew:1997:MLC

[1562] O. Balownew, T. Bode, A. B. Cremers, and J. Kalinski. Maintaining library catalogues with an RDBMS: A performance study. Lecture Notes in Computer Science, 1271:116–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berti:1997:MAD

[1563] L. Berti. Merging an active database and a reflective system: Modelling a new several active meta-levels architecture. Lecture Notes in Computer Science, 1271:119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1997:FDM

[1564] H. Gupta, I. McLaren, and A. Vella. A framework for database mining. Lecture Notes in Computer Science, 1271: 121–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haase:1997:QPT

[1565] O. Haase and A. Henrich. Query processing techniques for partly inaccessible distributed databases. Lecture Notes in Computer Science, 1271:123– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winiwarter:1997:DDO

[1566] W. Winiwarter and Y. Kambayashi.

DOA — the deductive object-oriented approach to the development of adaptive natural language interfaces. Lecture Notes in Computer Science, 1271: 137–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faulstich:1997:WWI

[1567] L. C. Faulstich, M. Spiliopoulou, and V. Linnemann. WIND: A warehouse for Internet data. Lecture Notes in Computer Science, 1271:169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conrad:1997:SII

[1568] S. Conrad, M. Hoeding, G. Saake, and I. Schmitt. Schema integration with integrity constraints. Lecture Notes in Computer Science, 1271:200–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1997:MID

[1569] L. Xu and A. Poulovassilis. A method for integrating deductive databases. Lecture Notes in Computer Science, 1271:215–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chazelle:1997:DTC

[1570] B. Chazelle. Discrepancy theory and computational geometry. Lecture Notes in Computer Science, 1272:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berretty:1997:DMP

[1571] R.-P. Berretty, M. Overmars, and A. F. Van der Stappen. Dynamic motion planning in low obstacle density environments. Lecture Notes in Computer Science, 1272:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guibas:1997:VBP

[1572] L. J. Guibas, J.-C. Latombe, S. M.

LaValle, and D. Lin. Visibility-based pursuit-evasion in a polygonal environment. Lecture Notes in Computer Science, 1272:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1997:MEM

[1573] P. K. Agarwal, L. J. Guibas, J. Hershberger, and E. Veach. Maintaining the extent of a moving point set. Lecture Notes in Computer Science, 1272: 31–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alstrup:1997:FCL

[1574] S. Alstrup, P. W. Lauridsen, P. Sommerlund, and M. Thorup. Finding cores of limited length. Lecture Notes in Computer Science, 1272:45– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shahrokhi:1997:BCL

[1575] F. Shahrokhi, O. Sykora, L. Szekely, and I. Vrto. On bipartite crossings, largest biplanar subgraphs, and the linear arrangement problem. Lecture Notes in Computer Science, 1272:55– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kosaraju:1997:AAG

[1576] S. Rao Kosaraju, A. A. Schaeffer, and L. G. Biesecker. Approximation algorithms for a genetic diagnostics problem. Lecture Notes in Computer Science, 1272:69–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hershberger:1997:CLS

[1577] J. Hershberger and J. Snoeyink. Cartographic line simplification and polygon CSG formulae in O(n log n) time. Lecture Notes in Computer Science, 1272: 93–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charikar:1997:CTL

[1578] M. Charikar, R. Motwani, P. Raghavan, and C. Silverstein. Constrained TSP and low-power computing. Lecture Notes in Computer Science, 1272: 104–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berman:1997:LLB

[1579] P. Berman, M. Charikar, and M. Karpinski. On-line load balancing for related machines. Lecture Notes in Computer Science, 1272:116–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stojanovic:1997:LTA

[1580] N. Stojanovic, P. Berman, D. Gumucio, and R. Hardison. A linear-time algorithm for the 1-Mismatch problem. Lecture Notes in Computer Science, 1272:126–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Majhi:1997:SGO

[1581] J. Majhi, R. Janardan, M. Smid, and P. Gupta. On some geometric optimization problems in layered manufacturing. Lecture Notes in Computer Science, 1272:136–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okano:1997:NTC

[1582] H. Okano, S. Misono, and K. Iwano.

New TSP construction heuristics and their relationships to the 2-Opt. Lecture Notes in Computer Science, 1272: 150–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amir:1997:PMH

[1583] A. Amir, M. Lewenstein, and N. Lewenstein. Pattern matching in hypertext. Lecture Notes in Computer Science, 1272:160–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeza-Yates:1997:MAS

[1584] R. Baeza-Yates and G. Navarro. Multiple approximate string matching. Lecture Notes in Computer Science, 1272: 174–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobkin:1997:ACG

[1585] D. P. Dobkin. Applied computational geometry. Lecture Notes in Computer Science, 1272:185–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devillers:1997:CCP

[1586] O. Devillers, G. Liotta, F. P.

Preparata, and R. Tamassia. Checking the convexity of polytopes and the planarity of subdivisions. Lecture Notes in Computer Science, 1272:186–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barequet:1997:VDP

[1587] G. Barequet, M. T. Dickerson, and M. T. Goodrich. Voronoi diagrams for polygon-offset distance functions. Lecture Notes in Computer Science, 1272: 200–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seiden:1997:RAA

[1588] S. Seiden. Randomized algorithms for that ancient scheduling problem. Lecture Notes in Computer Science, 1272: 210–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayashi:1997:OPA

[1589] T. Hayashi, K. Nakano, and S. Olariu. Optimal parallel algorithms for proximate points, with applications. Lecture Notes in Computer Science, 1272: 224–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eppstein:1997:EAS

[1590] D. Eppstein and D. W. Hart. An efficient algorithm for shortest paths in vertical and horizontal segments. Lecture Notes in Computer Science, 1272: 234–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:GPQ

[1591] D. Z. Chen, O. Daescu, and K. S. Klenk. On geometric path query problems. Lecture Notes in Computer Science, 1272:248–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldman:1997:LSH

[1592] S. A. Goldman, J. Parwatikar, and S. Suri. On-line schedufing with hard deadlines. Lecture Notes in Computer Science, 1272:258–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1997:LBM

[1593] S. K. Das and M. C. Pinotti. Load balanced mapping of data structures in parallel memory modules for fast and conflict-free templates access. Lecture Notes in Computer Science, 1272:272– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mulmuley:1997:PVP

[1594] K. Mulmuley. Parallel vs. parametric complexity. Lecture Notes in Computer Science, 1272:282–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez-Ortiz:1997:PIN

[1595] A. Lopez-Ortiz and S. Schuierer.

Position-independent near optimal searching and on-line recognition in star polygons. Lecture Notes in Computer Science, 1272:284–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Efrat:1997:DDS

[1596] A. Efrat, M. J. Katz, F. Nielsen, and M. Sharir. Dynamic data structures for fat objects and their appfications. Lecture Notes in Computer Science, 1272:297–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldwasser:1997:IAS

[1597] M. Goldwasser and R. Motwani. Intractability of assembly sequencing: Unit disks in the plane. Lecture Notes in Computer Science, 1272:307–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narasimhan:1997:HTS

[1598] G. Narasimhan. On Hamiltonian triangulations in simple polygons. Lecture Notes in Computer Science, 1272: 321–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertolazzi:1997:COD

[1599] P. Bertolazzi, G. Di Battista, and W. Didimo. Computing orthogonal drawings with the minimum number of bends. Lecture Notes in Computer Science, 1272:331–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwek:1997:SDF

[1600] S. Kwek. On a simple depth-first search strategy for exploring unknown graphs. Lecture Notes in Computer Science, 1272:345–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakostas:1997:ODH

[1601] A. Papakostas and I. G. Tollis. Orthogonal drawing of high degree graphs with small area and few bends. Lecture Notes in Computer Science, 1272: 354–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramnath:1997:CLT

[1602] S. Ramnath and H. Shen. A centroid labeling technique and its application to path selection in trees. Lecture Notes in Computer Science, 1272:368–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barequet:1997:OPA

[1603] G. Barequet, A. J. Briggs, M. T. Dickerson, and M. T. Goodrich. Offsetpolygon annulus placement problems. Lecture Notes in Computer Science, 1272:378–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBerg:1997:CCM

[1604] M. De Berg, P. Bose, D. Bremner, and S. Ramaswami. Computing constrained minimum-width annuli of point sets. Lecture Notes in Computer Science, 1272:392–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Segal:1997:GAP

[1605] M. Segal and K. Kedem. Geometric applications of posets. Lecture Notes in Computer Science, 1272:402–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1997:CPD

[1606] H. Gupta and R. Wenger. Constructing pairwise disjoint paths with few links. Lecture Notes in Computer Science, 1272:416–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodnik:1997:TDA

[1607] A. Brodnik, P. B. Miltersen, and J. I. Munro. Trans-dichotomous algorithms without multiplication — some upper and lower bounds. Lecture Notes in Computer Science, 1272:426–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rethman:1997:AAS

[1608] J. Rethman and E. Wanke. An approximation algorithm for stacking up bins from a conveyer onto pallets. Lecture Notes in Computer Science, 1272:440–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1997:RBT

[1609] K. S. Larsen, E. Soisalon-Soininen, and P. Widmayer. Relaxed balance through standard rotations. Lecture Notes in Computer Science, 1272:450–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hershberger:1997:EBR

[1610] J. Hershberger and S. Suri. Efficient breakout routing in printed circuit boards. Lecture Notes in Computer Science, 1272:462–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:PR

[1611] Z.-Z. Chen, M. Grigni, and C. Papadimitriou. Planarity revisited. Lecture Notes in Computer Science, 1272: 472–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arehart:1997:BSH

[1612] A. B. Arehart and W. A. Wolovich.

Bumpless switching in hybrid systems. Lecture Notes in Computer Science, 1273:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branicky:1997:TPM

[1613] M. S. Branicky, E. Dolginova, and N. Lynch. A toolbox for proving and maintaining hybrid specifications. Lecture Notes in Computer Science, 1273: 18–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branicky:1997:SHS

[1614] M. S. Branicky and S. E. Mattsson.

Simulation of hybrid systems. Lecture Notes in Computer Science, 1273:31– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brayman:1997:AKN

[1615] V. Brayman and J. Vagners. Application of the Kohn-Nerode control law extraction procedure to the inverted pendulum problem. Lecture Notes in Computer Science, 1273:57–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broucke:1997:DHS

[1616] M. Broucke and P. Varaiya. Decidability of hybrid systems with linear and nonlinear differential inclusions. Lecture Notes in Computer Science, 1273:77–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cofer:1997:RIH

[1617] D. D. Cofer. Reliable implementation of hybrid control systems for advanced avionics. Lecture Notes in Computer Science, 1273:93–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deshpande:1997:SFP

[1618] A. Deshpande, A. Goellue, and P. Varaiya. SHIFT: A formalism and a programming language for dynamic networks of hybrid automata. Lecture Notes in Computer Science, 1273:113– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heymann:1997:SMR

[1619] M. Heymann, F. Lin, and G. Meyer. Synthesis of minimally restrictive legal controllers for a class of hybrid systems. Lecture Notes in Computer Science, 1273:134–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knight:1997:CTM

[1620] J. F. Knight and B. Luense. Control theory, modal logic, and games. Lecture Notes in Computer Science, 1273: 160–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohn:1997:ABV

[1621] W. Kohn, A. Nerode, and J. B. Remmel. Agent based velocity control of highway systems. Lecture Notes in Computer Science, 1273:174–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolen:1997:CAR

[1622] J. F. Kolen and F. Zhao. A computational analysis of the reachability problem for a class of hybrid dynamical systems. Lecture Notes in Computer Science, 1273:215–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kourjanski:1997:CRH

[1623] M. Kourjanski and P. Varaiya. A class of rectangular hybrid systems with computable reach set. Lecture Notes in Computer Science, 1273:228–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lemmon:1997:SIS

[1624] M. Lemmon and C. J. Bett. Safe implementations of supervisory commands. Lecture Notes in Computer Science, 1273:235–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nerode:1997:HSG

[1625] A. Nerode, J. B. Remmel, and A. Yakhnis. Hybrid system games: Extraction of control automata with small topologies. Lecture Notes in Computer Science, 1273:248–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oezguener:1997:HCD

[1626] O. Oezguener, C. Hatipoglu, A. Iftar, and K. Redmill. Hybrid control desgin for a three vehicle scenario demonstration using overlapping decompositions. Lecture Notes in Computer Science, 1273:294–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pappas:1997:TCA

[1627] G. J. Pappas and S. Sastry. Towards continuous abstractions of dynamical and control systems. Lecture Notes in Computer Science, 1273:329–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raisch:1997:TOS

[1628] J. Raisch and S. O’Young. A totally ordered set of discrete abstractions for a given hybrid or continuous system. Lecture Notes in Computer Science, 1273:342–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stursberg:1997:CTH

[1629] O. Stursberg, S. Kowalewski, I. Hoffmann, and J. Preussig. Comparing timed and hybrid automata as approximations of continuous systems. Lecture Notes in Computer Science, 1273: 361–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomlin:1997:HCM

[1630] C. Tomlin, G. Pappas, J. Lygeros, and D. Godbole. Hybrid control models of next generation air traffic management. Lecture Notes in Computer Science, 1273:378–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Urquhart:1997:JET

[1631] K. Urquhart. Java and the emerging technology. Lecture Notes in Computer Science, 1274:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nelson:1997:TSL

[1632] T. H. Nelson. Transcopyright: A simple legal arrangement for sharing, reuse, and republication of copyrighted material on the net. Lecture Notes in Computer Science, 1274:7–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirano:1997:HDE

[1633] S. Hirano. HORB: Distributed execution of Java programs. Lecture Notes in Computer Science, 1274:29–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fukuda:1997:PMA

[1634] M. Fukuda, L. F. Bic, and M. B. Dillencourt. Performance of the MES-SENGERS autonomous-objects-based system. Lecture Notes in Computer Science, 1274:43–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallnau:1997:ECB

[1635] K. Wallnau, E. Morris, P. Feiler, and A. Earl. Engineering component-based systems with distributed object technology. Lecture Notes in Computer Science, 1274:58–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishioka:1997:LCN

[1636] T. Nishioka, Y. Hamazaki, and M. Tsukamoto. Local class name space facility for worldwide object-oriented distributed systems environment. Lecture Notes in Computer Science, 1274: 74–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuentes:1997:COA

[1637] L. Fuentes and J. M. Troya. A component-oriented architecture to design multimedia services on a distributed platform. Lecture Notes in Computer Science, 1274:90–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cavedon:1997:TTP

[1638] L. Cavedon, A. Rao, L. Sonenberg, and G. Tidhar. Teamwork via team plans in intelligent autonomous agent systems. Lecture Notes in Computer Science, 1274:106–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizuno:1997:WWI

[1639] H. Mizuno and H. Fukuoka. WISH: A Web information sharing system accessible via the WWW. Lecture Notes in Computer Science, 1274:122–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakashima:1997:OPC

[1640] H. Nakashima. Organic programming for cooperative computation. Lecture Notes in Computer Science, 1274:133– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grudin:1997:EGS

[1641] J. Grudin and L. Palen. Emerging groupware successes in major corporations: Studies of adoption and adaptation. Lecture Notes in Computer Science, 1274:142–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamakami:1997:ATF

[1642] T. Yamakami. An approach toward fits evaluation between organizational culture and groupware. Lecture Notes in Computer Science, 1274:154–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1997:RRP

[1643] K. Fujita and S. Kunifuji. A realization of a reflection of personal information on distributed brainstorming environment. Lecture Notes in Computer Science, 1274:166–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1997:HAM

[1644] W.-S. Li, K. S. Candan, K. Hirata, and Y. Hara. A hybrid approach to multimedia database systems through integration of semantics and media-based search. Lecture Notes in Computer Science, 1274:182–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bayer:1997:UBT

[1645] R. Bayer. The universal B-tree for multidimensional indexing: General concepts. Lecture Notes in Computer Science, 1274:198–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamiura:1997:SVL

[1646] M. Kamiura, H. Oiso, K. Tajima, and K. Tanaka. Spatial views and LODbased access control in VRML-object databases. Lecture Notes in Computer Science, 1274:210–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Messina:1997:MDI

[1647] P. Messina. Metacomputing and dataintensive applications. Lecture Notes in Computer Science, 1274:226–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishio:1997:AID

[1648] S. Nishio and M. Tsukamoto. Architectural issues of database systems in advanced computer networks. Lecture Notes in Computer Science, 1274:237– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1997:APM

[1649] D. B. Lange, M. Oshima, G. Kargoth, and K. Kosaka. Aglets: Programming mobile agents in Java. Lecture Notes in Computer Science, 1274:253–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamaguchi:1997:WCM

[1650] S. Yamaguchi, K.-I. Chinen, and H. Inoue. WWW cache management and its international deployment. Lecture Notes in Computer Science, 1274:267– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mansfield:1997:VDA

[1651] G. Mansfield, E. P. Duarte, M. Kitahashi, and S. Noguchi. VINES: Distributed algorithms for a Web based distributed network management system. Lecture Notes in Computer Science, 1274:281–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radoslavov:1997:ACD

[1652] P. I. Radoslavov and T. Tokuda. An algorithm for cost-delay balanced trees and its application to wide area network multicast. Lecture Notes in Computer Science, 1274:294–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kummerfeld:1997:SID

[1653] R. J. Kummerfeld. Server initiated delivery of multimedia information. Lecture Notes in Computer Science, 1274: 309–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moriai:1997:DLO

[1654] S. Moriai and H. Tokuda. Dynamic loadable object support for realtime Mach kernels. Lecture Notes in Computer Science, 1274:318–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwasaki:1997:MKI

[1655] M. Iwasaki, T. Takeuchi, M. Nakahara, and S. Nakamura. A microkernel for isochronous video-data transfer. Lecture Notes in Computer Science, 1274:334–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manning:1997:MMC

[1656] E. G. Manning, S. Khan, R. Lea, and G. C. Shoja. Metaspaces and mobile computing: Promises and challenges. Lecture Notes in Computer Science, 1274:350–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Budiarto:1997:PLH

[1657] Budiarto, Kaname Harumoto, Masahiko Tsukamoto, and Shojiro Nishio. Position locking: Handling location dependent queries in mobile computing environment. Lecture Notes in Computer Science, 1274:363–378, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hokimoto:1997:RHM

[1658] A. Hokimoto and T. Nakajima. Robust host mobility supports for adaptive mobile applications. Lecture Notes in Computer Science, 1274:379–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Terada:1997:ACI

[1659] M. Terada, Y. Murayama, and G. Mansfield. Access control for interorganizational computer network environment. Lecture Notes in Computer Science, 1274:394–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shinjo:1997:WWS

[1660] Y. Shinjo. The world wide shell based on the object-stacking model. Lecture Notes in Computer Science, 1274:410– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1997:DGP

[1661] Y. Sato and Y. Hamazaki. DeleGate: A general purpose application level gateway. Lecture Notes in Computer Science, 1274:426–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agerholm:1997:IBT

[1662] S. Agerholm and J. Frost. An Isabellebased theorem prover for VDM-SL. Lecture Notes in Computer Science, 1275:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrews:1997:EFS

[1663] J. H. Andrews. Executing formal specifications by translation to higher order logic programming. Lecture Notes in Computer Science, 1275:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Archer:1997:HST

[1664] M. Archer and C. Heitmeyer. Humanstyle theorem proving using PVS. Lecture Notes in Computer Science, 1275: 33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camilleri:1997:HAV

[1665] A. J. Camilleri. A hybrid approach to verifying liveness in a symmetric multiprocessor. Lecture Notes in Computer Science, 1275:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chetali:1997:FVC

[1666] B. Chetali and B. Heyd. Formal verification of concurrent programs in LP and in CoQ: A comparative analysis. Lecture Notes in Computer Science, 1275:69–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Constable:1997:MPC

[1667] R. L. Constable. ML programming in constructive type theory. Lecture Notes in Computer Science, 1275:87– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devillers:1997:PIS

[1668] M. Devillers, D. Griffioen, and O. Mueller. Possibly infinite sequences in theorem provers: A comparative study. Lecture Notes in Computer Science, 1275:89–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dowek:1997:PNF

[1669] G. Dowek. Proof normalization for a first-order formulation of higher-order logic. Lecture Notes in Computer Science, 1275:105–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dutertre:1997:UPE

[1670] B. Dutertre and S. Schneider. Using a PVS embedding of CSP to verify authentication protocols. Lecture Notes in Computer Science, 1275:121– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1997:VAP

[1671] J. Harrison. Verifying the accuracy of polynomial approximations in HOL. Lecture Notes in Computer Science, 1275:137–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirschkoff:1997:FFC

[1672] D. Hirschkoff. A full formalisation of πcalculus theory in the calculus of constructions. Lecture Notes in Computer Science, 1275:153–169, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1997:RDP

[1673] D. Kapur. Rewriting, decision procedures and lemma speculation for automated hardware verification. Lecture Notes in Computer Science, 1275:171– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laangbacka:1997:RRS

[1674] T. Laangbacka and J. Von Wright.

Refining reactive systems in HOL using action systems. Lecture Notes in Computer Science, 1275:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohri:1997:FBT

[1675] T. Mohri. On formalization of bicategory theory. Lecture Notes in Computer Science, 1275:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naraschewski:1997:TOO

[1676] W. Naraschewski. Towards an objectoriented progification language. Lecture Notes in Computer Science, 1275: 215–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peled:1997:VRS

[1677] D. Peled. Verification for robust specification. Lecture Notes in Computer Science, 1275:231–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santen:1997:TSM

[1678] T. Santen. A theory of structured model-based specifications in isabelle/ HOL. Lecture Notes in Computer Science, 1275:243–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simons:1997:PPI

[1679] M. Simons. Proof presentation for isabelle. Lecture Notes in Computer Science, 1275:259–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slind:1997:DUI

[1680] K. Slind. Derivation and use of induction schemes in higher-order logic. Lecture Notes in Computer Science, 1275: 275–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slotosch:1997:HOQ

[1681] O. Slotosch. Higher order quotients and their implementation in Isabelle HOL. Lecture Notes in Computer Science, 1275:291–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wenzel:1997:TCO

[1682] M. Wenzel. Type classes and overloading in higher-order logic. Lecture Notes in Computer Science, 1275:307– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zammit:1997:CSC

[1683] V. Zammit. A comparative study of Coq and HOL. Lecture Notes in Computer Science, 1275:323–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1997:CFA

[1684] S. K. Das and M. C. Pinotti. Conflictfree access to templates of trees and hypercubes in parallel memory systems. Lecture Notes in Computer Science, 1276:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krizanc:1997:MOP

[1685] D. Krizanc and L. Zhang. Many-to-one packet routing via matchings. Lecture Notes in Computer Science, 1276: 11–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Annexstein:1997:MTG

[1686] F. S. Annexstein, K. A. Berman, and R. Swaminathan. A multi-tree generating routing scheme using acyclic orientations. Lecture Notes in Computer Science, 1276:18–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertram-Kretzberg:1997:AHP

[1687] C. Bertram-Kretzberg, T. Hofmeister, and H. Lefmann. An algorithm for Heilbronn’s problem. Lecture Notes in Computer Science, 1276:23–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1997:OMP

[1688] K. W. Chan and F. Y. L. Chin. Optimal multiresolution polygonal approximation. Lecture Notes in Computer Science, 1276:32–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:CSC

[1689] W. Chen, Xiao Wen Deng, K. Wada, and K. Kawaguchi. Constructing a strongly convex superhull of points. Lecture Notes in Computer Science, 1276:42–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:1997:FRS

[1690] B. Zhu. Fast range searching with Delaunay triangulations. Lecture Notes in Computer Science, 1276:52–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1997:RHC

[1691] J.-Y. Cai and D. Sivakumar. Resolution of Hartmanis’ conjecture for NLhard sparse sets. Lecture Notes in Computer Science, 1276:62–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reischuk:1997:CLF

[1692] R. Reischuk. Can large fanin circuits perform reliable computations in the presence of noise? Lecture Notes in Computer Science, 1276:72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grolmusz:1997:SSR

[1693] V. Grolmusz. On set systems with restricted intersections modulo a composite number. Lecture Notes in Computer Science, 1276:82–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takenaga:1997:SVO

[1694] Y. Takenaga, M. Nouzoe, and S. Yajima. Size and variable ordering of OBDDs representing threshold functions. Lecture Notes in Computer Science, 1276:91–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atteson:1997:PNJ

[1695] K. Atteson. The performance of neighbor-joining algorithms of phylogeny reconstruction. Lecture Notes in Computer Science, 1276:101–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berry:1997:IET

[1696] V. Berry and O. Gascuel. Inferring evolutionary trees with strong combinatorial evidence. Lecture Notes in Computer Science, 1276:111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rice:1997:PHB

[1697] K. Rice and T. Warnow. Parsimony is hard to beat. Lecture Notes in Computer Science, 1276:124–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasieniec:1997:CCE

[1698] L. Gasieniec, J. Jansson, A. Lingas, and A. Oestlin. On the complexity of computing evolutionary trees. Lecture Notes in Computer Science, 1276:134– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brattka:1997:CI

[1699] V. Brattka. Computable invariance. Lecture Notes in Computer Science, 1276:146–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1997:SCR

[1700] Q. Zhou. Subclasses of computable real valued functions. Lecture Notes in Computer Science, 1276:156–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weihrauch:1997:CCL

[1701] K. Weihrauch and X. Zheng. Computability on continuous, lower semicontinuous and upper semi-continuous real functions. Lecture Notes in Computer Science, 1276:166–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmeister:1997:COS

[1702] T. Hofmeister, M. Krause, and H. U. Simon. Contrast-optimal k out of n secret sharing schemes in visual cryptography. Lecture Notes in Computer Science, 1276:176–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chua:1997:RTS

[1703] Seng Kiat Chua and S. Ling. A Rabin-type scheme based on y^{2}x^{3} + bx^{2} mod n. Lecture Notes in Computer Science, 1276:186–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kao:1997:OBA

[1704] M.-Y. Kao, J. Qi, and L. Tan. Optimal bidding algorithms against cheating in multiple-object auctions. Lecture Notes in Computer Science, 1276: 192–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:SAS

[1705] P. Eades and X. Lin. Spring algorithms and symmetry. Lecture Notes in Computer Science, 1276:202–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kashem:1997:GVR

[1706] M. Abul Kashem, X. Zhou, and T. Nishizeki. Generalized vertexrankings of partial k-trees. Lecture Notes in Computer Science, 1276:212– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:STT

[1707] L. Chen. A selected tour of the theory of identification matrices. Lecture Notes in Computer Science, 1276:222– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liang:1997:EDP

[1708] Y. D. Liang, Chin Lung Lu, and Chuan Yi Tang. Efficient domination on permutation graphs and trapezoid graphs. Lecture Notes in Computer Science, 1276:232–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1997:PMA

[1709] W.-L. Hsu. On physical mapping algorithms — an error-tolerant test for the consecutive ones property. Lecture Notes in Computer Science, 1276:242– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sankoff:1997:MPB

[1710] D. Sankoff and M. Blanchette. The median problem for breakpoints in comparative genomics. Lecture Notes in Computer Science, 1276:251–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chao:1997:FAA

[1711] K.-M. Chao. Fast algorithms for aligning sequences with restricted affine gap penalties. Lecture Notes in Computer Science, 1276:264–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qiao:1997:MRP

[1712] S. Qiao and W. S.-Y Wang. A matrix representation of phylogenetic trees. Lecture Notes in Computer Science, 1276:274–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peng:1997:ENS

[1713] S.-L. Peng, C.-W. Ho, T.-S. Hsu, and M.-T Ko. Edge and node searching problems on trees. Lecture Notes in Computer Science, 1276:284–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1997:CNO

[1714] P. Becker. Construction of nearly optimal multiway trees. Lecture Notes in Computer Science, 1276:294–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsai:1997:OAC

[1715] K.-H. Tsai and D.-W Wang. Optimal algorithms for circle partitioning. Lecture Notes in Computer Science, 1276:304–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1997:HRM

[1716] X. Tan and X. Song. Hexagonal routing of multiterminal nets. Lecture Notes in Computer Science, 1276:311–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lim:1997:DAP

[1717] S.-J. Lim and Y.-K Ng. Design and analysis of parallel set-term unification. Lecture Notes in Computer Science, 1276:321–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawabe:1997:MCO

[1718] Y. Kawabe and N. Ishii. On modularity of the completeness in ordersorted term rewriting systems. Lecture Notes in Computer Science, 1276:331– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Su:1997:TVQ

[1719] K. Su, D. Ding, and H. Chen. A three-valued quantificational logic of context. Lecture Notes in Computer Science, 1276:343–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poulakidas:1997:CSS

[1720] A. S. Poulakidas, A. Srinivasan, O. Egecioğlu, and O. Ibarra. A compact storage scheme for fast waveletbased subregion retrieval. Lecture Notes in Computer Science, 1276:353– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:DAM

[1721] Y. Zhang, J. Xiao, and X. Zhou. A declustering algorithm for minimizing spatial join cost. Lecture Notes in Computer Science, 1276:363–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:FTS

[1722] L. Wang and X. Jia. Fixed topology Steiner trees and spanning forests with applications in network communications. Lecture Notes in Computer Science, 1276:373–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xue:1997:BBA

[1723] G. Xue. A branch-and-bound algorithm for computing node weighted Steiner minimum trees. Lecture Notes in Computer Science, 1276:383–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1997:CCR

[1724] Y. Kobayashi, F. Otto, and P. Seedbold. A complete characterization of repetitive morphisms over the twoletter alphabet. Lecture Notes in Computer Science, 1276:393–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1997:AIS

[1725] S. M. Kim. An algorithm for identifying spliced languages. Lecture Notes in Computer Science, 1276:403–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hertrampf:1997:ST

[1726] U. Hertrampf. The shapes of trees. Lecture Notes in Computer Science, 1276: 412–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geidmanis:1997:TLA

[1727] D. Geidmanis, J. Kaneps, K. Apsitis, and D. Taimina. Tally languages accepted by alternating multitape finite automata. Lecture Notes in Computer Science, 1276:422–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwasaki:1997:IST

[1728] Y. Iwasaki, Y. Kajiwara, K. Obokata, and Y. Igarashi. Independent spanning trees of chordal rings. Lecture Notes in Computer Science, 1276:431–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1997:RG

[1729] Q. Hu, Y. Zhang, and X. Shen. Rearrangeable graphs. Lecture Notes in Computer Science, 1276:441–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sung:1997:PTN

[1730] T.-Y. Sung, T.-Y. Ho, and L.-H. Hsu. Projective torus networks. Lecture Notes in Computer Science, 1276:451– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1997:RSA

[1731] P. Cai and J.-Y. Cai. On the 100% rule of sensitivity analysis in linear programming. Lecture Notes in Computer Science, 1276:460–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takeuchi:1997:ETP

[1732] F. Takeuchi and H. Imai. Enumerating triangulations for products of two simplices and for arbitrary configurations of points. Lecture Notes in Computer Science, 1276:470–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1997:TSG

[1733] Z. Han and R. Lu. Tree structure genetic algorithm with a nourishment mechanism. Lecture Notes in Computer Science, 1276:482–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1997:ASP

[1734] X. Wu. An approach to scalability of parallel matrix multiplication algorithms. Lecture Notes in Computer Science, 1276:492–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1997:LAM

[1735] H. Liu, E. Song, and R. Sotudeh. Limited acceleration mechanism for cell loss free flow control in ATM networks. Lecture Notes in Computer Science, 1276:502–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:TJB

[1736] Y. Chen and Q. Li. They are just butterflies. Lecture Notes in Computer Science, 1276:512–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitanyi:1997:MS

[1737] P. Vitanyi. Mutual search. Lecture Notes in Computer Science, 1276:514– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teng:1997:EEG

[1738] S.-H Teng. Eigenvalues, eigenvectors, and graph partitioning. Lecture Notes in Computer Science, 1276:515–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agibalov:1997:PCF

[1739] G. P. Agibalov. Parallel computations on finite partially ordered sets. Lecture Notes in Computer Science, 1277:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:TLB

[1740] L. Chen. Tight lower bounds for computing shortest paths on proper interval and bipartite permutation graphs. Lecture Notes in Computer Science, 1277:7–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietz:1997:URT

[1741] R. D. Dietz, T. L. Casavant, T. E. Scheetz, and T. A. Braun. Using runtime uncertainty to robustly schedule parallel computation. Lecture Notes in Computer Science, 1277:13–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartmann:1997:TBD

[1742] P. Hartmann. A tuple-based data structure for distributed parallel processing of 3D dynamic meshes. Lecture Notes in Computer Science, 1277: 25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khenner:1997:APC

[1743] M. V. Khenner, D. V. Lyubimov, B. Roux, and S. V. Shklyaev. The application of parallel computations technique to the solution of certain hydrodynamic stability problems. Lecture Notes in Computer Science, 1277:40– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kushnarenko:1997:FFA

[1744] O. Kushnarenko and P. Schnoebelen.

A formal framework for the analysis of recursive-parallel programs. Lecture Notes in Computer Science, 1277:45– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Likhoded:1997:SDD

[1745] N. A. Likhoded and A. A. Tiountchik.

Systematic design of 3-dimensional fixed-size array processors. Lecture Notes in Computer Science, 1277:60– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lomazova:1997:PLD

[1746] I. Lomazova. On proving large distributed systems: Petri net modules verification. Lecture Notes in Computer Science, 1277:70–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pudov:1997:ISC

[1747] S. Pudov. Influence of self-connection weights on cellular-neural network stability. Lecture Notes in Computer Science, 1277:76–82, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakellariou:1997:EPS

[1748] R. Sakellariou. Estimating the parallel start — up overhead for parallelizing compilers. Lecture Notes in Computer Science, 1277:83–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seredynski:1997:PDE

[1749] F. Seredynski, P. Bouvry, and F. Arbab. Parallel and distributed evolutionary computation with MAN-IFOLD. Lecture Notes in Computer Science, 1277:94–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Severyanov:1997:PCF

[1750] V. Severyanov. Parallel computation of fractal sets with the help of neural networks and cellular automata. Lecture Notes in Computer Science, 1277:109–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sokolov:1997:TDA

[1751] V. A. Sokolov, E. E. Roubtsova, and S. A. Roubtsov. On a technology of design and analysis of dataflow programs. Lecture Notes in Computer Science, 1277:115–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoutchinin:1997:ILP

[1752] A. Stoutchinin. An integer linear programming model of software pipelining for the MIPS R8000 processor. Lecture Notes in Computer Science, 1277: 121–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valkovskii:1997:CCA

[1753] V. Valkovskii and D. Zerbino. Computations on cellular automata with defects. Lecture Notes in Computer Science, 1277:136–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:EIIb

[1754] T. Yang. Efficient implementation of the improved unsymmetric Lanczos process on massively distributed memory computers. Lecture Notes in Computer Science, 1277:145–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alpkocak:1997:SGF

[1755] A. Alpkocak and E. Ozkarahan. A spatial grid file for multimedia data representation. Lecture Notes in Computer Science, 1277:156–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bessonov:1997:OTP

[1756] O. Bessonov and B. Roux. Optimization techniques and performance analysis for different serial and parallel RISC-based computers. Lecture Notes in Computer Science, 1277:168– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borshchev:1997:ADE

[1757] A. V. Borshchev, Y. G. Karpov, V. V.

Roudakov, and A. Filippov. Analysis of a distributed election algorithm using COVERS 3.0 — A case study. Lecture Notes in Computer Science, 1277:175–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeDinechin:1997:USP

[1758] B. D. De Dinechin. A unified software pipeline construction scheme for modulo scheduled loops. Lecture Notes in Computer Science, 1277:189–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Germain:1997:HCS

[1759] C. Germain, J. Laminie, M. Pallud, and D. Etiemble. An HPF case study of a domain-decomposition based irregular application. Lecture Notes in Computer Science, 1277:201–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hluchy:1997:HAT

[1760] L. Hluchy, M. Dobrucky, and J. Astalos. Hybrid approach to task allocation in distributed systems. Lecture Notes in Computer Science, 1277:210– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwak:1997:VMN

[1761] H. Kwak, B. Lee, and A. R. Hurson. Viability of multithreading on networks of workstations. Lecture Notes in Computer Science, 1277:216–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jegou:1997:TMF

[1762] Y. Jegou. Task migration and fine grain parallelism on distributed memory architectures. Lecture Notes in Computer Science, 1277:226–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

On:1997:SBV

[1763] G.-W. On, B.-S. Lee, C.-E. Hong, and D.-H Chi. A scheme for building visual debugging environment with dynamic debugging method for parallel systems. Lecture Notes in Computer Science, 1277:241–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadopoulos:1997:CDC

[1764] G. A. Papadopoulos and F. Arbab.

Control-driven coordination programming in shared dataspace. Lecture Notes in Computer Science, 1277:247– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ten:1997:PAG

[1765] S. V. Ten and K. Otsuyama. Performance analysis of geometric modeling algorithm. Lecture Notes in Computer Science, 1277:262–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trichina:1997:VTS

[1766] E. Trichina and J. Oinonen. 3D visual tool supporting derivation of parallel programs for MIMD systems. Lecture Notes in Computer Science, 1277: 268–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:SAP

[1767] J. Wang, J. Li, and H. Kameda.

Scheduling algorithms for parallel transaction processing systems. Lecture Notes in Computer Science, 1277: 283–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cappello:1997:CPA

[1768] F. Cappello and D. Etiemble. Communications in parallel architectures and networks of workstations: From standardisation to new standards. Lecture Notes in Computer Science, 1277:298– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goossens:1997:MVC

[1769] B. Goossens. A multithreaded vector co-processor. Lecture Notes in Computer Science, 1277:311–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1997:HSC

[1770] R. Hoffmann and K.-P Voelkmann. Hardware support for 3D cellular processing. Lecture Notes in Computer Science, 1277:322–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Achasova:1997:CNL

[1771] S. M. Achasova. Cellular neural-like algorithms with heuristics for solving combinatorial optimization problems. Lecture Notes in Computer Science, 1277:330–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balandin:1997:AMS

[1772] M. Balandin, O. Chernyshev, and E. Shurina. Analysis of methods for solving large-scale non-symmetric linear systems with sparsed matrices. Lecture Notes in Computer Science, 1277: 336–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fabero:1997:PSN

[1773] J. C. Fabero, A. Bautista, and L. Casasus. Parallel simulation of non-linear phenomena with cellular automata. Lecture Notes in Computer Science, 1277:344–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khajdukov:1997:MSW

[1774] V. G. Khajdukov, V. D. Korneev, V. I.

Kostin, and V. V. Kovalevsky. Modelling of seismic wave propagation for 2D media (direct and inverse problems). Lecture Notes in Computer Science, 1277:350–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klimova:1997:DGP

[1775] O. V. Klimova. Decomposition on a group and parallel convolution and Fast Fourier Transform algorithms. Lecture Notes in Computer Science, 1277:358– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iiin:1997:PIS

[1776] V. P. Ii’in. Parallel implementation of symmetric alternating direction implicit methods. Lecture Notes in Computer Science, 1277:364–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markova:1997:CCA

[1777] V. Markova and S. Piskunov. Construction of composed ALT-models of cellular architectures. Lecture Notes in Computer Science, 1277:371–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Onuphre:1997:PCU

[1778] E. Onuphre and A. Chambarel. Parallel computation of an unsteady compressible flow. Lecture Notes in Computer Science, 1277:377–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sinitsyn:1997:PST

[1779] I. N. Sinitsyn. Parallel simulation technologies for stochastic systems. Lecture Notes in Computer Science, 1277: 383–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akimova:1997:PDA

[1780] E. N. Akimova. Parallel direct algorithms for solution of sparse linear systems. Lecture Notes in Computer Science, 1277:389–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dikenelli:1997:SPP

[1781] O. Dikenelli, O. Ozkasap, and E. Ozkarahan. Scheduling parallel programs involving parallel database interactions. Lecture Notes in Computer Science, 1277:391–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaliaev:1997:BMM

[1782] A. Kaliaev, I. Kaliaev, and I. Levin. The base module of multiprocessor system with structural-procedural organization of computing. Lecture Notes in Computer Science, 1277:394–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koch:1997:BPP

[1783] W. Koch. Bitwise processing — a paradigm for deriving parallel algorithms. Lecture Notes in Computer Science, 1277:396–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno:1997:TSN

[1784] E. D. Moreno, S. T. Kofuji, M. Stumm, and T. Abdelrahman. Tuning shared network cache size vs. second-level cache size in clusters-based multiprocessors. Lecture Notes in Computer Science, 1277:398–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeOliveira:1997:SMA

[1785] P. P. B. De Oliveira. A simple, multi-architecture, parallel procedure for generating combinations. Lecture Notes in Computer Science, 1277:400– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Samofalov:1997:PDP

[1786] V. Samofalov and A. Konovalov. Processing and debugging of parallel programs on the level of task model. Lecture Notes in Computer Science, 1277: 402–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiountchik:1997:SME

[1787] A. A. Tiountchik. Systolic modular exponentiation. Lecture Notes in Computer Science, 1277:404–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:HPI

[1788] T. Yang and H.-X. Lin. The highly parallel incomplete Gram–Schmidt preconditioner. Lecture Notes in Computer Science, 1277:406–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borshchev:1997:CCB

[1789] A. V. Borshchev, Y. G. Karpov, V. V. Roudakov, and A. Filippov. COV-ERS 3.0 — A C++ based graphical modeling and simulation tool. Lecture Notes in Computer Science, 1277:409– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pogudin:1997:SCC

[1790] Y. Pogudin and O. Bandman. Simulating cellular computations with ALT. A tutorial. Lecture Notes in Computer Science, 1277:424–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vazhenin:1997:SFS

[1791] A. Vazhenin and V. Morozov. SPARTH: a family of systems for parallel highaccuracy computations. Lecture Notes in Computer Science, 1277:436–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conrad:1997:MCC

[1792] M. Conrad and K.-P. Zauner. Molecular computing: From conformational pattern recognition to complex processing networks. Lecture Notes in Computer Science, 1278:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prusinkiewicz:1997:LVM

[1793] P. Prusinkiewicz. A look at the visual modeling of plants using Lsystems. Lecture Notes in Computer Science, 1278:11–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lim:1997:BCD

[1794] H. Lim. Bioinformatics and cheminformatics in the drug discovery cycle. Lecture Notes in Computer Science, 1278:30–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apweiler:1997:NDL

[1795] R. Apweiler, V. Junker, A. Gateau, and C. O’Donovan. New developments in linking of biological databases and computer-generation of annotation: SWISS-PROT and its computerannotated supplement TREMBL. Lecture Notes in Computer Science, 1278: 44–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salas:1997:EEG

[1796] F. Salas, J. Haas, G. Overton, and C. Stoeckert. EpoDB: An erythropoiesis gene expression database in progress. Lecture Notes in Computer Science, 1278:52–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Havel:1997:RAM

[1797] T. Havel, S. Hyberts, and I. Najfeld.

Recent advances in molecular distance geometry. Lecture Notes in Computer Science, 1278:62–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collado-Vides:1997:TMG

[1798] J. Collado-Vides, A. Huerta, and K. Klose. Three models of gene regulation in E. coli. Lecture Notes in Computer Science, 1278:72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frech:1997:NMD

[1799] K. Frech, K. Quandt, and T. Werner. A new method to develop highly specific models for regulatory DNA-Regions. Lecture Notes in Computer Science, 1278:79–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breuel:1997:TOO

[1800] G. Breuel and E. Gilles. Towards an object-oriented framework for the modeling of integrated metabolic processes. Lecture Notes in Computer Science, 1278:88–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kel:1997:TCD

[1801] A. Kel, O. Kel, O. Vishnevsky, and M. Ponomarenko. TRRD and COM-PEL databases on transcription linked to TRANSFAC as tools for analysis and recognition of regulatory sequences. Lecture Notes in Computer Science, 1278:99–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grigoriev:1997:IHD

[1802] A. Grigoriev, H. Lehrach, and J. Kumlien. Integrating heterogeneous datasets in genomic mapping: Radiation hybrids, YACS, genes and STS markers over the entire human chromosome X. Lecture Notes in Computer Science, 1278:106–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwikowski:1997:CAG

[1803] B. Schwikowski and M. Vingron. A clustering approach to generalized tree alignment with application to ALU repeats. Lecture Notes in Computer Science, 1278:115–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bornberg-Bauer:1997:SFM

[1804] E. Bornberg-Bauer. Simple folding model for HP lattice proteins. Lecture Notes in Computer Science, 1278:125– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zimmer:1997:FPF

[1805] R. Zimmer and R. Thiele. Fast protein fold recognition and accurate sequencestructure alignment. Lecture Notes in Computer Science, 1278:137–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonderLieth:1997:CSC

[1806] C.-W. Von der Lieth, E. Lang, and T. Kozar. Carbohydrates: Second-class citizens in biomedicine and bioinformatics ? Lecture Notes in Computer Science, 1278:147–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goebel:1997:SCN

[1807] U. Goebel, C. Forst, and P. Schuster. Structural constraints and neutrality in RNA. Lecture Notes in Computer Science, 1278:156–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwab:1997:SAF

[1808] C. Schwab, S. Handschuh, A. Teckentrup, and M. Wagener. A systematic approach to finding new lead structures having biological activity. Lecture Notes in Computer Science, 1278: 166–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaps:1997:VAC

[1809] A. Kaps, K. Heumann, D. Frishman, and M. Baehr. Visualization and analysis of the complete yeast genome. Lecture Notes in Computer Science, 1278:178–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suehnel:1997:VRM

[1810] J. Suehnel. Virtual reality modeling for structural biology. Lecture Notes in Computer Science, 1278:189–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dassow:1997:EGG

[1811] J. Dassow and V. Mitrana. Evolutionary grammars: A grammatical model for genome evolution. Lecture Notes in Computer Science, 1278:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paun:1997:DRD

[1812] G. Paun and A. Salomaa. From DNA recombination to DNA computing, via formal languages. Lecture Notes in Computer Science, 1278:210–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiter:1997:CCA

[1813] T. Eiter and G. Gottlob. The complexity class aitch theta0p2: Recent results and applications in AI and modal logic. Lecture Notes in Computer Science, 1279:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hennicker:1997:PSS

[1814] R. Hennicker and M. Wirsing. Proof systems for structured algebraic specifications: An overview. Lecture Notes in Computer Science, 1279:19–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1997:ACA

[1815] M. Li and P. Vitanyi. Average-case analysis via incompressibility. Lecture Notes in Computer Science, 1279:38– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mazurkiewicz:1997:LCE

[1816] A. Mazurkiewicz. Locally computable enumerations. Lecture Notes in Computer Science, 1279:51–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spielman:1997:CEC

[1817] D. A. Spielman. The complexity of error-correcting codes. Lecture Notes in Computer Science, 1279:67–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Upfal:1997:SAD

[1818] E. Upfal. Stochastic analysis of dynamic processes. Lecture Notes in Computer Science, 1279:85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avermiddig:1997:SMM

[1819] A. Avermiddig, M. Kunde, and A. Osterloh. k-k sorting on the multimesh. Lecture Notes in Computer Science, 1279:93–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1997:RCP

[1820] E. Best and T. Thielke. Refinement of coloured Petri nets. Lecture Notes in Computer Science, 1279:105–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badouel:1997:SPN

[1821] E. Badouel and P. Darondeau. Stratified Petri nets. Lecture Notes in Computer Science, 1279:117–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calabrese:1997:DAO

[1822] A. Calabrese. Distributed acyclic orientation of asynchronous anonymous networks. Lecture Notes in Computer Science, 1279:129–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choffrut:1997:GRR

[1823] C. Choffrut and L. Guerra. Generalized rational relations and their logical definability. Lecture Notes in Computer Science, 1279:138–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciebiera:1997:NBL

[1824] K. Ciebiera and A. Malinowski. A note on broadcasting with linearly bounded transmission faults in constant degree networks. Lecture Notes in Computer Science, 1279:150–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cucker:1997:LWC

[1825] F. Cucker and K. Meer. Logics which capture complexity classes over the reals. Lecture Notes in Computer Science, 1279:157–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drewes:1997:CDC

[1826] F. Drewes, H.-J. Kreowski, and D. Lapoire. Criteria to disprove context-freeness of collage languages. Lecture Notes in Computer Science, 1279:169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frid:1997:SCF

[1827] A. E. Frid. The subword complexity of fixed points of binary uniform morphisms. Lecture Notes in Computer Science, 1279:179–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasieniec:1997:EPC

[1828] L. Gasieniec and P. Indyk. Efficient parallel computing with memory faults. Lecture Notes in Computer Science, 1279:188–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruska:1997:BC

[1829] D. P. Gruska. Bounded concurrency. Lecture Notes in Computer Science, 1279:198–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hammar:1997:CTB

[1830] M. Hammar and B. J. Nilsson. Concerning the time bounds of existing shortest watchman route algorithms. Lecture Notes in Computer Science, 1279:210–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemaspaandra:1997:QOP

[1831] E. Hemaspaandra, L. A. Hemaspaandra, and H. Hempel. Query order in the polynomial hierarchy. Lecture Notes in Computer Science, 1279:222–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hertrampf:1997:PTM

[1832] U. Hertrampf. Polynomial time machines equipped with word problems over algebraic structures as their acceptance criteria. Lecture Notes in Computer Science, 1279:233–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karhumaeki:1997:PMP

[1833] J. Karhumaeki, W. Plandowski, and W. Rytter. Pattern-matching problems for 2-dimensional images described by finite automata. Lecture Notes in Computer Science, 1279:245–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koppenhagen:1997:CCC

[1834] U. Koppenhagen and E. W. Mayr. The complexity of the coverability, the containment, and the equivalence problems for commutative semigroups. Lecture Notes in Computer Science, 1279: 257–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kudlek:1997:CGD

[1835] M. Kudlek, S. Marcus, and A. Mateescu. Contextual grammars with distributed catenation and shuffle. Lecture Notes in Computer Science, 1279: 269–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuehnemann:1997:TDH

[1836] A. Kuehnemann. A two-dimensional hierarchy for attributed tree transducers. Lecture Notes in Computer Science, 1279:281–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LaTorre:1997:SWC

[1837] S. La Torre, M. Napoli, and M. Parente. Synchronization of 1-Way connected processors. Lecture Notes in Computer Science, 1279:293–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levcopoulos:1997:LTH

[1838] C. Levcopoulos and J. Gudmundsson. A linear-time heuristic for minimum rectangular coverings. Lecture Notes in Computer Science, 1279:305–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lomazova:1997:ONS

[1839] I. A. Lomazova. On occurrence net semantics for Petri nets with contacts. Lecture Notes in Computer Science, 1279:317–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1997:CAU

[1840] B. Martin. Cellular automata universality revisited. Lecture Notes in Computer Science, 1279:329–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mavronicolas:1997:TRC

[1841] M. Mavronicolas and N. Papadakis. Trade-off results for connection management. Lecture Notes in Computer Science, 1279:340–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nebel:1997:ACM

[1842] M. E. Nebel. On the average complexity of the membership problem for a generalized dyck language. Lecture Notes in Computer Science, 1279:352– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niedermeier:1997:TOL

[1843] R. Niedermeier, K. Reinhardt, and P. Sanders. Towards optimal locality in mesh-indexings. Lecture Notes in Computer Science, 1279:364–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okolnishnikova:1997:HNB

[1844] E. A. Okol’nishnikova. On the hierarchy of nondeterministic branching kprograms. Lecture Notes in Computer Science, 1279:376–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otto:1997:FUF

[1845] F. Otto and A. Sattler-Klein. FDT is undecidable for finitely presented monoids with solvable word problems. Lecture Notes in Computer Science, 1279:388–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petersen:1997:EPS

[1846] H. Petersen. The equivalence of pebbles and sensing heads for finite automata. Lecture Notes in Computer Science, 1279:400–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabinovich:1997:FAT

[1847] A. Rabinovich and B. A. Trakhtenbrot.

From finite automata toward hybrid systems. Lecture Notes in Computer Science, 1279:411–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadowski:1997:OQP

[1848] Z. Sadowski. On an optimal quantified propositional proof system and a complete language for NP intersection co-NP. Lecture Notes in Computer Science, 1279:423–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuierer:1997:LBL

[1849] S. Schuierer. Lower bounds in on-line geometric searching. Lecture Notes in Computer Science, 1279:429–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stephan:1997:CUT

[1850] F. C. Stephan and S. A. Terwijn. The complexity of universal text-learners. Lecture Notes in Computer Science, 1279:441–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verma:1997:UNF

[1851] R. M. Verma. Unique normal forms for nonlinear term rewriting systems: Root overlaps. Lecture Notes in Computer Science, 1279:452–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Virbitskaite:1997:BCP

[1852] I. Virbitskaite and A. Votintseva. Behavioural characterisations of partial order logics. Lecture Notes in Computer Science, 1279:463–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hand:1997:IDA

[1853] David J. Hand. Intelligent data analysis: Issues and opportunities. Lecture Notes in Computer Science, 1280: 1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800001.
pdf.

Apte:1997:DHC

[1854] Chidanand Apte, Se June Hong, Jonathan R. M. Hosking, Jorge Lepre, Edwin P. D. Pednault, and Barry K. Rosen. Decomposition of heterogeneous classification problems. Lecture Notes in Computer Science, 1280:17– 28, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800017.
pdf.

Muller:1997:MDS

[1855] Michael William Muller. Managing dialogue in a statistical expert assistant with a cluster-based user model. Lecture Notes in Computer Science, 1280: 29–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800029.
pdf.

McGeoch:1997:HFB

[1856] C. C. McGeoch, D. Precup, and P. R. Cohen. How to find big-oh in your data set (and how not to). Lecture Notes in Computer Science, 1280:41– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800041.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800041.
                                        pdf.

Berry:1997:DCU

[1857] Ian Berry and Paul Gough. Data classification using a W.I.S.E. toolbox. Lecture Notes in Computer Science, 1280: 53–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800053.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800053.
                                        pdf.

Cornish:1997:MMC

[1858] Tremaine A. O. Cornish. Mill’s methods for complete intelligent data analysis. Lecture Notes in Computer Science, 1280:65–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1280/12800065.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800065.
                                        pdf.

Gregory:1997:IMT

[1859] Dawn E. Gregory and Paul R. Cohen. Integrating many techniques for discovering structure in data. Lecture Notes in Computer Science, 1280:77– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800077.
pdf.

Levinson:1997:MRD

[1860] Robert Levinson and Jeff Wilkinson. Meta-reasoning for data analysis tool allocation. Lecture Notes in Computer Science, 1280:89–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800089.
pdf.

StAmant:1997:NDA

[1861] Robert St. Amant. Navigation for data analysis systems. Lecture Notes in Computer Science, 1280:101–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800101.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800101.
pdf.

Tsien:1997:ADC

[1862] Christine L. Tsien and James C.

Fackle. An annotated data collection system to support intelligent analysis of intensive care unit data. Lecture Notes in Computer Science, 1280:111– 121, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800111.
                                        pdf.

Yu:1997:CAU

[1863] Hairong Yu and Arthur Ramer. A combined approach to uncertain data analysis. Lecture Notes in Computer Science, 1280:123–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800123.
                                        pdf.

Schadler:1997:CAD

[1864] Kristina Schädler and Fritz Wysotzki. A connectionist approach to the distance-based analysis of relational data. Lecture Notes in Computer Science, 1280:137–148, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1280/12800137.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1280/12800137.pdf.

Schaedler:1997:CAD

[1865] K. Schaedler and F. Wysotzki. A connectionist approach to the distancebased analysis of relational data. Lecture Notes in Computer Science, 1280:

137–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glover:1997:EGB

[1866] Robin Glover and Peter Sharpe.

Efficient GA based techniques for automating the design of classification models. Lecture Notes in Computer Science, 1280:149–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800149.
pdf.

Lam:1997:DRM

[1867] C. P. Lam, G. A. W. West, and T. M.

Caelli. Data representations and machine learning techniques. Lecture Notes in Computer Science, 1280:161– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800161.
pdf.

Lo:1997:DKD

[1868] Suzanne Lo and A. Famili. Development of a knowledge-driven constructive induction mechanism. Lecture Notes in Computer Science, 1280:173– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800173.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800173.
                                        pdf.

Gama:1997:OLT

[1869] João Gama. Oblique linear tree. Lecture Notes in Computer Science, 1280: 187–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800187.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800187.
                                        pdf.

Haring:1997:FSN

[1870] S. Haring, J. N. Kok, and M. C. Van Wezel. Feature selection for neural networks through functional links found by evolutionary computation. Lecture Notes in Computer Science, 1280:199– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wezel:1997:FSN

[1871] S. Haring, J. N. Kok, and M. C. van Wezel. Feature selection for neural networks through functional links found by evolutionary computation. Lecture Notes in Computer Science, 1280:199– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800199.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800199.
                                        pdf.

Jensen:1997:BSM

[1872] David Jensen, Tim Oates, and Paul R.

Cohen. Building simple models: A case study with decision trees. Lecture Notes in Computer Science, 1280: 211–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800211.
pdf.

Piccardi:1997:ESL

[1873] M. Piccardi, R. Cucchiara, M. Bariani, and P. Mello. Exploiting symbolic learning in visual inspection. Lecture Notes in Computer Science, 1280:223– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800223.
pdf.

Scott:1997:FCE

[1874] P. D. Scott, R. J. Williams, and K. M.

Ho. Forming categories in exploratory data analysis and data mining. Lecture Notes in Computer Science, 1280: 235–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800235.
pdf.

Someren:1997:SDG

[1875] Maarten van Someren, Cristina Torres, and Floor Verdenius. A systematic description of greedy optimisation algorithms for cost sensitive generalisation. Lecture Notes in Computer Science, 1280:247–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800247.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800247.
                                        pdf.

VanSomeren:1997:SDG

[1876] M. Van Someren, C. Torres, and F. Verdenius. A systematic description of greedy optimisation algorithms for cost sensitive generalisation. Lecture Notes in Computer Science, 1280:247–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valtchev:1997:DMC

[1877] Petko Valtchev and Jérôme Euzenat. Dissimilarity measure for collections of objects and values. Lecture Notes in Computer Science, 1280:259–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800259.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800259.
                                        pdf.

Vullings:1997:ESU

[1878] H. J. L. M. Vullings, M. H. G. Verhaegen, and H. B. Verbruggen. ECG segmentation using time-warping. Lecture Notes in Computer Science, 1280: 275–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800275.
pdf.

Bellazzi:1997:ILD

[1879] Riccardo Bellazzi, Cristiana Larizza, and Alberto Riva. Interpreting longitudinal data through temporal abstractions: An application to diabetic patients monitoring. Lecture Notes in Computer Science, 1280:287– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800287.
pdf.

Kamp:1997:ISM

[1880] Vera Kamp and Frank Wietek. Intelligent support for multidimensional data analysis in environmental epidemiology. Lecture Notes in Computer Science, 1280:299–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800299.
pdf.

Gunn:1997:NPA

[1881] Steve R. Gunn, Martin Brown, and Kev M. Bossley. Network performance assessment for neurofuzzy data modelling. Lecture Notes in Computer Science, 1280:313–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800313.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800313.
                                        pdf.

Nascimento:1997:GAF

[1882] Susana Nascimento and Fernando Moura-Pires. A genetic approach to fuzzy clustering with a validity measure fitness function. Lecture Notes in Computer Science, 1280:325–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800325.
                                        pdf.

Roadknight:1997:AAN

[1883] C. M. Roadknight, D. Palmer-Brown, and G. E. Mills. The analysis of artificial neural network data models. Lecture Notes in Computer Science, 1280: 337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800337.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800337.
                                        pdf.

Huber:1997:SDA

[1884] Klaus-Peter Huber and Michael R.

Berthold. Simulation data analysis using fuzzy graphs. Lecture Notes in Computer Science, 1280:347–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800347.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800347.
pdf.

Klawonn:1997:MAF

[1885] Frank Klawonn and Erich-Peter Klement. Mathematical analysis of fuzzy classifiers. Lecture Notes in Computer Science, 1280:359–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800359.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800359.
pdf.

Lapp:1997:NFD

[1886] A. Lapp and H.-G. Kranz. Neuro-fuzzy diagnosis system with a rated diagnosis reliability and visual data analysis. Lecture Notes in Computer Science, 1280: 371–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800371.
pdf.

Turhan:1997:GFC

[1887] Meltem Turhan. Genetic fuzzy clustering by means of discovering membership functions. Lecture Notes in Computer Science, 1280:383–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800383.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800383.
pdf.

McSherry:1997:SIE

[1888] David McSherry. A strategy for increasing the efficiency of rule discovery in data mining. Lecture Notes in Computer Science, 1280:397–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800397.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800397.
pdf.

Schnattinger:1997:ITA

[1889] Klemens Schnattinger and Udo Hahn. Intelligent text analysis for dynamically maintaining and updating domain knowledge bases. Lecture Notes in Computer Science, 1280:409–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800409.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800409.
pdf.

Schlosser:1997:KDE

[1890] Michael Schlosser. Knowledge discovery in endgame databases. Lecture Notes in Computer Science, 1280:423– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800423.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800423.
pdf.

Darlington:1997:PIA

[1891] John Darlington, Yi ke Guo, Janjao Sutiwaraphun, and Hing Wing To. Parallel induction algorithms for data mining. Lecture Notes in Computer Science, 1280:437–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800437.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800437.
pdf.

Robinson:1997:DAQ

[1892] J. Robinson and B. G. T. Lowden. Data analysis for query processing. Lecture Notes in Computer Science, 1280:447– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800447.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800447.
pdf.

Siklossy:1997:DD

[1893] Laurent Siklóssy and Marc Ayel. Datum discovery. Lecture Notes in Computer Science, 1280:459–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800459.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800459.
                                        pdf.

Yuanhui:1997:CAE

[1894] Zhou Yuanhui, Lu Yuchang, and Shi Chunyi. A connectionist approach to extracting knowledge from databases. Lecture Notes in Computer Science, 1280:465–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1280/12800465.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800465.
                                        pdf.

Zhou:1997:CAE

[1895] Y. Zhou, Y. Lu, and C. Shi. A connectionist approach to extracting knowledge from databases. Lecture Notes in Computer Science, 1280:465–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vandenEijkel:1997:MPW

[1896] G. C. van den Eijkel, J. C. A. Van der Lubbe, and E. Backer. A modulated Parzen-windows approach for probability density estimation. Lecture Notes in Computer Science, 1280:479– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800479.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800479.
pdf.

Garcia:1997:IEQ

[1897] M. Rueda García, A. Arcos Cebrián, and E. Artés Rodríguez. Improvement on estimating quantiles in finite population using indirect methods of estimation. Lecture Notes in Computer Science, 1280:491–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800491.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800491.
pdf.

Kharin:1997:RCU

[1898] Yurij Kharin. Robustness of clustering under outliers. Lecture Notes in Computer Science, 1280:501–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800501.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800501.
pdf.

Schikuta:1997:BCS

[1899] Erich Schikuta and Martin Erhart.

The BANG-clustering system: Gridbased data analysis. Lecture Notes in Computer Science, 1280:513–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800513.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800513.
                                        pdf.

Liu:1997:TDM

[1900] W. Z. Liu, A. P. White, S. G. Thompson, and M. A. Bramer. Techniques for dealing with missing values in classification. Lecture Notes in Computer Science, 1280:527–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800527.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800527.
                                        pdf.

Ramoni:1997:UEK

[1901] Marco Ramoni and Paola Sebastiani. The use of exogenous knowledge to learn Bayesian networks from incomplete databases. Lecture Notes in Computer Science, 1280:537–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800537.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800537.
                                        pdf.

Wu:1997:RAO

[1902] John X. Wu, Gongxian Cheng, and Xiaohui Liu. Reasoning about outliers by modelling noisy data. Lecture Notes in Computer Science, 1280:549– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800549.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800549.
pdf.

Bradley:1997:RAS

[1903] Elizabeth Bradley and Matthew Easley. Reasoning about sensor data for automated system identification. Lecture Notes in Computer Science, 1280:561–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1280/12800561.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1280/12800561.
pdf.

Howe:1997:MDE

[1904] Adele E. Howe and Gabriel Somlo.

Modelling discrete event sequences as state transition diagrams. Lecture Notes in Computer Science, 1280:573– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1280/12800573.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1280/12800573.
pdf.

Boyd:1997:DDP

[1905] Sarah Boyd. Detecting and describing patterns in time-varying data using wavelets. Lecture Notes in Computer Science, 1280:585–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800585.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800585.
                                        pdf.

Schieffer:1997:DTB

[1906] Björn Schieffer and Günter Hotz. Diagnosis of tank ballast systems. Lecture Notes in Computer Science, 1280: 597–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1280/12800597.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800597.
                                        pdf.

Wolkenhauer:1997:QUM

[1907] Olaf Wolkenhauer. Qualitative uncertainty models from random set theory. Lecture Notes in Computer Science, 1280:609–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1280/12800609.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1280/12800609.
                                        pdf.

Milner:1997:SNH

[1908] R. Milner. Strong normalisation in higher-order action calculi. Lecture Notes in Computer Science, 1281:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garrigue:1997:EMS

[1909] J. Garrigue and D. Remy. Extending ML with semi-explicit higherorder polymorphism. Lecture Notes in Computer Science, 1281:20–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aiken:1997:ORP

[1910] A. Aiken, E. L. Wimmers, and J. Palsberg. Optimal representations of polymorphic types with subtyping. Lecture Notes in Computer Science, 1281:47– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ariola:1997:CLC

[1911] Z. M. Ariola and S. Blom. Cyclic lambda calculi. Lecture Notes in Computer Science, 1281:77–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohori:1997:TSS

[1912] A. Ohori. Type system for specializing polymorphism. Lecture Notes in Computer Science, 1281:107–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coppo:1997:ST

[1913] M. Coppo, F. Damiani, and P. Giannini. On strictness and totality. Lecture Notes in Computer Science, 1281: 138–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lawall:1997:SSP

[1914] J. L. Lawall and P. Thiemann. Sound specialization in the presence of computational effects. Lecture Notes in Computer Science, 1281:165–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kinoshita:1997:AAB

[1915] Y. Kinoshita, P. W. O’Hearn, A. J. Power, and M. Takeyama. An axiomatic approach to binary logical relations with applications to data refinement. Lecture Notes in Computer Science, 1281:191–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonsangue:1997:IDL

[1916] M. M. Bonsangue and J. N. Kok. Infinitary domain logic for finitary transition systems. Lecture Notes in Computer Science, 1281:213–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levy:1997:SRJ

[1917] J.-J. Levy. Some results in the joincalculus. Lecture Notes in Computer Science, 1281:233–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rathke:1997:LMC

[1918] J. Rathke and M. Hennessy. Local model checking for value-passing processes. Lecture Notes in Computer Science, 1281:250–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sere:1997:DRR

[1919] K. Sere and M. Walden. Data refinement of remote procedures. Lecture Notes in Computer Science, 1281:267– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abramsky:1997:TTA

[1920] S. Abramsky, S. Gay, and R. Nagarajan. A type-theoretic approach to deadlock-freedom of asynchronous systems. Lecture Notes in Computer Science, 1281:295–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Talcott:1997:CSM

[1921] C. L. Talcott. Composable semantic models for actor theories. Lecture Notes in Computer Science, 1281:321– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbanera:1997:CPP

[1922] F. Barbanera, S. Berardi, and M. Schivalocchi. “classical” programmingwith-proofs in lambda0s0y0mPA: An analysis of non-confluence. Lecture Notes in Computer Science, 1281:365– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Power:1997:ECS

[1923] J. Power and H. Thielecke. Environments, continuation semantics and indexed categories. Lecture Notes in Computer Science, 1281:391–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruce:1997:COE

[1924] K. B. Bruce, L. Cardelli, and B. C.

Pierce. Comparing object encodings. Lecture Notes in Computer Science, 1281:415–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poll:1997:SFW

[1925] E. Poll. System F with widthsubtyping and record updating. Lecture Notes in Computer Science, 1281: 439–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birkedal:1997:RIR

[1926] L. Birkedal and R. Harper. Relational interpretations of recursive types in an operational setting. Lecture Notes in Computer Science, 1281:458–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matthews:1997:ELF

[1927] S. Matthews. Extending a logical framework with a modal connective for validity. Lecture Notes in Computer Science, 1281:491–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boutin:1997:URB

[1928] S. Boutin. Using reflection to build efficient and certified decision procedures. Lecture Notes in Computer Science, 1281:515–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werner:1997:STT

[1929] B. Werner. Sets in types, types in sets. Lecture Notes in Computer Science, 1281:530–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayr:1997:CPN

[1930] R. Mayr. Combining Petri nets and PA-Processes. Lecture Notes in Computer Science, 1281:547–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamane:1997:SMC

[1931] S. Yamane and K. Nakamura. Symbolic model-checking method based on approximations and binary decision diagrams for real-time systems. Lecture Notes in Computer Science, 1281:562– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gardner:1997:TMH

[1932] P. Gardner and M. Hasegawa. Types and models for higher-order action calculi. Lecture Notes in Computer Science, 1281:583–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huet:1997:FST

[1933] G. Huet and H. Laulhere. Finitestate transducers as regular Boehm trees. Lecture Notes in Computer Science, 1281:604–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:1997:STS

[1934] M. Abadi. Secrecy by typing in security protocols. Lecture Notes in Computer Science, 1281:611–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjornson:1997:WTU

[1935] R. Bjornson, N. Carriero, and D. Gelernter. From weaving threads to untangling the Web: A view of coordination from Linda’s perspective. Lecture Notes in Computer Science, 1282:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kramer:1997:ESC

[1936] J. Kramer and J. Magee. Exposing the skeleton in the coordination closet. Lecture Notes in Computer Science, 1282:18–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lea:1997:DOS

[1937] D. Lea. Design for open systems in Java. Lecture Notes in Computer Science, 1282:32–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inverardi:1997:CAC

[1938] P. Inverardi, A. L. Wolf, and D. Yankelevich. Checking assumptions in component dynamics at the architectural level. Lecture Notes in Computer Science, 1282:46–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidan:1997:SBS

[1939] C. Bidan and V. Issarny. Security benefits from software architecture. Lecture Notes in Computer Science, 1282: 64–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minsky:1997:RCO

[1940] N. H. Minsky and V. Ungureanu. Regulated coordination in open distributed systems. Lecture Notes in Computer Science, 1282:81–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olivier:1997:DDA

[1941] P. A. Olivier. Debugging distributed applications using a coordination architecture. Lecture Notes in Computer Science, 1282:98–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nunes:1997:CDA

[1942] I. Nunes, J. L. Fiadeiro, and W. M.

Turski. Coordinating durative actions. Lecture Notes in Computer Science, 1282:115–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jagannathan:1997:CPS

[1943] S. Jagannathan. Communicationpassing style for coordination languages. Lecture Notes in Computer Science, 1282:131–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeJong:1997:SAL

[1944] E. De Jong. Software architecture for large control systems: A case study description. Lecture Notes in Computer Science, 1282:150–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stuurman:1997:ESA

[1945] S. Stuurman and J. Van Katwijk. Evaluation of software architectures for a control system: A case study. Lecture Notes in Computer Science, 1282:157– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzbacher:1997:MRC

[1946] A. A. Holzbacher, M. Perin, and M. Suedholt. Modeling railway control systems using graph grammars: A case study. Lecture Notes in Computer Science, 1282:172–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gelernter:1997:WLF

[1947] D. Gelernter and L. Zuck. On what Linda is: Formal description of Linda as a reactive system. Lecture Notes in Computer Science, 1282:187–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busi:1997:TSO

[1948] N. Busi, R. Gorrieri, and G. Zavattaro. Three semantics of the output operation for generative communication. Lecture Notes in Computer Science, 1282:205–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNicola:1997:CMA

[1949] R. De Nicola, G. L. Ferrari, and R. Pugliese. Coordinating mobile agents via blackboards and access rights. Lecture Notes in Computer Science, 1282:220–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1997:MCA

[1950] A. Brogi and J.-M Jacquet. Modeling coordination via asynchronous communication. Lecture Notes in Computer Science, 1282:238–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1997:POS

[1951] E. Best, F. S. De Boer, and C. Palamidessi. Partial order and SOS semantics for linear constraint programs. Lecture Notes in Computer Science, 1282:256–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denti:1997:PCM

[1952] E. Denti, A. Natali, and A. Omicini.

Programmable coordination media. Lecture Notes in Computer Science, 1282:274–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDerGoot:1997:STS

[1953] R. Van Der Goot, J. Schaeffer, and G. V. Wilson. Safer tuple spaces. Lecture Notes in Computer Science, 1282: 289–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hedman:1997:CAS

[1954] E. J. Hedman, J. N. Kok, and K. Sere.

Coordinating actions systems. Lecture Notes in Computer Science, 1282:302– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dingel:1997:AU

[1955] J. Dingel. Approximating UNITY. Lecture Notes in Computer Science, 1282: 320–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCann:1997:MUC

[1956] P. J. McCann and G.-C. Roman. Mobile UNITY coordination constructs applied to packet forwarding for mobile hosts. Lecture Notes in Computer Science, 1282:338–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Granoe:1997:LLO

[1957] K. Granoe and J. Paakki. From layer to layer: Object-oriented protocol refinement in kannel. Lecture Notes in Computer Science, 1282:355–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amadio:1997:AML

[1958] R. M. Amadio. An asynchronous model of locality, failure, and process mobility. Lecture Notes in Computer Science, 1282:374–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vion-Dury:1997:CCM

[1959] J.-Y. Vion-Dury, L. Bellissard, and V. Marangozov. A component calculus for modeling the olan configuration language. Lecture Notes in Computer Science, 1282:392–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buffo:1997:CMD

[1960] M. Buffo and D. Buchs. A coordination model for distributed object systems. Lecture Notes in Computer Science, 1282:410–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freisleben:1997:CPP

[1961] B. Freisleben and T. Kielmann. Coordination patterns for parallel computing. Lecture Notes in Computer Science, 1282:414–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kellett:1997:CMC

[1962] A. Kellett and M. Fisher. Concurrent MetateM as a coordination language. Lecture Notes in Computer Science, 1282:418–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadopoulos:1997:CBC

[1963] G. A. Papadopoulos and F. Arbab.

Control-based coordination of human and other activities in cooperative information systems. Lecture Notes in Computer Science, 1282:422–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rowstron:1997:UAT

[1964] A. Rowstron. Using asynchronous tuple-space access primitives (BONITA primitives) for process co-ordination. Lecture Notes in Computer Science, 1282:426–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tolksdorf:1997:BOO

[1965] R. Tolksdorf. Berlinda: An objectoriented platform for implementing coordination languages in Java. Lecture Notes in Computer Science, 1282:430– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller-Olm:1997:MCV

[1966] M. Müller-Olm. Modular compiler verification. A refinement-algebraic approach advocating stepwise abstraction. Lecture Notes in Computer Science, 1283:xii + 250, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amoura:1997:SIM

[1967] A. K. Amoura, E. Bampis, C. Kenyon, and Y. Manoussakis. Scheduling independent multiprocessor tasks. Lecture Notes in Computer Science, 1284:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arkin:1997:LSW

[1968] E. M. Arkin and R. Hassin. On local search for weighted k-set packing. Lecture Notes in Computer Science, 1284:13–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Azar:1997:LMC

[1969] Y. Azar and L. Epstein. On-line machine covering. Lecture Notes in Computer Science, 1284:23–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biedl:1997:AES

[1970] T. C. Biedl and M. Kaufmann. Areaefficient static and incremental graph drawings. Lecture Notes in Computer Science, 1284:37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloemer:1997:DBD

[1971] J. Bloemer. Denesting by bounded degree radicals. Lecture Notes in Computer Science, 1284:53–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandes:1997:LTA

[1972] U. Brandes and D. Wagner. A linear time algorithm for the arc disjoint Menger problem in planar directed graphs. Lecture Notes in Computer Science, 1284:64–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandstaedt:1997:DAT

[1973] A. Brandstaedt, V. Chepoi, and F. Dragan. Distance approximating trees for chordal and dually chordal graphs. Lecture Notes in Computer Science, 1284:78–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cornuejols:1997:DIP

[1974] G. Cornuejols, R. Urbaniak, R. Weismantel, and L. Wolsey. Decomposition of integer programs and of generating sets. Lecture Notes in Computer Science, 1284:92–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czumaj:1997:BDS

[1975] A. Czumaj and W. B. Strothmann. Bounded degree spanning trees. Lecture Notes in Computer Science, 1284: 104–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diks:1997:OAB

[1976] K. Diks and A. Pelc. Optimal adaptive broadcasting with a bounded fraction of faulty nodes. Lecture Notes in Computer Science, 1284:118–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djidjev:1997:WGS

[1977] H. N. Djidjev. Weighted graph separators and their applications. Lecture Notes in Computer Science, 1284:130– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fekete:1997:NEA

[1978] S. P. Fekete and J. Schepers. A new exact algorithm for general orthogonal d-dimensional knapsack problems. Lecture Notes in Computer Science, 1284:144–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1997:DDS

[1979] M. Fischer, F. Meyer auf der Heide, and W. B. Strothmann. Dynamic data structures for realtime management of large geometric scenes. Lecture Notes in Computer Science, 1284: 157–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1997:SRS

[1980] U. Foessmeier and M. Kaufmann. Solving rectilinear Steiner tree problems exactly in theory and practice. Lecture Notes in Computer Science, 1284:171– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frigioni:1997:DSV

[1981] D. Frigioni and G. F. Italiano. Dynamically switching vertices in planar graphs. Lecture Notes in Computer Science, 1284:186–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garefalakis:1997:NFR

[1982] T. Garefalakis. A new family of randomized algorithms for list accessing. Lecture Notes in Computer Science, 1284:200–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giancarlo:1997:LCT

[1983] R. Giancarlo and D. Guaiana. On-line construction of two-dimensional suffix trees. Lecture Notes in Computer Science, 1284:217–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glazebrook:1997:SMQ

[1984] K. D. Glazebrook and J. Niño-Mora.

Scheduling multiclass queueing networks on parallel servers: approximate and heavy-traffic optimality of Klimov’s priority rule. Lecture Notes in Computer Science, 1284:232–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grebinski:1997:ORG

[1985] V. Grebinski and G. Kucherov. Optimal reconstruction of graphs under the additive model. Lecture Notes in Computer Science, 1284:246–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helmberg:1997:FVS

[1986] C. Helmberg. Fixing variables in semidefinite relaxations. Lecture Notes in Computer Science, 1284:259–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henk:1997:TSK

[1987] M. Henk and R. Weismantel. Test sets of the knapsack problem and simultaneous Diophantine approximation. Lecture Notes in Computer Science, 1284:271–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwama:1997:TDM

[1988] K. Iwama and E. Miyano. Threedimensional meshes are less powerful than two-dimensional ones in oblivious routing. Lecture Notes in Computer Science, 1284:284–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalyanasundaram:1997:FTR

[1989] B. Kalyanasundaram and K. Pruhs. Fault-tolerant real-time scheduling. Lecture Notes in Computer Science, 1284:296–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kogan:1997:CGP

[1990] D. Kogan and A. Schuster. Collecting garbage pages in a distributed shared memory with reduced memory and communication overhead. Lecture Notes in Computer Science, 1284:308– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korupolu:1997:QFD

[1991] M. R. Korupolu and V. Ramachandran. Quasi-fully dynamic algorithms for two-connectivity, cycle equivalence and related problems. Lecture Notes in Computer Science, 1284:326–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krznaric:1997:MST

[1992] D. Krznaric, C. Levcopoulos, and B. J.

Nilsson. Minimum spanning trees in d dimensions. Lecture Notes in Computer Science, 1284:341–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1997:RBS

[1993] K. S. Larsen, T. Ottmann, and E. Soisalon-Soininen. Relaxed balance for search trees with local rebalancing. Lecture Notes in Computer Science, 1284:350–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller-Hannemann:1997:IAM

[1994] M. Mueller-Hannemann and K. Weihe.

Improved approximations for minimum cardinality quadrangulations of finite element meshes. Lecture Notes in Computer Science, 1284:364–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naor:1997:DSA

[1995] S. Naor, A. Orda, and Y. Petruschka. Dynamic storage allocation with known durations. Lecture Notes in Computer Science, 1284:378–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nolte:1997:CST

[1996] A. Nolte and R. Schrader. Coloring in sublinear time. Lecture Notes in Computer Science, 1284:388–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rethmann:1997:CAL

[1997] J. Rethmann and E. Wanke. Competitive analysis of on-line stack-up algorithms. Lecture Notes in Computer Science, 1284:402–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1997:SLB

[1998] A. S. Schulz and M. Skutella. Scheduling-LPs bear probabilities: randomized approximations for min-sum criteria. Lecture Notes in Computer Science, 1284:416–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Segal:1997:PSA

[1999] M. Segal. On piercing sets of axisparallel rectangles and rings. Lecture Notes in Computer Science, 1284:430– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sevastianov:1997:SPD

[2000] S. Sevastianov. Seven problems: so different yet close. Lecture Notes in Computer Science, 1284:443–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snoeyink:1997:LTR

[2001] J. Snoeyink and M. Van Kreveld.

Linear-time reconstruction of Delaunay triangulations with applications. Lecture Notes in Computer Science, 1284: 459–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trevisan:1997:ASS

[2002] L. Trevisan. Approximating satisfiable satisfiability problems. Lecture Notes in Computer Science, 1284:472– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vallee:1997:ACS

[2003] B. Vallee. Algorithms for computing signs of 2 x 2 determinants: Dynamics and average case analysis. Lecture Notes in Computer Science, 1284:486– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weihe:1997:RTC

[2004] K. Weihe and T. Willhalm. Reconstructing the topology of a CAD model — A discrete approach. Lecture Notes in Computer Science, 1284:500–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casti:1997:EPC

[2005] J. L. Casti. Emergent phenomena and computer worlds. Lecture Notes in Computer Science, 1285:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fogel:1997:TEE

[2006] L. J. Fogel. Top-evolutionary engineering. Lecture Notes in Computer Science, 1285:11–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chung:1997:FOU

[2007] C. Chung and R. G. Reynolds. Function optimization using evolutionary programming with self-adaptive cultural algorithms. Lecture Notes in Computer Science, 1285:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pan:1997:AEA

[2008] Z. Pan and L. Kang. An adaptive evolutionary algorithm for numerical optimization. Lecture Notes in Computer Science, 1285:27–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myung:1997:LBE

[2009] H. Myung and J.-H. Kim. Lagrangianbased evolutionary programming for constrained optimization. Lecture Notes in Computer Science, 1285:35– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsushita:1997:SIV

[2010] S. Matsushita, T. Furuhashi, H. Tsutsui, and Y. Uchikawa. Selection of input variables of fuzzy model using genetic algorithm with quick fuzzy inference. Lecture Notes in Computer Science, 1285:45–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1997:ESG

[2011] C.-Y. Lee and S. K. Han. Entropic sampling in genetic-entropic algorithm. Lecture Notes in Computer Science, 1285:54–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:CLS

[2012] J.-C. Chen. Computational and learning synergies with a coevolving multilevel architecture. Lecture Notes in Computer Science, 1285:63–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raik:1997:ESM

[2013] S. E. Raik and D. G. Browne. Evolving state and memory in genetic programming. Lecture Notes in Computer Science, 1285:73–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakao:1997:ECI

[2014] Z. Nakao, M. Takashibu, and Y.-W.

Chen. Evolutionary CT image reconstruction by image partitioning. Lecture Notes in Computer Science, 1285: 81–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stonier:1997:GLI

[2015] R. Stonier and D. Sturgess. Genetic learning of the irrigation cycle for water flow in cropped soils. Lecture Notes in Computer Science, 1285:89–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Song:1997:OPC

[2016] W.-K. Song and Z. Bien. Optimization of parameters of color image segmentation using evolutionary programming. Lecture Notes in Computer Science, 1285:97–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsujimura:1997:GAS

[2017] Y. Tsujimura and M. Gen. Genetic algorithms for solving multiprocessor scheduling problems. Lecture Notes in Computer Science, 1285:106–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhao:1997:SCE

[2018] Q. Zhao. A study on co-evolutionary learning of neural networks. Lecture Notes in Computer Science, 1285:116– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoshikawa:1997:KAF

[2019] T. Yoshikawa, T. Furuhashi, and Y. Uchikawa. Knowledge acquisition of fuzzy control rules for mobile robots using DNA coding method and pseudo-bacterial GA. Lecture Notes in Computer Science, 1285:126–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1997:ELA

[2020] M. W. Hwang and J. Y. Choi. Evolutionary learning algorithm for projection neural networks. Lecture Notes in Computer Science, 1285:136–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1997:ECT

[2021] X. Yao and Y. Liu. EPNet for chaotic time-series prediction. Lecture Notes in Computer Science, 1285:146–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:WSG

[2022] S.-H Chen. Would and should government lie about economic statistics: Simulations based on evolutionary cellular automata. Lecture Notes in Computer Science, 1285:157–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1997:TIC

[2023] K. P. Wong and A. Li. A technique for improving the convergence characteristic of genetic algorithms and its application to a genetic-based load flow algorithm. Lecture Notes in Computer Science, 1285:167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makita:1997:KEU

[2024] Y. Makita and M. Hagiwara. Knowledge extraction using neural network by an artificial life approach. Lecture Notes in Computer Science, 1285:177– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takeuchi:1997:IMU

[2025] I. Takeuchi and T. Furuhashi. An inference method using multiple patterns and modification of pattern space. Lecture Notes in Computer Science, 1285: 187–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kil:1997:RSB

[2026] R. M. Kil and Y. Song. Random search based on genetic operators. Lecture Notes in Computer Science, 1285:196– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1997:HEL

[2027] S.-B. Cho and S.-I. Lee. Hybrid evolutionary learning of fuzzy logic and genetic algorithm. Lecture Notes in Computer Science, 1285:206–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1997:FIU

[2028] J.-H. Lee and L. K. Hyung. Fuzzy identification of unknown systems based on GA. Lecture Notes in Computer Science, 1285:216–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nerome:1997:CCE

[2029] M. Nerome, K. Yamada, S. Endo, and H. Miyagi. Competitive co-evolution model on the acquisition of game strategy. Lecture Notes in Computer Science, 1285:224–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glaser:1997:CME

[2030] N. Glaser. The CoMoMAS methodology and environment for multi-agent system development. Lecture Notes in Computer Science, 1286:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kendall:1997:DPD

[2031] E. A. Kendall and M. T. Malkoun. Design patterns for the development of multi-agent systems. Lecture Notes in Computer Science, 1286:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lukose:1997:MEM

[2032] D. Lukose. Modelling extendible mobile agents. Lecture Notes in Computer Science, 1286:32–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DInverno:1997:MBE

[2033] M. D’Inverno and M. Luck. Making and breaking engagements: An operational analysis of agent relationships. Lecture Notes in Computer Science, 1286:48–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Decker:1997:CIA

[2034] K. Decker, K. Sycara, and M. Williamson.

Cloning for intelligent adaptive information agents. Lecture Notes in Computer Science, 1286:63–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jamison:1997:AAB

[2035] W. C. Jamison. ACACIA: An agency based collaboration framework for heterogeneous multi-agent systems. Lecture Notes in Computer Science, 1286: 76–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1997:MAC

[2036] L. He, Y. Chao, S. Kato, and T. Araki. A multi-agent cooperative reasoning system for amalgamated knowledge bases. Lecture Notes in Computer Science, 1286:92–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ekenberg:1997:DCM

[2037] L. Ekenberg. Detecting conflicts in multi-agent systems. Lecture Notes in Computer Science, 1286:106–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reed:1997:PFI

[2038] C. Reed, D. Long, M. Fox, and M. Garagnani. Persuasion as a form of inter-agent negotiation. Lecture Notes in Computer Science, 1286:120– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:MSS

[2039] M. Zhang and C. Zhang. Methodologies of solution synthesis in distributed expert systems. Lecture Notes in Computer Science, 1286:137–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1997:IAR

[2040] L. He, Y. Chao, S. Kato, and T. Araki.

Implementing an automated reasoning system for multi-agent knowledge and time. Lecture Notes in Computer Science, 1286:152–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:AMI

[2041] H. Yang and C. Zhang. Application of MAS in implementing rational IP routers on the priced Internet. Lecture Notes in Computer Science, 1286:166–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haegg:1997:SAF

[2042] S. Haegg. A sentinel approach to fault handling in multi-agent systems. Lecture Notes in Computer Science, 1286: 181–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kropf:1997:FHV

[2043] T. Kropf. Formal hardware verification: Methods and systems in comparison. Lecture Notes in Computer Science, 1287:xii + 367, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1997:SDT

[2044] M. Schneider. Spatial data types for database systems. finite resolution geometry for geographic information systems. Lecture Notes in Computer Science, 1288:xiii + 275, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1997:PSS

[2045] L. Bachmair. Paramodulation, superposition, and simplification. Lecture Notes in Computer Science, 1289:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchholz:1997:EGC

[2046] W. Buchholz. Explaining Gentzen’s consistency proof within infinitary proof theory. Lecture Notes in Computer Science, 1289:4–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buss:1997:AAT

[2047] S. R. Buss. Alogtime algorithms for tree isomorphism, comparison, and canonization. Lecture Notes in Computer Science, 1289:18–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carnielli:1997:ULG

[2048] W. A. Carnielli and P. R. S. Veloso. Ultrafilter logic and generic reasoning. Lecture Notes in Computer Science, 1289:34–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robinson:1997:IRM

[2049] J. A. Robinson. Informal rigor and mathematical understanding. Lecture Notes in Computer Science, 1289:54– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tammet:1997:RIM

[2050] T. Tammet. Resolution, inverse method and the sequent calculus. Lecture Notes in Computer Science, 1289: 65–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiuryn:1997:SL

[2051] J. Tiuryn. Subtyping over a lattice. Lecture Notes in Computer Science, 1289:84–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basin:1997:NMB

[2052] D. Basin, S. Matthews, and L. Vigano.

A new method for bounding the complexity of modal logics. Lecture Notes in Computer Science, 1289:89– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beklemishev:1997:PFI

[2053] L. D. Beklemishev. Parameter free induction and reflection. Lecture Notes in Computer Science, 1289:103–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borchert:1997:LAR

[2054] B. Borchert and F. Stephan. Looking for an analogue of Rice’s theorem in circuit complexity theory. Lecture Notes in Computer Science, 1289:114– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciabattoni:1997:TCB

[2055] A. Ciabattoni and D. Luchi. Two connections between linear logic and Lukasiewicz logics. Lecture Notes in Computer Science, 1289:128–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egly:1997:SCG

[2056] U. Egly and K. Genther. Structuring of computer-generated proofs by cut introduction. Lecture Notes in Computer Science, 1289:140–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilmore:1997:NSA

[2057] P. C. Gilmore. NaDSyl and some applications. Lecture Notes in Computer Science, 1289:153–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khakhanian:1997:MRA

[2058] V. K. Khakhanian. Markov’s rule is admissible in the set theory with intuitionistic logic. Lecture Notes in Computer Science, 1289:167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lisitsa:1997:BHT

[2059] A. Lisitsa and V. Sazonov. Bounded hyperset theory and Web-like data bases. Lecture Notes in Computer Science, 1289:172–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makowsky:1997:ID

[2060] J. A. Makowsky. Invariant definability. Lecture Notes in Computer Science, 1289:186–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matzinger:1997:CCR

[2061] R. Matzinger. Comparing computational representations of Herbrand models. Lecture Notes in Computer Science, 1289:203–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pape:1997:RTS

[2062] C. Pape and R. Haehnle. Restart tableaux with selection function. Lecture Notes in Computer Science, 1289: 219–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plisko:1997:TSL

[2063] V. Plisko. Two semantics and logics based on the goedel interpretation. Lecture Notes in Computer Science, 1289:233–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskevicius:1997:CDR

[2064] R. Pliuskevicius. On the completeness and decidability of a restricted first order linear temporal logic. Lecture Notes in Computer Science, 1289: 241–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Polacik:1997:PQI

[2065] T. Polacik. Propositional quantification in intuitionistic logic. Lecture Notes in Computer Science, 1289:255– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preining:1997:SP

[2066] N. Preining. Sketch-as-proof. Lecture Notes in Computer Science, 1289:264– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Setzer:1997:TST

[2067] A. Setzer. Translating set theoretical proofs into type theoretical programs. Lecture Notes in Computer Science, 1289:278–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

TortoraDeFalco:1997:DSP

[2068] L. Tortora De Falco. Denotational semantics for polarized (but nonconstrained) LK by means of the additives. Lecture Notes in Computer Science, 1289:290–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veanes:1997:USR

[2069] M. Veanes. The undecidability of simultaneous rigid E-unification with two variables. Lecture Notes in Computer Science, 1289:305–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Willard:1997:TRP

[2070] D. E. Willard. The tangibility reflection principle for self-verifying axiom systems. Lecture Notes in Computer Science, 1289:319–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xi:1997:UBS

[2071] H. Xi. Upper bounds for standardizations and an application. Lecture Notes in Computer Science, 1289:335– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baez:1997:IC

[2072] J. C. Baez. An introduction to ncategories. Lecture Notes in Computer Science, 1290:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bird:1997:ABA

[2073] R. S. Bird. Allegories as a basis for algorithmics. Lecture Notes in Computer Science, 1290:34–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jay:1997:SSD

[2074] C. B. Jay. Separating shape from data. Lecture Notes in Computer Science, 1290:47–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mellies:1997:FTR

[2075] P.-A Mellies. A factorisation theorem in rewriting theory. Lecture Notes in Computer Science, 1290:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lueth:1997:MMT

[2076] C. Lueth and N. Ghani. Monads and modular term rewriting. Lecture Notes in Computer Science, 1290:69– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1997:CPT

[2077] A. Corradini and F. Gadducci. A 2categorical presentation of term graph rewriting. Lecture Notes in Computer Science, 1290:87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattani:1997:PMCb

[2078] G. L. Cattani, I. Stark, and G. Winskel. Presheaf models for the π-calculus. Lecture Notes in Computer Science, 1290:106–126, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turi:1997:CMS

[2079] D. Turi. Categorical modelling of structural operational rules: Case studies. Lecture Notes in Computer Science, 1290:127–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pavlovic:1997:SIC

[2080] D. Pavlovic and S. Abramsky. Specifying interaction categories. Lecture Notes in Computer Science, 1290:147– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolter:1997:SNL

[2081] U. Wolter and A. Martini. Shedding new light in the world of logical systems. Lecture Notes in Computer Science, 1290:159–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossakowski:1997:CRL

[2082] T. Mossakowski, A. Tarlecki, and W. Pawlowski. Combining and representing logical systems. Lecture Notes in Computer Science, 1290:177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreev:1997:DAL

[2083] A. Andreev and S. Soloviev. A deciding algorithm for linear isomorphism of types with complexity O(n log_{2}(n)). Lecture Notes in Computer Science, 1290:197–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weihrauch:1997:EGM

[2084] K. Weihrauch and X. Zheng. Effectiveness of the global modulus of continuity on metric spaces. Lecture Notes in Computer Science, 1290:210–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hensel:1997:PPD

[2085] U. Hensel and B. Jacobs. Proof principles for datatypes with iterated recursion. Lecture Notes in Computer Science, 1290:220–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogendijk:1997:WDD

[2086] P. Hoogendijk and R. Backhouse.

When do datatypes commute? Lecture Notes in Computer Science, 1290: 242–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lellahi:1997:CCA

[2087] K. Lellahi and V. Tannen. A calculus for collections and aggregates. Lecture Notes in Computer Science, 1290: 261–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bucalo:1997:L

[2088] A. Bucalo and G. Rosolini. Lifting. Lecture Notes in Computer Science, 1290:281–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reus:1997:GSD

[2089] B. Reus and T. Streicher. General synthetic domain theory — A logical approach. Lecture Notes in Computer Science, 1290:293–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feitelson:1997:TPP

[2090] D. G. Feitelson, L. Rudolph, U. Schwiegelshohn, and K. C. Sevcik. Theory and practice in parallel job scheduling. Lecture Notes in Computer Science, 1291:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Downey:1997:UQT

[2091] A. B. Downey. Using queue time predictions for processor allocation. Lecture Notes in Computer Science, 1291: 35–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gibbons:1997:HAP

[2092] R. Gibbons. A historical application profiler for use by parallel schedulers. Lecture Notes in Computer Science, 1291:58–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feitelson:1997:MUL

[2093] D. G. Feitelson. Memory usage in the LANL CM-5 workload. Lecture Notes in Computer Science, 1291:78– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jann:1997:MWM

[2094] J. Jann, P. Pattnaik, H. Franke, and F. Wang. Modeling of workload in MPPs. Lecture Notes in Computer Science, 1291:95–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lagerstrom:1997:PPS

[2095] R. N. Lagerstrom and S. K. Gipp.

PScheD: Political scheduling on the CRAY T3E. Lecture Notes in Computer Science, 1291:117–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brecht:1997:EEP

[2096] T. B. Brecht. An experimental evaluation of processor pool-based scheduling for shared-memory NUMA multiprocessors. Lecture Notes in Computer Science, 1291:139–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parsons:1997:IMS

[2097] E. W. Parsons and K. C. Sevcik. Implementing multiprocessor scheduling disciplines. Lecture Notes in Computer Science, 1291:166–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanh:1997:OOA

[2098] Pham Hong Hanh and V. Simonenko. Objective-oriented algorithm for job scheduling in parallel heterogeneous systems. Lecture Notes in Computer Science, 1291:193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1997:IGS

[2099] W. Lee, M. Frank, V. Lee, and K. Mackenzie. Implications of I/O for gang scheduled workloads. Lecture Notes in Computer Science, 1291:215– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feitelson:1997:IUR

[2100] D. G. Feitelson and M. A. Jette. Improved utilization and responsiveness with gang scheduling. Lecture Notes in Computer Science, 1291:238–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hori:1997:GSD

[2101] A. Hori, H. Tezuka, and Y. Ishikawa. Global state detection using network preemption. Lecture Notes in Computer Science, 1291:262–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:PEG

[2102] F. Wang, M. Papaefthymiou, and M. Squillante. Performance evaluation of gang scheduling for parallel and distributed multiprogramming. Lecture Notes in Computer Science, 1291:277– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoy:1997:NNS

[2103] S. Antoy and Z. M. Ariola. Narrowing the narrowing space. Lecture Notes in Computer Science, 1292:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Compagnoni:1997:OCA

[2104] A. Compagnoni and M. Fernandez. An object calculus with algebraic rewriting. Lecture Notes in Computer Science, 1292:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gallardo:1997:ISM

[2105] J. E. Gallardo, F. Gutierrez, and B. C. Ruiz. Inheritance and selective method dispatching in concurrent haskell. Lecture Notes in Computer Science, 1292:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ancona:1997:OOM

[2106] D. Ancona and E. Zucca. Overriding operators in a mixin-based framework. Lecture Notes in Computer Science, 1292:47–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bowers:1997:RLP

[2107] A. F. Bowers, P. M. Hill, and F. Ibañez.

Resolution for logic programming with universal quantifiers. Lecture Notes in Computer Science, 1292:63–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gregory:1997:DAC

[2108] S. Gregory. A declarative approach to concurrent programming. Lecture Notes in Computer Science, 1292:79– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herrmann:1997:TDC

[2109] C. A. Herrmann and C. Lengauer. Transformation of divide and conquer to nested parallel loops. Lecture Notes in Computer Science, 1292:95– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paterson:1997:TLF

[2110] R. Paterson. Transforming lazy functions using comportment properties. Lecture Notes in Computer Science, 1292:111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucas:1997:TEE

[2111] S. Lucas. Transformations for efficient evaluations in functional programming. Lecture Notes in Computer Science, 1292:127–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serrano:1997:IEW

[2112] M. Serrano. Inline expansion: When and how? Lecture Notes in Computer Science, 1292:143–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossin:1997:HOV

[2113] C. Mossin. Higher-order value flow graphs. Lecture Notes in Computer Science, 1292:159–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giacobazzi:1997:IIA

[2114] R. Giacobazzi and F. Scozzari. Intuitionistic implication in abstract interpretation. Lecture Notes in Computer Science, 1292:175–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1997:OEF

[2115] M. Carlsson, G. Ottosson, and B. Carlson. An open-ended finite domain constraint solver. Lecture Notes in Computer Science, 1292:191–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giavitto:1997:SCR

[2116] J.-L. Giavitto, D. De Vito, and O. Michel. Semantics and compilation of recursive sequential streams in 81/2. Lecture Notes in Computer Science, 1292:207–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ogata:1997:ITR

[2117] K. Ogata and K. Futatsugi. Implementation of term rewritings with the evaluation strategy. Lecture Notes in Computer Science, 1292:225–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1997:RR

[2118] G. Barthe, J. Hatcliff, and M. H. Soerensen. Reflections on reflections. Lecture Notes in Computer Science, 1292: 241–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xi:1997:EUL

[2119] H. Xi. Evaluation under lambda abstraction. Lecture Notes in Computer Science, 1292:259–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1997:SRH

[2120] Z. Huang, C. Sun, and A. Sattar. Selective recomputation for handling sideeffects in parallel logic programs. Lecture Notes in Computer Science, 1292: 275–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sparud:1997:TLF

[2121] J. Sparud and C. Runciman. Tracing lazy functional computations using redex trails. Lecture Notes in Computer Science, 1292:291–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cousineau:1997:FPG

[2122] G. Cousineau. Functional programming and geometry. Lecture Notes in Computer Science, 1292:309–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thompson:1997:WDB

[2123] S. Thompson. Where do I begin? A problem solving approach in teaching functional programming. Lecture Notes in Computer Science, 1292:323– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1997:TFL

[2124] M. Hanus. Teaching functional and logic programming with a single computation model. Lecture Notes in Computer Science, 1292:335–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eisinger:1997:LIF

[2125] N. Eisinger, T. Geisler, and S. Panne.

Logic implemented functionally. Lecture Notes in Computer Science, 1292: 351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Findler:1997:DPP

[2126] R. B. Findler, C. Flanagan, M. Flatt, and S. Krishnamurthi. DrScheme: A pedagogic programming environment for scheme. Lecture Notes in Computer Science, 1292:369–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whittle:1997:EHN

[2127] J. Whittle, A. Bundy, and H. Lowe. An editor for helping novices to learn standard ML. Lecture Notes in Computer Science, 1292:389–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeRoure:1997:IDD

[2128] D. DeRoure. Introducing the declarative dungeon. Lecture Notes in Computer Science, 1292:407–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johansson:1997:PAC

[2129] E. Johansson, C. Jonsson, T. Lindgren, and J. Bevemyr. A pragmatic approach to compilation of erlang. Lecture Notes in Computer Science, 1292:419–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janousek:1997:FTD

[2130] J. Janousek and B. Melichar. Formal translations described by translation grammars with LR(kappa) input grammars. Lecture Notes in Computer Science, 1292:421–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujinami:1997:ART

[2131] N. Fujinami. Automatic run-time code generation in object-oriented languages. Lecture Notes in Computer Science, 1292:423–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munson:1997:TOT

[2132] E. V. Munson. Toward an operational theory of media. Lecture Notes in Computer Science, 1293:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandkuhl:1997:FSC

[2133] K. Sandkuhl. First steps to cross media publishing and multimodal documents. Lecture Notes in Computer Science, 1293:15–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahonen:1997:DSC

[2134] H. Ahonen. Disambiguation of SGML content models. Lecture Notes in Computer Science, 1293:27–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kilpelainen:1997:SE

[2135] P. Kilpeläinen and D. Wood. SGML and exceptions. Lecture Notes in Computer Science, 1293:39–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schroff:1997:GCS

[2136] T. Schroff and A. Brüggemann-Klein.

Grammar-compatible stylesheets. Lecture Notes in Computer Science, 1293: 51–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heather:1997:OAM

[2137] M. A. Heather and B. N. Rossiter.

Object awareness in multimedia documents. Lecture Notes in Computer Science, 1293:59–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

King:1997:LBF

[2138] P. R. King. A logic based formalism for temporal constraints in multimedia documents. Lecture Notes in Computer Science, 1293:87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myka:1997:TAH

[2139] A. Myka, H. Argenton, and U. Guentzer.

Towards automatic hypertextual representation of linear texts. Lecture Notes in Computer Science, 1293:103–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kosmynin:1997:UBC

[2140] A. Kosmynin and I. Davidson. Using background contextual knowledge for documents representation. Lecture Notes in Computer Science, 1293:123– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dharap:1997:TSD

[2141] C. Dharap and C. Mic Bowman. Typed structured documents for information retrieval. Lecture Notes in Computer Science, 1293:135–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murata:1997:TDS

[2142] M. Murata. Transformation of documents and schemas by patterns and contextual conditions. Lecture Notes in Computer Science, 1293:153–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:TFP

[2143] X. Wang and D. Wood. Tabular formatting problems. Lecture Notes in Computer Science, 1293:171–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minas:1997:VDV

[2144] M. Minas and L. Shklar. Visual definition of virtual documents for the World-Wide-Web. Lecture Notes in Computer Science, 1293:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldmann:1997:CCH

[2145] Mikael Goldmann and Mats Näslund.

The complexity of computing hard core predicates. Lecture Notes in Computer Science, 1294:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940001.
pdf.

Fujisaki:1997:SZK

[2146] Eiichiro Fujisaki and Tatsuaki Okamoto.

Statistical zero knowledge protocols to prove modular polynomial relations. Lecture Notes in Computer Science, 1294:16–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1294/12940016.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1294/12940016.
pdf.

Crescenzo:1997:KSV

[2147] Giovanni Di Crescenzo, Tatsuaki Okamoto, and Moti Yung. Keeping the SZK-verifier honest unconditionally. Lecture Notes in Computer Science, 1294:31–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1294/12940031.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1294/12940031.
pdf.

DiCrescenzo:1997:KSV

[2148] G. Di Crescenzo, T. Okamoto, and M. Yung. Keeping the SZK-Verifier honest unconditionally. Lecture Notes in Computer Science, 1294:31–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldreich:1997:FMC

[2149] Oded Goldreich. On the foundations of modern cryptography. Lecture Notes in Computer Science, 1294:46– 74, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1294/12940046.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940046.
                                        pdf.

Beaver:1997:PPE

[2150] Donald Beaver. Plug and play encryption. Lecture Notes in Computer Science, 1294:75–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1294/12940075.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1294/12940075.pdf.

Canetti:1997:DE

[2151] Ran Canetti, Cynthia Dwork, Moni Naor, and Rafail Ostrovsky. Deniable encryption. Lecture Notes in Computer Science, 1294:90–104, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL ftp://theory.lcs.mit.edu/
                                        pub/tcryptol/96-02r.ps;   http:/

/link.springer-ny.com/link/service/
series/0558/bibs/1294/12940090.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1294/12940090.pdf.

Goldreich:1997:EDE

[2152] Oded Goldreich, Shafi Goldwasser, and Shai Halevi. Eliminating decryption errors in the Ajtai– Dwork cryptosystem. Lecture Notes in Computer Science, 1294:105–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940105.
pdf.

Goldreich:1997:PKC

[2153] Oded Goldreich, Shafi Goldwasser, and Shai Halevi. Public-key cryptosystems from lattice reduction problems. Lecture Notes in Computer Science, 1294: 112–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940112.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940112.
pdf.

Gennaro:1997:RBU

[2154] Rosario Gennaro, Hugo Krawczyk, and Tal Rabin. RSA-based undeniable signatures. Lecture Notes in Computer Science, 1294:132–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1294/12940132.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940132.
                                        pdf.

Juels:1997:SBD

[2155] Ari Juels, Michael Luby, and Rafail Ostrovsky. Security of blind digital signatures. Lecture Notes in Computer Science, 1294:150–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1294/12940150.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940150.
                                        pdf.

Zheng:1997:DSH

[2156] Yuliang Zheng. Digital signcryption or how to achieve cost (signature & encryption) << cost(signature) + cost(encryption). Lecture Notes in Computer Science, 1294:165–179, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1294/12940165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940165.
                                        pdf.

Gennaro:1997:HSD

[2157] Rosario Gennaro and Pankaj Rohatgi. How to sign digital streams. Lecture Notes in Computer Science, 1294:180– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1294/12940180.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940180.
pdf.

Nguyen:1997:MHR

[2158] Phong Nguyen and Jacques Stern.

Merkle–Hellman revisited: A cryptanalysis of the Qu-Vanstone cryptosystem based on group factorizations. Lecture Notes in Computer Science, 1294: 198–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940198.
pdf.

Berson:1997:FMP

[2159] Thomas A. Berson. Failure of the McEliece public-key cryptosystem under message-resend and relatedmessage attack. Lecture Notes in Computer Science, 1294:213–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940213.
pdf.

Misarsky:1997:MAU

[2160] Jean-François Misarsky. A multiplicative attack using LLL algorithm on RSA signatures with redundancy. Lecture Notes in Computer Science, 1294: 221–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1294/12940221.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940221.
                                        pdf.

Bleichenbacher:1997:SKP

[2161] Daniel Bleichenbacher. On the security of the KMOV public key cryptosystem. Lecture Notes in Computer Science, 1294:235–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1294/12940235.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940235.
                                        pdf.

Lim:1997:KRA

[2162] Chae Hoon Lim and Pil Joong Lee. A key recovery attack on discrete log-based schemes using a prime order subgroup. Lecture Notes in Computer Science, 1294:249–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1294/12940249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940249.
                                        pdf.

Young:1997:PKA

[2163] Adam Young and Moti Yung. The prevalence of kleptographic attacks on discrete-log based cryptosystems. Lecture Notes in Computer Science, 1294: 264–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1294/12940264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940264.
pdf.

Bellare:1997:PRN

[2164] Mihir Bellare, Shafi Goldwasser, and Daniele Micciancio. “pseudo-random” number generation within cryptographic algorithms: The DSS case. Lecture Notes in Computer Science, 1294: 277–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940277.
pdf.

Cachin:1997:USA

[2165] Christian Cachin and Ueli M. Maurer. Unconditional security against memory-bounded adversaries. Lecture Notes in Computer Science, 1294:292– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940292.
pdf.

Maurer:1997:PAS

[2166] Ueli M. Maurer and Stefan Wolf.

Privacy amplification secure against active adversaries. Lecture Notes in Computer Science, 1294:307–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1294/12940307.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940307.
                                        pdf.

Naor:1997:VAI

[2167] Moni Naor and Benny Pinkas. Visual authentication and identification. Lecture Notes in Computer Science, 1294: 322–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1294/12940322.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940322.
                                        pdf.

Brassard:1997:QIP

[2168] Gilles Brassard. Quantum information processing: The good, the bad and the ugly. Lecture Notes in Computer Science, 1294:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1294/12940337.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940337.
                                        pdf.

Guajardo:1997:EAE

[2169] Jorge Guajardo and Christof Paar. Efficient algorithms for elliptic curve cryptosystems. Lecture Notes in Computer Science, 1294:342–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/series/0558/bibs/1294/12940342.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
1294/12940342.pdf.

Solinas:1997:IAA

[2170] Jerome A. Solinas. An improved algorithm for arithmetic on a family of elliptic curves. Lecture Notes in Computer Science, 1294:357–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940357.
pdf.

Takagi:1997:FRT

[2171] Tsuyoshi Takagi. Fast RSA-type cryptosystems using n-adic expansion. Lecture Notes in Computer Science, 1294: 372–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940372.
pdf.

Buchmann:1997:OWF

[2172] Johannes A. Buchmann and Sachar Paulus. A one way function based on ideal arithmetic in number fields. Lecture Notes in Computer Science, 1294: 385–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940385.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940385.
pdf.

Dolev:1997:EAM

[2173] Shlomi Dolev and Rafail Ostrovsky. Efficient anonymous multicast and reception. Lecture Notes in Computer Science, 1294:395–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1294/12940395.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940395.
                                        pdf.

Camenisch:1997:EGS

[2174] Jan L. Camenisch and Markus A. Stadler. Efficient group signature schemes for large groups. Lecture Notes in Computer Science, 1294:410– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1294/12940410.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940410.
                                        pdf.

Boneh:1997:EGS

[2175] Dan Boneh and Matthew Franklin. Efficient generation of shared RSA keys. Lecture Notes in Computer Science, 1294:425–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1294/12940425.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1294/12940425.
                                        pdf.

Frankel:1997:PR

[2176] Yair Frankel, Peter Gemmell, Philip D.

MacKenzie, and Moti Yung. Proactive RSA. Lecture Notes in Computer Science, 1294:440–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1294/12940440.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1294/12940440.pdf.

Canetti:1997:TRR

[2177] Ran Canetti. Towards realizing random oracles: Hash functions that hide all partial information. Lecture Notes in Computer Science, 1294:455– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940455.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940455.
pdf.

Bellare:1997:CRH

[2178] Mihir Bellare and Phillip Rogaway.

Collision-resistant hashing: Towards making UOWHFs practical. Lecture Notes in Computer Science, 1294:470– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940470.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940470.
pdf.

Knudsen:1997:FSH

[2179] Lars Ramkilde Knudsen and Bart Preneel. Fast and secure hashing based on codes. Lecture Notes in Computer Science, 1294:485–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940485.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940485.
pdf.

Golic:1997:EDC

[2180] Jovan Dj. Golic and Renato Menicocci. Edit distance correlation attack on the alternating step generator. Lecture Notes in Computer Science, 1294: 499–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940499.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940499.
pdf.

Biham:1997:DFA

[2181] Eli Biham and Adi Shamir. Differential fault analysis of secret key cryptosystems. Lecture Notes in Computer Science, 1294:513–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940513.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940513.
pdf.

Wagner:1997:CCM

[2182] David Wagner, Bruce Schneier, and John Kelsey. Cryptanalysis of the Cellular Message Encryption Algorithm. Lecture Notes in Computer Science, 1294:526–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1294/12940526.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1294/12940526.
pdf;   http://www.counterpane.com/
cmea-abstract.html.

Anonymous:1997:AIc

[2183] Anonymous. Author index. Lecture Notes in Computer Science, 1294:539– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1294/1294indx.pdf.

Abramsky:1997:GSP

[2184] S. Abramsky. Game semantics for programming languages. Lecture Notes in Computer Science, 1295:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babai:1997:CC

[2185] L. Babai. Communication complexity. Lecture Notes in Computer Science, 1295:5–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1997:TAT

[2186] H. L. Bodlaender. Treewidth: Algorithmic techniques and results. Lecture Notes in Computer Science, 1295:19– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1997:WTR

[2187] N. Dershowitz. When are two rewrite systems more than none? Lecture Notes in Computer Science, 1295:37– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dwork:1997:PAL

[2188] C. Dwork. Positive applications of lattices to cryptography. Lecture Notes in Computer Science, 1295:44–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1997:TBC

[2189] G. L. Ferrari and U. Montanari. A tile-based coordination view of asynchronous pi-calculus. Lecture Notes in Computer Science, 1295:52–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1997:CCS

[2190] J. Hromkovic and G. Schnitger. Communication complexity and sequential computation. Lecture Notes in Computer Science, 1295:71–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krajicek:1997:LBP

[2191] J. Krajicek. Lower bounds for a proof system with an exponential speedup over constant-depth Frege systems and over polynomial calculus. Lecture Notes in Computer Science, 1295:85– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liskiewicz:1997:CLS

[2192] M. Liskiewicz and R. Reischuk. Computational limitations of stochastic Turing machines and Arthur–Merlin games with small space bounds. Lecture Notes in Computer Science, 1295: 91–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roth:1997:LPK

[2193] D. Roth. Learning to perform knowledge-intensive inferences. Lecture Notes in Computer Science, 1295:108– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoening:1997:RPE

[2194] U. Schoening. Resolution proofs, exponential bounds, and Kolmogorov complexity. Lecture Notes in Computer Science, 1295:110–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Afrati:1997:EDC

[2195] F. Afrati, L. Guessarian, and M. De Rougemont. The expressiveness of datalog circuits (DAC). Lecture Notes in Computer Science, 1295:119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mundhenk:1997:CPE

[2196] M. Mundhenk, J. Goldsmith, and E. Allender. The complexity of policy evaluation for finite-horizon partiallyobservable Markov decision processes. Lecture Notes in Computer Science, 1295:129–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernardinello:1997:CTS

[2197] L. Bernardinello and L. Pomello. A category of transition systems and its relations with orthomodular posets. Lecture Notes in Computer Science, 1295: 139–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berard:1997:AZW

[2198] B. Bérard and C. Picaronny. Accepting Zeno words without making time stand still. Lecture Notes in Computer Science, 1295:149–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bollig:1997:CTR

[2199] B. Bollig and I. Wegener. Complexity theoretical results on partitioned (nondeterministic) binary decision diagrams. Lecture Notes in Computer Science, 1295:159–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonsangue:1997:SCU

[2200] M. M. Bonsangue and J. N. Kok. Specifying computations using hyper transition systems. Lecture Notes in Computer Science, 1295:169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattaneo:1997:SIM

[2201] G. Cattaneo, E. Formenti, L. Margara, and J. Mazoyer. A shift-invariant metric on S0Z inducing a non-trivial topology. Lecture Notes in Computer Science, 1295:179–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:SCC

[2202] G. Chen. Subtyping calculus of construction. Lecture Notes in Computer Science, 1295:189–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choffrut:1997:DBL

[2203] C. Choffrut and G. Pighizzini. Distances between languages and reflexivity of relations. Lecture Notes in Computer Science, 1295:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ryl:1997:PCS

[2204] I. Ryl, Y. Roos, and M. Clerbout. Partial characterization of synchronization languages. Lecture Notes in Computer Science, 1295:209–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heckel:1997:IST

[2205] R. Heckel, H. Ehrig, U. Wolter, and A. Corradini. Integrating the specification techniques of graph transformation and temporal logic. Lecture Notes in Computer Science, 1295:219– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drewes:1997:GTH

[2206] F. Drewes. On the generation of trees by hyperedge replacement. Lecture Notes in Computer Science, 1295:229– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernau:1997:RV

[2207] H. Fernau and R. Stiebe. Regulation by valences. Lecture Notes in Computer Science, 1295:239–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fokkink:1997:SCT

[2208] W. Fokkink and J. Van de Pol. Simulation as a correct transformation of rewrite systems. Lecture Notes in Computer Science, 1295:249–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gavoilie:1997:DIR

[2209] C. Gavoilie. On the dilation of interval routing. Lecture Notes in Computer Science, 1295:259–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khasidashvili:1997:RCF

[2210] Z. Khasidashvili and J. Glauert. Relating conflict-free stable transition and event models. Lecture Notes in Computer Science, 1295:269–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goerdt:1997:GCT

[2211] A. Goerdt. The giant component threshold for random regular graphs with edge faults. Lecture Notes in Computer Science, 1295:279–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzer:1997:MHF

[2212] M. Holzer. Multi-head finite automata: Data-independent versus data-dependent computations. Lecture Notes in Computer Science, 1295:299– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwama:1997:CFS

[2213] K. Iwama. Complexity of finding short resolution proofs. Lecture Notes in Computer Science, 1295:309–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jukna:1997:PVN

[2214] S. Jukna, A. Razborov, P. Savicky, and I. Wegener. On P versus NP intersection co-NP for decision trees and read-once branching programs. Lecture Notes in Computer Science, 1295:319– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karner:1997:CAF

[2215] G. Karner and W. Kuich. A characterization of abstract families of algebraic power series. Lecture Notes in Computer Science, 1295:327–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1997:RDL

[2216] Y. Kobayashi and F. Otto. Repetitiveness of D0L-Languages is decidable in polynomial time. Lecture Notes in Computer Science, 1295:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolpakov:1997:MLF

[2217] R. Kolpakov and G. Kucherov. Minimal letter frequency in n-th powerfree binary words. Lecture Notes in Computer Science, 1295:347–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korec:1997:RTG

[2218] I. Korec. Real-time generation of primes by a one-dimensional cellular automaton with 11 states. Lecture Notes in Computer Science, 1295:358– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krznaric:1997:OAC

[2219] D. Krznaric and C. Levcopoulos. Optimal algorithms for complete linkage clustering in d dimensions. Lecture Notes in Computer Science, 1295:368– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manzini:1997:ILC

[2220] G. Manzini and L. Margara. Invertible linear cellular automata over zm: Algorithmic and dynamical aspects. Lecture Notes in Computer Science, 1295:378–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin-Vide:1997:TLC

[2221] C. Martin-Vide, J. Miquel-Verges, and G. Paun. Two-level contextual grammars: The internal case. Lecture Notes in Computer Science, 1295:388– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meer:1997:CPR

[2222] K. Meer. Counting problems over the reals. Lecture Notes in Computer Science, 1295:398–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinel:1997:ISE

[2223] C. Meinel and T. Theobald. On the influence of the state encoding on OBDD-Representations of finite state machines. Lecture Notes in Computer Science, 1295:408–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1997:DTF

[2224] R. Meyer and A. Petit. Decomposition of TrPTL formulas. Lecture Notes in Computer Science, 1295:418–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mundhenk:1997:NHS

[2225] M. Mundhenk. NP-hard sets have many hard instances. Lecture Notes in Computer Science, 1295:428–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nickelsen:1997:DVL

[2226] A. Nickelsen. Deciding verbose languages with linear advice. Lecture Notes in Computer Science, 1295:438– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petersen:1997:HIS

[2227] H. Petersen. Homomorphic images of sentential forms and terminating grammars. Lecture Notes in Computer Science, 1295:448–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plump:1997:SOT

[2228] D. Plump. Simplification orders for term graph rewriting. Lecture Notes in Computer Science, 1295:458–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rensink:1997:DBA

[2229] A. Rensink and H. Wehrheim. Dependency-based action refinement. Lecture Notes in Computer Science, 1295:468–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Savicky:1997:HKB

[2230] P. Savicky and S. Zak. A hierarchy for (1, +kappa)-branching programs with respect to kappa. Lecture Notes in Computer Science, 1295:478–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sibeyn:1997:RFS

[2231] J. F. Sibeyn. Routing with finite speeds of memory and network. Lecture Notes in Computer Science, 1295: 488–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tyszkiewicz:1997:QAC

[2232] J. Tyszkiewicz. Queries and algorithms computable by polynomial time existential reflective machines. Lecture Notes in Computer Science, 1295:498– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1997:POS

[2233] W. Vogler. Partial order semantics and read arcs. Lecture Notes in Computer Science, 1295:508–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leonardis:1997:CCR

[2234] A. Leonardis and H. Bischof. Computational complexity reduction in eigenspace approaches. Lecture Notes in Computer Science, 1296:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruske:1997:AID

[2235] J. Bruske and G. Sommer. An algorithm for intrinsic dimensionality estimation. Lecture Notes in Computer Science, 1296:9–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pauwels:1997:FUC

[2236] E. J. Pauwels, P. Fiddelaers, and F. Mindru. Fully unsupervised clustering using center-surround receptive fields with applications to colorsegmentation. Lecture Notes in Computer Science, 1296:17–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1997:MSF

[2237] M. L. Williams, E. R. Hancock, and R. C. Wilson. Multi-sensor fusion with Bayesian inference. Lecture Notes in Computer Science, 1296:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lanser:1997:MVB

[2238] S. Lanser, C. Zierl, O. Munkelt, and B. Radig. MORAL — A visionbased object recognition system for autonomous mobile systems. Lecture Notes in Computer Science, 1296:33– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denzler:1997:RTP

[2239] J. Denzler and H. Niemann. Realtime pedestrian tracking in natural scenes. Lecture Notes in Computer Science, 1296:42–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Surendro:1997:NRO

[2240] K. Surendro and Y. Anzai. Nonrigid object recognition using principal component analysis and geometric hashing. Lecture Notes in Computer Science, 1296:50–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mustafa:1997:OIS

[2241] A. A. Y. Mustafa, L. G. Shapiro, and M. A. Ganter. Object identification with surface signatures. Lecture Notes in Computer Science, 1296:58– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Csurka:1997:CPP

[2242] G. Csurka and O. Faugeras. Computing projective and permutation invariants of points and lines. Lecture Notes in Computer Science, 1296:66–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suk:1997:PPP

[2243] T. Suk and J. Flusser. Point projective and permutation invariants. Lecture Notes in Computer Science, 1296: 74–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lasenby:1997:CPI

[2244] J. Lasenby and E. Bayro-Corrochano.

Computing 3D projective invariants from points and lines. Lecture Notes in Computer Science, 1296:82–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muresan:1997:GTI

[2245] L. Muresan. 2D → 2D geometric transformation invariant to arbitrary translations, rotations and scales. Lecture Notes in Computer Science, 1296:90– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aksak:1997:EFD

[2246] I. Aksak, C. Feist, V. Kijko, and R. Knoefel. Extraction of filled-in data from color forms. Lecture Notes in Computer Science, 1296:98–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buzug:1997:IVS

[2247] T. M. Buzug, C. Lorenz, and J. Weese.

Improvement of vessel segmentation by elastically compensated patient motion in digital subtraction angiography images. Lecture Notes in Computer Science, 1296:106–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herzog:1997:TDQ

[2248] A. Herzog, G. Krell, B. Michaelis, and J. Wang. Three-dimensional quasibinary image restoration for confocal microscopy and its application to dendritic trees. Lecture Notes in Computer Science, 1296:114–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bors:1997:MFI

[2249] A. G. Bors, W. Puech, I. Pitas, and J.-M. Chassery. Mosaicing of flattened images from straight homogeneous generalized cylinders. Lecture Notes in Computer Science, 1296:122–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kozera:1997:WPL

[2250] R. Kozera and R. Klette. Wellposedness of linear shape-from-shading problem. Lecture Notes in Computer Science, 1296:130–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuzikov:1997:CCS

[2251] A. Tuzikov and H. J. A. M. Heijmans. Comparing convex shapes using Minkowski addition. Lecture Notes in Computer Science, 1296:138–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kenmochi:1997:DDO

[2252] Y. Kenmochi and A. Imiya. Deformation of discrete object surfaces. Lecture Notes in Computer Science, 1296: 146–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chernov:1997:NAN

[2253] V. M. Chernov and A. V. Shabashev.

Non-archimedean normalized fields in texture analysis tasks. Lecture Notes in Computer Science, 1296:154–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kazantsev:1997:RTB

[2254] I. G. Kazantsev. The radon transformbased analysis of bidirectional structural textures. Lecture Notes in Computer Science, 1296:162–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chetverikov:1997:TSD

[2255] D. Chetverikov and K. Gede. Textures and structural defects. Lecture Notes in Computer Science, 1296:167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pollefeys:1997:SCA

[2256] M. Pollefeys and L. Van Cool. Selfcalibration from the absolute conic on the plane at infinity. Lecture Notes in Computer Science, 1296:175–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Svoboda:1997:BCC

[2257] T. Svoboda and P. Sturm. A badly calibrated camera in ego-motion estimation-propagation of uncertainty. Lecture Notes in Computer Science, 1296:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blake:1997:CCR

[2258] S. Blake. 6DOF calibration of a camera with respect to the wrist of a 5axis machine tool. Lecture Notes in Computer Science, 1296:191–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koller:1997:ACC

[2259] D. Koller, G. Klinker, E. Rose, and D. Breen. Automated camera calibration and 3D egomotion estimation for augmented reality applications. Lecture Notes in Computer Science, 1296:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farid:1997:ORE

[2260] H. Farid and E. P. Simoncelli. Optimally rotation-equivariant directional derivative kernels. Lecture Notes in Computer Science, 1296:207–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stammberger:1997:HFS

[2261] T. Stammberger, M. Michaelis, M. Reiser, and K.-H. Englmeier. A hierarchical filter scheme for efficient corner detection. Lecture Notes in Computer Science, 1296:215–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branca:1997:DDL

[2262] A. Branca, F. P. Lovergine, G. Attolico, and A. Distante. Defect detection on leather by oriented singularities. Lecture Notes in Computer Science, 1296:223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quan:1997:UAR

[2263] L. Quan and R. Mohr. Uniqueness of 3D affine reconstruction of lines with affine cameras. Lecture Notes in Computer Science, 1296:231–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanatani:1997:SEA

[2264] K. Kanatani. Self-evaluation for active vision by the geometric information criterion. Lecture Notes in Computer Science, 1296:247–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mendelsohn:1997:DTR

[2265] J. Mendelsohn, E. Simoncelli, and R. Bajcsy. Discrete-time rigidityconstrained optical flow. Lecture Notes in Computer Science, 1296:255–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiemker:1997:ISS

[2266] R. Wiemker. An iterative spectralspatial Bayesian labeling approach for unsupervised robust change detection on remotely sensed multispectral imagery. Lecture Notes in Computer Science, 1296:263–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smolka:1997:CEB

[2267] B. Smolka and K. W. Wojciechowski.

Contrast enhancement of badly illuminated images based on Gibbs distribution and random walk model. Lecture Notes in Computer Science, 1296:271– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hlavac:1997:ANL

[2268] V. Hlavac and J. Fojtik. Adaptive nonlinear predictor for lossless image compression. Lecture Notes in Computer Science, 1296:279–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:BSR

[2269] Z. Yang and S. Ma. Beyond standard regularization theory. Lecture Notes in Computer Science, 1296:289–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henkel:1997:FSC

[2270] R. D. Henkel. Fast stereovision by coherence detection. Lecture Notes in Computer Science, 1296:297–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menard:1997:SMU

[2271] C. Menard and A. Leonardis. Stereo matching using M-estimators. Lecture Notes in Computer Science, 1296:305– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lan:1997:RLB

[2272] Z.-D. Lan and R. Mohr. Robust location based partial correlation. Lecture Notes in Computer Science, 1296: 313–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hansen:1997:OSD

[2273] M. Hansen, K. Daniilidis, and O. Sommer. Optimization of stereo disparity estimation using the instantaneous frequency. Lecture Notes in Computer Science, 1296:321–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiskott:1997:SMC

[2274] L. Wiskott. Segmentation from motion: Combining Gabor- and Mallatwavelets to overcome aperture and correspondence problem. Lecture Notes in Computer Science, 1296:329–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weitzel:1997:CSR

[2275] L. Weitzel, K. Kopecz, C. Spengler, and R. Eckhorn. Contour segmentation with recurrent neural networks of pulse-coding neurons. Lecture Notes in Computer Science, 1296:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czuni:1997:MMB

[2276] L. Czuni, T. Sziranyi, and J. Zerubia. Multigrid MRF based picture segmentation with cellular neural networks. Lecture Notes in Computer Science, 1296:345–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1997:CSC

[2277] L. Williams, T. Wang, and K. Thornber. Computing stochastic completion fields in linear-time using a resolution pyramid. Lecture Notes in Computer Science, 1296:353–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krebs:1997:BNO

[2278] B. Krebs, M. Burkhardt, and F. M. Wahl. A Bayesian network for 3D object recognition in range data. Lecture Notes in Computer Science, 1296:361– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalocsai:1997:ISR

[2279] P. Kalocsai. Improving the shape recognition performance of a model with Gabor filter representation. Lecture Notes in Computer Science, 1296: 369–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohr:1997:BDV

[2280] R. Mohr, S. Picard, and C. Schmid. Bayesian decision versus voting for image retrieval. Lecture Notes in Computer Science, 1296:376–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kroener:1997:SNN

[2281] S. Kroener. A structured neural network invariant to cyclic shifts and rotations. Lecture Notes in Computer Science, 1296:384–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heijmans:1997:MGO

[2282] H. J. A. M. Heijmans. Morphological grain operators for binary images. Lecture Notes in Computer Science, 1296:392–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palagyi:1997:PST

[2283] K. Palagyi and A. Kubas. A parallel 12subiteration 3D thinning algorithm to extract medial lines. Lecture Notes in Computer Science, 1296:400–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodehorst:1997:AIS

[2284] V. Rodehorst. Architectural image segmentation using digital watersheds. Lecture Notes in Computer Science, 1296:408–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vavoulidis:1997:MIC

[2285] C. P. Vavoulidis and I. Pitas. Morphological iterative closest point algorithm. Lecture Notes in Computer Science, 1296:416–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pernus:1997:PMV

[2286] F. Pernus, A. Leonardis, and S. Kovacic. Planning multiple views for 3-D object recognition and pose determination. Lecture Notes in Computer Science, 1296:424–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christy:1997:FRO

[2287] S. Christy and R. Horaud. Fast and reliable object pose estimation from line correspondences. Lecture Notes in Computer Science, 1296:432–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poesl:1997:SDO

[2288] J. Poesl and H. Niemann. Statistical a-D object localization without segmentation using wavelet analysis. Lecture Notes in Computer Science, 1296:440–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1997:RTM

[2289] S. F. Li, W. Hsu, and H. K. Pung. A real-time monocular vision-based 3D mouse. Lecture Notes in Computer Science, 1296:448–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiskott:1997:FRE

[2290] L. Wiskott, J.-M. Fellous, N. Krueger, and C. Von der Malsburg. Face recognition by elastic bunch graph matching. Lecture Notes in Computer Science, 1296:456–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feraud:1997:CMN

[2291] R. Feraud, O. Bernier, J.-E. Viallet, and M. Collobert. A conditional mixture of neural networks for face detection, applied to locating and tracking an individual speaker. Lecture Notes in Computer Science, 1296:464–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1997:LUF

[2292] K. Yu, X. Jiang, and H. Bunke. Lipreading using Fourier transform over time. Lecture Notes in Computer Science, 1296:472–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiskott:1997:PFF

[2293] L. Wiskott. Phantom faces for face analysis. Lecture Notes in Computer Science, 1296:480–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasteratos:1997:NHS

[2294] A. Gasteratos, I. Andreadis, and P. Tsalides. A new hardware structure for implementation of soft morphological filters. Lecture Notes in Computer Science, 1296:488–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kovalev:1997:MAA

[2295] V. A. Kovalev and Y. S. Bondar. A method for anisotropy analysis of 3D images. Lecture Notes in Computer Science, 1296:495–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lagunovsky:1997:FLR

[2296] D. Lagunovsky and S. Ablameyko. Fast line and rectangle detection by clustering and grouping. Lecture Notes in Computer Science, 1296:503–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1997:ORO

[2297] L.-X. Zhou and W.-K. Gu. 1st and 2nd order recursive operators for adaptive edge detection. Lecture Notes in Computer Science, 1296:511–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasinski:1997:SNI

[2298] A. J. Kasinski. Smoothing noisy images without destroying predefined feature carriers. Lecture Notes in Computer Science, 1296:519–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skarbek:1997:LSM

[2299] W. Skarbek, M. Ghuwar, and K. Ignasiak. Local subspace method for pattern recognition. Lecture Notes in Computer Science, 1296:527–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paschos:1997:TEN

[2300] G. Paschos. Testing the effectiveness of non-linear rectification on Gabor energy. Lecture Notes in Computer Science, 1296:535–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chudy:1997:NLT

[2301] L. Chudy and V. Chudy. Neural-like thinning processing. Lecture Notes in Computer Science, 1296:543–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aksak:1997:DOG

[2302] I. Aksak, C. Feist, V. Kijko, and R. Knoefel. Detection of the objects with given shape on the grey-valued pictures. Lecture Notes in Computer Science, 1296:551–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aherne:1997:APS

[2303] F. Aherne, P. Rockett, and N. Thacker.

Automatic parameter selection for object recognition using a parallel multiobjective genetic algorithm. Lecture Notes in Computer Science, 1296:559– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Negi:1997:UTS

[2304] A. Negi, P. Sreedevi, and P. Debbarma. Unsupervised texture segmentation using Hermite transform filters. Lecture Notes in Computer Science, 1296:567–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarukhanyan:1997:DHM

[2305] H. G. Sarukhanyan. Decomposition of the Hadamard matrices and Fast Hadamard Transform. Lecture Notes in Computer Science, 1296:575–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zunic:1997:CDD

[2306] J. Zunic and N. Sladoje. A characterization of digital disks by discrete moments. Lecture Notes in Computer Science, 1296:582–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chernov:1997:OSS

[2307] V. M. Chernov and M. A. Chicheva.

One-step short-length DCT algorithms with data representation in the direct sum of the associative algebras. Lecture Notes in Computer Science, 1296: 590–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chutatape:1997:CES

[2308] O. Chutatape and L. Li. Character extraction from scene image using fuzzy entropy and rule-based technique. Lecture Notes in Computer Science, 1296:597–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carreno:1997:FIR

[2309] D. Carreño and X. Ginesta. Facial image recognition using neural networks and genetic algorithms. Lecture Notes in Computer Science, 1296:605– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kovalev:1997:EMA

[2310] V. Kovalev and M. Petrou. An energy minimisation approach to the registration, matching and recognition of images. Lecture Notes in Computer Science, 1296:613–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chernov:1997:EFC

[2311] V. M. Chernov and M. V. Pershina. “error-free” calculation of the convolution using generalized Mersenne and Fermat transforms over algebraic fields. Lecture Notes in Computer Science, 1296:621–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smolka:1997:NMT

[2312] B. Smolka and K. W. Wojciechowski. A new method of texture binarization. Lecture Notes in Computer Science, 1296:629–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nickolay:1997:POI

[2313] B. Nickolay, B. Schneider, and S. Jacob. Parameter optimisation of an image processing system using evolutionary algorithms. Lecture Notes in Computer Science, 1296:637–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogs:1997:ALU

[2314] A. Hoogs. Analysis of learning using segmentation models. Lecture Notes in Computer Science, 1296:645–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reulke:1997:SPI

[2315] R. Reulke. Stereo processing of image data from the airborne CCD scanner WAAC. Lecture Notes in Computer Science, 1296:653–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fan:1997:AMC

[2316] C. F. Fan, Z. M. Li, X. Q. Ye, and W. K. Gu. An adaptive method of color road segmentation. Lecture Notes in Computer Science, 1296:661–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohta:1997:OFD

[2317] N. Ohta. Optical flow detection using a general noise model for gradient constraint. Lecture Notes in Computer Science, 1296:669–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolesnik:1997:ASS

[2318] M. Kolesnik and G. Paar. Algorithmic solution and simulation results for vision-based autonomous mode of a planetary rover. Lecture Notes in Computer Science, 1296:677–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchiz:1997:FFB

[2319] J. M. Sanchiz, F. Pla, and J. A. Marchant. A framework for featurebased motion recovery in ground plane vehicle navigation. Lecture Notes in Computer Science, 1296:686–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gimelfarb:1997:TRM

[2320] G. L. Gimel’farb and R. M. Haralick. Terrain reconstruction from multiple views. Lecture Notes in Computer Science, 1296:694–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boluda:1997:DMI

[2321] J. A. Boluda, J. Domingo, F. Pardo, and J. Pelechano. Detecting motion independent of the camera movement through a log-polar differential approach. Lecture Notes in Computer Science, 1296:702–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnspang:1997:CFC

[2322] J. Arnspang, K. Henriksen, and N. Olsen. Coordinate-free camera calibration. Lecture Notes in Computer Science, 1296:710–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klingspohr:1997:PRT

[2323] H. Klingspohr, T. Block, and R.-R.

Grigat. A passive real-time gaze estimation system for human-machine interfaces. Lecture Notes in Computer Science, 1296:718–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fellenz:1997:AVS

[2324] W. Fellenz, K. Schluens, A. Koschan, and M. Teschner. An active vision system for obtaining high resolution depth information. Lecture Notes in Computer Science, 1296:726–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fayyad:1997:KDD

[2325] U. Fayyad. Knowledge discovery in databases: An overview. Lecture Notes in Computer Science, 1297:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gottlob:1997:CSI

[2326] G. Gottlob, N. Leone, and F. Scarcello.

On the complexity of some inductive logic programming problems. Lecture Notes in Computer Science, 1297:17– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puget:1997:ILP

[2327] J.-F. Puget. Inductive logic programming and constraint logic programming (abstract). Lecture Notes in Computer Science, 1297:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexin:1997:LPR

[2328] Z. Alexin, J. Csirik, T. Gyimothy, and M. Jelasity. Learning phonetic rules in a speech recognition system. Lecture Notes in Computer Science, 1297: 37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anthony:1997:CII

[2329] S. Anthony and A. M. Frisch. Cautious induction in inductive logic programming. Lecture Notes in Computer Science, 1297:45–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anthony:1997:GNL

[2330] S. Anthony and A. M. Frisch. Generating numerical literals during refinement. Lecture Notes in Computer Science, 1297:61–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blockeel:1997:LDI

[2331] H. Blockeel and L. De Raedt. Lookahead and discretization in ILP. Lecture Notes in Computer Science, 1297: 77–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bryant:1997:DMI

[2332] C. H. Bryant. Data mining via ILP: The application of progol to a database of enantioseparations. Lecture Notes in Computer Science, 1297:85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cussens:1997:PST

[2333] J. Cussens. Part-of-speech tagging using progol. Lecture Notes in Computer Science, 1297:93–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dehaspe:1997:MEM

[2334] L. Dehaspe. Maximum entropy modeling with clausal constraints. Lecture Notes in Computer Science, 1297:109– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dehaspe:1997:MAR

[2335] L. Dehaspe and L. De Raedt. Mining association rules in multiple relations. Lecture Notes in Computer Science, 1297:125–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeRaedt:1997:ULD

[2336] L. De Raedt and H. Blockeel. Using logical decision trees for clustering. Lecture Notes in Computer Science, 1297:133–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzeroski:1997:ISN

[2337] S. Dzeroski and T. Erjavec. Induction of slovene nominal paradigms. Lecture Notes in Computer Science, 1297:141– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flach:1997:NFI

[2338] P. A. Flach. Normal forms for inductive logic programming. Lecture Notes in Computer Science, 1297:149–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furukawa:1997:SCE

[2339] K. Furukawa, T. Murakami, K. Ueno, and T. Ozaki. On a sufficient condition for the existence of most specific hypothesis in progol. Lecture Notes in Computer Science, 1297:157–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furusawa:1997:ILP

[2340] M. Furusawa, N. Inuzuka, H. Seki, and H. Itoh. Induction of logic programs with more than one recursive clause by analyzing saturations. Lecture Notes in Computer Science, 1297:165–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geibel:1997:LFG

[2341] P. Geibel and F. Wysotzki. A logical framework for graph theoretical decision tree learning. Lecture Notes in Computer Science, 1297:173–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kakas:1997:LA

[2342] A. C. Kakas and F. Riguzzi. Learning with abduction. Lecture Notes in Computer Science, 1297:181–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1997:SPI

[2343] L. Martin and C. Vrain. Systematic predicate invention in inductive logic programming. Lecture Notes in Computer Science, 1297:189–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moyle:1997:LPE

[2344] S. Moyle and S. Muggleton. Learning programs in the event calculus. Lecture Notes in Computer Science, 1297: 205–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nienhuys-Cheng:1997:DBH

[2345] S.-H Nienhuys-Cheng. Distance between Herbrand interpretations: A measure for approximations to a target concept. Lecture Notes in Computer Science, 1297:213–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ozaki:1997:RPF

[2346] T. Ozaki, K. Furukawa, T. Murakami, and K. Ueno. Realizing progol by forward reasoning. Lecture Notes in Computer Science, 1297:227–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pompe:1997:PFO

[2347] U. Pompe and I. Kononenko. Probabilistic first-order classification. Lecture Notes in Computer Science, 1297: 235–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reddy:1997:LHD

[2348] C. Reddy and P. Tadepalli. Learning horn definitions with equivalence and membership queries. Lecture Notes in Computer Science, 1297:243–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadohara:1997:UAS

[2349] K. Sadohara and M. Haraguchi. Using abstraction schemata in inductive logic programming. Lecture Notes in Computer Science, 1297:256–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebag:1997:DIF

[2350] M. Sebag. Distance induction in first order logic. Lecture Notes in Computer Science, 1297:264–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srinivasan:1997:CPU

[2351] A. Srinivasan, R. D. King, S. H. Muggleton, and M. J. E. Sternberg. Carcinogenesis predictions using ILP. Lecture Notes in Computer Science, 1297: 273–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1997:DFO

[2352] I. Weber. Discovery of first-order regularities in a relational database using offline candidate determination. Lecture Notes in Computer Science, 1297: 288–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamamoto:1997:WHC

[2353] A. Yamamoto. Which hypotheses can be found with inverse entailment? Lecture Notes in Computer Science, 1297: 296–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alpuente:1997:SFU

[2354] M. Alpuente, M. Falaschi, G. Moreno, and G. Vidal. Safe folding/unfolding with conditional narrowing. Lecture Notes in Computer Science, 1298:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoy:1997:OND

[2355] S. Antoy. Optimal non-deterministic functional logic computations. Lecture Notes in Computer Science, 1298:16– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codish:1997:SBT

[2356] M. Codish and C. Taboch. A semantic basis for termination analysis of logic programs and its realization using symbolic norm constraints. Lecture Notes in Computer Science, 1298:31– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geser:1997:PFP

[2357] A. Geser and S. Gorlatch. Parallelizing functional programs by generalization. Lecture Notes in Computer Science, 1298:46–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirchner:1997:HOE

[2358] G. Kirchner and C. Ringeissen. Higherorder equational unification via explicit substitutions. Lecture Notes in Computer Science, 1298:61–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steggles:1997:PHO

[2359] L. J. Steggles. Parameterised higherorder algebraic specifications. Lecture Notes in Computer Science, 1298:76– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suzuki:1997:HOL

[2360] T. Suzuki, K. Nakagawa, and T. Ida. Higher-order lazy narrowing calculus: A computation model for a higherorder functional logic language. Lecture Notes in Computer Science, 1298: 99–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aoto:1997:CPT

[2361] T. Aoto and Y. Toyama. On composable properties of term rewriting systems. Lecture Notes in Computer Science, 1298:114–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucas:1997:NRC

[2362] S. Lucas. Needed reductions with context-sensitive rewriting. Lecture Notes in Computer Science, 1298:129– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohlebusch:1997:CTG

[2363] E. Ohlebusch. Conditional term graph rewriting. Lecture Notes in Computer Science, 1298:144–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Almendros-Jimenez:1997:LNP

[2364] J. M. Almendros-Jimenez and A. Gil-Luezas. Lazy narrowing with parametric order sorted types. Lecture Notes in Computer Science, 1298:159–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1997:TAT

[2365] G. Barthe and F. Van Raamsdonk.

Termination of algebraic type systems: The syntactic approach. Lecture Notes in Computer Science, 1298:174–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caires:1997:PNS

[2366] L. Caires and L. Monteiro. Proof net semantics of proof search computation. Lecture Notes in Computer Science, 1298:194–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1997:ESL

[2367] G. Barthe, F. Kamareddine, and A. Rios. Explicit substitutions for the lambdacap delta calculus. Lecture Notes in Computer Science, 1298:209– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MunozH:1997:LLV

[2368] Cesar A. Muñoz H. A left-linear variant of lambda-sigma. Lecture Notes in Computer Science, 1298:224–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khasidashvili:1997:PUN

[2369] Z. Khasidashvili and M. Ogawa. Perpetuality and uniform normalization. Lecture Notes in Computer Science, 1298:240–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdennadher:1997:MGE

[2370] S. Abdennadher and H. Schuetz. Model generation with existentially quantified variables and constraints. Lecture Notes in Computer Science, 1298:256– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nedjah:1997:OLR

[2371] N. Nedjah, C. D. Walter, and S. E. Eldridge. Optimal left-to-right patternmatching automata. Lecture Notes in Computer Science, 1298:273–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilks:1997:IEC

[2372] Y. Wilks. Information extraction as a core language technology. Lecture Notes in Computer Science, 1299:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grishman:1997:IET

[2373] R. Grishman. Information extraction: Techniques and challenges. Lecture Notes in Computer Science, 1299:10– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaizauskas:1997:CVL

[2374] R. Gaizauskas, K. Humphreys, S. Azzam, and Y. Wilks. Concepticons vs. lexicons: An architecture for multilingual information extraction. Lecture Notes in Computer Science, 1299:28– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basili:1997:LAI

[2375] R. Basili and M. T. Pazienza. Lexical acquisition for information extraction. Lecture Notes in Computer Science, 1299:44–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boguraev:1997:TTD

[2376] B. Boguraev and C. Kennedy. Technical terminology for domain specification and content characterisation. Lecture Notes in Computer Science, 1299: 73–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grefenstette:1997:SQL

[2377] G. Grefenstette. Short query linguistic expansion techniques: Palliating oneword queries by providing intermediate structures to text. Lecture Notes in Computer Science, 1299:97–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smeaton:1997:IRS

[2378] A. F. Smeaton. Information retrieval:

Still butting heads with natural language processing ? Lecture Notes in Computer Science, 1299:115–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guarino:1997:SMF

[2379] N. Guarino. Semantic matching: Formal ontological distinctions for information organization, extraction, and integration. Lecture Notes in Computer Science, 1299:139–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neri:1997:MLI

[2380] F. Neri and L. Saitta. Machine learning for information extraction. Lecture Notes in Computer Science, 1299:171– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cluet:1997:MQS

[2381] S. Cluet. Modeling and querying semistructured data. Lecture Notes in Computer Science, 1299:192–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlechta:1997:NLB

[2382] Karl Schlechta. Nonmonotonic logics: basic concepts, results, and techniques, volume 1187 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-62482-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 243 pp. LCCN Q339 .S35 1997.

Sipper:1997:EPC

[2383] Moshe Sipper. Evolution of parallel cellular machines: the cellular programming approach, volume 1194 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997.

CODEN LNCSD9. ISBN 3-540-62613-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 198 pp. LCCN QA76.58.S575 1997.

Nienhuys-Cheng:1997:FIL

[2384] S.-H. (Shan-Hwei) Nienhuys-Cheng and R. Wolf. Foundations of inductive logic programming, volume 1228 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-62927-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xvii + 404 pp. LCCN QA76.63 .N54 1997.

Kraetzschmar:1997:DRM

[2385] Gerhard Kraetzschmar. Distributed reason maintenance for multiagent systems, volume 1229 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63606-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 296 pp. LCCN QA76.76.I58 K73 1997.

Taft:1997:ARM

[2386] S. Tucker Taft and Robert A. Duff.

Ada 95 reference manual: language and standard libraries, volume 1246 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63144-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 526 pp. LCCN QA76.73.A35 A22 1997. International standard ISO/IEC 8652:1995(E).

Mycielski:1997:SLC

[2387] Jan Mycielski, Grzegorz Rozenberg, and Arto Salomaa. Structures in logic and computer science: a selection of essays, volume 1261 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63246-8. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 370 pp. LCCN QA76.S844 1997.